A Complete Bibliography of Publications in Biometrics: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

19 May 2021 Version 1.03

Title word cross-reference

2 × 2 [267, 484]. $69.95/£ [530, 524]. $79.95/£ [523, 529]. $84.95/£ [528]. $89.95/Euro [526, 522]. $94.95/£ [527]. ^{2} [855]. A [639]. D [184, 459]. F [1429]. H [1458]. K [225, 1552]. m [1341]. N [675]. P [1316, 488, 904, 1343, 83, 1195, 1344]. R × C [1655]. × [900, 1451, 1743, 725, 1685]. U [1044, 1638].

-Component [1552]. -Estimates [1341]. -Function [225]. -likelihood [1458]. -Mixture [675]. -Optimal [184, 459, 639]. -Statistics [1638, 1044]. -Studies [855]. -Test [1429]. -Value [83, 1316, 488]. -Values [1195, 1344, 904, 1343].

0 [530, 523, 525, 532, 522, 529, 527, 524, 528]. 0-19-925231-9 [531]. 0-387-95350-7 [524]. 0-387-95353-1 [523]. 0-387-95370-1 [529]. 0-387-95435-X [530]. 0-471-34772-8 [527]. 0-471-39685-0 [528].

0-521-66578-7 [532]. 0-521-78165-5 [525]. 0-8176-4209-9 [522].

1 [1141, 523, 42, 384, 1375, 1374, 529, 517, 1248]. 1-58488-469-X [1141]. 1-59047-080-X [1248]. 1997. [559]. 1999 [521].

2001 [560]. 2003 [561]. 2004 [1360]. 2005 [1012, 1916, 1571]. 2006 [1258]. 20th [820]. 22.99/$34.95 [531]. 24-Hours [620]. 2nd [1267, 1415, 1269, 1320, 1473, 1266, 1317, 1318, 1704, 1812, 1759, 1760, 1915, 1405, 1310, 1410, 1699, 1698, 1261, 1404, 1758, 1218, 1220, 1223, 1413].

3 [526]. 3-7643-6514-5 [526]. 32.95/$48.00 [532]. 3rd [1584, 1816, 1911, 1815, 1749, 1260].

4th [1268, 1471, 1702, 1811].

5 [526, 525]. 53 [559]. 55 [521]. 55.00/$80.00 [525]. 57 [560]. 59 [561].

60 [1360]. 60th [1272]. 61 [1012, 1916, 1571]. 62 [1258]. 63 [1744].

7 [532, 524].

8 [527].

9 [522].

A. [1584, 1643, 1585, 1701, 1704, 1868, 1760, 1913, 1316, 1905, 1903, 1746, 1752, 1754, 1907]. Aalen [560, 1803, 1497, 634]. Abalone [258]. Aberrant [603, 1550]. Ability [1357, 1617]. Abnormality [998]. Abrams [1078]. Absence [705, 911]. Absolute [1788, 559, 2]. Abstract [928]. Abundance [1236, 436, 435, 1098, 676, 1005, 1569, 1151, 724, 1734, 564, 568, 1641, 627, 6, 1836]. Abundance-Based [1098]. Accelerated [1769, 309, 1303, 999, 1167, 1723]. Accidents [64]. Accommodating [1099]. Account [719]. Accounting [708, 1342, 1161, 515, 9, 288, 571]. Accumulation [107, 1375, 890, 1877]. Accuracy [1399, 1124, 623, 837, 567, 1479, 36, 88, 1664, 177, 1886, 857, 1069, 1327]. Accurate [1199, 448, 481]. Acid [1501, 281]. Acknowledgement [1477, 1474]. Acknowledgements [195, 367, 496]. Across [1682, 1109]. Activation [1875]. Active [35]. Activity [1383, 1782, 177]. Actually [508]. Acute [1614]. Adaptive [839, 1373, 1408, 1785, 72, 344, 1841, 884, 715, 217, 313, 472, 1726, 1605, 316, 438, 1209, 883, 1119, 525, 1522, 1808]. Additional [271, 1843]. Additive [332, 424, 268, 1774, 590, 1371, 1049, 1060, 1497, 577, 1460, 750, 21, 1180, 1187, 560]. Additive-Risk [590]. Address [1321]. Addressing [239, 242]. Adequacy [309, 1448]. Adjust [56, 434]. Adjusted [1334, 311, 246, 419, 26, 859, 1831, 1600, 1287, 310, 1846]. Adjusting [541, 66, 902, 218]. Adjustment [332, 959, 488, 305, 203, 1661, 1657, 1532]. Adjustments [1130, 83, 356, 349, 560, 1079]. Adjuvant [1849]. Administrative [1304]. Adults [1662]. Advanced [1874, 1236, 1075, 1248, 1641]. Advances [1912, 1696, 1575, 1813]. Advantages [313]. Affect [62, 183]. Affecting [298]. AFLP [1196]. after [588]. Against [1775, 224, 351]. Age [959, 1876, 1882, 427, 1068, 1668, 221, 249, 385, 339, 626, 76, 606, 419, 908, 10, 1001]. Age-Adjusted [419]. Age-at-Harvest [1668]. Age-Dependent [1882, 76, 1001]. Age-Specific [959, 339, 626, 10]. Age-Structured [606]. Ageing [1525]. Agent [880, 753, 1515]. Agents [681, 565, 1840]. Aggregate [476, 121]. Aggregation [1040]. Aging [40]. Agreement [275, 490, 1070, 1722, 1906]. Agresti [1473]. Agriculture [1220]. AIDS [1891, 372, 361, 112]. Air [1863, 1452, 1739]. Akaike [210]. Akushevich [1869]. al [492, 1919, 1920]. Alaska [627]. Albert [961]. Alcoholism [117]. Algorithm [1866, 895, 324, 1787, 685]. Algorithms [1129, 1578, 1041, 1441, 1844, 1374]. Alho [1221]. All-or-None [1796]. All-or-Nothing [508]. Allele [702, 703, 296]. Allele-Frequency [702]. Allele-Sharing [703, 296]. Alleles [299]. Allelic [158, 1278]. Allelic-Loss [158]. Allowing [291, 1353, 554]. Alterations [1454]. Alternating [131]. Alternative [267, 1729, 856]. Alternatives [1429, 224]. Always [534]. Alzheimer [1776]. Amino [1501, 281]. Among [490, 275, 1916, 833, 122, 549, 862, 580, 832]. Amstrup [1222]. Analyses [1225, 1227, 1268, 1911, 1130, 181, 1121, 1097, 556, 1849, 1374, 622, 696, 1095, 1396, 498, 501]. analysi [1920]. Analysing [1517]. Analysis [765, 208, 937, 1090, 1141, 1317, 1318, 1644, 1812, 1816, 1814, 1914, 1238, 241, 1486, 410, 64, 503, 1634, 291, 1458, 1730, 1609, 141, 1291, 649, 801, 1039, 1435, 1498, 1768, 1160, 107, 1366, 791, 1668, 772, 1450, 58, 56, 1351, 214, 793, 1301, 128, 66, 819, 1692, 279, 372, 28, 1790, 246, 407, 143, 1771, 773, 707, 1456, 78, 617, 1195, 570, 137, 1603, 165, 1756, 1412, 1803, 507, 1533, 277, 1282, 1465, 687, 160, 1568, 1864, 1343, 299, 620, 1799, 200, 1854, 492, 1191, 943, 829, 1630, 1659, 1577, 1295, 1479, 1502, 907, 582, 1447, 1625]. Analysis [1758, 587, 1695, 1260, 1107, 1061, 1723, 678, 147, 1652, 720, 1064, 1184, 1792, 1136, 672, 724, 749, 1452, 1776, 1767, 345, 606, 832, 1330, 1623, 1592, 462, 1200, 1118, 1716, 798, 948, 906, 1312, 1861, 1549, 1385, 1593, 26, 1185, 706, 452, 1111, 1601, 859, 538, 941, 166, 361, 1873, 529, 722, 1377, 916, 169, 426, 1378, 513, 1137, 411, 168, 1793, 1566, 1341, 951, 1737, 1344, 1830, 577, 824, 1073, 1413, 281, 1831, 1909, 432, 1367, 666, 1490, 1782, 1640, 1715, 588, 1264, 750, 1430, 1487, 897, 94, 1754, 1168, 1835, 1802, 1570, 197, 1162, 1259, 113, 70, 888]. Analysis [934, 1193, 1379, 1368, 1483, 1101, 416, 110, 1696, 917, 428, 408, 207, 1283, 1481, 1779, 1827, 117, 1198, 1212, 1277, 1558, 341, 1380, 1681, 413, 623, 504, 1919, 1085, 1473, 1271, 1868, 1521, 1310, 1810, 1910, 1747, 1075, 1222, 1752, 1865, 1584, 1757, 1806, 1468]. Analyte [273]. Analytic [760, 236, 1610]. Analyze [849, 1848]. Analyzing [1008, 699, 892, 258, 766, 1670, 580, 692, 939, 1391, 349, 751, 1009, 159, 415]. and/or [487]. Andersen [1271]. Anderson [1236, 1701, 1641]. Andres [1909]. Angular [1346]. Animal [436, 292, 83, 1569, 1151, 1625, 1789, 347, 996, 842, 1607]. Animals [568, 6, 1636]. Ankerst [1223]. Anniversary [1272]. Annual [1171, 35]. ANOVA [523, 1727, 609]. Answer [362]. Answer-Until-Correct [362]. Antibody [212]. Antiretroviral [1532]. Antiviral [42]. Applicability [491, 492]. Application [1741, 603, 424, 1486, 1173, 1053, 1045, 1874, 1882, 152, 1736, 1891, 372, 28, 896, 219, 60, 105, 1159, 665, 1568, 998, 1454, 1455, 1871, 1792, 1621, 1776, 580, 1449, 1677, 1826, 1623, 1599, 369, 1339, 1514, 723, 480, 361, 67, 1662, 893, 1378, 182, 1729, 680, 1831, 307, 1095, 1367, 1782, 1117, 252, 1392, 725, 260, 638, 1379, 1488, 1545, 748, 416, 1532, 585, 905, 942, 610, 1212, 1069]. Applications [278, 306, 1470, 175, 615, 1821, 1730, 586, 1342, 1195, 1910, 1753, 1371, 1758, 1853, 1614, 1618, 1710, 1666, 632, 1604, 822, 1873, 1115, 347, 1423, 634, 758, 1204, 1440, 932, 1249, 1277, 1846, 1267, 1761, 1811, 1578, 1407, 1410, 1577, 1083, 1459, 1076]. Applied [1638, 116, 226, 555, 1246, 824, 1528, 1812, 1810, 1261, 1799, 1583]. Applying [861]. Approach [332, 760, 785, 744, 489, 881, 918, 739, 919, 308, 541, 1876, 1798, 1669, 300, 228, 1882, 1160, 1543, 912, 569, 1425, 1097, 8, 1672, 741, 1854, 1188, 1659, 1885, 1742, 555, 548, 1447, 1414, 1049, 442, 740, 237, 1297, 1339, 256, 1671, 458, 1434, 29, 361, 421, 182, 1210, 640, 1661, 666, 460, 936, 742, 1157, 673, 88, 1059, 743, 828, 197, 1302, 894, 1368, 1546, 1300, 54, 408, 380, 550, 1676, 1840, 1194, 1293, 560, 528, 1318, 1869, 1406, 1466, 1219, 1801]. Approaches [1078, 278, 1399, 1356, 586, 215, 226, 1232, 313, 1111, 1557, 1681]. Approximate [1766, 252, 448, 31]. Approximating [130]. Approximation [1917, 1199, 1918, 944, 1887]. Approximations [350, 1426, 1208]. Arancio [1258]. Arbitrarily [1558]. Archival [1671]. Arctic [199]. Area [1735, 1068, 1062, 1125]. Areal [990]. Areas [1912]. Arise [1156]. Arising [980, 981, 982, 983, 984, 985, 979, 986, 987, 172]. Arithmetic [1572]. Arm [1533, 715, 798]. Arm-in-Cage [1533]. Armed [1622]. Arms [314]. Arnason [502]. Arnason-Schwarz [502]. Arnold [1813]. Array [1902, 766, 1844, 1377]. Array-Based [1844]. Arrays [1381, 1279, 571, 897]. Article [240]. Articles [927, 975, 1034, 1091, 1142, 1229, 1255, 1273, 1322, 1419, 1475]. Ascertainment [22]. Aspects [773, 896, 456]. Assays [278, 1554, 707, 360]. Assemblages [1877]. Assess [1202]. Assessing [278, 275, 711, 708, 946, 1195, 1783, 1050, 1279, 1894, 359, 1295, 1183, 1540, 1900, 1357, 1651, 993, 725, 1448, 1546, 857]. Assessment [297, 1130, 1044, 4, 1070, 1736, 1897, 623, 570, 1568, 1479, 515, 859, 67, 74, 1640, 172, 942]. Associated [912, 1434, 691, 1480, 1113]. Association [1123, 300, 1039, 1435, 1336, 1768, 129, 103, 608, 198, 267, 1172, 299, 1901, 1499, 832, 386, 745, 1834, 1835, 1547, 847]. Associations [1798, 1764]. Assumed [1659]. Assuming [1171]. Assumptions [1570]. Asymptomatic [994]. Asymptotic [1683, 1643]. Asymptotically [1848]. Asymptotics [175, 1583]. Asynchronous [1792]. Atkinson [1082]. Attributable [952, 219, 893, 855]. Attrition [1370]. AUC [578]. Augmentation [182, 1737]. Augmented [1202, 251]. Aureus [1612]. Author [593, 138, 776, 1235, 726]. Authors [140, 365, 1686, 1401, 188, 326, 328, 400, 1309, 1233, 1403, 1127, 1688, 1684, 1627, 1629, 808, 1745, 1307]. Autocorrelation [1487]. Autoregressive [446, 1654, 545, 548, 321]. Auxiliary [271, 372, 1791, 200, 1890, 1201, 471, 1596, 413]. Average [1780, 1171, 1282]. Averaging [556, 1805]. Axon [135]. Axon-Scaled [135].

B [559, 521, 1012, 1225, 1141, 1085, 1268, 1318, 1416, 1813, 1868, 1911, 1405, 1310, 1221, 1809, 1695, 1260, 1526, 1222, 834]. B-Spline [834]. B. [1915, 1217, 1800, 1858, 1691, 1751]. Background [285]. Baddeley [1217]. Bailey [1903]. Baker [1916]. Balaena [469]. Balakrishnan [1813, 522]. Balancing [587]. Balding [1639]. Band [791, 232]. Band-Recovery [791]. Bands [636]. Barbu [1910]. Barker [1270]. Barnett [1815]. Basal [1447]. Base [255]. Based [297, 1065, 937, 1701, 1559, 615, 1098, 676, 1039, 1435, 1882, 453, 161, 935, 679, 1429, 23, 1005, 128, 66, 279, 717, 1354, 384, 139, 304, 1854, 1885, 567, 1841, 575, 1844, 852, 537, 546, 245, 368, 1166, 1666, 1592, 1503, 1549, 144, 1439, 398, 1893, 628, 171, 1720, 673, 1287, 317, 756, 1663, 828, 1180, 1835, 1547, 177, 16, 1545, 298, 1300, 1620, 408, 1889, 1532, 415, 1443, 1198, 1764, 318, 1873, 1574, 1582]. Basel [526]. Baseline [44, 1661, 383]. Baseline-Risk [44]. Basford [94]. Basic [1267]. Basis [385]. Batch [1372]. Bayes [52, 723, 1851, 1593, 1542, 1766, 932, 1347, 1277]. Bayes-Type [1593]. Bayesian [744, 1585, 1645, 918, 919, 1079, 741, 1747, 1414, 740, 821, 1466, 1579, 742, 743, 528, 900, 839, 1486, 81, 881, 1373, 263, 1667, 1730, 739, 1609, 141, 586, 535, 1850, 1106, 959, 1876, 212, 1798, 98, 1874, 1345, 1721, 1543, 405, 912, 1668, 1884, 1829, 215, 705, 213, 1616, 214, 340, 467, 377, 569, 621, 542, 543, 506, 755, 840, 1154, 1425, 477, 853, 831, 134, 1880, 707, 8, 944, 505, 270, 1036, 1548, 618, 507, 576, 665, 997, 1284, 339, 626, 1332, 1674, 943, 1103, 247, 548, 582, 1447, 1384, 1061, 1608, 5, 442, 147, 237, 1037, 1896]. Bayesian [504, 1852, 544, 787, 148, 1851, 564, 1549, 1385, 32, 1305, 706, 762, 668, 1189, 265, 426, 182, 1246, 1210, 854, 1830, 805, 771, 666, 667, 669, 1895, 1395, 1487, 1112, 897, 830, 1438, 31, 1485, 1879, 37, 1739, 110, 1001, 1389, 1675, 54, 1875, 905, 850, 887, 1163, 1293, 910, 639, 1114, 1515, 1078, 1415, 1868, 1905, 1748, 1906, 1414, 1800, 1311, 1576]. Bayesian-Frequentist [1395]. Be [385, 697]. Behavior [1353, 680, 1607]. Behavioral [1128, 62, 1461, 996]. Behrens [902]. Beijing [1614]. Below [199]. Below-Threshold [199]. Benchmark [289, 859]. Benefit [1843]. Benefits [280]. Berk [1907]. Berkson [369]. Bernoulli [1392]. Bertoin [1690]. Best [126, 702, 1440]. Beta [122, 139, 1083, 19, 625]. Beta-Binomial [122, 139]. Beta-Binomial-Poisson [625]. Beta-Logistic [19]. Betensky [1917]. Better [1843]. Between [374, 1798, 475, 336, 221, 215, 392, 384, 1501, 299, 1333, 16, 264, 690, 1007, 1172, 838, 1051, 1159, 1295, 1871, 558, 1593, 625]. between- [1159]. between-Study [1295]. Beyond [1695]. Bi [858]. Bi-Factorial [858]. Biallelic [1197, 757]. Bias [1164, 1130, 950, 959, 679, 66, 198, 1343, 1295, 515, 345, 1397, 1898, 764, 906, 18, 1593, 951, 846, 1796, 1595, 934, 114, 857, 1919, 1920]. Bias-Corrected [1397, 846]. Biased [839, 387, 814, 1201]. Biased-Coin [839]. Bidirectional [794]. Binary [85, 440, 446, 521, 603, 1550, 64, 486, 1500, 956, 772, 320, 793, 103, 1007, 457, 198, 102, 955, 131, 687, 998, 116, 259, 1765, 1823, 555, 1806, 672, 1000, 50, 1597, 860, 1514, 82, 1185, 452, 1391, 166, 1003, 1299, 358, 426, 347, 3, 411, 1566, 316, 1326, 499, 383, 1168, 789, 992, 416, 1056, 770]. Binding [485, 281]. Binomial [1680, 1358, 12, 122, 139, 151, 30, 558, 261, 584, 224, 846, 21, 260, 625]. Bio [827]. Bio-X [827]. Bioassay [182]. Bioassays [175]. Bioavailability [1816]. Bioconductor [1214]. Bioequivalence [1816, 1243, 1407]. Bioinformatics [1228, 1129, 1346, 827, 1703, 1527, 1750, 1696, 1214]. Biologic [1505]. Biological [1236, 1155, 652, 456, 1641]. Biologists [1587]. Biology [1214, 526, 1144, 252, 526, 1260, 1260, 1262]. Biomarker [440, 1554, 881, 1544]. Biomarkers [834, 1290, 1617, 1069]. Biomedical [1762, 1696]. Biometric [1530, 1588, 1647, 1705, 1531, 1589, 1648, 1706, 1325]. Biometrical [96, 1513]. Biometrics [48, 1744, 827, 95, 187, 189, 325, 327, 329, 399, 807, 1402, 1687, 1126, 774, 775, 454, 1072, 592]. Biometry [1709, 827, 1086]. Bioprocessing [1094]. Biosciences [1637]. Biostatistics [1215, 1748, 827, 1760]. Biphasic [42]. Biplot [916, 917]. Biplots [720]. Bird [959, 626]. Birds [362]. Birkhäuser [526, 522]. Births [1738]. Bishop [1639]. Biswas [1696]. Bivand [1799]. Bivariate [1123, 300, 651, 1660, 1495, 1159, 613, 886, 630, 1853, 1346, 1185, 166, 1663, 650, 942]. Bladder [1454]. Blanke [1693]. Block [56, 551, 1516, 393, 1391]. Blocking [137]. Blood [620]. Boca [1141]. Boelen [1905]. Böhning [1806]. Bolker [1809]. Bolstad [1645]. Bone [1599]. Book [965, 1078, 1014, 1236, 46, 93, 495, 972, 973, 974, 1228, 1088, 1089, 1086, 1225, 1227, 1087, 1090, 1141, 1226, 1085, 1267, 1415, 1268, 1470, 1471, 1269, 1320, 1473, 1319, 1266, 1317, 1472, 1270, 1417, 1271, 1318, 1416, 1528, 1584, 1643, 1646, 1642, 1644, 1585, 1703, 1527, 1586, 1645, 1587, 1701, 1702, 1704, 1529, 1757, 1812, 1813, 1868, 1759, 1762, 1869, 1816, 1866, 1911, 1815, 1760, 1761, 1915, 1870, 1867, 1912, 1913, 1811, 1814, 1914, 864, 1213, 1128, 1523, 1693, 1572, 1405, 530, 925, 777, 865, 1639]. Book [1574, 1573, 1634, 594, 1521, 1074, 1021, 778, 1019, 1080, 1316, 866, 728, 596, 1310, 1408, 867, 1409, 1458, 523, 1519, 868, 1863, 1463, 1077, 1079, 1905, 1904, 1755, 809, 1578, 1580, 1407, 1862, 969, 1018, 1810, 654, 1749, 1081, 1697, 1016, 1632, 729, 811, 924, 1214, 730, 1518, 1859, 1082, 526, 922, 1215, 1692, 657, 1221, 779, 1633, 1631, 1809, 1410, 1024, 780, 1638, 1522, 963, 1699, 1216, 1855, 966, 1698, 1903, 1756, 1412, 1803, 1910, 1690, 1516, 1525, 1261, 967, 1748, 1906, 1465, 781, 964, 1524, 1023, 1020, 595, 1864, 1747, 782, 1753, 1799, 1520, 731, 1217, 1854]. Book [732, 1025, 1406, 1808, 1630, 1577, 1404, 1758, 1414, 1695, 1806, 733, 1260, 1265, 869, 1083, 659, 812, 1800, 783, 597, 1459, 1637, 1750, 1636, 1218, 1526, 1805, 962, 821, 1075, 1411, 598, 525, 1312, 1466, 1469, 1467, 1861, 1807, 1022, 656, 1313, 532, 1860, 1314, 522, 1464, 735, 810, 870, 1575, 1311, 529, 871, 1219, 1462, 1801, 734, 872, 1220, 1076, 1576, 1858, 142, 193, 238, 283, 331, 366, 401, 429, 455, 562, 599, 1015, 531, 1746, 527, 1579, 1222, 1223, 970, 1248, 1073, 1635, 1413, 1517, 1641, 1857, 1909, 968, 1752, 923, 1640, 1856, 1460, 524, 873, 1581, 1468]. Book [1264, 1582, 1315, 1754, 921, 1689, 658, 1700, 1908, 1694, 1802, 1013, 1865, 736, 920, 874, 1259, 1583, 655, 1461, 1224, 1804, 737, 1262, 528, 1017, 1696, 875, 1907, 1263, 1691, 1751, 876, 877]. Boosted [1355]. Boosting [1887, 1180, 685]. Bootstrap [297, 72, 464, 914, 1260, 516, 915, 814, 227, 1646]. Borchers [1236, 1641]. Borgan [560, 559, 1803]. Bosc [1693]. Boston [522]. Both [23, 761]. Bound [679]. Boundaries [605, 1622]. Bounded [1838]. Bounds [1785, 1594, 680]. Bové [1918]. Bowhead [627, 469]. Brain [1846]. Branching [1213]. Braun [1413]. Brazzale [1583]. BRCA1 [1882]. Breast [1916, 833, 615, 1792, 1339, 1675]. Breastfeeding [1445]. Breastmilk [517]. Breeding [889, 347, 10, 94]. Bregman [1012, 1011, 1010]. Breslow [450]. Bridge [1662]. Bridging [1395]. Brief [533, 600, 660, 738, 784, 813, 825, 878, 926, 971, 1026, 1027, 1028, 1029, 1030, 1031, 1032, 1084, 1250, 1140]. Briggs [1465]. Broemeling [1748, 1906]. Brown [1261]. Buck [1632]. Buckeridge [1801]. Buckland [1236, 1641]. Buckley [1493]. Buckley-James [1493]. building [1854]. Bulmer [1086]. Buncher [1749]. Burgess [1642]. Burnham [1236, 1641].

C [521, 1266, 1528, 1529, 1757, 1813, 1816, 1870, 523, 1463, 1862, 1749, 1632, 1215, 1410, 1522, 1698, 1525, 1414, 1218, 1469, 532, 522, 529, 1219, 1729, 1222, 1073, 1641, 1752, 1264, 1583]. C. [1870, 1912, 1639]. C.-D [1525]. Cage [1533]. Cai [1917]. Calculating [1232, 1010, 1011, 1796, 1012]. Calculation [289, 803, 1457, 769]. Calculations [1698, 681, 1503, 26, 173]. Calibration [152, 273, 1676]. Cambridge [1903, 525, 532]. Campbell [1702]. Campolongo [1909]. Canal [905]. Cancer [1916, 157, 833, 615, 586, 1874, 28, 1494, 1849, 1454, 431, 1792, 1682, 715, 248, 288, 805, 307, 994, 1480, 1389, 1675, 1228, 1869]. Candidate [1882, 1665]. Cannings [1639]. Cannot [385]. Canonical [785, 1776]. Can’t [1572, 128]. Capture [291, 1353, 1538, 62, 427, 931, 549, 958, 1425, 1152, 1234, 1151, 643, 548, 346, 1670, 290, 644, 832, 1786, 63, 607, 957, 502, 891, 1737, 1222, 10, 179, 1300, 1506, 1394, 996, 1004, 652]. capture-recapture [652]. CAR [990]. Carcinogenesis [603, 1486, 478]. Carcinogenicity [83, 477]. Care [1078]. Carey [1214]. Cariboni [1909]. Carlin [1868]. Carlo [1699, 1260, 1041, 1710, 1789, 1249]. Carol [522]. Carolina [1248]. Carrier [1831]. Carroll [1266]. Cary [1248]. Cascades [1624]. Case [561, 1042, 1863, 803, 1457, 212, 1040, 1039, 1435, 1386, 1882, 1721, 1884, 1351, 28, 1672, 277, 151, 1288, 795, 1885, 510, 1883, 648, 1232, 1441, 1723, 1203, 1677, 345, 1592, 462, 1200, 493, 767, 203, 1385, 1593, 1111, 1010, 696, 1011, 993, 909, 1219, 689, 855, 432, 667, 1490, 1287, 260, 1162, 1583, 751, 757, 942, 1198, 559, 1012, 2, 821]. Case-Cohort [803, 1457, 1232, 1723, 1677, 1200, 993]. Case-Control [1042, 212, 1040, 1039, 1435, 1386, 1882, 1884, 1351, 28, 1672, 277, 1288, 795, 1885, 510, 1883, 648, 345, 1592, 462, 493, 203, 1385, 1593, 1111, 1010, 696, 1011, 909, 432, 667, 1490, 1287, 260, 751, 1198, 561, 559, 1012, 2]. Case-Only [1040, 689, 1162]. Casella [1694]. Cases [25]. Catch [234, 257]. Categorical [1473, 690, 1874, 618, 412, 272, 1049, 500, 1311, 364, 55, 1308, 1681, 1073]. Categorizing [718]. Causal [1472, 357, 991, 1571, 336, 1817, 1719, 1780, 370, 570, 1177, 1784, 514, 1096, 1797, 508, 746, 1246, 689, 1818, 1396]. Cause [355, 1657, 40]. Cause-Specific [1657]. Caused [764]. Causes [208, 1450]. Cautionary [709, 772, 558]. Caveats [104]. Cay [6]. CD4 [105]. cDNA [766]. Cell [105, 849, 480, 317]. Censored [937, 357, 424, 693, 447, 374, 574, 309, 378, 1683, 286, 935, 1350, 1660, 23, 945, 755, 68, 1724, 1495, 463, 165, 136, 1412, 49, 464, 829, 697, 1107, 1136, 500, 111, 245, 144, 244, 423, 206, 24, 29, 361, 375, 73, 1565, 750, 1168, 37, 1120, 54, 312, 1558, 333]. Censoring [208, 845, 1725, 465, 566, 236, 377, 407, 239, 242, 617, 1292, 433, 1659, 204, 1331, 907, 1330, 798, 112, 250, 758, 444, 1212]. Census [436]. Center [989]. Center-Specific [989]. Century [96, 820]. Cerebellum [1379]. Cerioli [1082]. Certificates [168]. Cessation [1852]. CGH [1902, 1377]. Chain [567, 147, 255, 996, 842]. Chains [332, 1910, 560]. Challenge [1894]. Challenges [1088]. Chambers [1914]. Chan [1758]. Chance [1245]. Chang [1808, 1522]. Change [1171, 385, 90, 1484, 450]. Changepoint [131, 1066]. Changes [692, 1542]. Chao [1423]. Chaoqun [1578]. Character [385, 298]. Characteristic [1062, 164]. Characteristics [161]. Characterization [1349]. Characterizing [384, 105, 1375, 1352]. Charge [162]. Chaudhary [1584]. Checking [1291, 223, 835, 368, 1605]. Checks [1081]. Chemical [318]. Chemoinformatics [1301]. Chemoprevention [155]. Chemotherapy [1874]. Chernick [1646]. Cheung [1318]. Chevret [1464]. Chi [475, 58]. Chi-square [58]. Chi-Squared [475]. Chiasmata [415]. Child [1367]. Childhood [219]. Children [1662]. ChIP [1880, 1548, 1380]. ChIP-Chip [1880, 1548, 1380]. Choice [1642, 1795, 516]. Choose [1757]. Choosing [575]. Chop [1842]. Chop-Lump [1842]. Chow [1816, 1522, 1698]. Chromosomal [491, 254, 492]. Chronic [25, 89, 848, 116]. Chuang [1463]. Chuang-Stein [1463]. Circadian [1008, 359, 609, 1009]. Circular [1619, 717, 1354]. Circular-Circular [1354]. Circular-Linear [1619, 1354]. Claeskens [1805]. Class [278, 709, 1353, 1489, 1902, 519, 154, 1191, 1431, 1781, 612, 1555, 1287, 1886, 380, 850, 887, 520]. Classes [672, 1357]. Classical [369, 388, 313]. Classification [384, 629, 1855, 30, 1479, 1637, 1666, 1376, 82, 1062, 1726]. Classified [1068]. Clinical [1078, 785, 1470, 1269, 1320, 839, 699, 1408, 337, 1821, 1500, 1850, 1580, 170, 378, 1622, 488, 1697, 1859, 1387, 372, 143, 1522, 86, 1698, 276, 576, 1243, 314, 1406, 1338, 1841, 678, 20, 715, 1797, 91, 692, 373, 882, 313, 1662, 1219, 1462, 112, 854, 1223, 324, 554, 1718, 861, 1162, 677, 1696, 883, 1163, 1119, 1596].

Clinicians [1317, 1317]. Clipped [1788]. Clonal [849]. Clonality [1348, 1651]. Clones [1156]. Closed [350, 549, 958, 1608, 1786, 63, 957, 1423, 841]. Closed-Form [350]. Closure [858]. Cluster [956, 72, 344, 1683, 1366, 1350, 793, 569, 28, 955, 1159, 1678, 87, 274, 1299, 411, 1108, 499, 92, 501, 1443, 1764, 1904, 1637]. Cluster-Correlated [92]. Cluster-Randomized [87]. Clustered [1768, 1160, 710, 1551, 236, 556, 1603, 1772, 500, 462, 692, 1200, 939, 452, 1601, 1003, 411, 1211, 1058, 992, 1483, 501]. Clustering [1491, 1578, 134, 998, 254, 1763, 1822, 1377, 1115, 640, 1367, 1287, 828, 1545, 1620, 1194]. Clusters [5, 252, 1872, 1427, 430, 1908]. Co [661, 826]. Co-Editors [661, 826]. Coagulation [1690]. Coarsened [1511]. Coarsening [1212]. Codominant [115]. Coefficient [490, 614, 1292, 1722, 1161, 786, 1371, 1652, 1452, 82, 908, 1100, 16, 770]. Coefficients [133, 1159, 100, 1553, 355, 1557, 663]. Cognition [1245]. Cognitive [1066, 1186]. Cohort [803, 1457, 228, 1039, 1770, 1232, 1723, 580, 1677, 462, 1200, 993]. Coin [839, 387]. Collapsing [1727]. Collision [1196, 1156]. Colon [603, 1486]. Colorectal [1916, 833, 1874]. Column [129, 1048]. Combination [1725, 831, 1338, 38]. Combinations [157, 1840]. Combine [1736]. Combined [1199]. Combining [940, 59, 77, 249, 1288, 1849, 1375, 1374, 232, 1376, 449, 313, 1062, 222, 725, 380, 1544]. Comment [518, 1917, 1918, 1744, 1509, 1919, 186, 1920]. Comments [956, 955]. Common [1043, 688, 1533, 537, 790]. Community [1669, 416, 806]. Community-Intervention [806]. Companion [1085]. Comparability [802]. Comparative [315, 1903, 1841, 1661, 645, 1277]. Compare [1853, 164]. Compared [716]. Comparing [900, 1680, 1042, 489, 1301, 838, 1501, 474, 1682, 1900, 1599, 1877, 1514, 1095, 1664, 1059, 1617]. Comparison [135, 1500, 308, 27, 566, 42, 721, 829, 1441, 580, 1797, 1397, 1786, 305, 1176, 1851, 1439, 1651, 1893, 588, 856]. Comparisons [940, 862, 50, 1296, 244, 1211, 1058, 1729, 1440]. Compartment [269]. Compatibility [1299, 757]. Competing [208, 1498, 1768, 637, 536, 204, 852, 355, 305, 939, 554, 1168]. Competition [56]. Complete [1124, 56, 380]. Completed [835]. Completed-Data [835]. Completely [1817, 1818]. Complex [556, 952, 1670, 1881, 118, 702, 552]. Compliance [374, 315, 1781, 508]. Complicated [1298]. Component [916, 893, 110, 917, 1552, 585]. Components [818, 614, 844, 1398, 76, 539]. Composite [378]. Composition [255]. Compound [893, 1108]. Compton [1634]. Computational [1214, 1577, 1144, 1414, 1010, 1575, 1011, 1012, 1759, 1762, 1264]. Computations [913]. Computer [1260]. Computer-Intensive [1260]. Concentration [582, 1447]. Concentrations [199]. Concept [1176]. Conception [840]. Concepts [1410, 532, 861, 1074]. Concordance [319, 490, 614, 1292]. Condition [404]. Conditional [291, 391, 759, 1435, 1386, 710, 215, 1654, 1603, 662, 1610, 148, 1385, 73, 55, 1095, 15, 579, 885].

Conditional-Cumulant-of-Exposure [15]. Conditionally [545, 394]. Conditioning [1743]. Conditions [1070, 848, 385]. Conducting [208]. Confidence [323, 900, 1680, 636, 1785, 792, 72, 1343, 575, 1615, 421, 1378, 1442, 1341, 627, 725, 260, 1120, 162, 770]. Confounded [1594]. Confounders [1780]. Confounding [541, 1161, 383, 1570, 1595]. Congdon [1311, 1415]. Conjugate [285, 507]. Conjunction [1673]. Connections [104]. Consequences [127]. Conservative [1439]. Considerations [1244, 611, 395, 727]. Consortium [1883]. Constant [1171, 1614]. Constants [448]. Constrained [605, 185, 542, 1650, 1381]. Constraints [354, 745]. Construct [1757]. Constructed [1354]. Construction [1642, 680, 1048]. Consultant [929, 976, 1092, 1143, 1230, 1256, 1274, 1323, 1420, 1476, 1590]. Consumed [1871]. Contact [748]. Contain [227]. Containing [162]. Contaminated [1830]. Contingency [900, 129, 77, 1685, 1840]. Continual [451, 398, 557, 1293]. Continuation [104]. Continuous [57, 1053, 1730, 1507, 1838, 956, 955, 474, 786, 1854, 1444, 1885, 1331, 903, 1742, 1000, 301, 1716, 1671, 261, 718, 1658, 153, 71, 1837, 552, 751, 1300, 1394, 390, 1004, 942, 1119, 420, 1691]. Continuous-Time [1671, 1300]. Continuum [1689]. Contour [226]. Contralateral [1792]. Contrast [463]. Contributions [497]. Control [1042, 759, 212, 1040, 1039, 1435, 1386, 1882, 1622, 1884, 1351, 28, 1672, 277, 1288, 795, 902, 1885, 510, 1883, 648, 1203, 345, 1592, 462, 493, 478, 203, 1385, 1593, 1111, 1010, 1893, 696, 1011, 909, 432, 667, 1490, 830, 1287, 260, 751, 1198, 847, 561, 559, 1012, 2]. Control-Versus-Treatment [1893]. Controlled [1269, 1320]. Controlling [280, 1504]. Controls [28, 18]. Controversy [397, 591, 1395]. Conventional [1671]. Convex [123]. Convolution [684]. Cook [1697, 1692, 1630]. Copula [1659, 1434]. Copulas [1715]. Copy [1711]. Core [1414]. Coregionalized [1173]. Cormack [1053, 607, 1536]. Coronary [1782]. Correct [1164, 362]. Corrected [1397, 846, 936]. Correcting [53, 764, 112, 16]. Correction [194, 284, 402, 560, 561, 559, 521, 863, 1012, 1916, 1571, 363, 475, 1672, 1360, 1258, 124, 934, 114, 1919, 1920]. Corrections [47]. Correlated [1584, 1550, 1486, 489, 1498, 1366, 320, 476, 687, 1770, 1890, 544, 1439, 1715, 484, 92, 1283, 883, 933]. Correlating [360]. Correlation [785, 521, 603, 319, 490, 614, 1292, 1776, 500, 628, 3, 1831, 434, 16, 1886, 857, 770]. Correlations [52, 1776, 75, 650]. Correspondence [43, 282, 978, 526, 1864]. Corresponding [1600]. Cost [1559, 447, 1121, 1465, 1623, 483, 642, 727, 1120, 333]. Cost-Effectiveness [1121, 1465, 1120, 333]. Costs [111, 422]. Count [97, 418, 1207, 98, 405, 912, 1897, 1184, 1851, 666, 1430, 641]. Counterfactual [1006]. Counts [48, 675]. Course [1345, 1116, 1713]. Courses [1387]. Covariables [30]. Covariance [1757, 1106, 353, 56, 844, 773, 1652, 1397, 211, 448, 1878]. Covariate [332, 560, 79, 1057, 1656, 1728, 474, 200, 949, 1432, 1328, 1900, 90, 305, 203, 1658, 993, 1661, 1831, 936, 1600, 1655, 473, 1657, 1168, 1162, 15, 251, 789, 380, 1532, 1537, 1329]. Covariate-Adjusted [1600]. Covariate-Dependent [789]. Covariates [424, 1053, 950, 201, 800, 379, 943, 1167, 643, 1670, 53, 500, 662, 712, 1890, 938, 1060, 1898, 622, 183, 951, 1737, 55, 1565, 1095, 1888, 646, 1663, 70, 1201, 1493, 1394, 898, 1004, 1596, 413, 698]. Cowling [518]. Cowpea [1455]. Cows [747]. Cox [1417, 621, 209, 1000, 671, 1337, 1497, 908, 29, 375, 146, 634, 36, 33, 251]. Crainiceanu [1266]. CRC [1141]. Creel [59]. Cretaceous [186]. Criteria [1586, 989, 1071]. Criterion [1667, 1181, 39, 210, 37, 1277, 639]. Critical [351, 448]. Cross [467, 1375, 1710, 691]. Cross-Sectional [467, 1375, 691]. Crossclassified [1487]. Crosses [1777]. Crossing [1599]. Crossover [1853]. Crossovers [253]. Crowley [1223]. Crude [493]. Cryer [1758]. Crypt [603, 1550]. CT [479]. Cumulant [15]. Cumulative [683, 1040, 852, 368, 1598, 752]. Cure [693, 233, 586, 300, 651, 201, 1450, 247, 1190, 1852, 34, 33, 650, 905, 887]. Cure-Mixture [300]. Cure-Rate [586]. Current [1169, 249, 377]. Curve [1735, 1173, 128, 1410, 86, 890, 398, 51, 1062, 1125, 797, 183]. Curve-Free [86, 398]. Curves [1356, 79, 396, 107, 799, 239, 242, 49, 837, 1652, 14, 1621, 123, 1599, 1877, 423, 229, 1793, 164, 1069, 911, 652]. Curving [293]. Cutpoint [718]. Cycle [340]. Cycles [7]. Cycling [842]. cygnus [427]. Cytogenetic [442]. Cytotoxic [880].

D [1078, 1236, 521, 1090, 1470, 1471, 1266, 1472, 1270, 1417, 1318, 1642, 1701, 1702, 1812, 1762, 1693, 1639, 1316, 1863, 1580, 1407, 1862, 1697, 1632, 1518, 1215, 1692, 1699, 1525, 1524, 1577, 1404, 1758, 1806, 1260, 1218, 1360, 1219, 1801, 1258, 1858, 527, 1223, 1248, 1641, 1909, 1865, 1259, 1262, 1263, 1751]. D. [1693, 1692, 1221, 1748, 1861, 1579, 1640]. Dabo [1860]. Dabo-Niang [1860]. D’Agostino [1463]. Dahm [492]. Daily [177, 1739]. Dairy [747]. Daneshkhah [1632]. Daniels [1747]. Dasgupta [1643]. Data [765, 636, 97, 85, 306, 446, 1008, 1554, 342, 1237, 937, 521, 1227, 1141, 1251, 1473, 1584, 1868, 1814, 1914, 357, 603, 1550, 418, 424, 1042, 1486, 699, 410, 693, 447, 991, 1571, 271, 64, 503, 291, 1353, 563, 57, 615, 1316, 436, 374, 1489, 1373, 1310, 1437, 79, 583, 141, 535, 289, 1199, 616, 574, 212, 258, 292, 27, 1040, 98, 25, 309, 1057, 1435, 1386, 1498, 1882, 1817, 1768, 1160, 378, 901, 1105, 405, 1656, 912, 180, 1683, 1358, 1366, 1068, 791, 1668, 772, 935, 1350, 1660, 1810, 320, 1884, 1736, 1897, 23, 58]. Data [1169, 1303, 1551, 1902, 766, 945, 223, 1351, 1556, 214, 793, 1859, 1082, 236, 249, 467, 377, 755, 701, 1154, 819, 608, 1602, 1692, 279, 68, 372, 1790, 1354, 262, 1809, 407, 457, 556, 1361, 1771, 1880, 835, 629, 617, 1195, 1495, 414, 1603, 165, 136, 1756, 618, 800, 219, 1242, 1159, 476, 1288, 121, 1849, 49, 412, 1375, 1568, 1747, 794, 1799, 998, 786, 199, 200, 1191, 1188, 1280, 1281, 464, 829, 1286, 204, 76, 1151, 1371, 849, 247, 272, 35, 555, 907, 990, 548, 1625, 1839, 1806, 1770, 1132, 100, 704, 1260, 346, 1107, 470, 1670]. Data [1871, 852, 1775, 1049, 431, 147, 1304, 1055, 409, 1184, 1136, 724, 301, 749, 1431, 1452, 406, 500, 1449, 1677, 1826, 1881, 1767, 345, 245, 695, 1653, 232, 13, 644, 832, 75, 167, 1190, 1067, 1330, 1433, 1623, 1666, 1890, 504, 462, 692, 1200, 553, 860, 1060, 1604, 1346, 1716, 1610, 1075, 1851, 261, 1549, 1385, 1305, 144, 939, 1185, 1512, 452, 1111, 158, 423, 1601, 859, 101, 206, 1738, 11, 1391, 352, 166, 696, 24, 29, 361, 375, 1102, 1825, 1003, 1299, 1873, 67, 1732, 358, 746, 1285, 1658, 1189, 628, 691, 1377, 916, 394, 426, 1186, 1115, 1714, 1378, 182, 513, 1100, 364, 1513, 73]. Data [1122, 1246, 790, 1605, 1137, 3, 1566, 1211, 1058, 153, 612, 951, 1737, 1567, 1830, 577, 1326, 824, 1073, 805, 1565, 10, 281, 1308, 1831, 1887, 1372, 432, 1367, 1453, 666, 1490, 1818, 349, 71, 460, 1117, 1600, 1557, 1715, 588, 499, 750, 1430, 1487, 1713, 94, 444, 1287, 1832, 40, 1485, 260, 1139, 638, 828, 861, 172, 754, 888, 992, 1009, 934, 1120, 1379, 1368, 1493, 1488, 1535, 1545, 1483, 1787, 1101, 416, 92, 501, 645, 917, 1174, 312, 1889, 1676, 1794, 850, 843, 1481, 253, 415, 1779, 1778, 469, 842, 7, 302, 641, 663, 1277, 1558, 333, 341, 910, 885, 1380, 1681, 162]. Data [420, 1114, 933, 770, 559, 2, 652, 1919, 1920, 1812, 1633, 1412, 1465, 1312, 1311, 1746, 1517, 1315, 1754, 1088, 1578, 1524, 1577, 1219, 1413]. Database [397, 546, 591, 1395]. Databases [1304]. Datasets [1523, 1774, 222]. Datta [1750, 1696]. Davey [1757]. Davies [529]. Davis [1073]. Davison [1583]. Day [1580]. Dealing [682, 1046]. Dean [1226]. Death [120, 1780, 1361, 1724, 78, 168]. Deaths [1738]. December [560, 1012, 1916, 1571]. Deciphering [1144]. Decision [61, 237, 343, 554, 438]. Decision-Theoretic [343, 554]. Decline [105, 1066, 186]. Declining [234]. Decoding [318]. Decompose [1109]. Decompositions [609]. Deconvolution [582, 248, 1382]. Defensible [1559]. Defined [1211]. Definition [176]. Delayed [258, 22]. Dementia [1066]. Demétrio [521]. DeMets [1697, 1219]. Demidenko [1076]. Demographic [436, 606, 832]. Demography [1221]. Demystifying [1340]. Dense [1667]. Density [571]. Deonier [1264]. Dependence [291, 374, 1451, 1902, 215, 1007, 1525, 1839, 625]. Dependent [1882, 161, 58, 465, 566, 407, 617, 49, 1659, 76, 907, 712, 445, 1623, 938, 798, 1296, 206, 628, 112, 1178, 1565, 666, 1430, 1139, 172, 789, 1001, 1212, 1069, 420]. Depth [1577]. Derived [892, 1304]. Describing [256]. Descriptive [417]. Design [440, 1124, 1228, 1317, 1521, 1850, 280, 1616, 1522, 125, 1533, 665, 1808, 1338, 630, 1502, 1853, 346, 611, 678, 1901, 232, 462, 993, 909, 674, 642, 1752, 387, 482, 1175, 883, 471, 639, 1515, 1644, 1411, 1694, 1644, 1816, 1903, 1220]. Design-Unbiased [125]. Designed [1534]. Designing [324, 422]. Designs [1399, 1741, 839, 441, 1199, 1145, 649, 228, 1040, 1882, 170, 753, 56, 793, 1387, 1202, 315, 1361, 1094, 952, 576, 536, 999, 858, 17, 184, 716, 50, 216, 715, 217, 1337, 1200, 373, 118, 313, 206, 393, 335, 398, 941, 459, 1505, 316, 1370, 386, 745, 1834, 343, 554, 473, 677, 1048, 1516]. Despite [1848]. Detect [1451, 1357, 9, 352, 1872]. Detectability [294]. Detected [1170]. Detecting [221, 1619, 1484, 1434, 487, 1063]. Detection [1554, 1683, 1350, 1654, 134, 1436, 5, 1539, 644, 360, 767, 274, 371, 804, 1108, 1047, 287, 1908, 1617, 1833, 1764]. Determination [427, 705, 126, 237, 633, 1370, 482]. Determining [880, 1564, 1427]. Deutsch [1862]. Development [1521, 1407, 1581, 1314]. Developmental [214, 1097, 1003]. Developments [4, 143]. Deviation [1788]. Devlin [190, 191]. Diagnosis [249, 519, 154, 1765, 1481, 520]. Diagnostic [278, 709, 1399, 1507, 161, 215, 705, 1619, 623, 1748, 49, 515, 50, 1357, 146, 527, 334, 725, 1664, 481, 163, 1886, 162, 1081]. Diagnostics [1155, 1721, 602, 1847, 671, 1720]. Diagonal [1873]. DIC [1851]. Dichotomization [474]. Dichotomous [1010, 668, 1011, 1012]. Dieckmann [525]. Dietary [1514]. Differ [417]. Difference [44, 336, 392, 45, 558]. Differences [22, 1828, 159, 487]. Different [1554, 1387, 1116, 580]. Differential [1755, 1195, 1036, 1279, 1037, 1740, 1370, 1832, 1546, 1194, 1711]. Differentially [493, 512, 571, 1120, 1833]. Diffusion [944]. Diggle [1574]. Dilution [707, 567]. Dimension [829, 1621]. Dimensional [1167, 1844, 14, 1666, 1060, 1873, 851, 1493, 1545, 416]. Dimensionality [513, 1804]. Dimensions [77, 1693]. Dinosaur [186]. Direct [1594, 1782]. Directed [536]. Directions [4, 143, 1088]. Dirichlet [1569, 339]. Discontinuities [5]. Discovery [1226, 1880, 1381, 830, 1438, 638, 1263, 1382]. Discrete [323, 1251, 57, 475, 1722, 474, 786, 1132, 147, 500, 1312, 624, 146, 513, 250, 552, 851, 1537, 698]. Discrete-Valued [475, 698]. Discriminant [417, 629, 1716, 538, 1873, 513]. Discriminatory [722, 1617]. Discussion [1033, 1253, 1418, 241, 918, 919, 980, 981, 982, 983, 984, 985, 240, 1818, 986, 987]. Discussions [1148, 242, 1149, 741, 1560, 1146, 740, 1561, 1508, 1147, 1362, 1562, 742, 1363, 743, 1819, 1364]. Disease [1124, 561, 1042, 25, 453, 221, 213, 467, 376, 1654, 1171, 134, 243, 1533, 105, 476, 299, 116, 510, 5, 431, 546, 1800, 1776, 787, 1357, 684, 1206, 386, 745, 771, 667, 1782, 163, 1481, 430, 1482, 1801]. Disease-Mapping [1206]. Disease-Prevalence [453]. Disease-Specific [561, 510]. Diseases [300, 1177, 1636, 525, 1059, 650]. Disequilibrium [1881]. Disomy [341]. Dispersion [846]. Dispersions [1065]. Dissimilarity [255]. Distance [1236, 1065, 913, 57, 1099, 294, 1733, 720, 1064, 632, 538, 1641, 1552]. Distance-Based [1065]. Distinct [1005]. Distinguishing [155]. Distributed [1391, 1739]. Distribution [1042, 1725, 1683, 12, 279, 796, 886, 1331, 1083, 1542, 130, 1108, 1170, 252, 947, 677, 757, 1481, 253, 415, 1]. Distributional [1570]. Distributions [323, 1207, 566, 717, 1354, 1650, 721, 1196, 285, 230, 245, 373, 1346, 1002, 265, 627, 428, 302, 625, 1518]. Diurnal [220]. Diversity [838, 1709, 186]. Diving [568]. Dmitrienko [1463]. DN [397]. DNA [744, 1867, 530, 918, 919, 1910, 741, 740, 742, 743, 739, 485, 591, 456, 1137, 1395, 416, 1836, 255, 1711, 1519, 1752]. Do [1316, 1576]. Dobson [1815]. Domain [1384]. Domains [1879, 1298]. Dominance [1774, 1621]. Dominici [1863]. Doncaster [1757]. Don’t [128]. Dose [636, 881, 1730, 940, 289, 687, 536, 858, 1674, 1894, 630, 1742, 1792, 32, 335, 941, 1464, 1615, 1505, 1440, 387, 317, 565, 756, 1663, 851, 1163, 1840, 1515, 1314]. Dose-Escalation [1515]. Dose-Finding [881, 630, 335, 941, 1464, 1505, 317, 565, 756, 1163]. Dose-Response [1730, 940, 536, 1615, 1440]. Doses [1387]. Dosimetry [442, 288]. Double [1741, 241, 1099, 239, 242, 568, 171]. Double-Observer [1099]. Double-Platform [568]. Double-Sampling [1741]. Doubly [991, 1571, 1107, 361, 750, 1493]. Dougherty [1700]. Down [387]. Dramatic [1442]. Drawbacks [272, 1308]. Drop [198, 40, 383]. Drop-Out [40, 383]. Drop-Outs [198]. Dropout [97, 181, 602, 786, 612, 1555, 197, 1056, 1778]. Dropouts [612, 951, 1794]. Drs [492]. Drug [1226, 1521, 708, 1407, 1375, 1183, 1540, 1314, 1532]. Dual [1176]. Dual-Endpoint [1176]. Dudoit [1214, 1753]. Due [286, 234, 1780, 1361]. Duncan [1410]. Dunson [1858]. Duration [340, 697, 994]. During [39, 415, 253]. Durrett [1867, 530]. Dynamic [765, 1500, 1692, 1103, 1340, 1189, 394, 1315]. Dynamical [1624, 1446]. Dynamics [1838, 42, 853, 665, 1847, 525, 1204].

E. [815, 1215]. E.-J. [815]. Early [157, 1539, 371, 684, 487]. Early- [684]. East [6]. Ebola [1203]. Ecological [1080, 262, 1288, 1568, 1517, 498, 1809]. Ecologists [1238]. Ecology [1857]. Edgington [1467]. Edited [1236, 1228, 1225, 1226, 1128, 1215, 1216, 1217, 1220, 1222, 1223, 1224, 1088, 1227, 1320, 1319, 1317, 1271, 1318, 1813, 1762, 1912, 1213, 1639, 1074, 1316, 1310, 1463, 1905, 1749, 1214, 1221, 1633, 1261, 1404, 1260, 1083, 1218, 1312, 1313, 1860, 1314, 1464, 1575, 1311, 1219, 1801, 1858, 1315, 1259, 1461, 1262, 1696, 1263]. Edition [1251, 1320, 1318, 1267, 1415, 1268, 1471, 1269, 1473, 1266, 1317, 1584, 1702, 1704, 1812, 1759, 1816, 1911, 1815, 1760, 1915, 1811, 1405, 1310, 1749, 1410, 1699, 1698, 1261, 1404, 1758, 1260, 1218, 1220, 1223, 1413]. editor [532, 960, 187, 189, 325, 327, 329, 330, 399, 533, 600, 660, 601, 738, 784, 813, 825, 878, 926, 971, 1026, 1027, 1028, 1029, 1030, 1031, 1032, 1084, 1250, 1402, 1687, 1744, 1126, 774, 775, 454, 1140, 1072, 494, 961, 592]. Editors [525, 522, 807, 1254, 815, 653, 661, 826, 1478, 816, 1708, 817]. Eeden [1469]. Effect [950, 1719, 150, 1780, 1342, 243, 1765, 1110, 389, 635, 1797, 1597, 508, 908, 218, 158, 700, 1400, 1858, 1178, 1006, 466, 1889, 1778, 1329]. Effective [611, 1564]. Effectiveness [1121, 1465, 1120, 333]. Effects [1356, 357, 64, 503, 1353, 563, 1458, 79, 1594, 604, 895, 1899, 1817, 427, 152, 935, 12, 221, 1556, 155, 506, 109, 796, 570, 381, 474, 1282, 151, 664, 665, 1568, 794, 116, 1177, 1894, 1369, 53, 514, 409, 1792, 132, 1428, 1067, 1328, 90, 1610, 305, 11, 459, 24, 1658, 993, 364, 1206, 229, 855, 640, 1536, 1766, 1857, 1782, 1818, 439, 222, 646, 40, 383, 31, 70, 609, 789, 888, 1224, 1598, 1739, 1675, 1552, 54, 312, 408, 1347, 1063, 207, 1827, 302, 413, 1472]. Efficacy [1841, 1176, 756, 1163]. Efficacy-Toxicity [756]. Efficacy/Toxicity [1841]. Efficiency [1559, 271, 17, 1479, 395, 1593, 642, 88, 176, 1596]. Efficient [119, 1199, 801, 611, 1618, 1482]. Effort [257]. Efron [1456]. Efron-Type [1456]. EIC [516]. Eight [1176]. Eisenberg [1634]. Eiser [1632]. election [58]. Electromagnetic [219]. Elementary [1585]. Eliciting [1632]. Ellison [1080]. EM-Boosting [685]. EMA [1568]. Emerging [1614]. Emigration [346, 232]. Emphasis [474]. Empirical [1736, 1631, 1244, 1593, 1125, 1201, 1347, 420]. Encompassing [1497]. Encounter [1733]. End [1043]. Endpoint [1500, 1824, 567, 1432, 1176, 1505]. Endpoints [760, 337, 1495, 1665, 505, 259, 678, 1071, 694, 1360, 854, 250, 1247, 205]. Engineers [1529]. Enhanced [1280, 443, 1546]. Entropy [1710]. Entry [699, 39, 906]. Environment [1385, 1593, 349, 1547, 892, 1573]. Environmental [48, 1572, 1634, 1863, 385, 1592, 222, 1259, 1175]. Epidemic [1613, 1614, 1203]. Epidemics [388]. Epidemiologic [1074, 133, 1802]. Epidemiological [289, 286]. Epidemiology [1131, 1863, 1079, 1045, 290, 824, 179, 1696, 1270, 1405, 1800]. Epigenetic [1454]. Episode [1332]. Episodically [1871]. Epistatic [1347]. eQTL [1436, 1038]. Equal [82]. Equality [23, 354, 310]. Equation [1416, 60, 1466, 1740, 1720, 88]. Equations [1755, 185, 457, 511, 1603, 1397, 1391, 210, 266, 1488, 550]. Equilibrium [1881, 757]. Equivalence [1207, 259, 716, 1900, 1662, 482]. Equivalence/Noninferiority [482]. Eradication [89]. Erosion [700]. Error [278, 709, 1266, 759, 1079, 1045, 41, 682, 1456, 1289, 800, 1791, 1672, 902, 1823, 1441, 1871, 53, 949, 1449, 1553, 1428, 1205, 90, 369, 764, 203, 1197, 1102, 1825, 202, 936, 473, 1204, 1290, 897, 124, 1488]. Error-in-Variable [897]. Errors [303, 294, 109, 952, 643, 389, 500, 712, 1328, 1592, 1355, 425]. Errors-in-Variables [303]. Erythrocytes [480]. Escalation [1515]. Essays [1912]. Essentials [1089]. Establish [716]. Establishing [482, 757]. Estimate [357, 436, 1361, 346, 691, 312]. Estimated [332, 895, 1890, 448, 469, 560]. Estimates [1155, 453, 1429, 1005, 109, 1295, 1304, 63, 1341, 1157, 1443]. Estimating [709, 1399, 1356, 1124, 1236, 521, 1042, 563, 57, 44, 1624, 614, 435, 228, 1040, 1054, 1882, 453, 185, 213, 549, 128, 1171, 1733, 457, 198, 385, 239, 242, 511, 1603, 121, 567, 697, 174, 514, 1499, 1776, 1797, 355, 1397, 493, 890, 1734, 11, 1391, 210, 266, 893, 3, 675, 1178, 1641, 1170, 932, 88, 1480, 631, 1547, 1302, 992, 934, 1120, 1193, 1488, 1598, 6, 1739, 1836, 1300, 752, 1347, 1394, 1773, 550, 471, 1004, 253, 333, 1511, 1382, 1919, 1920]. Estimation [636, 559, 338, 991, 1571, 233, 1730, 616, 1769, 959, 1098, 25, 311, 403, 676, 845, 1541, 1817, 792, 1719, 1725, 108, 161, 59, 1728, 120, 465, 1169, 234, 257, 637, 578, 392, 844, 1154, 1780, 1342, 156, 315, 1724, 125, 1650, 276, 80, 219, 1791, 1446, 1292, 52, 1375, 304, 116, 795, 1674, 886, 1177, 1103, 1165, 1167, 204, 1331, 76, 1151, 643, 1608, 1055, 2, 301, 389, 635, 443, 662, 75, 445, 632, 373, 553, 860, 787, 87, 767, 702, 1469, 882, 564, 889, 82, 572, 423, 1179, 1738, 568, 375, 1825, 1828, 34, 248, 1714, 1258, 1210, 149, 183, 1729]. Estimation [10, 627, 1349, 994, 669, 296, 19, 1818, 719, 1657, 1393, 1837, 33, 1438, 638, 177, 16, 754, 92, 1158, 1506, 1552, 1889, 550, 114, 585, 231, 883, 1537, 382, 641, 1482, 911, 770]. Estimator [350, 1052, 862, 508, 211, 1593, 1423, 846, 677, 16, 251, 1535, 295]. Estimators [353, 356, 479, 516, 1397, 148, 723]. Estrous [842]. Ette [1520].

Evaluate [319, 451]. Evaluating [61, 490, 1507, 1665, 1444, 1885, 1871, 216, 722, 788, 1886, 1329]. Evaluation [760, 1294, 1905, 235, 989, 831, 226, 1205, 1176, 31, 1247, 205, 806, 1078]. Event [332, 1237, 1271, 1812, 699, 410, 615, 27, 935, 1450, 701, 414, 1803, 1332, 672, 500, 763, 1190, 1623, 419, 908, 1512, 1601, 1828, 1189, 1108, 250, 1138, 460, 1676, 1558, 560]. Events [1457, 378, 128, 78, 617, 1333, 245, 763, 798, 1170, 1600, 1283, 1827, 1630, 1404]. Everitt [1225, 1141, 1911, 1915, 1085, 1268]. Eveson [653]. Evett [191]. Evidence [1316, 286, 1244, 1701, 1526]. Evidential [591]. Evolution [1241, 1575, 1186, 1468, 530, 1867]. Exact [1785, 616, 792, 901, 320, 1451, 687, 647, 299, 259, 1597, 419, 558, 1893, 351, 484, 1312]. Exam [371]. Examination [1339]. Example [1123]. Examples [1755]. Excel [1075]. Exceptional [1519]. Excess [1871, 1566]. Excessive [692]. Exchangeable [1768, 1299, 499]. Existing [1095]. Expanding [468]. Expectancy [35]. Expected [1426, 1796, 422, 1488]. Experienced [699]. Experiment [583, 1516, 1339]. Experimental [1644, 1777, 1613, 665, 1411, 1178, 1752, 1175, 1515]. Experimentation [1226]. Experiments [1642, 1644, 616, 1345, 1534, 155, 551, 590, 137, 1094, 1548, 768, 1533, 1894, 1502, 1503, 1740, 1671, 456, 512, 1220, 674, 1344, 1300, 996, 1004, 1903, 1464]. Expert [1632]. Explained [1216, 36, 1006, 466]. Explanatory [682]. Explicit [1538]. Exploitation [59]. Exploiting [1728, 1593]. Exploration [79, 916, 645, 917]. Exploratory [910]. Exploring [1839, 261, 1344, 1082]. Exponentially [1846]. Exposure [286, 506, 219, 277, 160, 90, 493, 1010, 1011, 1490, 222, 1290, 287, 177, 15, 1012]. Exposure-Time-Response [160]. Exposures [1592]. Expressed [512, 571, 1833]. Expression [881, 1437, 766, 819, 1436, 629, 1195, 1036, 1116, 1279, 1280, 1502, 1038, 1037, 916, 1115, 1576, 1832, 645, 917, 1194, 1114, 1711]. Expressions [1546]. Extended [457, 746, 1404]. Extending [1667, 860, 1224]. Extension [1053, 516, 891, 1211]. Extensions [1853, 1231, 1423, 634, 1857, 1845, 1866]. Extinction [1213]. Extra [418, 261]. Extra-Binomial [261]. Extra-Poisson [418]. Extrapolation [805]. Extrema [1117]. Extremes [1404].

F [1408, 1519, 1077, 1516, 1260, 1222, 1517, 1857, 1909, 524]. F. [1863, 1630, 1312, 1860, 1856, 1581]. Factor [1727, 236, 1383, 421]. Factorial [858]. Factors [1164, 701, 1276, 1061, 1851, 1766, 932]. Faenza [1339]. Failure [765, 374, 1769, 1498, 1366, 1660, 931, 1303, 637, 1602, 407, 136, 1412, 800, 536, 1281, 1167, 1770, 1107, 1723, 1890, 355, 462, 1200, 146, 854, 1367, 750, 1168, 850, 843]. Failure-Time [462, 146, 854]. Falk [1404]. Fallout [1449, 369]. False [1504, 830, 725, 1438, 638, 752, 1382]. False-Negative [725]. False-Positive [752]. Familial [300, 1040, 457, 1776, 650, 430, 117]. Families [457, 516, 552, 1193]. Family [1040, 116, 795, 1110, 1111, 1513, 432, 1547, 751, 1675]. Family-Based [1547]. Family-Specific [1111]. Far [137]. Faraway [1224]. Farrell [815]. Fast [1821]. Fatal [1059]. Fatality [1614]. Fayers [1640]. FDR [1158]. Feasibility [317]. Feasible [1124]. Feature [1191, 1666]. Feature-Specific [1191]. Features [1381]. Fecundability [150]. Federer [1516]. Feedback [99, 1607]. Fekjær [560]. Female [1831]. Ferraty [1860]. Fertility [41, 340]. Fertilization [61]. Fetal [1895]. Fever [388]. Field [1587, 137, 349]. Fielding [1637]. Fields [219]. Fill [66]. Final [588]. Finding [881, 17, 630, 1742, 1314, 335, 941, 1464, 1505, 387, 317, 565, 756, 1663, 851, 1163, 1840]. Fine [1172, 1696]. Fingerprints [416]. Finite [632, 1671, 182, 1443, 1409]. Finite-State [1671]. First [363, 427, 747, 249]. First-Order [363]. Firth [521]. Fish [1155, 676]. Fisher [902]. Fisheries [723]. Fit [711, 58, 1619, 799, 139, 590, 1182, 1335, 995, 502, 73, 1605, 797, 522]. Fitness [768, 1740]. Fitted [219]. Fitting [185, 42, 1104, 1610, 227, 652]. Fixed [1899, 221, 1369, 1206]. Fjeld [1634]. Fleas [853]. Flexibility [1282]. Flexible [834, 713, 605, 1145, 1891, 249, 1548, 1672, 613, 1165, 1055, 1828, 1834, 1290, 487, 302]. Florida [1141]. Flowgraph [1237]. Fluorescence [1793]. FMR1 [1831]. fMRI [1875]. Foci [603, 1550]. Focused [1181]. Folate [585]. Follow [561, 647, 35, 510, 445, 168, 1686, 1684]. Follow-Up [647, 35, 445, 168, 561, 510]. Following [1785, 792, 91, 288]. Follows [140, 365, 138, 1401, 188, 326, 328, 400, 1309, 1233, 1403, 1127, 1688, 776, 1627, 1629, 1235, 726, 808, 1745, 1307]. Food [1044, 1736]. Foods [1871]. Forcing [1847]. Forecasting [1221]. Foreman [191]. Forestry [60]. Form [350, 1165, 671]. Forms [65]. Formula [836, 1600]. Formulation [1536]. Forum [929, 976, 1092, 1143, 1230, 1256, 1274, 1323, 1420, 1476, 1590]. Forward [1082]. Four [1853]. Fraction [651, 679, 247, 650]. Fractional [1854, 118]. Fractions [219, 893, 887]. Fragile [491, 254, 492]. Fragment [1196]. Fragmentation [1690]. Frailties [747]. Frailty [1882, 379, 433, 1333, 1107, 406, 395, 763, 1189, 790, 149, 673, 650, 1300, 1394, 905, 850, 117, 641]. Frailty-Model-Based [1882]. Framework [1068, 1625, 1662]. Frameworks [1784]. Francis [1141, 1086]. Francis/CRC [1141]. Frangakis [240]. Free [1735, 86, 398]. Free-Response [1735]. Freedman [1012]. Freedom [681]. Frequencies [1197]. Frequency [702, 1278, 253]. Frequentist [900, 519, 1395, 520]. Frequently [1450]. Friedman [1219]. FROC [1735]. Frühwirth [1409]. Frühwirth-Schnatter [1409]. FSR [1821]. Full [895]. Function [1882, 80, 160, 1847, 795, 1742, 1496, 697, 852, 1666, 225, 1302, 888, 992, 1448]. Functional [1486, 270, 381, 1116, 671, 1549, 1002, 1793, 1830, 1372, 1832, 40, 383, 1485, 1754, 761, 1773, 1712, 585, 1878, 1329, 1310, 1860]. Functions [1730, 759, 683, 23, 463, 148, 1100, 934, 1773, 310, 1919, 1920]. Fundamentals [1462]. Fungal [226]. Funnel [66]. Funnel-Plot-Based [66]. Furberg [1219]. Furosemide [235]. Further [157]. Future [1088, 48, 4, 143, 1275, 1572].

G [521, 1089, 1087, 1090, 1141, 1270, 1869, 1866, 1911, 1814, 1916, 1573, 1578, 1859, 526, 1406, 1265, 1218, 1805, 1807, 1073, 1635, 1517, 1857]. G. [1586, 1815, 653, 1859, 1750, 1694]. Gain [271, 553, 1400]. Gaining [280]. Gains [18]. Galecki [1695]. Galton [1086]. Galwey [1695]. Game [680]. Gamma [1567, 850, 641]. Gamma-Frailty [641]. Gan [1578]. Gap [1333, 245, 1395, 1427]. Garthwaite [1632]. Gasca [526]. Gascuel [1575]. Gatelli [1909]. Gaussian [1489, 1160, 75, 1822, 1372, 1715]. GE [916, 917]. GE-Biplot [916, 917]. Gear [234]. GEE [319, 75, 211]. Geisser [1417]. Gene [1437, 1882, 766, 819, 629, 1036, 1279, 1280, 1502, 1144, 1037, 1385, 1593, 916, 1115, 1576, 386, 745, 1112, 1713, 638, 1835, 1547, 1546, 645, 917, 1833, 1194, 1751, 1114, 1711]. Gene-Disease [386, 745]. Gene-Environment [1385, 1593, 1547]. Gene-Trait [1835]. GeneChip [897]. General [1160, 791, 637, 354, 1625, 1046, 749, 393, 746, 421, 1555, 727, 1886, 887]. Generalizability [861]. Generalization [992]. Generalizations [822]. Generalized [1815, 424, 1044, 417, 1106, 899, 614, 1777, 405, 185, 542, 457, 1398, 385, 109, 1603, 507, 1381, 1679, 902, 943, 1765, 990, 1183, 537, 1055, 1459, 662, 13, 1428, 1397, 458, 352, 210, 995, 1714, 1567, 577, 1888, 173, 178, 540, 1180, 1224, 1187, 806, 382, 663, 1764, 933, 1458, 1460]. Generating [1297]. Generation [1088, 1426]. Genes [1882, 1116, 158, 1434, 512, 1278, 571, 1113, 298, 1833]. Genetic [297, 1916, 833, 1199, 1774, 385, 838, 116, 587, 1433, 553, 256, 1157, 159, 847, 1077]. Genetically [551]. Genetics [96, 114, 910, 1226, 1639]. Genome [1501, 896, 1381, 1883, 1276, 1901, 1117, 1264]. Genome-Wide [1883, 1276, 1901]. Genomes [1144]. Genomewide [1834]. Genomic [1880, 1195, 998, 1191, 1192, 1277, 847, 1700]. Genomics [1753]. Genotype [1040, 384, 1197, 757, 1836]. Genotypes [1455, 1592]. Genotypic [342]. Genotyping [1197]. Gentleman [1214]. Geoadditive [1795]. Geochemistry [1731]. Geocoding [1511]. Geodetic [913]. Geographic [1908]. Geometric [1059]. Geometry [1577]. George [192]. Geostatistics [1574, 1862, 1582]. Gerbils [853]. Germination [108]. Germination-Growth [108]. Get [128]. GGobi [1692]. Gibbs [182]. Giovanna [1920]. Gjessing [1803]. Glidden [1215]. GLM [51]. Global [1909]. GM [896]. Gold [278, 709, 1356, 705, 911]. Gollery [1703]. Gomez [1799]. Gomez-Rubio [1799]. Gompertzian [156]. Gone [137]. Gönen [1582]. Good [680]. Goodness [711, 58, 799, 139, 590, 1182, 1335, 995, 502, 797, 522]. Goodness-of-Fit [711, 58, 799, 139, 590, 1182, 995, 502, 522]. Gotelli [1080]. Grace [1416]. Grades [1293]. Grafen [531]. Graphical [79, 1507, 17, 916, 94, 645, 917]. Graphics [1692, 1413, 1523]. Graphs [1357]. Great [853]. Green [1580]. Greenacre [1864]. Gregoire [1573]. Groesen [1689]. Group [937, 64, 1785, 605, 1899, 904, 276, 30, 1441, 1844, 678, 884, 20, 1296, 313, 144, 1211, 1661, 324, 439, 1664, 1379]. Grouped [180, 1677, 1620]. Grouping [286]. Groups [336, 1501, 1853, 580, 557]. Growth [619, 1213, 1173, 258, 108, 799, 155, 755, 156, 1410, 1455, 1652, 1484, 804, 1895, 754, 761, 652]. Gu [523]. Guide [1317, 1270, 1646, 1238, 1216, 1695, 823, 824, 1463]. Guidelines [235]. Gupta [1083]. Gustafson [1079].

H [560, 1088, 1317, 1523, 1572, 1571, 1573, 1080, 1905, 1632, 1803, 1261, 1465, 1861, 527]. H. [1317, 1587, 1757, 1761, 1904, 1698, 1637, 1315]. Habitat [547]. Habits [1455]. Haccou [1213]. Haenszel [992]. Hails [531]. Handbook [1225, 1268, 1759, 1131, 1240, 1227, 1470, 1471, 1703, 1911, 1639, 1083, 1222, 1223]. Handle [1717]. Haplotype [1039, 1435, 299, 1592, 1197, 1198]. Haplotype-Based [1039, 1435, 1592, 1198]. Hardy [1041, 1881, 532, 757]. Hardy-Weinberg [1041, 1881, 757]. Harlamov [1691]. Harman [1635]. Harmonic [1008, 1009]. Harvest [1668]. Hauschke [1407]. Hayes [1904]. Hazard [1917, 574, 1918, 463, 80, 1282, 795, 395, 148, 1657]. Hazards [309, 180, 1450, 465, 1829, 376, 68, 1494, 904, 136, 800, 613, 1165, 1770, 1000, 671, 1166, 1190, 508, 461, 1497, 624, 375, 936, 750, 33, 145, 1598, 698]. Health [1078, 1702, 980, 1068, 981, 982, 983, 984, 985, 1871, 1452, 1604, 544, 1801, 979, 986, 987, 1634, 1863]. Health-Care [1078]. Healthy [1515]. Heart [1782]. Heavy [1756]. Heavy-Tailed [1756]. Heckman [19]. Held [1918]. Helge [1912]. Hellinger [632, 1552]. Hematopoiesis [269]. Hemodialysis [1826]. Hemophiliacs [424]. Herald [6]. Herbicide [175]. Heredity [1086]. Heritability [350]. Herpesvirus [691]. Hesketh [1268, 1704]. Heterogeneity [1353, 44, 1099, 1569, 701, 65, 1608, 395, 607, 547, 719]. Heterogeneous [958, 644, 957, 1047]. Heteroscedastic [1448]. Heumann [1912]. Hidden [711, 1703, 1902, 1880, 106, 1910, 105, 147, 341]. Hierarchical [1486, 708, 1373, 535, 1876, 1345, 1774, 1548, 618, 1381, 476, 1288, 52, 1678, 1116, 1352, 254, 1103, 990, 1800, 1374, 544, 1851, 668, 1370, 1830, 1453, 1112, 1875, 231]. Hierarchical-Clustering [254]. Hierarchically [749]. High [1167, 1666, 1060, 1873, 513, 571, 1493, 1545, 416]. High-Density [571]. High-Dimensional [1167, 1666, 1060, 1873, 1493, 1545, 416]. Highly [299, 1439, 481]. Hinkelmann [1644]. Hirji [1312]. Hispanic [1776]. Historical [478, 168]. Histories [332, 560]. History [1271, 1074, 1040, 1803, 116, 1110, 514, 1275, 1109, 684, 1675]. HIV [338, 42, 384, 570, 1446, 1445, 665, 1375, 1281, 1103, 514, 580, 1677, 1374, 256, 1176, 1740, 335, 1662, 517, 1095]. HIV-1 [42, 384, 1375, 1374, 517]. HIV-Infected [1662]. Hjort [1805]. Hofmann [1523]. Hogan [1747]. Hoijtink [1905]. Holm [1916]. Homeostasis [359]. Homogeneity [1065, 1775, 20, 1881, 45, 633, 393]. Hong [1371, 1614]. Honour [1912]. Hormone [582, 1447, 1297, 1063, 7]. Hormones [1798]. Horses [235]. Hoshmand [1220]. Hosmer [1812]. Hospital [1123, 1612]. Host [122]. Hothorn [1225]. Hours [620]. Household [563]. Housing [222]. HPD [1685]. Hu [1408]. Huber [1214, 522]. Huber-Carol [522]. Huesler [1404]. Human [424, 1634, 41, 691, 114]. Humans [362, 1636]. Hunting [52]. Hwang [1744]. Hybrid [1055, 1056, 1778]. Hybridization [1277]. Hyper [723]. Hyper-Parameters [723]. Hyphae [226]. Hypotheses [342, 1673, 621, 30, 91, 1905]. Hypothesis [1043, 1156, 1848, 856].

I. [1908]. I/II [881, 1338, 1163]. Iacus [1755]. IBS [1272, 1709]. Identifiability [1817, 1234, 1818]. Identification [491, 492, 273, 1441, 1740, 1349, 571, 1194]. Identified [126, 1453]. Identify [771]. Identifying [157, 1116, 254, 1843, 828, 1113]. Idiosyncrasy [239, 242]. Ieno [1517, 1857]. II [881, 378, 1616, 1338, 468, 903, 237, 1428, 343, 554, 550, 1163]. IIA [715]. III [468]. Illian [1861]. Illness [120, 1724]. Illness-Death [120]. Illnesses [116]. Illustrated [94]. Image [797]. Images [479]. Imaging [270]. Immigration [1738]. Immune [1202]. Immunodeficiency [424]. Immunohistochemical [278]. Impact [1164, 769, 39, 861]. Impacts [1079]. Imperfect [1445]. Implementation [1808, 588, 1515]. Implementations [605]. Implications [1295, 1479]. Implicit [1180]. Important [1780]. Imprecise [686]. Imprinting [1192]. Improve [1849, 1375]. Improved [211, 171, 1833]. Improving [1005, 1157, 1596]. Imputation [1399, 772, 372, 835, 199, 167, 355, 29, 361, 1718, 898, 1845]. Incidence [683, 467, 376, 219, 852, 431, 450, 1480]. Incomplete [1044, 1489, 1543, 1656, 710, 772, 551, 457, 200, 75, 1246, 1887, 841, 1393, 444, 16, 380, 1511]. Incomplete-Data [1246]. Incorporating [258, 622, 588, 1655, 1836]. Incorporation [268, 128, 553, 640, 1833]. Incorrectly [1848]. Increased [65]. Increasing [177]. Incremental [1479, 1120, 333]. Independence [264, 391, 690, 475, 1728, 223, 1051, 1385, 1593, 334]. Independent [556, 558, 1391, 1170, 260, 428, 1681]. Independent-Block [1391]. Index [1735, 1424, 176]. Indicators [672, 1000]. Indices [1098, 1669, 676, 262]. Individual [838, 999, 116, 1670, 53, 1536, 1737, 177]. Individual-Level [53]. Individualizing [1874]. Individuals [549, 1734]. Indrayan [1760]. Induced [376, 1711]. Induction [1635]. Industry [1226, 1749]. Infant [1445]. Infected [42, 1662, 517]. Infection [424, 563, 691]. Infections [89, 748]. Infectious [681, 1636, 525]. Infectivity [517]. Inference [275, 1701, 1761, 991, 1571, 81, 615, 263, 1437, 1755, 212, 269, 98, 336, 1386, 1719, 901, 542, 556, 370, 102, 1650, 1495, 721, 414, 944, 270, 1036, 1533, 412, 304, 1772, 1165, 1496, 1455, 1203, 443, 1096, 1618, 217, 1328, 1592, 1852, 860, 1604, 558, 1740, 32, 1305, 1789, 1102, 51, 1125, 1100, 1576, 1246, 951, 250, 1117, 499, 444, 484, 481, 1685, 539, 1302, 789, 1001, 1389, 430, 698, 1417, 1089, 1417, 1631, 1693]. Inferences [621, 840, 463, 997, 575, 537, 859, 746, 1832, 1596]. Inferring [1671, 158, 1278]. Inflated [151, 914, 1184, 224, 915, 947]. Influence [1155, 1721, 555, 352, 1888, 197]. Influencing [1061]. Influential [352]. Informatics [1801]. Information [1294, 1667, 1543, 1181, 1154, 137, 285, 1849, 200, 1280, 1144, 587, 1384, 1792, 553, 478, 703, 210, 1400, 1344, 1655, 1843, 37, 579, 380, 1277, 1586]. Informative [97, 446, 1905, 1366, 181, 236, 377, 433, 769, 1767, 695, 1330, 1623, 951, 1830, 250, 758, 383, 501]. Informatively [697]. Infusion [317, 438]. Inheritance [115]. Inhomogeneous [1351, 1772, 1302]. Initial [90]. Initiated [1156]. Initiative [980, 981, 982, 983, 984, 985, 986, 987, 979]. Injury [64]. Input [1361]. Insecticide [182]. Insemination [747]. Instability [158]. Instability-Selection [158]. Institute [1248]. Instrument [174]. Instruments [1514]. Integer [272]. Integrated [465, 60, 470]. Integrating [436]. Integration [1710]. Intensity [541, 1511]. Intensity-Score [541]. Intensive [1633, 1260]. Inter [1875]. Inter-Subject [1875]. Interacting [853, 1384]. Interaction [221, 127, 1183, 1540, 118, 351, 610]. Interactions [268, 122, 1547, 1162]. Interactive [1128, 1692]. Interacts [395]. Interdisciplinary [1869]. Interest [299]. Interesting [1587]. Interim [882, 169]. Intermediate [1814, 1594, 1724]. Intermittent [695]. Internal [280]. International [1325]. Interobserver [275]. Interplot [56]. Interpolation [831]. Interpretation [1227, 51, 1640]. Interpreting [132]. Interval [424, 693, 374, 1667, 574, 25, 895, 1777, 792, 286, 1660, 392, 755, 68, 407, 1724, 165, 136, 1412, 204, 1331, 575, 1190, 82, 423, 29, 375, 73, 725, 770]. Interval-Based [575]. Interval-Censored [424, 693, 374, 574, 286, 1660, 755, 68, 165, 136, 1412, 423, 29, 375]. Intervals [323, 900, 1680, 72, 1343, 1615, 1378, 1442, 1341, 260, 1120, 162]. Intervention [157, 1203, 1367, 806]. Intestinal [1156]. Intra [838]. Intra-Individual [838]. Intraclass [521, 392, 633, 3, 770]. Intrareader [857]. Intrauterine [804]. Introduction [1473, 1644, 1815, 1129, 1410, 1138, 1460, 1264, 1269, 1320, 1527, 1645, 1697, 1631, 1695, 1750, 1865]. Introductory [1085, 1216]. Intuitive [1316]. Invariant [342, 609, 1187]. Invasions [1352]. Inverse [1628, 344, 1674, 1492, 112, 251]. Investigate [861]. Investigating [485]. Investigation [1532]. Investigators [1462]. invited [1254]. IPCW [112]. Irizarry [1214]. Irradiation [288]. Irregular [626]. Irreversible [120]. ISBN [1141, 530, 523, 526, 525, 532, 522, 529, 531, 527, 1248, 524, 528]. Isotonic [706, 387]. Isotope [480]. Isotropy [1050]. Israel [526]. Issues [980, 981, 982, 1410, 983, 984, 985, 979, 642, 986, 987, 652]. Item [1454, 1443]. Iterated [1764]. Iterative [829, 1374]. IV [821]. ix [526]. Izenman [1855].

J [1078, 1236, 1085, 1267, 1471, 1269, 1320, 1266, 1416, 1642, 1702, 1757, 1813, 1866, 1815, 1761, 1870, 1914, 1639, 1574, 1571, 1521, 1080, 1310, 1458, 1904, 653, 1632, 1221, 1633, 1638, 1698, 1412, 1747, 1799, 1520, 1406, 1630, 1404, 1758, 1414, 1260, 1265, 1459, 1636, 525, 1861, 1801, 1223, 1413, 1641, 1857, 1909, 1856, 94, 1259, 1461, 1224, 1804, 1262, 528, 1696]. J. [1702, 1869, 1816, 1814, 1574, 815, 1578, 1749, 1855, 1690, 1753, 1313, 1222, 1689, 1224]. J.-P. [1816]. J.-Y. [1749]. Jaakko [1916]. Jagers [1213]. James [1493]. Jarvis [1572]. Jenkinson [1632]. Jensen [1217]. Jiang [1459]. Joffe [561]. Joint [303, 586, 535, 956, 1874, 1105, 602, 569, 608, 1602, 1104, 246, 1880, 955, 1188, 1281, 1333, 1824, 1331, 1066, 1841, 672, 1776, 1432, 1826, 1767, 1330, 790, 1566, 460, 1557, 1715, 1827]. Jointly [881, 1556]. Jolly [1053, 607, 1536, 192, 1153, 1425]. Joshi [1088]. Journal [1164]. Journals [196]. Judgement [1534, 686]. Judgements [1632]. Judgment [1535]. Julious [817]. June [559]. Jureckova [1313]. Jurisica [1263]. Just [1453]. Just-Identified [1453].

K. [1318, 1803, 1758]. K.-A [1576]. K.-S. [1758]. Kaplan [937, 1500, 128, 244, 307]. Kappa [489, 81, 392, 1722, 82, 633]. Kappas [1604]. Kaprio [1916]. Kargupta [1088]. Katz [1317]. Kazakhstan [853]. Keeling [1636]. Keiding [1271]. Kempthorne [1644]. Kennedy [1814]. Kenward [1859]. Kernel [1433, 628]. Kernel-Based [628]. Khan [1862]. Khattree [1762]. Killup [1216]. Kin [228]. Kin-Cohort [228]. Kinetic [944]. Kinetics [1383, 480, 585]. King [1516]. Kitagawa [1586]. Klein [1090]. Kleinbaum [1090, 1270]. Klugkist [1905]. Knowledge [277, 362, 1263]. Known [1554, 360]. Koch [1073]. Kodlin [863]. Kokko [1587]. Kong [1371, 1614]. Konishi [1586]. Kosorok [1631]. Kowalski [1638]. KPSS [1606]. Kravchenko [1869]. Krishnan [1866]. Kroese [1699]. Kroonenberg [1810]. Kuhnert [1806]. Kulkarni [1635].

L [1236, 1012, 1089, 1266, 1585, 1759, 1316, 1904, 1410, 1748, 1360, 1219, 1462, 1222, 1641, 1101]. L-Splines [1101]. L. [1642, 1572, 1634, 1697, 1633, 1906, 1577, 1805, 1801, 1641]. Laake [1236, 1641]. Laan [1753]. Labeled [480]. Labeled-Cell [480]. Labeling [480]. Lack [1619]. Lack-of-Fit [1619]. Lag [1739]. Lagged [357]. Lai [1525]. Lan [1406]. Langholz [559]. Large [1523, 1693, 953, 1774, 954, 1514, 640]. Large-Sample [640]. Larose [1524]. Laslett [653]. LASSO [1303]. Late [170, 820, 906, 684, 487]. Late-Onset [170]. Late-Stage [684]. Latent [278, 446, 709, 1353, 1489, 161, 1902, 543, 1410, 519, 154, 835, 794, 1191, 1824, 1823, 672, 1000, 301, 749, 695, 1653, 1852, 589, 1186, 1858, 153, 612, 1895, 54, 1872, 1886, 380, 1840, 520, 933, 1607]. Latent-Class [1489]. Latent-Model [1824]. Latent-State [1607]. Latent-Variable [1823]. Lattice [178]. Lattice-Ordered [178]. Lattices [1839]. Lawless [1630]. Laws [1404]. Lawson [1800, 1405]. Lazar [1754]. Lead [39, 222, 1480, 1389]. Lead-in [39]. Learning [1090, 1750, 1726, 1635, 685, 1907, 1855]. Least [489, 1303, 829, 1433, 434]. Least-Squares [489, 1433]. Lecture [1272]. Lee [1458, 1804, 1466]. Left [907, 375]. Left-Truncated [375]. Lemeshow [1812, 1811]. Length [1196, 1381]. Length-Constrained [1381]. Lesion [755]. Letter [960, 1402, 1687, 1744, 1126, 774, 775, 816, 454, 1072, 961, 592, 817, 653]. Leuangthong [1862]. Leukemia [219]. Level [708, 340, 1678, 53, 301, 90, 461, 32, 1185, 1211, 1453]. Levels [1727, 212, 421, 434]. Levene [65, 393]. Levie [1075]. Levy [1811]. Lewis [1226]. Li [1081]. Liberty [616]. Library [1379]. Lichtenstein [1916]. Life [1334, 1701, 1529, 1757, 845, 1216, 35, 1496, 1892, 26, 1109, 531, 1640]. Life-History [1109]. Lifetime [336, 1002, 310]. Likelihood [175, 1538, 1155, 228, 201, 1435, 1736, 747, 279, 139, 106, 230, 1446, 613, 412, 209, 304, 200, 1188, 464, 204, 301, 749, 1677, 1297, 1890, 69, 879, 63, 1125, 1714, 149, 846, 202, 173, 460, 499, 1157, 646, 673, 425, 448, 481, 1180, 1617, 789, 894, 1201, 298, 428, 1300, 1293, 420, 1846, 1806, 1458]. Likelihood-Based [279, 139, 1180, 298, 1300]. Likelihood-Ratio [106]. Likelihoods [1359, 627]. Limit [287, 1617]. Limitation [383]. Limits [1554, 360, 571]. Limnios [1910]. Line [518, 1422, 294, 403, 953, 293, 802, 622, 764, 348, 954, 568, 472]. Linear [1815, 1912, 1458, 350, 1106, 899, 604, 1777, 108, 405, 1429, 1619, 542, 354, 1215, 1354, 1398, 109, 796, 1289, 80, 943, 1765, 1369, 1459, 662, 84, 1428, 1433, 702, 458, 352, 24, 995, 394, 1714, 1567, 1766, 1248, 1888, 173, 1453, 1856, 177, 1224, 898, 806, 302, 382, 663, 1482, 1764, 933, 1695, 1319, 1459]. Linkage [703, 552, 1192, 1193]. Links [1006]. List [290, 389, 179, 841, 1537]. Lists [290, 1710, 179, 841]. Literature [128]. Literature-Based [128]. Littell [1218]. Liu [1816, 1577]. Lives [1529]. Load [570, 360]. Lobsters [427]. Local [264, 1155, 76, 555, 352, 628, 1888, 1117, 197, 227, 550, 1212]. Localization [8, 1276]. Location [1568, 1792, 158, 746, 1278, 1875]. Location-Specific [1792]. Locations [1833, 1511]. LOCF [772]. Loci [1777, 945, 1436, 299, 1038, 1190, 9, 669, 761, 1347, 1712, 1878]. Locus [8, 115, 894, 1113, 757]. Log [836, 721, 647, 112, 1248, 1600, 487, 428]. Log-Rank [836, 647, 112, 1600, 487]. Logistic [1543, 1181, 1215, 1182, 200, 184, 1369, 132, 1898, 762, 73, 19, 439, 1290, 40, 383, 124, 15, 761, 1544, 413]. Logit [12]. Logit-Normal [12]. Loglinear [674]. Lognormal [1155]. Lombardo [1801]. Long [1500, 308, 1326, 631]. Long- [631]. Long-Term [1500, 308]. Longitudinal [85, 446, 1554, 1252, 1704, 1489, 486, 586, 535, 834, 899, 1105, 772, 1556, 602, 608, 1602, 1633, 1104, 198, 1603, 1242, 999, 412, 1747, 786, 1188, 1103, 1824, 76, 1132, 100, 514, 1055, 1431, 662, 949, 1432, 1826, 1767, 695, 1653, 445, 167, 1067, 1330, 1623, 504, 1604, 1716, 1002, 1512, 452, 101, 459, 1102, 1825, 67, 804, 993, 394, 1186, 1714, 1100, 790, 1605, 1566, 153, 612, 951, 1370, 1326, 824, 1895, 71, 460, 1557, 444, 40, 383, 70, 888, 934, 1368, 1488, 1101, 1056, 1889, 1878, 1676, 1794, 1778, 7, 302, 663, 1329, 885, 1327, 1919, 1920, 1461]. Loss [616, 791, 1153, 1280, 158, 1278]. Louis [1868]. Low [636, 396, 1894, 1298, 1187]. Low-Dose [636, 1894]. Low-Rank [1298, 1187]. Lump [1842]. Lunardon [1920]. Lung [116]. Lymphocyte [438].

M [561, 521, 1086, 1090, 1266, 1317, 1270, 1318, 1584, 1646, 1702, 1813, 1814, 1523, 1639, 1571, 1080, 1578, 653, 1859, 1082, 526, 1221, 1631, 1522, 1525, 1747, 1753, 1406, 1404, 1414, 1265, 1636, 525, 522, 1360, 1073, 1517, 1909, 1640, 1264, 528, 1696]. M. [1703, 1645, 1914, 1755, 1810, 1809, 1638, 1864, 1808, 1575, 1219, 1576, 1857, 1582, 1804]. Machin [1318, 1702, 1580, 1640]. Machine [1750]. Machines [1433]. MacKinnon [1865]. Magnetic [270]. Magnusson [1087]. Mahalanobis [57]. Main [473]. Maindonald [1413]. Major [298]. Making [438]. Malaria [893]. Malthusian [1208].

Mammalian [1144]. Mammography [857]. Management [525]. Manifold [1855]. Manly [1222, 1260]. Mantel [992]. Manton [1869]. Many [1555]. Mapping [1667, 895, 1777, 1884, 221, 945, 896, 1038, 431, 1800, 1190, 321, 787, 1206, 1112, 894, 761, 1712, 1878, 1751]. Maps [5, 1157]. Marasinghe [1814]. March [561, 521, 1360]. Marginal [391, 690, 899, 228, 1366, 772, 1771, 800, 795, 1770, 1775, 514, 82, 1299, 71, 1655, 501]. Marginalized [412, 1326, 789]. Marginally [458]. Marin [1414]. Marine [220]. Mark [436, 258, 791, 59, 1771, 1331, 1734, 1738, 674, 1836]. Mark-Recapture [791, 59, 1738, 674]. Mark-Recapture-Recovery [436, 258]. Mark-Recapture-Type [1836]. Mark-Resight [791, 1734]. Marked [1306, 1771, 1051, 1734, 714]. Marker [1294, 49, 1444, 553, 1357, 908, 1400, 1006, 466]. Markers [297, 785, 157, 1667, 1725, 1196, 105, 1885, 1376, 788, 1327]. Markov [560, 1703, 1409, 1910, 1691, 332, 711, 1353, 161, 710, 465, 1902, 1612, 1724, 1880, 106, 414, 105, 1381, 1611, 147, 1431, 1781, 255, 1675, 996, 842, 341, 1910]. Markovich [1756]. Marks [1051]. Marrow [1599]. Martinez [1759]. Martinussen [1315]. Masking [145]. Mass [1549, 1122]. Matched [649, 1435, 648, 580, 345, 1597, 203, 1010, 696, 1011, 1341, 667, 1490, 482, 751, 1198, 1012, 1472]. Matched-Pair [649, 482]. Matching [1479, 938, 18]. Maternal [1445]. Maternal-Infant [1445]. Math [1087]. Mathematical [1813, 526, 1575]. Mathematics [1903]. MATLAB(R) [1759]. Matrices [353]. Matrix [1352, 448]. Matter [1739]. Matthews [1320, 1269]. Maximally [119, 1679, 130]. Maximin [17, 459]. Maximization [1666]. Maximum [1538, 880, 201, 1446, 613, 200, 204, 301, 749, 723, 63, 1714, 846, 202, 307, 1157, 646, 673, 425, 1617, 298, 856]. Maximum-Tolerated [880]. May [1812, 697]. mBIC [1667]. McClish [527]. McCulloch [1215]. Mcdonald [1222]. McLachlan [1866]. McNemar [1231]. McNemar’s [860, 130]. Mean [1334, 336, 845, 1070, 1279, 697, 537, 1621, 1179, 1258, 448, 287, 1535, 1174, 1443, 162]. Mean-Variance [1279]. Means [721, 1455, 111, 1828, 1378, 178, 428, 1607]. Measure [466]. Measured [950, 1281, 1000, 500, 712]. Measurement [1045, 294, 41, 682, 1289, 800, 1791, 1672, 1823, 643, 1871, 53, 1135, 949, 1449, 1553, 1205, 1328, 1592, 90, 764, 203, 1305, 1102, 1825, 202, 936, 473, 1204, 1290, 425, 124, 1488, 1266, 1079]. Measurements [785, 1602, 620, 199, 662, 70, 16, 1676, 529]. Measures [1215, 1456, 1906, 174, 722, 177, 255, 933, 437]. Measuring [1336, 1722, 703, 797]. Mechanisms [48, 99, 1297]. Med [65]. Median [447, 543, 1496, 1794]. Mediation [1396, 1865]. Medical [1141, 1164, 1316, 1849, 50, 111, 1623, 823, 1802, 333, 1702, 1760, 1915]. Medicine [1316, 497, 527, 1261, 1748, 1581]. Meier [937, 1500, 128, 244, 307]. Meiosis [253, 415]. Melanoma [450]. Membership [1211]. Memory [1776]. Menardi [1920]. Mendelian [1110]. Menopause [908]. Menstrual [340, 7]. Menten [1383]. Mesbah [522]. Meta [760, 1130, 1609, 128, 623, 66, 1343, 1295, 504, 1610, 1779, 1806].

Meta-Analyses [1130]. Meta-Analysis [1609, 128, 66, 1343, 1295, 1779, 623, 504, 1806]. Meta-Analytic [760, 1610]. Methicillin [1612]. Methicillin-Resistant [1612]. Method [1624, 451, 66, 682, 86, 17, 575, 537, 1046, 1677, 1890, 1484, 1010, 398, 557, 568, 1011, 1115, 1178, 719, 15, 751, 1193, 1201, 1493, 1788, 298, 1347, 1004, 1293, 1845, 420, 1012, 1699]. Methodology [1530, 1588, 1647, 1705, 1521, 69, 6]. Methods [518, 208, 1226, 1646, 1868, 1762, 1811, 1405, 1074, 1538, 1863, 1422, 201, 1435, 1407, 1697, 945, 1121, 793, 1215, 477, 1522, 590, 629, 1241, 1906, 1524, 613, 1375, 492, 1103, 1260, 1038, 1384, 1134, 1792, 1797, 1337, 355, 1176, 1807, 144, 1313, 206, 879, 1464, 995, 1299, 529, 527, 296, 588, 439, 646, 841, 1664, 1718, 124, 1139, 645, 1532, 415, 532, 1528, 1642, 1814]. Methylation [1752]. Metric [1064]. Metz [525]. MICE [1349]. Michaelis [1383]. Michailidis [1750]. Microarray [1345, 1303, 819, 1280, 704, 611, 1118, 1503, 456, 1434, 512, 1378, 1117, 1713, 638, 1545]. Microarrays [1036, 1502, 485, 1137, 1752]. Microbial [912, 416]. microRNA [1710]. Microtus [1053]. Migration [1671]. Miles [1078]. Milliken [1218]. Million [1523]. Mínguez [1813]. Minima [227]. Minimax [184]. Minimize [422]. Minimizing [227]. Minimum [632, 538, 1506, 1552]. Mining [1088, 1524, 298]. Mining-Next [1088]. Mink [610]. Misclassification [950, 1079, 102, 277, 1046, 452, 584, 1488]. Misclassified [1304, 493, 860, 426, 347, 698]. Mises [230, 1346]. Mismatches [290, 179]. Mismeasured [474, 624]. Mismeasurement [474]. Misreported [1110]. Missing [85, 991, 1571, 201, 1057, 1817, 901, 637, 77, 1780, 1361, 1724, 835, 379, 199, 943, 35, 695, 167, 1328, 355, 1898, 1716, 166, 746, 168, 1746, 951, 55, 577, 680, 1888, 1453, 1490, 1818, 646, 15, 251, 1488, 1174, 1889, 898, 1779, 1859, 1747]. Missingness [446, 696, 1717]. Misspecification [1765, 1823, 1428, 439, 888]. Mitra [1750]. Mixed [1008, 1319, 57, 350, 1106, 892, 604, 405, 152, 1429, 354, 1104, 1398, 796, 381, 1261, 664, 665, 1568, 1765, 1695, 1049, 53, 1055, 301, 749, 662, 84, 1428, 1067, 1328, 1433, 1218, 1610, 589, 458, 1549, 1390, 1601, 538, 11, 352, 459, 24, 995, 394, 1714, 1513, 229, 1766, 349, 40, 383, 789, 1009, 1224, 1101, 1187, 1063, 806, 1778, 302, 382, 663, 1329, 1482, 1764, 1076, 1857]. Mixed-Effect [1778]. Mixed-Effects [665, 1568, 1067, 1328, 1610, 24, 789]. Mixture [708, 1489, 486, 1207, 1409, 1609, 300, 1777, 181, 549, 958, 230, 304, 704, 1276, 695, 1190, 504, 1786, 69, 1822, 939, 67, 34, 957, 893, 182, 74, 1555, 675, 1895, 1392, 287, 1552, 1114, 625]. Mixtures [1207, 786, 632, 369, 890, 1346, 118, 261, 265, 63]. Mizon [1497]. MLE [1001]. MLEs [895]. MM [1429]. MM-Estimates [1429]. Model [85, 446, 1319, 410, 291, 708, 233, 1489, 486, 1053, 535, 834, 350, 1291, 1769, 1624, 269, 435, 309, 845, 1336, 1882, 427, 451, 108, 1543, 1068, 791, 710, 1660, 1153, 848, 1897, 1450, 181, 120, 1169, 1902, 1891, 122, 249, 376, 621, 1425, 354, 1602, 1654, 156, 581, 519, 477, 831, 139, 154, 1436, 1880, 835, 125, 796, 136, 1548, 618, 243, 800, 768, 1282, 476, 131, 837, 1568, 1854, 1279, 433, 1167, 1333, 1765, 1824, 1823, 1885, 849, 1066, 914, 704, 184, 470, 1670, 1608, 1049, 1795, 1183, 1540, 1723, 1652, 1499, 409, 132, 1000, 485, 749, 1203, 671, 443]. Model [662, 712, 949, 1553, 1826, 1374, 84, 368, 1653, 1781, 232, 1190, 1067, 1805, 355, 1335, 1740, 822, 82, 908, 939, 1390, 522, 398, 24, 266, 375, 1003, 67, 34, 746, 691, 1377, 684, 607, 502, 1186, 893, 1714, 1513, 74, 1605, 347, 1858, 224, 915, 1566, 612, 1536, 1766, 202, 634, 758, 19, 1895, 936, 439, 1655, 1204, 750, 1837, 1112, 33, 897, 287, 1485, 1717, 814, 70, 1224, 1545, 145, 1101, 650, 1056, 227, 159, 761, 1675, 1620, 1872, 1394, 1875, 1886, 390, 1443, 520, 842, 117, 641, 910, 1380, 639, 1114, 1607, 698, 1701, 1574, 1805, 1582]. Model-Based [1885, 398, 1545, 1620, 1443, 1574, 1582]. Model-building [1854]. Model-Checking [368]. Model-Unbiased [125]. Modeling [97, 278, 306, 619, 1356, 1554, 1252, 1470, 1416, 1704, 1812, 1550, 913, 1173, 319, 881, 303, 583, 940, 713, 1731, 1876, 1798, 300, 956, 1874, 1345, 1725, 1884, 12, 1303, 1829, 215, 1556, 602, 1569, 1518, 41, 340, 467, 569, 506, 755, 608, 279, 1007, 1104, 1774, 1410, 134, 1880, 1152, 955, 1242, 105, 1445, 339, 626, 1747, 794, 1332, 220, 1188, 1281, 1611, 1066, 100, 704, 1276, 1871, 852, 442, 14, 1055, 409, 1800, 1539, 1037, 406, 1767, 832, 1297, 544, 369, 906, 256, 1424, 388, 1466, 1740, 480, 158, 101, 668, 371, 1285, 265, 1513]. Modeling [1246, 790, 1605, 517, 612, 404, 1536, 1372, 667, 460, 1557, 1278, 1832, 1689, 1180, 609, 1379, 650, 1262, 1875, 996, 1878, 1676, 610, 625, 1607, 1586, 1813, 1636]. Modelling [1240, 1854, 1695, 1861, 1415, 1587]. Models [709, 1008, 711, 1237, 1334, 1251, 1266, 1703, 1757, 1867, 1912, 424, 530, 1239, 693, 991, 1571, 1353, 615, 1458, 1609, 1519, 586, 683, 1155, 1106, 892, 1838, 899, 1045, 292, 688, 651, 201, 604, 1057, 1899, 1777, 1105, 1721, 1070, 405, 180, 1358, 152, 161, 1429, 268, 491, 1619, 42, 549, 958, 1612, 799, 747, 236, 377, 542, 543, 150, 103, 68, 1354, 1809, 1104, 1398, 385, 682, 109, 831, 590, 106, 944, 1289, 80, 1533, 1446, 1910, 99, 381, 1381, 1261, 1159, 1282, 664, 665, 1288, 1284, 613, 104, 412, 379, 545, 786, 1234, 1116]. Models [1352, 200, 1454, 943, 1165, 1765, 1824, 1823, 1371, 643, 247, 990, 548, 1695, 1132, 1110, 346, 802, 1369, 1133, 1795, 53, 514, 1055, 301, 1459, 1431, 1452, 1776, 406, 395, 662, 1432, 1618, 695, 1166, 13, 1428, 763, 1328, 1433, 1900, 123, 504, 632, 1060, 1786, 461, 69, 764, 1497, 1610, 1822, 624, 1851, 458, 573, 1359, 1549, 1512, 1789, 1601, 206, 11, 352, 459, 995, 1102, 1825, 1311, 146, 1189, 804, 248, 63, 607, 957, 394, 891, 1115, 182, 1076, 1100, 364, 73, 1206, 229, 149, 674, 1848, 1729, 153, 1555, 675, 1047, 1737, 1567, 1720, 55, 577, 1326, 1248, 1565, 1857]. Models [771, 1888, 173, 1453, 349, 71, 1460, 646, 673, 932, 1392, 444, 40, 383, 1396, 1315, 21, 1879, 540, 163, 172, 37, 1461, 177, 789, 1009, 1448, 1739, 110, 1187, 1836, 255, 1506, 1889, 898, 996, 1063, 806, 207, 231, 1794, 905, 850, 887, 843, 1544, 1778, 1827, 302, 382, 610, 663, 1329, 1482, 1764, 341, 847, 885, 933, 523, 524, 1319, 1815, 1409, 1215, 1524, 1218, 1856, 1224, 1761, 1633]. Models-Statistics [1519]. Moderate [1326]. Modern [1266, 823, 531, 1638, 1855, 1411]. Modes [536, 1417]. Modification [1012, 423, 1010, 1011]. Modified [1667, 1722, 65, 1368, 1277]. Molecular [1241, 771, 252]. Molecules [1301]. Molenaar [1689]. Molenberghs [1859].

Moment [1423]. Momentary [1568]. Monahan [1856]. Monitor [1741]. Monitored [20]. Monitoring [1500, 378, 314, 1406, 903, 144, 1219, 1462, 854, 1388]. Monophyly [932]. Monotone [1730, 209, 858, 1742, 1615]. Monotonic [799]. Monte [1699, 1260, 1041, 1710, 1789, 1249]. Monte-Carlo [1249]. Montfort [1461]. Monthly [1371]. Morabia [1074]. Morbidity [1869]. Morphology [1846]. Mortality [561, 1869, 1068, 1172, 510, 1539, 1671, 1739]. Most [342]. Mothers [517]. Motif [1880]. Motor [64, 1210]. Moulton [1904]. Mourao [1087]. Mouse [1379]. Mover [1390]. Mover-Stayer [1390]. Moye [1316, 1585, 1462]. MRI [1754]. mRNA [1710, 564, 1711]. Muenchen [1913]. Muller [1319, 1576]. Multi [892, 1622, 849, 841, 1379]. Multi-Armed [1622]. Multi-environment [892]. Multi-Group [1379]. Multi-Library [1379]. Multi-List [841]. Multi-Type [849]. Multiallelic [670]. Multiarm [1841]. Multicategory [1726]. Multicenter [44]. Multiclass [1437]. Multidimensional [1433]. Multidimensionality [291]. Multievent [891]. Multifactorial [720]. Multilevel [1252, 1160, 1398, 664, 1305, 1367, 1837, 207, 1704]. Multilist [1393]. Multimarker [1835]. Multimarker-Based [1835]. Multimodel [1740]. Multinomial [1719, 1182, 771]. Multiple [1399, 1554, 785, 275, 490, 81, 615, 391, 690, 337, 1667, 940, 834, 27, 1669, 895, 1777, 753, 935, 12, 129, 766, 215, 543, 755, 1602, 1387, 372, 556, 835, 505, 1036, 131, 687, 314, 1332, 199, 678, 716, 672, 1000, 1432, 84, 167, 355, 1376, 1716, 274, 633, 1439, 29, 361, 1189, 694, 1360, 153, 1344, 627, 1349, 667, 719, 1440, 1718, 40, 383, 31, 1879, 638, 754, 1872, 205, 1532, 1293, 1544, 7, 1069, 1753]. Multiple-Agent [753]. Multiple-Cause [40]. Multiple-Changepoint [131]. Multiple-Community [1669]. Multiple-Interval [895]. Multiple-Peak [1349]. Multiple-Recapture [754]. Multiple-Record [719]. Multiple-Test [1344]. Multiple-Testing [766]. Multiplicative [1556, 764, 349]. Multiplicities [708]. Multiplicity [155, 859, 1287]. Multiplicity-Adjusted [859]. Multiply [580]. Multiresolution [1173]. Multiresponse [94]. Multiscale [1122]. Multisite [502]. Multistage [1169, 1205]. Multistate [1237, 710, 1660, 1425, 1653, 1512, 1658, 891]. Multivariable [1317, 1854]. Multivariate [765, 1065, 1227, 1085, 1319, 1105, 12, 1612, 844, 1082, 236, 377, 701, 150, 103, 1104, 1855, 136, 505, 800, 1722, 664, 104, 794, 199, 1577, 990, 273, 1231, 1775, 409, 724, 1776, 167, 1067, 504, 762, 24, 1186, 364, 149, 673, 1392, 416, 645, 1174, 380, 850, 843, 942, 933]. Multiway [1810]. Musings [820]. Muskrat [610]. Muskrat-Mink [610]. Mutation [1110, 572]. Mutations [1375, 1281]. Muth [1267]. Mutual [475]. mysticetus [469].

N [1269, 1320, 1270, 1271, 1528, 1813, 1634, 1695, 1805, 522, 527, 1579, 1517, 1857, 1460, 1581, 1754, 1259]. N. [1762, 1756, 1910, 1583]. Nadarajah [1083]. Nagl [1228]. Naik [1762]. Naitee [1314]. Naïve [814]. Nason [1807]. Natural [1416, 1573, 514, 684, 1220]. Nature [1244, 1526]. Near [292].

Near-Singularity [292]. Nearly [225]. Need [1741]. Needed [1451]. Negative [224, 846, 725, 21]. Nelder [1458]. Nepal [1367]. Nest [959, 1876]. Nest-Specific [1876]. Nested [1039, 1781, 1879, 559, 2]. Nesting [339, 626, 6]. Network [1832]. Neurochemical [135]. Neurotoxicology [942]. Nevada [1449, 369]. Newborns [116]. Next [1088]. Neyman [1302]. Neyman-Scott [1302]. Niang [1860]. Niche [1518]. Nicola [1920]. Niels [1916]. Nikulin [522]. Ninan [1258]. Nitrogen [61]. No [946, 1597, 692]. Noisy [229]. Non [1160, 465, 414]. Non-Gaussian [1160]. Non-Markov [465, 414]. Nonadherence [1342, 1796]. Noncompliance [1817, 769, 1781, 1797, 798, 948, 879, 67, 746, 112, 1818, 1718]. Nonconjugate [110]. Nonconstant [1447]. None [1796]. Nonhomogeneous [1611, 1601]. Nonidentifiability [644, 1626]. Nonignorable [85, 208, 486, 1817, 618, 786, 515, 696, 612, 1555, 577, 1490, 1818, 646, 1717, 1488, 1056, 1778, 1212]. Nonignorably [943, 166]. Noninferiority [509, 649, 259, 1900, 482]. Noninteger [1308]. Nonlinear [1008, 1266, 175, 152, 185, 799, 664, 665, 1847, 53, 1134, 1618, 13, 1328, 11, 1186, 425, 1009, 1175, 1063, 231, 1870, 1804]. Nonmelanoma [28]. Nonmonotone [1717]. Nonnegative [717, 1354]. Nonnormal [1482]. Nonoverlapping [252]. Nonparametric [1471, 1528, 233, 1730, 79, 1609, 1768, 396, 108, 120, 1169, 1829, 566, 1556, 1121, 1724, 78, 1292, 902, 886, 204, 1331, 1496, 1133, 442, 1134, 245, 1296, 32, 144, 244, 423, 206, 34, 1658, 1542, 265, 229, 1661, 994, 1117, 1655, 178, 1438, 1664, 1249, 814, 113, 1224, 1535, 1158, 1712, 1878, 295, 641, 911, 1711, 1382, 1756]. Nonproportional [1829, 904, 1598]. Nonrandom [197]. Nonrandomized [1177]. Nonrare [1457]. Nonresponse [486, 618, 879, 668, 1718, 1443]. Nonstandard [1070]. Nonstationary [603, 101]. Nonsuperiority [1900]. Nonzero [887]. Nordic [1916, 833]. Normal [839, 12, 721, 537, 448, 428]. Normality [16]. Normality-Based [16]. Normalization [1280, 1118]. Normalized [579]. North [1248, 6]. Note [709, 330, 601, 1306, 1628, 772, 1398, 904, 1504, 493, 558, 1359, 694, 1360, 92]. Nothing [508]. Novel [1124]. Novikov [1012]. Nuclear [552]. Nucleotide [485]. Nuisance [271, 534, 856]. Number [435, 1426, 128, 623, 1007, 1052, 1504, 1734, 18, 1210, 1427, 625, 1711]. Numbers [1358, 526, 1404, 1179, 1258]. Numerical [831]. Nutrient [16]. Nutritional [1045].

O [560, 1572, 1310, 1862, 896, 1803, 1575, 1258]. O. [1644, 1218]. Oakley [1632]. Oberman [1012]. Objective [1742]. Objectives [227]. O’Brien [902]. Observation [710, 1613, 848, 1533, 1767, 1430]. Observational [1780, 697, 1623, 1341, 1178]. Observations [1334, 311, 1724, 831, 463, 116, 1296, 883]. Observed [935, 385, 719, 1168, 70, 757]. Observer [1099]. Observers [490]. Obtained [1296]. Obuchowski [527]. Occupancy [1359, 1047]. Occupation [465, 1169]. Occupational [287]. Occurrence [1332]. Occurrences [419].

Odds [1336, 180, 1284, 409, 493, 73, 689, 260, 1163]. oeni [896]. Off [1593, 627]. Offs [756]. O’Hagan [1632]. Oligonucleotide [1279, 1117, 571]. Omnibus [305]. Oncology [903, 343, 565, 1223]. One [337, 89, 694, 1360, 448]. One-Sample [89]. One-Sided [337, 694, 1360, 448]. Onghena [1467]. Only [1040, 689, 1095, 1162, 1787, 856]. Onset [619, 1042, 170, 755, 1481]. Open [1151, 346, 832, 607, 550, 471]. Operating [161, 1062, 164]. Operatorial [1860]. Optimal [1642, 759, 1850, 1725, 1613, 56, 1616, 315, 576, 184, 216, 1340, 483, 459, 421, 316, 677, 1175, 487, 639]. Optimality [449]. Optimization [1199, 185, 371, 176]. Order [363, 1351, 621, 542, 1775, 1621, 358]. Order-Constrained [542]. Order-Preserving [1621]. Order-Restricted [621]. Ordered [27, 272, 557, 1308, 178, 1440, 1681]. Ordering [1455, 1357]. Orderings [997]. Orders [1265]. Ordinal [306, 1356, 608, 1742, 100, 1775, 1431, 1067, 1542, 680, 439, 40, 540, 54]. Organism [768]. Organisms [122]. Orthonormal [1109]. OSCAR [1491]. Other [1587, 829, 177]. Oud [1461]. Our [1275]. Outbreak [563, 1043]. Outcome [1123, 1719, 1350, 1452, 445, 860, 544, 1434, 577, 1843, 751, 1329, 420, 770]. Outcome-Dependent [445, 420]. Outcomes [956, 1874, 569, 543, 1097, 955, 687, 630, 1742, 1841, 1784, 1871, 84, 87, 624, 589, 244, 166, 411, 153, 680, 1095, 383, 1663, 1879, 380, 942, 1119, 1640]. Outcrossing [159]. Outputation [556]. Outputs [831]. Outs [198]. Overall [490]. Overdiagnosis [1480]. Overdispersed [1851, 19]. Overdispersion [914, 915]. Overidentified [1453]. Overlap [295]. Overlapping [1682]. Overviews [20]. Oxford [531].

P [1078, 1236, 1415, 1319, 1271, 1528, 1757, 1868, 1811, 1213, 1128, 1574, 1079, 1905, 653, 1810, 1632, 1520, 1854, 1414, 1411, 1467, 1576, 1641, 1640, 1581, 1908, 1696]. P. [1816, 1699, 1636, 1807, 1311, 1223, 961, 1865]. p53 [278]. Pair [649, 482]. Paired [937, 23, 259, 50, 860, 144, 244, 411, 1058]. Pairs [1384, 1597]. Pairwise [1437, 1104, 304]. Panel [410, 688, 848, 1601, 666, 1430, 1889, 641]. Panel-Count [666]. Panulims [427]. Paper [1033, 1253, 1418, 1254]. Paradigm [760, 1783]. Paradis [1468]. Parallel [1338]. Parameter [1624, 792, 156, 276, 1375, 464, 445, 553, 1469, 1208, 689, 846, 1717, 856]. Parameterization [364]. Parameters [323, 1680, 271, 436, 417, 534, 542, 198, 80, 1103, 132, 662, 832, 723, 70, 754, 1620]. Parametric [1471, 1417, 693, 583, 108, 249, 150, 768, 1885, 1134, 516, 1564, 758, 1717, 1158, 550, 390, 1443]. Parentage [587]. Parker [1012, 1010]. Parmar [1318]. Part [1566, 1502, 206]. Partial [1673, 1303, 578, 747, 277, 121, 464, 829, 362, 1455, 1714, 149, 1393, 1063]. Partially [1399, 1618, 1168, 550, 685]. Particle [1732]. Particles [1172]. Particular [299]. Particulate [1739]. Partitioning [1884, 213, 1114]. Partly [1773, 885]. Past [48, 4, 143]. Path [293]. Pathway [1433]. Patient [1850, 1303, 989, 1640, 1663, 1482, 788]. Patient-Reported [1640]. Patient-Specific [1663, 1482]. Patients [42]. Pattern [486, 181, 637, 695, 1555]. Pattern-Mixture [486]. Patterns [1869, 359, 256, 828, 1861]. Paul [1916]. Paulson [1470]. Pawitan [1458]. pbk [531]. Peak [1349]. Pearson [1199, 475]. Pebesma [1799]. Pedigree [297, 889, 1325]. Pedigrees [702, 894]. Penalized [574, 747, 511, 629, 999, 1191, 1454, 464, 149, 1493, 1620, 1845]. Penalty [464]. Penetrance [228, 1882, 1193]. Peng [1863]. pennsylvanicus [1053]. Penttinen [1861]. People [1587]. PEPFAR [1741]. Peptide [281]. Peptide-Binding [281]. Percent [1171]. Perfectly [950]. Performance [900, 958, 1161, 957, 1565]. Perinatal [338]. Period [39]. Periodic [311, 1533, 1001, 1389, 7]. Periodontal [1605]. Periods [935]. Perkins [1750]. Permanent [232]. Permutation [342, 1398, 9, 164]. Persistence [547]. Perspective [1294, 1266, 1579, 1907]. Perspectives [820, 1246]. Peto [423]. Petronis [1752]. Pharmaceutical [1267, 1470, 1749, 1463]. Pharmacodynamic [303]. Pharmacokinetic [303, 443, 712, 13, 1662, 1567, 38]. Pharmacokinetic/Pharmacodynamic [303]. Pharmacology [1243, 1520]. Pharmacometrics [1520]. Phase [1728, 1502, 767, 881, 170, 378, 396, 753, 1616, 1387, 86, 576, 1338, 468, 903, 237, 715, 343, 554, 565, 1163]. Phenotype [384, 281]. Philosophical [1244]. Photo [627]. Photoidentification [469]. Phylogenetic [1374, 1109]. Phylogenetics [1468]. Phylogeny [932]. Physical [896, 1782, 1689]. Picek [1313]. Pigeot [1407]. Pilkey [1572]. Pilkey-Jarvis [1572]. Pilot [440, 280]. Pioneer [1086]. Pioneers [820]. Placebo [39]. Placement [722]. Planning [551, 315, 1344]. Plant [1422, 94, 176]. Plasma [585]. Platform [568]. Plot [1507, 66, 1516]. Plots [1619, 835]. Plus [1085, 121]. Pocket [1270]. Point [264, 1306, 1155, 1351, 1301, 519, 1771, 1803, 1050, 1051, 1606, 1772, 254, 1896, 1861, 714, 669, 1287, 225, 520, 1511]. Points [1051, 1296]. Poisson [1917, 1918, 915, 418, 279, 151, 914, 632, 1601, 224, 1108, 947, 1552, 641, 625]. Polacheck [653]. Policies [373, 677]. Pollock [232]. Pollutants [199]. Pollution [1863, 1452, 1739]. Polya [1609]. Polyclonal [1156]. Polydesigns [1096]. Polymerase [567]. Polymorphic [299]. Polynomial [1165]. Polynomials [1854]. Polytomous [301]. Pools [163, 318]. Popular [1295]. Population [1869, 303, 417, 4, 1054, 1882, 453, 549, 1005, 958, 853, 219, 999, 1455, 587, 1608, 389, 443, 712, 606, 644, 832, 1626, 1786, 1424, 957, 1423, 675, 296, 1204, 1393, 1837, 38, 1300, 1394, 550, 585, 471, 1004, 1537, 1443]. Population-Based [1882, 453]. Populations [1811, 838, 220, 1151, 849, 681, 537, 1789, 1378, 841, 1236, 1213, 1077, 1641]. Positions [895]. Positive [1438, 752]. Positives [1504]. Possible [1555]. Possibly [1770, 1716]. Post [1534, 137, 686, 1095]. Post-Randomization [1095]. Post-Stratification [1534, 686]. Posterior [1437, 167, 1206]. Postmortem [1392]. Poststratified [1535]. Posttest [773, 635]. Potency [1183]. Potent [318]. Potential [1097]. Potentially [208, 577]. Pothole [404]. Power [759, 803, 1457, 1678, 65, 1232, 395, 1729, 173]. Powerful [342, 1548, 1597]. pp [1141, 530, 523, 526, 525, 532, 522, 529, 531, 527, 1248, 524]. Practical [1317, 1318, 1463, 127, 1414, 1695, 823, 824]. Practice [1531, 1589, 1648, 1706, 1239, 1135, 1756, 1864]. Practitioners [1646]. Pragmatic [1854]. Prairie [404]. Pre [343]. Pre-Phase [343]. Precision [545]. Predator [1624]. Predator-Prey [1624]. Predict [1572, 222]. Predicting [1303, 1384, 764, 1002]. Prediction [1529, 1693, 405, 1181, 152, 384, 1456, 60, 1110, 388, 248, 1729, 1843, 382, 1845]. Prediction-Based [384]. Prediction-Focused [1181]. Predictions [946, 1518, 664, 797, 524]. Predictive [98, 1897, 701, 837, 50, 36, 1565]. Predictiveness [1444, 1885]. Predictor [794, 718, 1542]. Predictors [1491, 840, 474, 1062, 631]. Pregnancy [580]. Premature [931]. Premutation [1831]. Prentice [760, 1071]. Prescott [1261]. Presence [1594, 845, 234, 566, 41, 1154, 1602, 239, 242, 617, 1445, 1441, 1670, 1608, 1781, 1623, 1599, 1763, 668, 1828, 1490, 473, 1168, 1192, 1787, 1598, 857]. Presence-Only [1787]. Present [48, 856]. Preserving [1621, 1396]. Presidential [1321]. Press [1141, 525, 532, 531, 528]. Pressure [620]. Pretest [773, 635]. Pretest-Posttest [773, 635]. Prevalence [1042, 453, 705, 1304, 767, 163]. Prevention [1130, 615, 307]. Prey [1624]. Primary [1824, 678, 1432]. Primer [1701, 1080, 1856, 1316, 1909]. Principal [370, 1361, 1665, 1339, 916, 1782, 917, 585]. Principles [1417]. Printing [177]. Prions [573]. Prior [1569, 1154, 285, 545, 1564]. Priors [507, 339, 504, 1605, 1830]. Probabilistic [1669, 1903, 1245]. Probabilities [62, 465, 1169, 1891, 840, 102, 414, 1196, 277, 121, 485, 644, 1047, 10, 1632]. Probability [1761, 1437, 491, 1007, 82, 538, 112, 251, 1001, 579, 176, 625, 530, 1643, 1867]. Proband [751]. Probit [771, 1886]. Problem [89, 1862, 209, 902, 591]. Problematic [95]. Problems [536, 1449, 1887, 1469]. Procedure [1541, 1621, 1010, 1011, 171, 1012]. Procedures [275, 59, 491, 42, 1753, 858, 1618, 51, 176, 1471]. Process [263, 1569, 1612, 581, 385, 1771, 1724, 1803, 1051, 105, 1824, 1601, 1484, 1186, 183, 1372, 1170, 1168, 1773, 1511, 1691]. Processes [264, 1213, 1306, 475, 710, 1613, 1351, 1240, 1631, 1050, 1606, 1772, 1611, 1432, 1896, 1390, 668, 714, 225, 1302, 1690, 1259]. Processing [1122, 1700]. Product [1187]. Production [135]. Profile [620, 1806, 481, 1773]. Profile-Likelihood [481]. Profiles [1104, 1116, 1280, 582, 1447, 1329]. Profiling [1113]. Progenies [115]. Prognostic [1434, 1565, 1069]. Program [752]. Programming [1914]. Programs [1539, 371]. Progression [467, 243, 105, 1281, 1675, 1482]. Progressive [710, 1533]. Proof [1176]. Proof-of-Concept [1176]. Propensity [357, 938, 1845]. Properties [670, 211, 1651]. Proportion [699, 1054, 304, 1006, 466, 1506]. Proportional [845, 180, 1450, 376, 68, 1494, 136, 243, 1284, 613, 1165, 409, 1000, 671, 111, 1166, 1190, 508, 461, 624, 375, 73, 71, 936, 33, 145, 698]. Proportions [900, 1500, 479, 1041, 558, 19, 482, 1685]. Proposal [502]. Proschan [1406].

Prospective [1375, 1514, 680, 994, 1327]. Prostate [1916, 833, 1480]. Protease [1666]. Protective [212]. Protein [1384, 1346]. Proteomic [1549]. Proteomics [1576, 1263]. Protocols [177]. Proximity [1436]. PSA [1480]. Pseudo [1334, 1297]. Pseudo-Likelihood [1297]. Pseudo-Observations [1334]. Pseudolikelihood [1368]. Pseudopartial [1890]. Pseudopartial-Likelihood [1890]. Pseudovalues [852]. Psychiatric [555]. Psychology [1245]. Public [1801]. Publication [1164, 1130, 679, 66, 1343, 1295]. Pulsatile [1798, 582, 1447, 1063]. Pursuit [525]. Putting [285].

Q [26]. Q-TWiST [26]. Qiu [1744]. QTL [895]. Quadrat [1154, 1052]. Quadratic [1100]. Qualification [1135]. Qualified [115]. Qualitative [351]. Quality [1334, 311, 246, 26, 1581, 310, 1640]. Quality-Adjusted [311, 246, 310]. Quality-of-Life-Adjusted [1334, 26]. Quantal [636, 48, 859]. Quantification [1071, 571]. Quantifying [912, 450, 1400, 1565, 1192]. Quantile [79, 1280, 1892, 1662, 1546, 843]. Quantitative [1777, 945, 1436, 8, 1520, 30, 1038, 1190, 9, 74, 669, 894, 1113, 159, 298, 761, 1347, 1712, 1878, 942, 1634]. Quantitative-Trait-Locus [894]. Quantities [677]. Quasi [1155, 434, 1293]. Quasi-Least [434]. Quasi-Likelihood [1155, 1293]. Questionnaire [794]. Queueing [848].

R [1078, 1236, 1089, 1085, 1266, 1472, 1417, 1762, 1913, 530, 1634, 1521, 1863, 815, 1463, 1904, 1755, 1749, 1632, 1214, 1692, 1699, 1903, 1261, 1465, 1799, 1630, 1577, 1404, 1806, 1218, 1075, 531, 1641, 1264, 1583, 1224, 1907, 1225, 1584, 1870, 1914, 1809, 1799, 1808, 1758, 1807, 1313, 1413, 1857, 1460, 1468, 1913]. R. [1227, 1646, 1813, 1759, 1867, 1631, 1220, 1700, 1696]. R.-D [1404]. Rabe [1268, 1704]. Rabe-Hesketh [1268, 1704]. Race [235]. Radiation [1792]. Radio [931, 547]. Radio-Tag [931]. Rakow [1632]. Ramsay [1310]. Random [619, 1356, 64, 503, 233, 263, 1458, 713, 604, 1899, 427, 1358, 12, 221, 1556, 150, 109, 796, 133, 1161, 151, 1765, 1066, 1369, 409, 132, 1428, 1156, 364, 1766, 439, 1204, 646, 70, 609, 789, 1675, 1552, 54, 1889, 207, 390, 1004, 1827, 302, 413, 1690, 1858]. Random-Effect [1765]. Random-Effect-Based [1889]. Random-Effects [1428, 439, 1675]. Randomised [1904]. Randomization [1408, 892, 257, 1841, 1337, 373, 1095, 677, 1119, 1388, 1260, 1467]. Randomization-Derived [892]. Randomized [785, 1269, 1320, 1500, 1817, 1622, 1728, 862, 1678, 468, 1797, 508, 87, 798, 948, 1339, 879, 1848, 250, 994, 1818, 1718, 883, 1596]. Rank [937, 1551, 836, 647, 304, 1853, 144, 112, 640, 1211, 1058, 1729, 1600, 1396, 1298, 1483, 1187, 487]. Rank-Based [937, 304, 144]. Ranked [441, 322, 801, 1054, 1710, 483, 1893, 727]. Ranked-Set [801]. Ranking [1503, 1713]. Rankings [686]. Rao [363]. Rapidly [1059]. Rare [1404, 419, 1138, 6]. Rasch [291]. Rasch-Type [291]. Ratcliff [815]. Rate [693, 586, 649, 201, 59, 234, 155, 1733, 1614, 90, 32, 34, 830, 725, 905, 1382]. Raters [275]. Rates [616, 959, 378, 549, 1171, 476, 52, 580, 1682, 321, 544, 1671, 572, 1738, 1370, 1438, 638, 748, 1482]. Ratio [350, 649, 1336, 106, 1791, 230, 104, 689, 173, 448, 484, 482, 725, 260, 1617, 428, 579, 114]. Rationale [492]. Ratios [219, 493, 532, 1657, 1120, 1163, 333]. Raton [1141]. Rattanasiri [1806]. Ratto [1909]. Ray [479]. Reaction [930, 977, 1035, 1093, 1257, 1324, 1421, 1591, 1649, 1707, 583, 567]. Reaction-Based [567]. Reader [930, 977, 1035, 1093, 1257, 1324, 1421, 1591, 1649, 1707]. Real [388]. Real-Time [388]. Really [1316]. Reanalysis [288]. Reasoning [1316, 1245, 1635]. Reassessment [451, 398, 557, 169, 1293]. Recapture [291, 1353, 436, 1538, 258, 62, 791, 931, 59, 958, 1425, 1152, 1234, 1151, 643, 548, 346, 470, 1670, 290, 644, 832, 1786, 1738, 63, 607, 957, 502, 891, 674, 1737, 1222, 10, 179, 754, 1836, 1300, 1506, 1394, 996, 390, 1004, 652]. Received [508, 588]. Receiver [1062, 164]. Receiving [1719]. Recombination [253, 415]. Reconstructing [1575]. Reconstruction [297, 1383]. Record [719]. Records [1849, 889]. Recovery [503, 436, 141, 616, 258, 292, 791, 137, 470, 232, 1671]. Recovery/Recapture [470]. Recruitment [1155]. Recurrence [886, 113]. Recurrent [410, 935, 848, 1771, 78, 617, 1177, 1333, 1630, 763, 1601, 1828, 1600, 1283, 1827]. Recurrent-Event [410]. Redetermination [437]. Reduce [1280]. Reduced [627]. Reduction [829, 1621, 1898, 18, 513, 1804]. Reestimation [792, 882]. Referees [1477]. Reference [1155]. Refinement [884]. Reflection [760]. Regimes [1340]. Region [221, 1900, 404]. Regional [1731, 1671]. Regions [953, 1682, 348, 954, 421]. Registration [797]. Registries [453]. Registry [1849, 805]. Regression [1334, 1470, 1812, 1870, 447, 175, 1628, 1373, 1491, 950, 574, 1291, 1918, 309, 801, 1057, 1899, 1543, 1070, 1181, 1619, 1891, 578, 799, 543, 150, 608, 246, 682, 1771, 1494, 1855, 1152, 1182, 165, 80, 60, 1159, 1161, 151, 999, 209, 200, 1854, 464, 1165, 1659, 914, 1892, 1695, 1770, 648, 802, 1107, 852, 1049, 1795, 1046, 132, 724, 1492, 1374, 111, 1067, 1433, 1900, 123, 1592, 1337, 355, 632, 1898, 369, 256, 1514, 589, 1002, 218, 706, 762, 700, 29, 361, 146, 584, 1285, 1115, 73, 224, 915, 1793, 1848, 55, 202, 36, 1355, 1372, 19, 1715, 1655]. Regression [473, 524, 387, 1290, 425, 750, 1430, 40, 383, 1315, 124, 1835, 163, 498, 15, 70, 251, 1224, 1448, 1483, 54, 408, 1773, 898, 1907, 1676, 1794, 843, 390, 1481, 1544, 7, 1558, 413, 1917, 1215]. Regressions [133, 507]. Regular [927, 975, 1034, 1091, 1142, 1229, 1255, 1273, 1322, 1419, 1475, 225]. Regularization [1620]. Regularizations [1628, 1492]. Regularized [1769, 1167]. Regulation [222]. Regulatory [235, 1144]. Reid [1583]. Reiss [1404]. Rejoinder [744, 1563, 1510, 815, 1150, 956, 1820, 653, 958, 242, 1365, 954, 1011, 988, 915, 961, 1009, 1920, 917, 520, 817]. Relapse [693]. Related [1912, 1735, 551, 1784, 677, 1461]. Relatedness [342]. Relating [281]. Relationship [384]. Relationships [160, 1871]. Relative [1124, 1040, 279, 285, 1165, 1183, 564, 703]. Relatives [116]. Reliability [1910, 1205, 642, 861, 1525, 1529, 1579]. Reliable [1506, 1635]. Remaining [1002]. Remission [376]. REML [350]. Removal [1005]. Removal-Based [1005]. Removals [390]. Renewal [1170]. Reparameterizing [181]. Repeated [97, 1356, 785, 1215, 102, 1894, 174, 798, 1305, 529, 752, 933, 437]. Replacement [1734]. Replicated [441, 675, 1713]. Replied [365, 138, 1686, 1401, 188, 326, 328, 400, 1309, 1233, 1403, 1127, 1688, 1684, 776, 1627, 1629, 1235, 726, 808, 1745, 1307]. Reply [593, 1918, 492, 190]. Report [738, 825, 661, 826, 1478, 1708]. Reported [1640]. Reporting [616]. Reports [533, 660, 784, 813, 878, 926, 971, 1026, 1027, 1028, 1029, 1030, 1031, 1032, 1084, 1250, 1140, 600]. Reproducibility [319]. Requirements [1678]. Requiring [271]. Resampling [411, 1139, 828, 1532]. Resampling-Based [828, 1532]. Research [676, 315, 1756, 1604, 532, 1762, 1698, 1802]. Researchers [1646]. Residual [1040, 845, 1161, 1496, 1892, 1206, 1720]. Residual-Based [1720]. Residuals [68, 368, 1764]. Resight [791, 1734]. Resistance [1375, 1281, 182]. Resistant [1612]. Resolvable [1048]. Resonance [270]. Resource [58, 1625]. Resources [1573]. Respect [989]. Respiratory [116, 1614]. Responded [140]. Response [636, 1735, 1408, 1730, 940, 62, 1057, 103, 1202, 1094, 687, 536, 160, 858, 1454, 1674, 1765, 1823, 697, 100, 1183, 1540, 14, 32, 452, 859, 1003, 1615, 1542, 421, 316, 1326, 1440, 40, 1175, 996, 1119, 191, 1]. Response-Adaptive [1408, 1119]. Responses [839, 690, 486, 12, 129, 198, 102, 35, 1328, 1716, 74, 347, 1442, 646, 383, 1056, 1889]. Responsiveness [174]. Restarting [227]. Restricted [1334, 336, 257, 621, 1289, 1900, 1469]. Restriction [804]. Restrictions [358]. Restrictive [39]. Results [440]. Retrospective [1770, 941]. Review [1078, 1236, 1228, 1088, 1089, 1086, 1225, 1227, 1087, 1090, 1141, 1226, 1267, 1415, 1268, 1470, 1471, 1269, 1320, 1473, 1319, 1266, 1317, 1472, 1270, 1271, 1318, 1416, 1528, 1584, 1643, 1646, 1642, 1644, 1585, 1703, 1527, 1586, 1645, 1587, 1701, 1702, 1704, 1529, 1757, 1812, 1813, 1868, 1759, 1762, 1869, 1816, 1866, 1911, 1815, 1760, 1761, 1915, 1870, 1867, 1912, 1913, 1811, 1814, 1914, 1213, 1128, 1523, 1693, 1572, 1405, 530, 1639, 1574, 1573, 1634, 1521, 1074, 1080, 1310, 1408, 1409, 1458, 523, 1519, 1863, 1463, 1077]. Review [1079, 1905, 1904, 1755, 196, 1578, 1580, 1407, 1862, 1810, 1749, 1081, 1697, 1632, 1214, 1518, 1859, 1082, 526, 1215, 1692, 1221, 1633, 1631, 1809, 1410, 1638, 1522, 1699, 1216, 1855, 1698, 1903, 1756, 1412, 1803, 1910, 1690, 1516, 1525, 1261, 1748, 1906, 1465, 1524, 1864, 1747, 1753, 1799, 1520, 1217, 1854, 1406, 1808, 1630, 1577, 1404, 1758, 1414, 1806, 1265, 1083, 1800, 1459, 1637, 1750, 1636, 1218, 1526, 1805, 821, 1075, 1411, 525, 1312, 1466, 1469, 1467, 1861, 1807, 1313, 532, 1860, 1314, 522, 1464, 1575, 1311, 529, 1219, 1462, 1801, 1220, 1076, 1576, 1858, 531, 1746, 527, 1579, 1222]. Review [1223, 1248, 1073, 1635, 1413, 1517, 1641, 1857, 1909, 1752, 1640, 1856, 1460, 524, 1581, 1468, 1264, 94, 1582, 1315, 1754, 1689, 1700, 1908, 1694, 1802, 1865, 1259, 1583, 1461, 1224, 1804, 1262, 528, 1696, 1907, 1263, 1691, 1751]. Reviews [965, 1014, 93, 495, 972, 973, 974, 1085, 1417, 864, 925, 777, 865, 594, 1021, 778, 1019, 1316, 866, 728, 596, 867, 868, 809, 969, 1018, 654, 1016, 729, 811, 924, 730, 922, 657, 779, 1024, 780, 963, 966, 967, 781, 964, 1023, 1020, 595, 782, 731, 732, 1025, 1695, 733, 1260, 869, 659, 812, 783, 597, 962, 598, 1022, 656, 735, 810, 870, 871, 734, 872, 193, 238, 283, 331, 366, 429, 562, 599, 1015, 970, 968, 923, 873, 921, 658, 1013, 736, 920, 874, 655, 737, 1017, 875, 876, 877, 46, 142, 401, 455]. Revised [1251, 1404]. Revisit [45]. Revisited [1188, 1395, 1158]. Reweighting [1374]. Rhythms [1008, 609, 1009]. Riani [1082]. Ribeiro [1574]. Richard [1497]. Richness [1154]. Ridout [521]. Right [239, 242, 464, 829, 907, 1330]. Right-Censored [464, 829]. Ring [503, 141]. Ring-Recovery [503, 141]. Risk [636, 1634, 1044, 44, 1040, 1736, 989, 213, 590, 952, 507, 1165, 884, 45, 1060, 449, 859, 691, 74, 1579, 484, 752, 942, 559, 2]. Risks [1498, 1768, 637, 1654, 1783, 285, 204, 852, 1499, 355, 305, 939, 1285, 1168]. Ritz [1870]. Robert [1414]. Robin [1519]. Robust [991, 1571, 935, 1429, 844, 414, 1036, 1577, 1883, 346, 1901, 232, 1919, 1118, 458, 674, 351, 1600, 934, 1920, 92, 231, 471, 639, 1313]. Robustified [1714]. Robustness [709, 1282, 1824, 1304, 69]. ROC [1356, 1735, 1291, 49, 837, 1885, 1666, 123, 1376, 1357, 51, 1125, 183, 1069, 911]. ROC-Based [1666]. Rock [427]. Rodent [477]. Rodolphe [1519]. Roff [1260]. Rogerson [1908]. Rohani [1636]. Role [760]. Root [905]. Rosenberger [1408]. Rotary [177]. Rousset [1077]. Row [1048]. Row-Column [1048]. Royston [1854]. Rubin [1472, 240]. Rubinsteina [1699]. Rubio [1799]. Rule [417]. Rules [61, 605]. Ruppert [1266]. Ryan [1411].

S [1236, 521, 1225, 1141, 1085, 1268, 1470, 1269, 1417, 1586, 1702, 1704, 1529, 1812, 1816, 1811, 1316, 1409, 1519, 815, 1755, 1580, 1215, 1410, 1522, 1698, 1910, 1753, 1799, 1083, 1750, 1466, 1467, 1860, 522, 1464, 529, 1801, 1222, 1073, 1641, 961, 1752, 1460, 1264, 528, 1696]. S-Plus [1085]. S. [1228, 1226, 1320, 1812, 1911, 1811, 1580, 1214, 1216, 1758, 1806, 1635, 1909, 1802]. S.-C [1816, 1522, 1698, 1752]. S.-Y [1466]. Sabanés [1918]. Sabelis [525]. Safety [708]. SAGE [564, 1379]. Saisana [1909]. Saltelli [1909]. Sample [323, 1559, 308, 803, 649, 27, 89, 792, 901, 1451, 705, 836, 356, 1342, 838, 721, 505, 952, 1678, 126, 1279, 769, 575, 681, 237, 217, 87, 211, 1503, 882, 1564, 26, 633, 1010, 1011, 169, 640, 1729, 1370, 745, 1796, 173, 1600, 1655, 444, 482, 1583, 1174, 437, 1012, 1698]. Sample-Size [633, 437]. Sampled [229, 1193]. Samples [322, 1098, 1036, 121, 1514, 1893, 260, 163, 1836, 1443]. Sampling [1236, 1741, 1472, 241, 441, 322, 1099, 801, 1054, 1386, 72, 344, 1005, 1154, 28, 953, 1733, 239, 242, 1289, 1052, 293, 764, 1734, 767, 348, 954, 483, 472, 171, 182, 1641, 1209, 727, 751, 1201, 420, 1811, 1573]. Sandwich [356]. Sarabia [1813]. SAS [1584, 1913, 1463, 1808, 1248, 1073, 1790, 1141, 1911, 1248, 1814]. SAS(R) [1218]. Satorra [1461]. Saturation [234]. Sauerbrei [1854]. Saunders [1529]. Saveliev [1857]. Scalar [1329]. Scale [1568, 1187].

Scale-Invariant [1187]. Scaled [135, 84]. Scaling [1064]. Scalp [288]. Scan [1286, 1883, 546]. Scanning [1501]. Scatter [1763]. Schabenberger [1218]. Schafer [1633]. Schbath [1519]. Schedule [880]. Scheike [1315]. Scheme [1055, 420]. Schemes [1201]. Scherer [363]. Schizophrenia [1392]. Schnatter [1409]. Schwarz [502]. Science [1520]. Sciences [1701, 1696, 531, 1470, 1702, 1757, 1128, 1220, 1689, 1461]. Scientific [676, 1244, 1075]. Scientists [1572, 528, 1529, 1216]. Score [541, 1683, 1776, 1166, 938, 449, 633, 224, 936, 539, 1681]. Score-Independent [1681]. Scores [357, 585]. Scoring [272, 1308]. Scott [363, 1302]. Screening [1124, 1673, 119, 25, 1616, 118, 449, 248, 307, 994, 343, 1480, 1001, 1389, 752, 1226]. Seabird [6]. Seamlessly [468]. Search [397, 591, 1395, 1112, 1082]. Seber [1053, 1153, 1425, 607, 1536]. Second [1351]. Second-Order [1351]. Secondary [276, 91]. Secretion [1297]. Secretion-Generating [1297]. Secretions [1063]. Sectional [467, 1375, 1732, 691]. See [128]. Seemingly [1045]. Segal [1696]. Segmentation [744, 918, 739, 919, 741, 226, 740, 1377, 742, 743]. Segmentation/Clustering [1377]. Segmented [802, 473]. Segments [131]. Segregation [114, 117]. SEIR [1203]. Selected [239, 242, 1679, 889, 1378, 130, 1095, 894]. Selecting [701, 512, 788]. Selection [1491, 1727, 1821, 1077, 1850, 1106, 899, 604, 1874, 1622, 1181, 1656, 989, 1891, 464, 1625, 470, 1369, 1795, 345, 1666, 1805, 1503, 1822, 1740, 158, 718, 266, 547, 577, 771, 1796, 1180, 1545, 416, 1620, 898, 176, 910, 1858]. Selective [1113]. Self [1090, 713, 239, 242, 14, 1339, 422]. Self-Designing [422]. Self-Examination [1339]. Self-Learning [1090]. Self-Modeling [14]. Self-Selected [239, 242]. Selvin [1802]. Semi [1910, 1717, 1691, 1910]. Semi-Markov [1910, 1691, 1910]. Semi-Parametric [1717]. Semiautomatic [226]. Semicompeting [1783, 1499, 1285]. Semicontinuous [949]. Semiparametric [1550, 410, 308, 535, 683, 616, 845, 1057, 1336, 1728, 945, 214, 1631, 617, 1495, 1152, 165, 1284, 795, 1151, 247, 907, 648, 1133, 1540, 409, 1184, 635, 406, 1826, 1374, 1653, 516, 1328, 1433, 1118, 369, 1385, 480, 1305, 939, 1185, 1825, 1189, 893, 202, 667, 460, 1557, 425, 1664, 113, 751, 1201, 312, 1773, 1283, 1676, 1004, 1443, 1827, 7, 1558, 420, 1746]. Semiparametrically [540]. Semivarying [1371]. Sensitivity [208, 451, 181, 570, 1747, 1659, 1064, 723, 1341, 1095, 1909, 1782, 1570, 197, 498, 1001, 1212]. Separability [1306, 714]. Separation [1898]. September [1258]. Sequence [744, 1867, 530, 918, 739, 919, 1880, 838, 741, 1381, 131, 1853, 740, 281, 742, 743, 255]. Sequences [1501, 998, 1177, 485]. Sequential [306, 937, 509, 1785, 605, 403, 792, 170, 1622, 488, 904, 276, 678, 91, 313, 144, 941, 1178, 324, 588, 439, 1664, 579, 1388]. Sequentially [20]. Serfling [1577]. Serial [707, 245, 910]. Series [98, 1903, 99, 199, 1758, 1134, 216, 1326, 524, 1607, 1081]. Series-Nonparametric [1134]. Serum [163]. Service [1529, 544]. Set [441, 322, 801, 1054, 483, 1893, 727]. Sets [1301]. Setting [1783]. Several [537, 290, 449]. Severe [1614]. Severity [64, 1332]. Sex [532]. Sexual [1604]. Shaked [1265]. Shalabh [1912]. Shantikumar [1265]. Shao [1698]. Shape [609]. Shape-Invariant [609]. Shared [1449, 763, 1003, 1717, 850]. Sharing [703, 296]. Sharp [1378]. Sheep [1061]. Sheskin [1471]. Shiboski [1215]. Shmulevich [1700]. Short [631]. Short-Term [631]. Shoukri [1584]. Should [133]. Shrinkage [1491, 353, 564, 1593, 1873, 585]. Shrinkage-based [1873]. Shrunken [1195]. Sided [337, 694, 1360, 448]. Sides [761]. Siegmund [1751]. Sigmund [525]. Sign [1595]. Signal [1700]. Signatures [771]. Signed [1551, 1058]. Signed-Rank [1551]. Significance [1118]. Significant [1654]. Silverman [1310]. SIMEX [1046]. Similarity [1098, 1669, 262, 1835]. Simon [1316]. Simple [1559, 451, 1622, 268, 129, 66, 952, 426, 324, 1212]. Simplex [118]. Simplifying [1359]. Simulation [1240, 167, 1138, 124, 1755, 1699]. Simultaneous [1680, 937, 1491, 1727, 463, 1328, 767, 1378, 1194]. Simultaneously [125]. Single [1173, 81, 753, 1783, 115, 715, 371]. Single- [1173, 753]. Single-Arm [715]. Single-Locus [115]. Singpurwalla [1579]. Singularity [292]. Site [1449, 1359, 1047, 369]. Sites [491, 254, 492]. Sivakumar [1088]. Size [803, 294, 956, 792, 280, 1451, 705, 549, 569, 836, 1342, 1650, 1698, 955, 1678, 126, 769, 575, 681, 587, 1608, 389, 237, 644, 217, 87, 1503, 882, 1564, 26, 483, 633, 1010, 1011, 169, 1423, 1729, 1370, 675, 745, 1796, 173, 296, 1600, 482, 501, 1174, 1300, 1394, 550, 471, 1004, 1537, 437, 1012]. Size/Power [803]. Sizes [1559, 1366, 1279, 87, 1626, 1299, 1732, 499, 1872]. Skill [946, 1507]. Skin [28]. Skrondal [1704]. Sliced [1628, 1492]. Small [323, 649, 356, 953, 721, 1279, 1404, 87, 211, 348, 954, 444, 1583, 1174]. Small-Sample [323, 649, 356, 721, 211, 444]. Smeeton [1528]. Smith [1857, 1089, 1517]. Smoking [1852]. Smooth [799, 796, 947, 1558]. Smoothers [628]. Smoothing [747, 1166, 321, 1424, 1299, 248, 1206, 609, 1298, 610, 523]. Smoothly [1788]. Smooths [1187]. SMR [168]. Soay [1061]. Society [1325]. Software [185, 1513, 1695, 1914]. Sojourn [25]. Solution [209, 1898]. Solutions [1214]. Solved [1862]. Some [1155, 773]. Sound [1145]. Source [1043]. Sources [1569, 861]. Souvaine [1577]. Space [1049, 1469, 1789, 1536, 805, 1259, 851]. Space-Time [1049, 805, 1259]. Sparse [1485, 992]. Spatia [262]. Spatial [518, 603, 1306, 1238, 1405, 693, 913, 1731, 475, 405, 1683, 1350, 1884, 223, 1351, 1654, 1774, 831, 1172, 134, 137, 1050, 1606, 52, 1772, 545, 1799, 1286, 1839, 1800, 406, 1896, 321, 1861, 1512, 1605, 797, 1108, 404, 714, 349, 1487, 1287, 1048, 1875, 415, 382, 1764, 1511]. Spatial-Temporal [1306, 714]. Spatially [1486, 1173, 1538, 1005, 476, 544, 1391, 1605, 675, 933]. Spatiotemporal [581, 270, 1352, 1390, 610]. Speciated [1172]. Species [435, 1098, 1154, 581, 853, 1052, 890, 1877, 159, 295]. Specific [561, 959, 1876, 956, 989, 569, 955, 339, 626, 1191, 510, 431, 1792, 662, 1111, 1828, 10, 1657, 1663, 1482]. Specificity [122]. Specified [458, 32, 1848].

Specifying [1605]. Spectrometry [1549, 1122]. Spencer [1221]. Spiegelhalter [1078]. Spline [615, 523, 834, 574, 160, 1165, 443, 1166, 321, 1601, 609, 1063, 1845]. Spline-Based [615, 1166]. Spline-Enhanced [443]. Splines [1373, 747, 999, 648, 1297, 1298, 1101]. Split [793, 1516]. Split-Cluster [793]. Spontaneous [572]. Spread [581]. Sprent [1528]. Springer [530, 523, 529, 524]. SPSS [1913, 1227]. Spurious [227]. square [58]. Squared [475]. Squares [489, 1303, 829, 1433, 434]. Stable [828]. Stage [119, 1199, 1669, 1541, 1160, 120, 465, 1169, 28, 1177, 431, 1901, 715, 217, 1337, 373, 684, 1505, 386, 745, 771, 422, 677, 1676, 231, 1443, 1329]. Stage-Specific [431]. Stain [591]. Standard [278, 709, 1356, 705, 109, 952, 716, 622, 1513, 911]. Staphylococcus [1612]. Stapleton [1761]. Stata [1252, 1268, 1704]. State [477, 1533, 1852, 1671, 1390, 1789, 1536, 1607]. State-Space [1789, 1536]. Stated [1642]. States [891, 667]. Stationarity [1606]. Statistic [363, 1199, 836, 1286, 1883, 546, 579, 1427]. Statistical [1089, 1225, 1267, 1415, 1268, 1471, 1317, 1417, 1528, 1586, 1813, 1911, 1761, 1814, 135, 1239, 1639, 1348, 497, 196, 980, 269, 649, 235, 491, 981, 1697, 982, 1121, 793, 1518, 1301, 819, 983, 1855, 1903, 1412, 896, 1241, 984, 1678, 985, 1630, 1695, 1244, 1038, 1383, 1792, 485, 1203, 1618, 1526, 1313, 1156, 1484, 1662, 979, 1635, 1752, 986, 1754, 987, 1865, 255, 430, 1198, 341, 318, 652, 1316, 529, 527, 1141, 1405, 1316, 1863, 1221, 1465, 1406, 1861, 1464, 1462, 1581, 1908, 1694, 1259, 1696, 1907]. Statisticians [1217]. Statistics [760, 1267, 1643, 1702, 1759, 1915, 1128, 1044, 1519, 1463, 1079, 1683, 392, 1240, 1243, 1679, 1853, 1807, 244, 633, 823, 1439, 130, 1729, 531, 1223, 1600, 1751, 856, 1085, 1585, 1645, 1080, 1638, 1414, 821, 1860, 1583, 1087, 1749, 1216]. Status [1169, 249, 377, 1724]. Stayer [1390]. Steel [1575]. Stefansky [1266]. Stein [1463]. Steinijans [1407]. Step [768]. Step-Stress [768]. Stepwise [1541, 1615, 1505, 176]. Stereology [1217]. Steven [817]. Stewart [1319]. Stochastic [1755, 435, 1240, 156, 581, 944, 820, 849, 1775, 1203, 1887, 1112, 1525, 1265, 1262]. Stock [1155, 222]. Stockwell [1518]. Stokes [1073]. Stopped [588]. Stopping [605]. Stoyan [1861]. Strain [896]. Strategies [1573, 1747]. Strategy [766]. Stratification [1534, 370, 1361, 686, 1339, 1178, 1393]. Stratified [948, 633, 113, 1388]. Stratum [1782]. Street [1642]. Streibig [1870]. Strength [1709, 822]. Stress [768, 822]. Stress-Strength [822]. Stroup [1218]. Structural [99, 1823, 514, 1484, 1102, 1720, 484, 725, 1416, 1466]. Structure [1077, 54, 1878, 318]. Structured [1049, 606, 540]. Structures [688, 352, 1886]. Strycker [1410]. Stuart [1916]. Studies [1741, 1164, 81, 44, 1538, 940, 803, 1457, 1039, 89, 1407, 286, 679, 931, 705, 623, 150, 1780, 477, 773, 1494, 1672, 339, 626, 1747, 795, 1885, 1479, 587, 1770, 611, 648, 1232, 514, 1901, 580, 1677, 695, 1710, 445, 167, 1592, 493, 87, 821, 1786, 203, 882, 1593, 633, 218, 703, 1010, 335, 1434, 361, 1003, 1011, 1662, 1615, 993, 1219, 547, 168, 855, 1341, 386, 745, 1661, 250, 577, 680, 1178, 1834, 1796, 432, 642, 667, 554, 1440, 1392, 1547, 38, 1583, 751, 390, 1779, 1198, 847, 437, 1012, 1816, 1859]. Study [440, 561, 1317, 157, 1863, 1850, 1882, 42, 28, 151, 1295, 555, 510, 697, 1792, 1776, 1203, 635, 1826, 1205, 1604, 1514, 700, 1782, 473, 40, 124, 1379, 645, 1394, 905, 1515]. Study/Validation [473]. Studying [768]. Stulajter [524]. Subdivided [1077]. Subject [97, 85, 446, 204, 1331, 643, 1652, 662, 1797, 452, 746, 1617, 1875]. Subject-Specific [662]. Subjectivity [528]. Subjects [699, 1387, 1843]. Subpopulations [1005]. Subsampling [315]. Subsequent [879]. Subset [998, 1095]. Substantial [18]. Substantiate [681]. Substudy [426]. Subunit [956, 569, 955, 1211]. Subunit-Specific [956, 569, 955]. Success [1007, 52, 625]. Suggestion [1112]. Suicide [1371]. Sullivan [1270]. Sum [1853, 640, 1211, 1729]. Summands [1179, 1258]. Summaries [606]. Summary [1735]. Sums [717, 1354, 1179, 1258]. Sun [1412]. Superiority [509, 716]. Superiority/Equivalence [716]. Supervised [1491, 685]. Supplemented [909]. Suppressor [158, 1278]. Supremum [836]. Surface [1094, 1183, 1540]. Surfaces [421]. Surrogacy [1783]. Surrogate [760, 785, 1294, 1495, 1665, 1784, 1071, 1400, 1006, 1247, 466, 205]. Surveillance [546, 804, 1801, 1908, 1537]. Survey [59, 257, 219, 1881]. Surveys [518, 1099, 1422, 294, 403, 676, 681, 627]. Survival [306, 1741, 937, 1334, 1812, 241, 79, 308, 535, 834, 959, 1876, 651, 309, 311, 1874, 1105, 1721, 180, 23, 1303, 1829, 566, 1891, 1556, 1602, 1215, 68, 372, 22, 246, 143, 239, 242, 1456, 80, 507, 1722, 1282, 339, 626, 49, 837, 1188, 1177, 1333, 1286, 1659, 204, 1331, 247, 769, 907, 1061, 147, 1499, 409, 1136, 406, 1677, 1826, 1599, 373, 1060, 87, 1610, 906, 144, 26, 244, 423, 700, 1434, 1658, 790, 1566, 634, 1565, 432, 1557, 673, 897, 1315, 37, 677, 1493, 1696, 487, 408, 207, 887, 469, 310, 885, 1090, 1271, 1318, 1803, 1802]. Survival-Time [566]. Survivor [1780]. Survivors [308, 862, 631]. Susceptibility [1916, 833, 1197, 1532]. Swayne [1692]. Swine [388]. Switching [1409]. Symmetrically [172]. Syndrome [1614]. Synthesis [286]. System [1103, 1537, 1073]. Systems [48, 1838, 465, 719, 1262, 1416].

T [1236, 1868, 1573, 653, 1697, 1633, 1410, 1516, 1524, 1406, 1695, 1750, 1411, 1222, 1641, 1909, 1315, 317]. T-cell [317]. T. [1225, 1866, 1573, 1632, 1695, 1314]. Table [484, 725, 1685, 1840]. Tables [900, 1451, 129, 77, 267, 1743, 1655]. Tag [616, 791, 931, 1153, 1671]. Tag-Recovery [616, 1671]. Tagging [1671]. Tail [1437]. Tailed [1756]. Takes [719]. Tanur [528]. Tarantola [1909]. Tasks [362]. Tavaré [1264]. Taylor [1141]. Technique [1507, 451]. Techniques [940, 1855, 829, 1136, 1637, 368]. Technological [456]. Technology [571]. Telemetry [931, 1625, 547]. Tell [1316]. Temporal [1306, 1670, 321, 274, 797, 714, 1773]. Temporary [346, 232]. Tensor [1187]. Teratology [261]. Term [1500, 308, 631]. Terminal [763, 1623, 1283, 1827]. Terminating [1332, 1828]. Termination [666]. Terms [1008, 1009]. Tessellation [263]. Test [417, 1043, 83, 280, 1429, 320, 670, 1551, 129, 705, 799, 623, 139, 1182, 1606, 1445, 65, 299, 902, 1441, 1901, 1614, 20, 1449, 1597, 860, 478, 369, 1497, 633, 1156, 1439, 1893, 1299, 502, 224, 640, 1211, 1058, 1729, 1344, 178, 448, 552, 164, 1174, 579, 752, 162, 1231]. Testing [264, 1529, 603, 1306, 418, 374, 391, 690, 475, 688, 1899, 23, 766, 66, 22, 904, 115, 505, 131, 1753, 126, 30, 1455, 1441, 1844, 1775, 716, 500, 1881, 91, 1335, 9, 1542, 224, 1766, 714, 334, 19, 186, 1657, 287, 638, 1547, 1532, 310, 856, 1711]. Tests [1399, 1356, 342, 1124, 1065, 937, 1348, 337, 509, 1785, 1507, 180, 161, 935, 215, 1619, 356, 1398, 1842, 267, 838, 1501, 106, 230, 1051, 647, 259, 567, 30, 914, 515, 50, 245, 1166, 45, 419, 1296, 449, 305, 522, 393, 1651, 358, 1658, 694, 1360, 1505, 915, 112, 1341, 1848, 307, 334, 173, 351, 932, 725, 1664, 1287, 947, 481, 1249, 163, 539, 1617, 757, 487, 579, 1886, 1388, 1467]. Text [1090]. Textbook [1702, 1580]. Their [417, 1098, 895, 1910, 1440, 1249, 861, 1120, 1459, 1618]. Theoretic [343, 554]. Theory [1294, 1319, 1643, 1642, 1761, 1239, 1408, 1578, 1454, 1808, 1135, 822, 1076, 1746, 861, 1635]. Therapies [42, 1849, 1338]. Therapy [1228, 905]. There [1098]. Theus [1523]. Thickness [450]. Thomas [1641, 1236]. Thompson [1526]. Three [708, 120, 477, 1533, 1678, 1844, 1357, 645]. Three-Dimensional [1844]. Three-Level [708]. Three-Stage [120]. Three-State [477, 1533]. Threshold [688, 199, 704, 347]. Thresholding [830]. Thresholds [623]. Thyroid [288]. Tidal [1621]. Tight [828]. Tiling [1381]. Tilted [1846]. Tilting [196]. Timber [664]. Time [765, 1237, 1334, 1812, 357, 424, 699, 693, 1353, 615, 374, 583, 713, 616, 1769, 62, 98, 25, 1498, 1345, 152, 1366, 1068, 1660, 1450, 1081, 1303, 566, 602, 747, 506, 701, 407, 136, 1412, 800, 99, 160, 49, 1116, 199, 1281, 886, 1167, 1611, 1331, 1770, 1107, 1049, 1723, 1134, 672, 1452, 500, 712, 245, 1781, 1190, 1890, 462, 938, 1200, 1296, 388, 1671, 1002, 146, 1189, 854, 250, 1178, 805, 1565, 1367, 460, 1557, 524, 750, 1657, 1837, 1713, 1480, 1168, 1259, 1389, 1300, 579, 408, 1394, 1676, 850, 843, 390, 1004, 1558, 1069, 1607, 1, 1758]. Time-at-Liberty [616]. Time-Course [1116]. Time-Dependent [49, 712, 938, 1178, 1565, 1069]. Time-Lagged [357]. Time-Series [199]. Time-to-Event [1237, 1812, 699, 615, 672, 500, 1190, 250, 460, 1676, 1558]. Time-to-Relapse [693]. Time-Varying [424, 506, 1452, 1781, 1002, 1557, 1657, 408]. Times [233, 308, 196, 1613, 1533, 1722, 1333, 1767, 1623, 798, 1555, 1170, 1430, 113]. Timing [338]. Tissue [479, 1392]. Tolerance [396]. Tolerated [880]. Toluene [177]. Too [137]. Tool [1212]. Tools [1605, 1137]. Topsoil [1731]. Toutenburg [1912]. Toxicities [170]. Toxicity [881, 396, 1097, 903, 1841, 1003, 317, 756, 1163, 1293]. Toxicology [48, 214, 1175]. Trade [1593, 756]. Trade-Off [1593]. Trade-Offs [756]. Training [417]. Trait [1777, 945, 1436, 8, 1038, 1190, 9, 1109, 669, 1835, 894, 1113, 761, 1347, 1712, 1878]. Traits [385, 1653, 552]. Trajectories [1002, 761]. Tramontano [1527]. Transcription [1276, 1383]. Transcriptional [1113]. Transcripts [564]. Transect [518, 1422, 294, 403, 953, 293, 622, 764, 348, 954, 568, 472]. Transform [1109, 513, 761]. Transform-Both-Sides [761]. Transformation [1611, 1610, 1827]. Transformations [713, 11]. Transformed [80, 1282]. Transfusion [691]. Transition [465, 1724, 121, 412, 1375, 1653, 589, 1102, 1825, 1001]. Transitional [85, 1512]. Translations [928]. Translators [1474]. Transmission [338, 563, 1612, 1445, 117]. Transplantation [1599]. Trawl [676]. Treated [42]. Treatment [357, 1874, 1719, 1622, 935, 989, 127, 243, 862, 126, 1177, 697, 514, 716, 635, 395, 1797, 1597, 508, 373, 1340, 879, 393, 1893, 1828, 1442, 1178, 788, 1006, 1718, 31, 1162, 677, 466, 1598, 312, 1228]. Treatment-Covariate [1162]. Treatment-Specific [1828]. Treatments [551, 1853, 716, 216, 1599, 1059]. Tree [1609, 115, 546, 1109, 408]. Tree-Based [546, 408]. Trees [724, 1355, 673]. Trend [418, 83, 320, 670, 478, 419, 706, 349, 113]. Trends [1682, 1424, 805, 1480]. Trial [699, 881, 615, 1500, 372, 1774, 1783, 276, 1338, 468, 508, 1176, 112, 680, 994, 1367, 588, 438, 861, 94]. Trials [1078, 785, 208, 839, 1316, 337, 1821, 586, 892, 649, 1817, 792, 170, 378, 396, 1622, 1358, 753, 488, 1728, 1616, 1007, 1387, 143, 1202, 1842, 86, 570, 576, 862, 314, 1678, 1406, 903, 1841, 769, 678, 237, 20, 1677, 715, 1797, 91, 692, 373, 798, 948, 313, 879, 1662, 1219, 1462, 1505, 316, 1848, 854, 324, 307, 1095, 1818, 343, 554, 1718, 422, 317, 565, 1162, 677, 1696, 806, 883, 1163, 1119, 1596, 625, 1269, 1320, 1408, 1904, 1580, 1697, 1522]. Triangular [265]. Triggers [157]. Trigonometric [717, 1354]. Trim [66]. Trophic [122]. Truncated [64, 1207, 1136, 1514, 375]. Truncation [204, 907, 1335, 444, 1168]. Trust [1883]. Tsay [1749]. Tsiatis [1746]. Tu [1638]. Tukey [94]. Tumor [619, 155, 156, 1650, 1652, 1792, 158, 1651, 1278, 444, 1675]. Tumorigenicity [590]. Tumors [278, 629, 1156]. Turner [1521]. Twin [1790, 1159, 76, 1513]. Twins [1916, 833]. TWiST [26]. Two [264, 119, 690, 308, 1199, 269, 475, 27, 1669, 336, 1541, 1160, 901, 1728, 566, 28, 463, 838, 505, 230, 1177, 1295, 1502, 1853, 174, 1901, 14, 301, 1621, 580, 715, 217, 1900, 1337, 1397, 373, 360, 461, 798, 1296, 767, 591, 558, 1185, 1390, 1505, 893, 1566, 1211, 386, 745, 1661, 179, 334, 1453, 434, 482, 725, 422, 565, 260, 677, 851, 428, 1174, 1676, 231, 1840, 1537, 1443, 1329, 1681]. Two-Arm [798]. Two-Compartment [269]. Two-Component [893]. Two-Dimensional [14, 851]. Two-Group [1296, 1211, 1661]. Two-Level [301, 461, 1185, 1453]. Two-List [1537]. Two-Part [1566]. Two-Phase [1728, 1502, 767]. Two-Sample [308, 27, 901, 838, 505, 1174]. Two-Stage [119, 1199, 1669, 1541, 1160, 28, 1177, 1901, 715, 217, 1337, 373, 1505, 386, 745, 422, 677, 1676, 231, 1443, 1329]. Two-Stain [591]. Two-State [1390]. Type [291, 356, 1456, 849, 1593, 1788, 1836, 902, 1428]. Types [637, 1602].

U [525]. U.S [1869]. U.S. [404]. Umbrella [997, 1357]. Umesh [363]. UMPU [267]. Unbalanced [1054, 716, 1726]. Unbiased [125, 702, 260]. Uncertain [891, 1632]. Uncertainty [912, 1281, 1670, 1608, 1449, 288, 1836]. Uncommon [1442]. Unconditional [259, 1597, 558, 55, 484]. Uncontrolled [218]. Underdispersed [279]. Underlying [761]. Understand [1159]. Understanding [117]. Undervalued [497]. Unequal [322, 647, 1299]. Unequally [229]. Unexamined [773]. Unidentifiable [1450]. Unified [1458, 1406, 1742, 63, 1300, 390, 1194]. Uniform [1335]. Uniparental [341]. Unit [127, 1210]. Unit-Treatment [127]. Unity [95]. Univariate [1227, 1319, 1756, 1292]. University [525, 532, 531]. Unknown [699, 427, 1652, 1734, 1179, 1258]. Unlabeled [1301]. Unmasked [145]. Unmeasured [1595]. Unmodeled [1099]. Unobserved [1569, 719]. Unpaired [23]. Unphased [1592]. Unpublished [679]. Unrecognized [395]. Unrelated [1045]. Unseen [1098]. Until [362]. Unusual [1123]. Unwin [1523]. Up-and-Down [387]. Upper [186]. US$34.95 [1248]. US$79.95/£ [1141]. Use [651, 396, 235, 223, 133, 1910, 30, 395, 478, 9, 916, 434, 539, 917, 471]. Useless [1572]. Users [1913]. Using [1008, 342, 1334, 1225, 1141, 1252, 1268, 1704, 1911, 1164, 503, 1609, 1199, 1463, 605, 616, 892, 1876, 212, 1435, 1882, 1070, 405, 1181, 12, 931, 1736, 1303, 1351, 1214, 1569, 1654, 372, 356, 1790, 239, 242, 1783, 115, 944, 768, 1722, 1159, 160, 1279, 1808, 1151, 567, 1695, 1806, 704, 346, 1383, 1384, 174, 514, 485, 1776, 406, 1374, 504, 1376, 890, 764, 1610, 1734, 261, 564, 1549, 889, 244, 1391, 51, 1062, 265, 628, 248, 63, 1505, 1186, 513, 1513, 73, 547, 74, 797, 411, 149, 1848, 1108, 612, 1737, 1766, 1248, 1073, 19, 349, 1715, 524, 387]. Using [750, 841, 1713, 1009, 1101, 1836, 761, 408, 1427, 550, 1063, 685, 1163, 1481, 1482, 1596, 341, 1382, 1881, 1851, 1413, 1085]. Utility [1666]. Utilization [544]. Utilizing [357].

V [1213, 1407, 1214, 1215, 1910, 1259]. V. [1862, 1799]. Vaccine [563, 586, 1202, 1842, 570, 1894, 1677, 1176, 1095]. Valentine [1573]. Valid [1848]. Validation [785, 426, 473]. Validity [522]. Value [83, 488, 1479, 591, 1316]. Valued [475, 698]. Values [1195, 904, 1343, 50, 889, 746, 722, 1344, 351]. Vannucci [1576]. Variability [297, 418, 1342, 9, 1875]. Variable [1491, 1821, 899, 1874, 1181, 1656, 543, 1410, 1823, 1331, 1795, 537, 1000, 301, 749, 90, 1822, 18, 1858, 153, 771, 1895, 897, 1180, 1545, 416, 1620, 898, 933]. Variables [303, 690, 1594, 372, 682, 538, 1355, 1843, 471, 1854]. Variance [1757, 818, 350, 614, 1070, 453, 1342, 1733, 1774, 1398, 65, 1279, 76, 1295, 882, 26, 393, 1109, 171, 539, 888, 1448, 92, 110, 1695]. Variance-Function [888]. Variances [534, 895]. Variate [1791]. Variation [1213, 220, 256, 261, 36, 861]. Variations [1516]. Varicella [748]. Varies [294]. Varieties [176]. Variety [892, 349]. Variogram [223]. Various [847]. Vary [989, 549]. Varying [424, 1173, 623, 506, 786, 100, 1183, 1652, 1452, 1553, 1781, 1296, 1002, 908, 1100, 1557, 499, 1657, 748, 1872, 408, 663]. Varying-Coefficient [908, 1100]. Vatutin [1213]. Vector [448]. Vedel [1217]. Vehicle [64]. Verification [1124, 515, 183, 857]. Verified [1399]. Verleysen [1804]. Version [633]. Versus [1893, 1453, 1123]. Very [513].

Verzani [1085]. vi [1141]. Via [372, 1877, 1830, 1620, 319, 1458, 8, 1611, 1055, 1710, 182, 790, 482, 1113, 806]. View [519, 254, 520, 1803]. vii [1248]. viii [530]. Viral [89, 42, 570, 360, 1740]. Virulence [525]. Virus [424]. Visit [374]. Visits [626]. Visual [172]. Visualizing [1523]. Vitamin [1367]. Vito [526]. Vittinghoff [1215]. Volterra [526]. Volume [1644, 1123, 1621, 821]. Volunteers [1515].

W [1087, 1645, 1812, 1759, 1814, 1310, 1408, 1081, 1214, 1516, 1695, 1218, 525, 532, 94]. W. [1319, 1747, 1854]. Wagenmakers [815]. Wald [356]. Wald-Type [356]. Walk [536, 1204]. Walker [1857]. Walls [1633]. Walters [1702]. Wang [1698, 1919, 1752]. Waning [506]. Ward [1612]. Waterman [1264]. Wavelet [1549, 513, 830, 1807]. Wavelet-Based [1549]. Web [1209]. WECARE [1792]. Weed [1173]. Weight [160]. Weighted [937, 489, 83, 1722, 1677, 69, 244, 1726, 112, 307, 16, 251, 1483, 1788, 487, 1427]. Weights [814]. Weinberg [1041, 1881, 757]. Weir [191]. Welch [1695]. Wellcome [1883]. West [1695]. Western [427]. Wetland [404]. Whales [627, 469]. Where [1199, 294, 697]. Whether [126]. which [1170]. While [280]. Who [1843]. Wide [1883, 1276, 1901]. Wigle [1263]. Wilcoxon [640, 1058, 1729, 1788]. Wilcoxon-Type [1788]. Wild [1789]. Wildlife [4, 1668]. Wiley [527, 528]. Wilkinson [1262]. Willan [1465]. Williams [1128, 1520]. Willis [19]. Windmills [196]. Winkel [1581]. Withdrawal [152]. Within [1652, 411, 1159, 1770]. Within-Cluster [411, 1159]. Within-Subject [1652]. Without [25, 278, 709, 1356, 1087, 1725, 1171, 500]. without/with [500]. Wittes [1406]. Wolfinger [1218]. Wombling [1896]. Women [980, 981, 982, 983, 984, 985, 580, 986, 987, 1675, 979]. Wood [1460]. Words [1519, 1519]. World [1135]. Wu [1578].

X [1638, 527, 528, 1141, 530, 479, 827, 1248]. X-Ray [479]. X.-H [527]. Xenograft [444]. Xie [1525]. xiii [523, 532]. xix [524]. xv [531, 527]. xvii [525]. xxiv [529]. xxxiii [522].

Y. [1749]. Yakir [1751]. Yamada [1908]. Yeast [573, 1117]. Yes [946]. Yes/No [946]. Yesha [1088]. Yield [664, 723, 176]. York [530, 523, 529, 527, 524, 528]. Young [1089].

Z [1915]. Zelterman [1248]. Zero [151, 914, 1184, 224, 915, 484, 725, 947, 887]. Zero-Inflated [151, 914, 1184, 224, 915, 947]. Zeros [1871, 1566, 162]. Zhang [1581]. Zhou [527]. Zidek [1259]. Zuur [1517, 1857].

References

Kodlin:1967:NRT

[1] D. Kodlin. A new response time distribution. Biometrics, 23(??):227– 239, 1967. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [863].

Langholz:1997:EAR

[2] B. Langholz and Ø. Borgan. Estimation of absolute risk from nested casecontrol data. Biometrics, 53(??):767–774, June 1997. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [559].

Ridout:1999:EIC

[3] Martin S. Ridout, Clarice G. B. Demétrio, and David Firth. Estimating intraclass correlation for binary data. Biometrics, 55(1):137–148, March 1999. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [521].

Buckland:2000:WPA

[4] S. T. Buckland, I. B. J. Goudie, and D. L. Borchers. Wildlife population assessment: Past developments and future directions. Biometrics, 56(1): 1–12, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Knorr-Held:2000:BDC

[5] Leonhard Knorr-Held and Günter Raßer. Bayesian detection of clusters and discontinuities in disease maps. Biometrics, 56(1):13–21, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Welsh:2000:MEA

[6] A. H. Welsh, R. B. Cunningham, and R. L. Chambers. Methodology for estimating the abundance of rare animals: Seabird nesting on North East herald cay. Biometrics, 56(1):22–30, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2000:SRP

[7] Daowen Zhang, Xihong Lin, and MaryFran Sowers. Semiparametric regression for periodic longitudinal hormone data from multiple menstrual cycles. Biometrics, 56(1):31–39, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

George:2000:LQT

[8] A. W. George, K. L. Mengersen, and G. P. Davis. Localization of a quantitative trait locus via a Bayesian approach. Biometrics, 56(1):40– 51, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nettleton:2000:AVU

[9] Dan Nettleton and R. W. Doerge. Accounting for variability in the use of permutation testing to detect quantitative trait loci. Biometrics, 56 (1):52–58, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwarz:2000:EAS

[10] Carl J. Schwarz and A. Neil Arnason. Estimation of age-specific breeding probabilities from capture–recapture data. Biometrics, 56(1):59–64, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oberg:2000:EDT

[11] Ann Oberg and Marie Davidian. Estimating data transformations in nonlinear mixed effects models. Biometrics, 56(1):65–72, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coull:2000:REM

[12] Brent A. Coull and Alan Agresti. Random effects modeling of multiple binomial responses using the multivariate binomial logit-normal distribution. Biometrics, 56(1):73–80, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lindsey:2000:GNM

[13] J. K. Lindsey, W. D. Byrom, J. Wang, P. Jarvis, and B. Jones. Generalized nonlinear models for pharmacokinetic data. Biometrics, 56(1): 81–88, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ladd:2000:SMT

[14] William M. Ladd and Mary J. Lindstrom. Self-modeling for twodimensional response curves. Biometrics, 56(1):89–97, March 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2000:CCE

[15] C. Y. Wang and Wen-Tao Huang. Conditional-cumulant-of-exposure method in logistic missing covariate regression. Biometrics, 56(1):98– 105, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2000:WNB

[16] C. Y. Wang. Weighted normality-based estimator in correcting correlation coefficient estimation between incomplete nutrient measurements. Biometrics, 56(1):106–112, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Imhof:2000:GMF

[17] Lorens Imhof and Weng Kee Wong. A graphical method for finding maximin efficiency designs. Biometrics, 56(1):113–117, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ming:2000:SGB

[18] Kewei Ming and Paul R. Rosenbaum. Substantial gains in bias reduction from matching with a variable number of controls. Biometrics, 56(1):118– 124, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Slaton:2000:ETO

[19] Terra L. Slaton, Walter W. Piegorsch, and Stephen D. Durham. Estimation and testing with overdispersed proportions using the beta-logistic regression model of Heckman and Willis. Biometrics, 56(1):125–133, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2000:HTO

[20] Zhengqing Li and Yan Li. A homogeneity test in overviews with group sequentially monitored clinical trials. Biometrics, 56(1):134–138, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thurston:2000:NBA

[21] Sally W. Thurston, M. P. Wand, and John K. Wiencke. Negative binomial additive models. Biometrics, 56(1):139–144, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fine:2000:TDS

[22] Jason P. Fine and Anastasios A. Tsiatis. Testing for differences in survival with delayed ascertainment. Biometrics, 56(1):145–153, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dallas:2000:TES

[23] Michael J. Dallas and P. V. Rao. Testing equality of survival functions based on both paired and unpaired censored data. Biometrics, 56(1):154– 159, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2000:LME

[24] Wei Pan and Thomas A. Louis. A linear mixed-effects model for multivariate censored data. Biometrics, 56(1):160–166, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2000:EST

[25] Tony H. H. Chen, H. S. Kuo, M. F. Yen, M. S. Lai, L. Tabar, and S. W. Duffy. Estimation of sojourn time in chronic disease screening without data on interval cases. Biometrics, 56(1):167–172, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murray:2000:VSS

[26] Susan Murray and Bernard Cole. Variance and sample size calculations in quality-of-life-adjusted survival analysis (Q-TWiST). Biometrics, 56 (1):173–182, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:2000:TSC

[27] Shu-Hui Chang. A two-sample comparison for multiple ordered event data. Biometrics, 56(1):183–189, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fears:2000:ATS

[28] Thomas R. Fears and Mitchell H. Gail. Analysis of a two-stage casecontrol study with cluster sampling of controls: Application to nonmelanoma skin cancer. Biometrics, 56(1):190–198, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2000:MIA

[29] Wei Pan. A multiple imputation approach to Cox regression with interval-censored data. Biometrics, 56(1):199–203, March 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hung:2000:UBG

[30] Ming-Chin Hung and William H. Swallow. Use of binomial group testing in tests of hypotheses for classification or quantitative covariables. Biometrics, 56(1):204–212, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2000:ABE

[31] Peter F. Thall, Richard M. Simon, and Yu Shen. Approximate Bayesian evaluation of multiple treatment effects. Biometrics, 56(1):213–219, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mukhopadhyay:2000:BNI

[32] Saurabh Mukhopadhyay. Bayesian nonparametric inference on the dose level with specified response rate. Biometrics, 56(1):220–226, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sy:2000:ECP

[33] Judy P. Sy and Jeremy M. G. Taylor. Estimation in a Cox proportional hazards cure model. Biometrics, 56(1):227–236, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peng:2000:NMM

[34] Yingwei Peng and Keith B. G. Dear. A nonparametric mixture model for cure rate estimation. Biometrics, 56(1):237–243, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Izmirlian:2000:ALE

[35] Grant Izmirlian, Dwight Brock, Luigi Ferrucci, and Caroline Phillips. Active life expectancy from annual follow-up data with missing responses. Biometrics, 56(1):244–248, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schemper:2000:PAE

[36] Michael Schemper and Robin Henderson. Predictive accuracy and explained variation in Cox regression. Biometrics, 56(1):249–255, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Volinsky:2000:BIC

[37] Chris T. Volinsky and Adrian E. Raftery. Bayesian information criterion for censored survival models. Biometrics, 56(1):256–262, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wakefield:2000:CPP

[38] Jonathan Wakefield and Nargis Rahman. The combination of population pharmacokinetic studies. Biometrics, 56(1):263–270, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Landin:2000:IRE

[39] R. Landin, D. J. DeBrota, T. A. DeVries, W. Z. Potter, and M. A. Demitrack. The impact of restrictive entry criterion during the placebo lead-in period. Biometrics, 56(1):271–278, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

TenHave:2000:MEL

[40] Thomas R. Ten Have, Michael E. Miller, Beth A. Reboussin, and Margaret K. James. Mixed effects logistic regression models for longitudinal ordinal functional response data with multiple-cause drop-out from the longitudinal study of aging. Biometrics, 56(1):279–287, March 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2000:MHF

[41] David B. Dunson and Clarice R. Weinberg. Modeling human fertility in the presence of measurement error. Biometrics, 56(1):288–292, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ding:2000:CSM

[42] A. Adam Ding and Hulin Wu. A comparison study of models and fitting procedures for biphasic viral dynamics in HIV-1 infected patients treated with antiviral therapies. Biometrics, 56(1):293–300, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:Ca

[43] Anonymous. Correspondence. Biometrics, 56(1):301–303, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohning:2000:ERD

[44] Dankmar Böhning and Jesus Sarol Jr. Estimating risk difference in multicenter studies under baseline-risk heterogeneity. Biometrics, 56(1):304– 308, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lui:2000:RTH

[45] Kung-Jong Lui and Colleen Kelly. A revisit on tests for homogeneity of the risk difference. Biometrics, 56(1):309–315, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:BRa

[46] Anonymous. Book reviews. Biometrics, 56(1):316–323, March 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:Cb

[47] Anonymous. Corrections. Biometrics, 56(1):324–325, March 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bailer:2000:QCM

[48] A. John Bailer and Walter W. Piegorsch. From quantal counts to mechanisms and systems: The past, present, and future of biometrics in environmental toxicology. Biometrics, 56(2):327–336, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heagerty:2000:TDR

[49] Patrick J. Heagerty, Thomas Lumley, and Margaret S. Pepe. Timedependent ROC curves for censored survival data and a diagnostic marker. Biometrics, 56(2):337–344, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leisenring:2000:CPV

[50] Wendy Leisenring, Todd Alono, and Margaret Sullivan Pepe. Comparisons of predictive values of binary medical diagnostic tests for paired designs. Biometrics, 56(2):345–351, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pepe:2000:IRC

[51] Margaret Sullivan Pepe. An interpretation for the ROC curve and inference using GLM procedures. Biometrics, 56(2):352–359, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2000:HBE

[52] Zhuoqiong He and Dongchu Sun. Hierarchical Bayes estimation of hunting success rates with spatial correlations. Biometrics, 56(2):360–367, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ko:2000:CME

[53] Hyejin Ko and Marie Davidian. Correcting for measurement error in individual-level covariates in nonlinear mixed effects models. Biometrics, 56(2):368–375, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xie:2000:REC

[54] Minge Xie, Douglas G. Simpson, and Raymond J. Carroll. Random effects in censored ordinal regression: Latent structure and Bayesian approach. Biometrics, 56(2):376–383, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Satten:2000:CUC

[55] Glen A. Satten and Raymond J. Carroll. Conditional and unconditional categorical regression models with missing covariates. Biometrics, 56 (2):384–388, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

David:2000:OCB

[56] Olivier David, Hervé Monod, and Joël Amoussou. Optimal complete block designs to adjust for interplot competition with a covariance analysis. Biometrics, 56(2):389–393, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bedrick:2000:EMD

[57] Edward J. Bedrick, Jodi Lapidus, and Joseph F. Powell. Estimating the Mahalanobis distance from mixed continuous and discrete data. Biometrics, 56(2):394–401, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dasgupta:2000:CSG

[58] Nairanjana Dasgupta and J. Richard Alldredge. A chi-square goodnessof-fit analysis of dependent resource election data. Biometrics, 56(2): 402–408, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dahiya:2000:EER

[59] Ram C. Dahiya, Shampa Saha, and Hassan Lakkis. Estimation of exploitation rate by combining mark-recapture procedures with a creel survey. Biometrics, 56(2):409–413, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gregoire:2000:PIR

[60] Timothy G. Gregoire, Oliver Schabenberger, and Fanzhi Kong. Prediction from an integrated regression equation: a forestry application. Biometrics, 56(2):414–419, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Antoniadou:2000:EDR

[61] T. Antoniadou and D. Wallach. Evaluating decision rules for nitrogen fertilization. Biometrics, 56(2):420–426, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chao:2000:CRW

[62] Anne Chao, Wenten Chu, and Chiu-Hsieh Hsu. Capture–recapture when time and behavioral response affect capture probabilities. Biometrics, 56 (2):427–433, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pledger:2000:UML

[63] Shirley Pledger. Unified maximum likelihood estimates for closed capture–recapture models using mixtures. Biometrics, 56(2):434–442, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barry:2000:AGT

[64] S. C. Barry and T. J. O’Neill. The analysis of group truncated binary data with random effects: Injury severity in motor vehicle accidents. Biometrics, 56(2):443–450, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hines:2000:IPM

[65] W. G. S. Hines and R. J. O’Hara Hines. Increased power with modified forms of the Levene (med) test for heterogeneity of variance. Biometrics, 56(2):451–454, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duval:2000:TFS

[66] Sue Duval and Richard Tweedie. Trim and fill: a simple funnel-plot-based method of testing and adjusting for publication bias in meta-analysis. Biometrics, 56(2):455–463, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pauler:2000:MML

[67] Donna K. Pauler and Nan M. Laird. A mixture model for longitudinal data with application to assessment of noncompliance. Biometrics, 56 (2):464–472, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Farrington:2000:RPH

[68] C. P. Farrington. Residuals for proportional hazards models with interval-censored survival data. Biometrics, 56(2):473–482, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Markatou:2000:MMR

[69] Marianthi Markatou. Mixture models, robustness, and the weighted likelihood methodology. Biometrics, 56(2):483–486, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2000:RAW

[70] C. Y. Wang, Naisyin Wang, and Suojin Wang. Regression analysis when covariates are regression parameters of a random effects model for observed longitudinal measurements. Biometrics, 56(2):487–495, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2000:MML

[71] Peter Xue-Kun Song and Ming Tan. Marginal models for longitudinal continuous proportional data. Biometrics, 56(2):496–502, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christman:2000:BCI

[72] Mary C. Christman and Jeffrey S. Pontius. Bootstrap confidence intervals for adaptive cluster sampling. Biometrics, 56(2):503–510, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rabinowitz:2000:UCL

[73] Daniel Rabinowitz, Rebecca A. Betensky, and Anastasios A. Tsiatis. Using conditional logistic regression to fit proportional odds models to interval censored data. Biometrics, 56(2):511–518, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Razzaghi:2000:RAQ

[74] Mehdi Razzaghi and Ralph L. Kodell. Risk assessment for quantitative responses using a mixture model. Biometrics, 56(2):519–527, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lipsitz:2000:GGE

[75] Stuart R. Lipsitz, Geert Molenberghs, Garrett M. Fitzmaurice, and Joseph Ibrahim. GEE with Gaussian estimation of the correlations when data are incomplete. Biometrics, 56(2):528–536, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huggins:2000:LEA

[76] Richard Huggins. Local estimation of age-dependent variance components from longitudinal twin data. Biometrics, 56(2):537–545, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dominici:2000:CCT

[77] Francesca Dominici. Combining contingency tables with missing dimensions. Biometrics, 56(2):546–553, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2000:NAR

[78] Debashis Ghosh and D. Y. Lin. Nonparametric analysis of recurrent events and death. Biometrics, 56(2):554–562, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bowman:2000:GEC

[79] A. W. Bowman and E. M. Wright. Graphical exploration of covariate effects on survival data through nonparametric quantile curves. Biometrics, 56(2):563–570, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gray:2000:ERP

[80] Robert J. Gray. Estimation of regression parameters and the hazard function in transformed linear survival models. Biometrics, 56(2):571– 576, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Basu:2000:BIK

[81] Sanjib Basu, Mousumi Banerjee, and Ananda Sen. Bayesian inference for kappa from single and multiple studies. Biometrics, 56(2):577–582, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nam:2000:IEK

[82] Jun-Mo Nam. Interval estimation of the kappa coefficient with binary classification and an equal marginal probability model. Biometrics, 56 (2):583–585, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2000:WVA

[83] James J. Chen, Karl K. Lin, Mohammad Huque, and Ramin B. Arani. Weighted p-value adjustments for animal carcinogenicity trend test. Biometrics, 56(2):586–592, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2000:SLM

[84] Xihong Lin, Louise Ryan, Mary Sammel, Daowen Zhang, Chantana Padungtod, and Xiping Xu. A scaled linear mixed model for multiple outcomes. Biometrics, 56(2):593–601, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2000:TML

[85] Paul S. Albert. A transitional model for longitudinal binary data subject to nonignorable missing data. Biometrics, 56(2):602–608, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gasparini:2000:CFM

[86] Mauro Gasparini and Jeffrey Eisele. A curve-free method for Phase I clinical trials. Biometrics, 56(2):609–615, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Manatunga:2000:SSE

[87] Amita K. Manatunga and Shande Chen. Sample size estimation for survival outcomes in cluster-randomized studies with small cluster sizes. Biometrics, 56(2):616–621, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sutradhar:2000:AEE

[88] Brajendra C. Sutradhar and Kalyan Das. On the accuracy of efficiency of estimating equation approach. Biometrics, 56(2):622–625, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2000:OSP

[89] Debbie M. Cheng and Stephen W. Lagakos. The one-sample problem from eradication studies of chronic viral infections. Biometrics, 56(2): 626–633, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lyles:2000:ECM

[90] Robert H. Lyles and Glen McFarlane. Effects of covariate measurement error in the initial level and rate of change of an exposure variable. Biometrics, 56(2):634–639, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2000:TSH

[91] Aiyi Liu, Ming Tan, James M. Boyett, and Xiaoping Xiong. Testing secondary hypotheses following sequential clinical trials. Biometrics, 56 (2):640–644, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:2000:NRV

[92] Rick L. Williams. A note on robust variance estimation for clustercorrelated data. Biometrics, 56(2):645–646, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:BRb

[93] Anonymous. Book reviews. Biometrics, 56(2):647–648, June 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Talbot:2000:RGA

[94] M. Talbot. Review: Graphical Analysis of Multiresponse Data: Illustrated with a Plant Breeding Trial, by K. E. Basford and J. W. Tukey. Biometrics, 56(2):649–650, June 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://www.jstor.org/

stable/2677019.

Stigler:2000:PUB

[95] Stephen M. Stigler. The problematic unity of biometrics. Biometrics, 56(3):653–658, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Elston:2000:CBG

[96] Robert C. Elston and Elizabeth A. Thompson. A century of biometrical genetics. Biometrics, 56(3):659–666, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2000:MRC

[97] Paul S. Albert and Dean A. Follmann. Modeling repeated count data subject to informative dropout. Biometrics, 56(3):667–677, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2000:BPI

[98] Ming-Hui Chen and Joseph G. Ibrahim. Bayesian predictive inference for time series count data. Biometrics, 56(3):678–685, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2000:STS

[99] Wensheng Guo and Morton B. Brown. Structural time series models with feedback mechanisms. Biometrics, 56(3):686–691, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kauermann:2000:MLD

[100] Göran Kauermann. Modeling longitudinal data with ordinal response by varying coefficients. Biometrics, 56(3):692–698, September 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nunez-Anton:2000:MNL

[101] Vicente Núñez-Antón and Dale L. Zimmerman. Modeling nonstationary longitudinal data. Biometrics, 56(3):699–705, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fujisawa:2000:IAM

[102] Hironori Fujisawa and Shizue Izumi. Inference about misclassification probabilities from repeated binary responses. Biometrics, 56(3):706–711, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ekholm:2000:AMM

[103] Anders Ekholm, John W. McDonald, and Peter W. F. Smith. Association models for a multivariate binary response. Biometrics, 56(3):712–718, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heagerty:2000:MCR

[104] Patrick J. Heagerty and Scott L. Zeger. Multivariate continuation ratio models: Connections and caveats. Biometrics, 56(3):719–732, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guihenneuc-Jouyaux:2000:MMD

[105] Chantal Guihenneuc-Jouyaux, Sylvia Richardson, and Ira M. Longini Jr. Modeling markers of disease progression by a hidden Markov process:

Application to characterizing CD4 cell decline. Biometrics, 56(3):733– 741, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giudici:2000:LRT

[106] Paolo Giudici, Tobias Ryden, and Pierre Vandekerkhove. Likelihoodratio tests for hidden Markov models. Biometrics, 56(3):742–747, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christen:2000:AAC

[107] J. Andrés Christen and Miguel Nakamura. On the analysis of accumulation curves. Biometrics, 56(3):748–754, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chiu:2000:NPE

[108] S. N. Chiu, M. P. Quine, and M. Stewart. Nonparametric and parametric estimation for a linear germination-growth model. Biometrics, 56(3): 755–760, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Friedl:2000:SEE

[109] Herwig Friedl and Göran Kauermann. Standard errors for EM estimates in generalized linear models with random effects. Biometrics, 56(3): 761–767, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wolfinger:2000:NBA

[110] Russell D. Wolfinger and Robert E. Kass. Nonconjugate Bayesian analysis of variance component models. Biometrics, 56(3):768–774, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2000:PMR

[111] D. Y. Lin. Proportional means regression for censored medical costs. Biometrics, 56(3):775–778, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robins:2000:CND

[112] James M. Robins and Dianne M. Finkelstein. Correcting for noncompliance and dependent censoring in an AIDS clinical trial with inverse probability of censoring weighted (IPCW) log-rank tests. Biometrics, 56(3):779–788, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2000:NST

[113] Mei-Cheng Wang and Ying-Qing Chen. Nonparametric and semiparametric trend analysis for stratified recurrence times. Biometrics, 56(3): 789–794, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yao:2000:BCS

[114] Yi-Ching Yao and John Jen Tai. Bias correction for segregation ratio estimation in human genetics. Biometrics, 56(3):795–800, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gillet:2000:QTS

[115] Elizabeth M. Gillet and Hans-Rolf Gregorius. Qualified testing of singlelocus codominant inheritance using single tree progenies. Biometrics, 56(3):801–807, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Houwing-Duistermaat:2000:EIG

[116] Jeanine J. Houwing-Duistermaat, Hans C. van Houwelingen, and J. Peter de Winter. Estimation of individual genetic effects from binary observations on relatives applied to a family history of respiratory illnesses and chronic lung disease of newborns. Biometrics, 56(3):808–814, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2000:FMS

[117] Heping Zhang and Kathleen Merikangas. A frailty model of segregation analysis: Understanding the familial transmission of alcoholism. Biometrics, 56(3):815–823, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McConkey:2000:FSD

[118] Brendan J. McConkey, Paul G. Mezey, D. George Dixon, and Bruce M. Greenberg. Fractional simplex designs for interaction screening in complex mixtures. Biometrics, 56(3):824–832, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berger:2000:MET

[119] Toby Berger, James W. Mandell, and P. Subrahmanya. Maximally efficient two-stage screening. Biometrics, 56(3):833–840, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Datta:2000:NET

[120] Somnath Datta, Glen A. Satten, and Susmita Datta. Nonparametric estimation for the three-stage irreversible illness-death model. Biometrics, 56(3):841–847, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hawkins:2000:ETP

[121] D. L. Hawkins and Chien-Pai Han. Estimating transition probabilities from aggregate samples plus partial transition data. Biometrics, 56(3): 848–854, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diserud:2000:BBM

[122] Ola H. Diserud and Rode Ødegaard. The beta-binomial model for host specificity among organisms in trophic interactions. Biometrics, 56(3): 855–861, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lloyd:2000:RMC

[123] Chris J. Lloyd. Regression models for convex ROC curves. Biometrics, 56(3):862–867, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thoresen:2000:SSM

[124] Magne Thoresen and Petter Laake. A simulation study of measurement error correction methods in logistic regression. Biometrics, 56(3):868– 872, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerow:2000:SMU

[125] Ken Gerow and Charles E. McCulloch. Simultaneously model-unbiased, design-unbiased estimation. Biometrics, 56(3):873–878, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Horn:2000:SSD

[126] Manfred Horn, Rudiger Vollandt, and Charles W. Dunnett. Sample size determination for testing whether an identified treatment is best. Biometrics, 56(3):879–881, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gadbury:2000:UTI

[127] Gary L. Gadbury and Hari K. Iyer. Unit-treatment interaction and its practical consequences. Biometrics, 56(3):882–885, September 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duchateau:2000:ENE

[128] Luc Duchateau, Laurence Collette, Richard Sylvester, and Jean-Pierre Pignon. Estimating number of events from the Kaplan–Meier curve for incorporation in a literature-based meta-analysis: What you don’t see you can’t get! Biometrics, 56(3):886–892, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Decady:2000:STA

[129] Y. J. Decady and D. R. Thomas. A simple test of association for contingency tables with multiple column responses. Biometrics, 56(3):893–896, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rabinowitz:2000:ADM

[130] Daniel Rabinowitz and Rebecca A. Betensky. Approximating the distribution of maximally selected McNemar’s statistics. Biometrics, 56(3): 897–902, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Halpern:2000:MCT

[131] Aaron L. Halpern. Multiple-changepoint testing for an alternating segments model of a binary sequence. Biometrics, 56(3):903–908, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Larsen:2000:IPL

[132] Klaus Larsen, Jørgen Holm Petersen, Esben Budtz-Jørgensen, and Lars Endahl. Interpreting parameters in the logistic regression model with random effects. Biometrics, 56(3):909–914, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenland:2000:WSE

[133] Sander Greenland. When should epidemiologic regressions use random coefficients? Biometrics, 56(3):915–921, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gangnon:2000:BDM

[134] Ronald E. Gangnon and Murray K. Clayton. Bayesian detection and modeling of spatial disease clustering. Biometrics, 56(3):922–935, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baileru:2000:SCA

[135] A. John Baileru, Siyu Liu, Michelle L. Smith, and Lori G. Isaacson. Statistical comparison of axon-scaled neurochemical production. Biometrics, 56(3):936–939, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goggins:2000:PHM

[136] William B. Goggins and Dianne M. Finkelstein. A proportional hazards model for multivariate interval-censored failure time data. Biometrics, 56(3):940–943, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilmour:2000:PBG

[137] Arthur R. Gilmour. Post blocking gone too far! recovery of information and spatial analysis in field experiments. Biometrics, 56(3):944–945, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:2000:ARF

[138] Walter T. Federer. The author replied as follows. Biometrics, 56(3): 945–946, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Garren:2000:LBG

[139] Steven T. Garren, Richard L. Smith, and Walter W. Piegorsch. On a likelihood-based goodness-of-fit test of the beta-binomial model. Biometrics, 56(3):947–949, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brooks:2000:ARF

[140] Steve Brooks, Byron Morgan, and Martin Ridout. The authors responded as follows. Biometrics, 56(3):950, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brooks:2000:BAR

[141] S. P. Brooks, E. A. Catchpole, B. J. T. Morgan, and S. C. Barry. On the Bayesian analysis of ring-recovery data. Biometrics, 56(3):951–956, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2000:BRa

[142] M. S. Ridout. Book reviews. Biometrics, 56(3):957–967, September 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fleming:2000:SAC

[143] Thomas R. Fleming and D. Y. Lin. Survival analysis in clinical trials: Past developments and future directions. Biometrics, 56(4):971–983, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murray:2000:NRB

[144] Susan Murray. Nonparametric rank-based methods for group sequential monitoring of paired censored survival data. Biometrics, 56(4):984–990, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:2000:MUP

[145] Wen Hsiang Wei and Michael R. Kosorok. Masking unmasked in the proportional hazards model. Biometrics, 56(4):991–995, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parker:2000:DCR

[146] Corette B. Parker and Elizabeth R. Belong. A diagnostic for Cox regression with discrete failure-time models. Biometrics, 56(4):996–1001, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kozumi:2000:BAD

[147] Hideo Kozumi. Bayesian analysis of discrete survival data with a hidden Markov chain. Biometrics, 56(4):1002–1006, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McKeague:2000:BEC

[148] Ian W. McKeague and Mourad Tighiouart. Bayesian estimators for conditional hazard functions. Biometrics, 56(4):1007–1015, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ripatti:2000:EMF

[149] Samuli Ripatti and Juni Palmgren. Estimation of multivariate frailty models using penalized partial likelihood. Biometrics, 56(4):1016–1022, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ecochard:2000:MPR

[150] René Ecochard and David G. Clayton. Multivariate parametric random effect regression models for fecundability studies. Biometrics, 56(4):1023– 1029, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hall:2000:ZIP

[151] Daniel B. Hall. Zero-inflated Poisson and binomial regression with random effects: a case study. Biometrics, 56(4):1030–1039, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Concordet:2000:CNM

[152] Didier Concordet and Olivier G. Nunez. Calibration for nonlinear mixed effects models: an application to the withdrawal time prediction. Biometrics, 56(4):1040–1046, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:2000:LVM

[153] Jason Roy and Xihong Lin. Latent variable models for longitudinal data with multiple continuous outcomes. Biometrics, 56(4):1047–1054, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Garrett:2000:LCM

[154] Elizabeth S. Garrett and Scott L. Zeger. Latent class model diagnosis. Biometrics, 56(4):1055–1067, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2000:DET

[155] David B. Dunson and Gregg E. Dinse. Distinguishing effects on tumor multiplicity and growth rate in chemoprevention experiments. Biometrics, 56(4):1068–1075, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ferrante:2000:PEG

[156] L. Ferrante, S. Bompadre, L. Possati, and L. Leone. Parameter estimation in a Gompertzian stochastic model for tumor growth. Biometrics, 56(4):1076–1081, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2000:ICC

[157] Stuart G. Baker. Identifying combinations of cancer markers for further study as triggers of early intervention. Biometrics, 56(4):1082–1087, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Newton:2000:ILE

[158] Michael A. Newton and Yoonjung Lee. Inferring the location and effect of tumor suppressor genes by instability-selection modeling of allelic-loss data. Biometrics, 56(4):1088–1097, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2000:QGM

[159] Rongling Wu and Bailian Li. A quantitative genetic model for analyzing species differences in outcrossing species. Biometrics, 56(4):1098–1104, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hauptmann:2000:AET

[160] Michael Hauptmann, Jürgen Wellmann, Jay H. Lubin, Philip S. Rosenberg, and Lothar Kreienbrock. Analysis of exposure-time-response relationships using a spline weight function. Biometrics, 56(4):1105–1108, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2000:EOC

[161] Richard J. Cook, Edmund T. M. Ng, and Maureen O. Meade. Estimation of operating characteristics for dependent diagnostic tests based on latent Markov models. Biometrics, 56(4):1109–1117, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2000:CIM

[162] Xiao-Hua Zhou and Wanzhu Tu. Confidence intervals for the mean of diagnostic test charge data containing zeros. Biometrics, 56(4):1118– 1125, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vansteelandt:2000:RMD

[163] S. Vansteelandt, E. Goetghebeur, and T. Verstraeten. Regression models for disease prevalence with diagnostic tests on pools of serum samples. Biometrics, 56(4):1126–1133, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Venkatraman:2000:PTC

[164] E. S. Venkatraman. A permutation test to compare receiver operating characteristic curves. Biometrics, 56(4):1134–1138, December 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goetghebeur:2000:SRA

[165] Els Goetghebeur and Louise Ryan. Semiparametric regression analysis of interval-censored data. Biometrics, 56(4):1139–1144, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Paik:2000:BBD

[166] Myunghee Cho Paik, Ralph Sacco, and I.-Feng Lin. Bivariate binary data analysis with nonignorably missing outcomes. Biometrics, 56(4): 1145–1156, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2000:MIP

[167] Minzhi Liu, Jeremy M. G. Taylor, and Thomas R. Belin. Multiple imputation and posterior simulation for multivariate missing data in longitudinal studies. Biometrics, 56(4):1157–1163, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rittgen:2000:SAH

[168] Werner Rittgen and Nikolaus Becker. SMR analysis of historical followup studies with missing death certificates. Biometrics, 56(4):1164–1169, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Posch:2000:IAS

[169] Martin Posch and Peter Bauer. Interim analysis and sample size reassessment. Biometrics, 56(4):1170–1176, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2000:SDP

[170] Ying Kuen Cheung and Rick Chappell. Sequential designs for Phase I clinical trials with late-onset toxicities. Biometrics, 56(4):1177–1182, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proschan:2000:IDS

[171] Michael A. Proschan and Janet Wittes. An improved double sampling procedure based on the variance. Biometrics, 56(4):1183–1187, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Viana:2000:SDM

[172] Marlos Viana and Ingram Olkin. Symmetrically dependent models arising in visual assessment data. Biometrics, 56(4):1188–1191, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shieh:2000:PSS

[173] Gwowen Shieh. On power and sample size calculations for likelihood ratio tests in generalized linear models. Biometrics, 56(4):1192–1196, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2000:ERI

[174] Madeleine King and Annette Dobson. Estimating the responsiveness of an instrument using more than two repeated measures. Biometrics, 56(4):1197–1203, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bellio:2000:ALA

[175] R. Bellio, J. E. Jensen, and P. Seiden. Applications of likelihood asymptotics for nonlinear regression in herbicide bioassays. Biometrics, 56(4): 1204–1212, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yonezawa:2000:DPE

[176] Katsuei Yonezawa, Takuro Ishii, and Kentaro Yano. Definition of the probability efficiency index for optimization of the stepwise yield selection procedures of plant varieties. Biometrics, 56(4):1213–1217, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wameling:2000:ITE

[177] A. Wameling, M. Schäper, J. Kunert, M. Blaszkewicz, C. van Thriel, M. Zupanic, and A. Seeber. Individual toluene exposure in rotary printing: Increasing accuracy of estimation by linear models based on protocols of daily activity and other measures. Biometrics, 56(4):1218–1221, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Strand:2000:GNT

[178] Matthew Strand. A generalized nonparametric test for lattice-ordered means. Biometrics, 56(4):1222–1226, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Seber:2000:CRE

[179] George A. F. Seber, John T. Huakau, and David Simmons. Capture– recapture, epidemiology, and list mismatches: Two lists. Biometrics, 56(4):1227–1232, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Colosimo:2000:TPH

[180] Enrico A. Colosimo, Liciana V. A. S. Chalita, and Clarice G. B. Demétrio. Tests of proportional hazards and proportional odds models for grouped survival data. Biometrics, 56(4):1233–1240, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daniels:2000:RPM

[181] Michael J. Daniels and Joseph W. Hogan. Reparameterizing the pattern mixture model for sensitivity analyses under informative dropout. Biometrics, 56(4):1241–1248, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2000:BAF

[182] Pingping Qu and Yinsheng Qu. A Bayesian approach to finite mixture models in bioassay via data augmentation and Gibbs sampling and its application to insecticide resistance. Biometrics, 56(4):1249–1255, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rodenberg:2000:RCE

[183] Cindy Rodenberg and Xiao-Hua Zhou. ROC curve estimation when covariates affect the verification process. Biometrics, 56(4):1256–1262, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2000:MOD

[184] Joy King and Weng-Kee Wong. Minimax D-optimal designs for the logistic model. Biometrics, 56(4):1263–1267, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Contreras:2000:FNC

[185] Martha Contreras and Louise M. Ryan. Fitting nonlinear and constrained generalized estimating equations with optimization software. Biometrics, 56(4):1268–1271, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Solow:2000:CTD

[186] Andrew R. Solow. A comment on testing for a decline in Upper Cretaceous dinosaur diversity. Biometrics, 56(4):1272–1273, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:EBa

[187] Anonymous. To the Editor of Biometrics. Biometrics, 56(4):1274–1275, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:ARF

[188] Anonymous. The authors replied as follows. Biometrics, 56(4):1275– 1276, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:EBb

[189] Anonymous. To the Editor of Biometrics. Biometrics, 56(4):1276, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:RD

[190] Anonymous. A reply by Devlin. Biometrics, 56(4):1277, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:RDE

[191] Anonymous. A response to Devlin by Evett, Foreman, and Weir. Biometrics, 56(4):1277, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cormack:2000:GJ

[192] Rachard M. Cormack. George Jolly. Biometrics, 56(4):1278, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2000:BRb

[193] M. S. Ridout. Book reviews. Biometrics, 56(4):1279–1294, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:Cc

[194] Anonymous. Correction. Biometrics, 56(4):1294, December 2000. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2000:A

[195] Anonymous. Acknowledgements. Biometrics, 56(4):1295, December 2000. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carroll:2001:RTS

[196] Raymond J. Carroll. Review times in statistical journals: Tilting at windmills? Biometrics, 57(1):1–6, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Verbeke:2001:SAN

[197] Geert Verbeke, Geert Molenberghs, Herbert Thijs, Emmanuel Lesaffre, and Michael G. Kenward. Sensitivity analysis for nonrandom dropout: a local influence approach. Biometrics, 57(1):7–14, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fitzmaurice:2001:BEA

[198] Garrett M. Fitzmaurice, Stuart R. Lipsitz, Geert Molenberghs, and Joseph G. Ibrahim. Bias in estimating association parameters for longitudinal binary responses with drop-outs. Biometrics, 57(1):15–21, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hopke:2001:MIM

[199] Philip K. Hopke, Chuanhai Liu, and Donald B. Rubin. Multiple imputation for multivariate data with missing and below-threshold measurements: Time-series concentrations of pollutants in the Arctic. Biometrics, 57(1):22–33, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Horton:2001:MLA

[200] Nicholas J. Horton and Nan M. Laird. Maximum likelihood analysis of logistic regression models with incomplete covariate data and auxiliary information. Biometrics, 57(1):34–42, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2001:MLM

[201] Ming-Hui Chen and Joseph G. Ibrahim. Maximum likelihood methods for cure rate models with missing covariates. Biometrics, 57(1):43–52, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schafer:2001:SML

[202] Daniel W. Schafer. Semiparametric maximum likelihood for measurement error model regression. Biometrics, 57(1):53–61, March 2001. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McShane:2001:CME

[203] Lisa M. McShane, Douglas N. Midthune, Joanne F. Dorgan, Laurence S. Freedman, and Raymond J. Carroll. Covariate measurement error adjustment for matched case-control studies. Biometrics, 57(1):62–73, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hudgens:2001:NML

[204] Michael G. Hudgens, Glen A. Satten, and Ira M. Longini Jr. Nonparametric maximum likelihood estimation for competing risks survival data subject to interval censoring and truncation. Biometrics, 57(1):74–80, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2001:EMS

[205] Jane Xu and Scott L. Zeger. The evaluation of multiple surrogate endpoints. Biometrics, 57(1):81–87, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OGorman:2001:NMM

[206] John T. O’Gorman and Michael G. Akritas. Nonparametric models and methods for designs with dependent censored data: Part I. Biometrics, 57(1):88–95, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yau:2001:MMS

[207] Kelvin K. W. Yau. Multilevel models for survival analysis with random effects. Biometrics, 57(1):96–102, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andrea:2001:MCS

[208] Rotnitzky Andrea, Daniel Scharfstein, Ting-Li Su, and James Robins. Methods for conducting sensitivity analysis of trials with potentially nonignorable competing causes of censoring. Biometrics, 57(1):103–113, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heinze:2001:SPM

[209] Georg Heinze and Michael Schemper. A solution to the problem of monotone likelihood in Cox regression. Biometrics, 57(1):114–119, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2001:AIC

[210] Wei Pan. Akaike’s information criterion in generalized estimating equations. Biometrics, 57(1):120–125, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mancl:2001:CEG

[211] Lloyd A. Mancl and Timothy A. DeRouen. A covariance estimator for GEE with improved small-sample properties. Biometrics, 57(1):126–134, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carey:2001:BIP

[212] Vincent J. Carey, Carol J. Baker, and Richard Platt. Bayesian inference on protective antibody levels using case-control data. Biometrics, 57 (1):135–142, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Denison:2001:BPE

[213] D. G. T. Denison and C. C. Holmes. Bayesian partitioning for estimating disease risk. Biometrics, 57(1):143–149, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dominici:2001:BSA

[214] Francesca Dominici and Giovanni Parmigiani. Bayesian semiparametric analysis of developmental toxicology data. Biometrics, 57(1):150–157, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dendukuri:2001:BAM

[215] Nandini Dendukuri and Lawrence Joseph. Bayesian approaches to modeling the conditional dependence between multiple diagnostic tests. Biometrics, 57(1):158–167, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leung:2001:ODE

[216] Denis Heng-Yan Leung and You-Gan Wang. Optimal designs for evaluating a series of treatments. Biometrics, 57(1):168–171, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2001:SSI

[217] Qing Liu and George Y. H. Chi. On sample size and inference for twostage adaptive designs. Biometrics, 57(1):172–177, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Naranjo:2001:ARE

[218] Joshua D. Naranjo and Joseph W. McKean. Adjusting for regression effect in uncontrolled studies. Biometrics, 57(1):178–181, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenland:2001:EPA

[219] Sander Greenland. Estimation of population attributable fractions from fitted incidence ratios and exposure survey data, with an application to electromagnetic fields and childhood leukemia. Biometrics, 57(1):182– 188, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hjellvik:2001:MDV

[220] Vidar Hjellvik, Olav Rune GodøS, and Dag TjøSstheim. Modeling diurnal variation of marine populations. Biometrics, 57(1):189–196, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dean:2001:DIB

[221] C. B. Dean, M. D. Ugarte, and A. F. Militino. Detecting interaction between random region and fixed age effects in disease mapping. Biometrics, 57(1):197–202, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Straws:2001:CDP

[222] Warren J. Straws, Raymond J. Carroll, Steven M. Bortnick, John R. Menkedick, and Bradley D. Schultz. Combining datasets to predict the effects of regulation of environmental lead exposure in housing stock. Biometrics, 57(1):203–210, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diblasi:2001:UVC

[223] A. Diblasi and A. W. Bowman. On the use of the variogram in checking for independence in spatial data. Biometrics, 57(1):211–218, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2001:STT

[224] Martin Ridout, John Hinde, and Clarice G. B. Demétrio. A score test for testing a zero-inflated Poisson regression model against zero-inflated negative binomial alternatives. Biometrics, 57(1):219–223, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:2001:FNR

[225] Charles C. Taylor, Ian L. Dryden, and Rahman Farnoosh. The Kfunction for nearly regular point processes. Biometrics, 57(1):224–231, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Inglis:2001:ESA

[226] Iain M. Inglis and Alison J. Gray. An evaluation of semiautomatic approaches to contour segmentation applied to fungal hyphae. Biometrics, 57(1):232–239, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wood:2001:MMF

[227] Simon N. Wood. Minimizing model fitting objectives that contain spurious local minima by bootstrap restarting. Biometrics, 57(1):240–244, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chatterjee:2001:MLA

[228] Nilanjan Chatterjee and Sholom Wacholder. A marginal likelihood approach for estimating penetrance from kin-cohort designs. Biometrics, 57(1):245–252, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rice:2001:NME

[229] John A. Rice and Colin O. Wu. Nonparametric mixed effects models for unequally sampled noisy curves. Biometrics, 57(1):253–259, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grimshaw:2001:LRT

[230] Scott D. Grimshaw, David G. Whiting, and Thomas H. Morris. Likelihood ratio tests for a mixture of two von Mises distributions. Biometrics, 57(1):260–265, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yeap:2001:RTS

[231] Beow Yong Yeap and Marie Davidian. Robust two-stage estimation in hierarchical nonlinear models. Biometrics, 57(1):266–272, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lindberg:2001:CBR

[232] M. S. Lindberg, W. L. Kendall, J. E. Hines, and Michael G. Anderson. Combining band recovery data and Pollock’s robust design to model temporary and permanent emigration. Biometrics, 57(1):273–281, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Betensky:2001:NEC

[233] Rebecca A. Betensky and David A. Schoenfeld. Nonparametric estimation in a cure model with random cure times. Biometrics, 57(1):282–286, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dauk:2001:CEP

[234] Philip C. Dauk and Carl J. Schwarz. Catch estimation in the presence of declining catch rate due to Gear saturation. Biometrics, 57(1):287– 293, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chu:2001:SER

[235] Karin K. Chu, Naisyin Wang, Scott Stanley, and Noah D. Cohen. Statistical evaluation of the regulatory guidelines for use of furosemide in race horses. Biometrics, 57(1):294–301, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2001:FAM

[236] David B. Dunson and Sally D. Perreault. Factor analytic models of clustered multivariate data with informative censoring. Biometrics, 57 (1):302–308, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leung:2001:BDA

[237] Denis Heng-Yan Leung and You-Gan Wang. A Bayesian decision approach for sample size determination in Phase II trials. Biometrics, 57 (1):309–312, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2001:BRa

[238] M. S. Ridout. Book reviews. Biometrics, 57(1):313–331, March 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frangakis:2001:AIE

[239] Constantine E. Frangakis and Donald B. Rubin. Addressing an idiosyncrasy in estimating survival curves using double sampling in the presence of self-selected right censoring. Biometrics, 57(2):333–342, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [240, 241] and rejoinder [242].

Robins:2001:DFR

[240] James Robins, Andrea Rotnitzky, and Marco Bonetti. Discussion of the Frangakis and Rubin article. Biometrics, 57(2):343–347, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [239, 242].

Baker:2001:DDS

[241] Stuart G. Baker. Discussion of double sampling for survival analysis. Biometrics, 57(2):348–350, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [239, 242].

Frangakis:2001:RDA

[242] Constantine E. Frangakis and Donald B. Rubin. Rejoinder to discussions on addressing an idiosyncrasy in estimating survival curves using double sampling in the presence of self-selected right censoring. Biometrics, 57 (2):351–353, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [239, 240, 241].

Greene:2001:MPT

[243] Tom Greene. A model for a proportional treatment effect on disease progression. Biometrics, 57(2):354–360, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murray:2001:UWK

[244] Susan Murray. Using weighted Kaplan–Meier statistics in nonparametric comparisons of paired censored survival outcomes. Biometrics, 57(2): 361–368, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2001:NTG

[245] D. Y. Lin and Zhiliang Ying. Nonparametric tests for the gap time distributions of serial events based on censored data. Biometrics, 57 (2):369–375, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fine:2001:JRA

[246] Jason P. Fine and Richard D. Gelber. Joint regression analysis of survival and quality-adjusted survival. Biometrics, 57(2):376–382, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ibrahim:2001:BSM

[247] Joseph G. Ibrahim, Ming-Hui Chen, and Debajyoti Sinha. Bayesian semiparametric models for survival data with a cure fraction. Biometrics, 57(2):383–388, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pinsky:2001:EPC

[248] Paul F. Pinsky. Estimation and prediction for cancer screening models using deconvolution and smoothing. Biometrics, 57(2):389–395, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2001:FPM

[249] David B. Dunson and Donna D. Baird. A flexible parametric model for combining current status and age at first diagnosis data. Biometrics, 57(2):396–403, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scharfstein:2001:IRS

[250] Daniel Scharfstein, James M. Robins, Wesley Eddings, and Andrea Rotnitzky. Inference in randomized studies with informative censoring and discrete time-to-event endpoints. Biometrics, 57(2):404–413, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2001:AIP

[251] C. Y. Wang and Hua Yun Chen. Augmented inverse probability weighted estimator for Cox missing covariate regression. Biometrics, 57(2):414– 419, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Su:2001:NCA

[252] Xiaoping Su, Sylvan Wallenstein, and David Bishop. Nonoverlapping clusters: Approximate distribution and application to molecular biology. Biometrics, 57(2):420–426, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2001:EFD

[253] Kai Yu and Eleanor Feingold. Estimating the frequency distribution of crossovers during meiosis from recombination data. Biometrics, 57(2): 427–434, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hou:2001:ICF

[254] Chia-Ding Hou, Jengtung Chiang, and John Jen Tai. Identifying chromosomal fragile sites from a hierarchical-clustering point of view. Biometrics, 57(2):435–440, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2001:SMD

[255] Tiee-Jian Wu, Ya-Ching Hsieh, and Lung-An Li. Statistical measures of DNA sequence dissimilarity under Markov chain models of base composition. Biometrics, 57(2):441–448, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mayer-Hamblett:2001:RMA

[256] Nicole Mayer-Hamblett and Steve Self. A regression modeling approach for describing patterns of HIV genetic variation. Biometrics, 57(2):449– 460, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dauk:2001:CER

[257] Philip C. Dauk and Carl J. Schwarz. Catch estimation with restricted randomization in the effort survey. Biometrics, 57(2):461–468, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Catchpole:2001:AAM

[258] E. A. Catchpole, S. N. Freeman, B. J. T. Morgan, and W. J. Nash. Abalone i: Analyzing mark-recapture-recovery data incorporating growth and delayed recovery. Biometrics, 57(2):469–477, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsueh:2001:UET

[259] Huey-Miin Hsueh, Jen-Pei Liu, and James J. Chen. Unconditional exact tests for equivalence or noninferiority for paired binary endpoints. Biometrics, 57(2):478–483, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Troendle:2001:UCI

[260] James F. Troendle and Jennifer Frank. Unbiased confidence intervals for the odds ratio of two independent binomial samples with application to case-control data. Biometrics, 57(2):484–489, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moore:2001:EEB

[261] Dirk F. Moore, Choon Keun Park, and Woollcott Smith. Exploring extra-binomial variation in teratology data using continuous mixtures. Biometrics, 57(2):490–494, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fewster:2001:SIS

[262] R. M. Fewster and S. T. Buckland. Similarity indices for spatia i ecological data. Biometrics, 57(2):495–501, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Blackwell:2001:BIR

[263] P. G. Blackwell. Bayesian inference for a random tessellation process. Biometrics, 57(2):502–507, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Allard:2001:TLI

[264] Denis Allard, Anders Brix, and Joël Chadoeuf. Testing local independence between two point processes. Biometrics, 57(2):508–517, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perron:2001:BNM

[265] F. Perron and K. Mengersen. Bayesian nonparametric modeling using mixtures of triangular distributions. Biometrics, 57(2):518–528, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2001:MSE

[266] Wei Pan. Model selection in estimating equations. Biometrics, 57(2): 529–534, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fuchs:2001:UAT

[267] Camil Fuchs. UMPU and alternative tests for association in 2 × 2 tables. Biometrics, 57(2):535–538, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coull:2001:SII

[268] Brent A. Coull, David Ruppert, and M. P. Wand. Simple incorporation of interactions into additive models. Biometrics, 57(2):539–545, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Catlin:2001:SIT

[269] Sandra N. Catlin, Janis L. Abkowitz, and Peter Guttorp. Statistical inference in a two-compartment model for hematopoiesis. Biometrics, 57(2):546–553, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gossl:2001:BSI

[270] C. Gössl, D. P. Auer, and L. Fahrmeir. Bayesian spatiotemporal inference in functional magnetic resonance imaging. Biometrics, 57(2):554– 562, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barker:2001:EGA

[271] Richard J. Barker and Laimonis Kavalieris. Efficiency gain from auxiliary data requiring additional nuisance parameters. Biometrics, 57(2):563– 566, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ivanova:2001:DIS

[272] Anastasia Ivanova and Vance W. Berger. Drawbacks to integer scoring for ordered categorical data. Biometrics, 57(2):567–570, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jones:2001:AIM

[273] Geoffrey Jones and David M. Rocke. Analyte identification in multivariate calibration. Biometrics, 57(2):571–576, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Molinari:2001:MTC

[274] Nicolas Molinari, Chistophe Bonaldi, and Jean-Pierre Daurés. Multiple temporal cluster detection. Biometrics, 57(2):577–583, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Altaye:2001:IPA

[275] Mekibib Altaye, Allan Dormer, and Neil Klar. Inference procedures for assessing interobserver agreement among multiple raters. Biometrics, 57 (2):584–588, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gorfine:2001:ESP

[276] Malka Gorfine. Estimation of a secondary parameter in a group sequential clinical trial. Biometrics, 57(2):589–597, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gustafson:2001:CCA

[277] Paul Gustafson, Nhu D. Le, and Refik Saskin. Case-control analysis with partial knowledge of exposure misclassification probabilities. Biometrics, 57(2):598–609, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2001:LCM

[278] Paul S. Albert, Lisa M. McShane, Joanna H. Shih, and The U.S.National Cancer Institute Bladder Tumor Marker Network. Latent class modeling approaches for assessing diagnostic error without a gold standard: With applications to p53 immunohistochemical assays in bladder tumors. Biometrics, 57(2):610–619, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Faddy:2001:LBM

[279] M. J. Faddy and R. J. Bosch. Likelihood-based modeling and analysis of data underdispersed relative to the Poisson distribution. Biometrics, 57(2):620–624, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coffey:2001:CTS

[280] Christopher S. Coffey and Keith E. Muller. Controlling test size while gaining the benefits of an internal pilot design. Biometrics, 57(2):625– 631, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Segal:2001:RAA

[281] Mark R. Segal, Michael P. Cummings, and Alan E. Hubbard. Relating amino acid sequence to phenotype: Analysis of peptide-binding data. Biometrics, 57(2):632–643, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:Ca

[282] Anonymous. Correspondence. Biometrics, 57(2):643, June 2001. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2001:BRb

[283] M. S. Ridout. Book reviews. Biometrics, 57(2):644–658, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:Cb

[284] Anonymous. Correction. Biometrics, 57(2):659–660, June 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenland:2001:PBI

[285] Sander Greenland. Putting background information about relative risks into conjugate prior distributions. Biometrics, 57(3):663–670, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2001:SEE

[286] Richard J. Cook, Babette B. Brumback, Melanie B. Wigg, and Louise M. Ryan. Synthesis of evidence from epidemiological studies with intervalcensored exposure due to grouping. Biometrics, 57(3):671–680, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:2001:MMO

[287] Douglas J. Taylor, Lawrence L. Kupper, Stephen M. Rappaport, and Robert H. Lyles. A mixture model for occupational exposure mean testing with a limit of detection. Biometrics, 57(3):681–688, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schafer:2001:TCF

[288] Daniel W. Schafer, Jay H. Lubin, Elaine Ron, Marilyn Stovall, and Raymond J. Carroll. Thyroid cancer following scalp irradiation: a reanalysis accounting for uncertainty in dosimetry. Biometrics, 57(3):689–697, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Budtz-Jorgensen:2001:BDC

[289] Esben Budtz-Jørgensen, Niels Keiding, and Philippe Grandjean. Benchmark dose calculation from epidemiological data. Biometrics, 57(3):698– 706, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2001:CRE

[290] Alan J. Lee, George A. F. Seber, Jennifer K. Holden, and John T. Huakau. Capture–recapture, epidemiology, and list mismatches: Several lists. Biometrics, 57(3):707–713, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartolucci:2001:ACR

[291] Francesco Bartolucci and Antonio Forcina. Analysis of capture–recapture data with a Rasch-type model allowing for conditional dependence and multidimensionality. Biometrics, 57(3):714–719, September 2001. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Catchpole:2001:NSM

[292] E. A. Catchpole, P. M. Kgosi, and B. J. T. Morgan. On the nearsingularity of models for animal recovery data. Biometrics, 57(3):720– 726, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hiby:2001:LTS

[293] Lex Hiby and M. B. Krishna. Line transect sampling from a curving path. Biometrics, 57(3):727–731, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2001:MEL

[294] Song Xi Chen and Ann Cowling. Measurement errors in line transect surveys where detectability varies with distance and size. Biometrics, 57(3):732–742, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yue:2001:NES

[295] Jack C. Yue, Murray K. Clayton, and Feng-Chang Lin. A nonparametric estimator of species overlap. Biometrics, 57(3):743–749, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Skaug:2001:ASM

[296] Hans Julius Skaug. Allele-sharing methods for estimation of population size. Biometrics, 57(3):750–756, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Almudevar:2001:BAV

[297] Anthony Almudevar. A bootstrap assessment of variability in pedigree reconstruction based on genetic markers. Biometrics, 57(3):757–763, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2001:MLB

[298] Rongling Wu, Bailian Li, Samuel S. Wu, and George Casella. A maximum likelihood-based method for mining major genes affecting a quantitative character. Biometrics, 57(3):764–768, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hirotsu:2001:ETA

[299] Chihiro Hirotsu, Satoshi Aoki, Toshiya Inada, and Yoshie Kitao. An exact test for the association between the disease and alleles at highly polymorphic loci with particular interest in the haplotype analysis. Biometrics, 57(3):769–778, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chatterjee:2001:BCM

[300] Nilanjan Chatterjee and Joanna Shih. A bivariate cure-mixture approach for modeling familial association in diseases. Biometrics, 57(3):779–786, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2001:MLE

[301] Sik-Yum Lee and Jian-Qing Shi. Maximum likelihood estimation of twolevel latent variable models with mixed continuous and polytomous data. Biometrics, 57(3):787–794, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2001:LMM

[302] Daowen Zhang and Marie Davidian. Linear mixed models with flexible distributions of random effects for longitudinal data. Biometrics, 57(3): 795–802, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bennett:2001:EVJ

[303] James Bennett and Jon Wakefield. Errors-in-variables in joint population pharmacokinetic/pharmacodynamic modeling. Biometrics, 57(3):803– 812, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heller:2001:PRB

[304] Glenn Heller and Jing Qin. Pairwise rank-based likelihood for estimation and inference on the mixture proportion. Biometrics, 57(3):813–817, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McKeague:2001:OTC

[305] Ian W. McKeague, Peter B. Gilbert, and Phyllis J. Kanki. Omnibus tests for comparison of competing risks with adjustment for covariate effects. Biometrics, 57(3):818–828, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2001:SOM

[306] James H. Albert and Siddhartha Chib. Sequential ordinal modeling with applications to survival data. Biometrics, 57(3):829–836, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2001:MWK

[307] Yu Shen and Jianwen Cai. Maximum of the weighted Kaplan–Meier tests with application to cancer prevention and screening trials. Biometrics, 57(3):837–843, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Broet:2001:SAT

[308] Philippe Broët, Yann De Rycke, Pascale Tubert-Bitter, Joseph Lellouch, Bernard Asselain, and Thierry Moreau. A semiparametric approach for the two-sample comparison of survival times with long-term survivors. Biometrics, 57(3):844–852, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2001:AHR

[309] Ying Qing Chen. Accelerated hazards regression model and its adequacy for censored survival data. Biometrics, 57(3):853–860, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2001:TES

[310] Hongwei Zhao and Anastasios A. Tsiatis. Testing equality of survival functions of quality-adjusted lifetime. Biometrics, 57(3):861–867, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2001:QAS

[311] Pai-Lien Chen and Pranab K. Sen. Quality-adjusted survival estimation with periodic observations. Biometrics, 57(3):868–874, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2001:SET

[312] Ronghui Xu and David P. Harrington. A semiparametric estimate of treatment effects with censored data. Biometrics, 57(3):875–885, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:2001:AGS

[313] Hans-Helge Müller and Helmut Schäfer. Adaptive group sequential designs for clinical trials: Combining the advantages of adaptive and of classical group sequential approaches. Biometrics, 57(3):886–891, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hellmich:2001:MCT

[314] Martin Hellmich. Monitoring clinical trials with multiple arms. Biometrics, 57(3):892–898, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frangakis:2001:CSD

[315] Constantine E. Frangakis and Stuart G. Baker. Compliance subsampling designs for comparative research: Estimation and optimal planning. Biometrics, 57(3):899–908, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenberger:2001:OAD

[316] William F. Rosenberger, Nigel Stallard, Anastasia Ivanova, Cherice N. Harper, and Michelle L. Ricks. Optimal adaptive designs for binary response trials. Biometrics, 57(3):909–913, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2001:DFB

[317] Peter F. Thall, Hsi-Guang Sung, and Aniruddha Choudhury. Dosefinding based on feasibility and toxicity in T-cell infusion trials. Biometrics, 57(3):914–921, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2001:SDP

[318] Lei Zhu, Jacqueline M. Hughes-Oliver, and S. Stanley Young. Statistical decoding of potent pools based on chemical structure. Biometrics, 57(3): 922–930, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barnhart:2001:MCC

[319] Huiman X. Barnhart and John M. Williamson. Modeling concordance correlation via GEE to evaluate reproducibility. Biometrics, 57(3):931– 940, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Corcoran:2001:ETT

[320] Chris Corcoran, Louise Ryan, Pralay Senchaudhuri, Cyrus Mehta, Nitin Patel, and Geert Molenberghs. An exact trend test for correlated binary data. Biometrics, 57(3):941–948, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

MacNab:2001:ASS

[321] Ying C. MacNab and C. B. Dean. Autoregressive spatial smoothing and temporal spline smoothing for mapping rates. Biometrics, 57(3):949–956, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhoj:2001:RSS

[322] Dinesh S. Bhoj. Ranked set sampling with unequal samples. Biometrics, 57(3):957–962, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:2001:SSC

[323] Alan Agresti and Yongyi Min. On small-sample confidence intervals for parameters in discrete distributions. Biometrics, 57(3):963–971, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schoenfeld:2001:SAD

[324] David A. Schoenfeld. A simple algorithm for designing group sequential clinical trials. Biometrics, 57(3):972–974, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:EBa

[325] Anonymous. To the Editor of Biometrics. Biometrics, 57(3):975, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:ARFa

[326] Anonymous. The authors replied as follows. Biometrics, 57(3):976, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:EBb

[327] Anonymous. To the Editor of Biometrics. Biometrics, 57(3):976–978, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:ARFb

[328] Anonymous. The authors replied as follows. Biometrics, 57(3):978–980, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:EBc

[329] Anonymous. To the Editor of Biometrics. Biometrics, 57(3):980, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:EN

[330] Anonymous. Editor’s note. Biometrics, 57(3):980, September 2001. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2001:BRc

[331] M. S. Ridout. Book reviews. Biometrics, 57(3):982–990, September 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aalen:2001:CAE

[332] Odd O. Aalen, Ørnulf Borgan, and Harald Fekjær. Covariate adjustment of event histories estimated from Markov chains: The additive approach.

Biometrics, 57(4):993–1001, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [560].

Zhao:2001:EMC

[333] Hongwei Zhao and Lili Tian. On estimating medical cost and incremental cost-effectiveness ratios with censored data. Biometrics, 57(4): 1002–1008, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2001:TIT

[334] Yu Shen, Dongfeng Wu, and Marvin Zelen. Testing the independence of two diagnostic tests. Biometrics, 57(4):1009–1017, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2001:DFD

[335] John O’Quigley, Michael D. Hughes, and Terry Fenton. Dose-finding designs for HIV studies. Biometrics, 57(4):1018–1029, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2001:CID

[336] Pei-Yun Chen and Anastasios A. Tsiatis. Causal inference on the difference of the restricted mean lifetime between two groups. Biometrics, 57(4):1030–1038, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bloch:2001:OST

[337] Daniel A. Bloch, Tze Leung Lai, and Pascale Tubert-Bitter. One-sided tests in clinical trials with multiple endpoints. Biometrics, 57(4):1039– 1047, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balasubramanian:2001:ETP

[338] Raji Balasubramanian and Stephen W. Lagakos. Estimation of the timing of perinatal transmission of HIV. Biometrics, 57(4):1048–1058, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2001:BMA

[339] Chong Z. He, Dongchu Sun, and Yolande Tra. Bayesian modeling of agespecific survival in nesting studies under Dirichlet priors. Biometrics, 57(4):1059–1066, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2001:BML

[340] David B. Dunson. Bayesian modeling of the level and duration of fertility in the menstrual cycle. Biometrics, 57(4):1067–1073, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2001:SAU

[341] H. Zhao, J. Li, and W. P. Robinson. Statistical analysis of uniparental disomy data using hidden Markov models. Biometrics, 57(4):1074–1079, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Almudevar:2001:MPP

[342] Anthony Almudevar. Most powerful permutation invariant tests for relatedness hypotheses using genotypic data. Biometrics, 57(4):1080–1088, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stallard:2001:DTD

[343] Nigel Stallard and Peter F. Thall. Decision-theoretic designs for pre-Phase II screening trials in oncology. Biometrics, 57(4):1089–1095, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christman:2001:IAC

[344] Mary C. Christman and Feng Lan. Inverse adaptive cluster sampling. Biometrics, 57(4):1096–1105, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2001:MCC

[345] I-Feng Lin and Myunghee Cho Paik. Matched case-control data analysis with selection bias. Biometrics, 57(4):1106–1112, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kendall:2001:UOR

[346] William L. Kendall and Rhema Bjorkland. Using open robust design models to estimate temporary emigration from capture–recapture data. Biometrics, 57(4):1113–1122, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rekaya:2001:TMM

[347] R. Rekaya, K. A. Weigel, and D. Gianola. Threshold model for misclassified binary responses with applications to animal breeding. Biometrics, 57(4):1123–1129, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Melville:2001:LTS

[348] G. J. Melville and A. H. Welsh. Line transect sampling in small regions. Biometrics, 57(4):1130–1137, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:2001:AVE

[349] Alison Smith, Brian Cullis, and Robin Thompson. Analyzing variety by environment data using multiplicative mixed models and adjustments for spatial field trend. Biometrics, 57(4):1138–1147, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burch:2001:CFA

[350] Brent D. Burch and Ian R. Harris. Closed-form approximations to the REML estimator of a variance ratio (or heritability) in a mixed linear model. Biometrics, 57(4):1148–1156, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Silvapulle:2001:TAQ

[351] Mervyn J. Silvapulle. Tests against qualitative interaction: Exact critical values and robust tests. Biometrics, 57(4):1157–1165, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ouwens:2001:LID

[352] Mario J. N. M. Ouwens, Frans E. S. Tan, and Martijn P. F. Berger. Local influence to detect influential data structures for generalized linear mixed models. Biometrics, 57(4):1166–1172, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Daniels:2001:SEC

[353] Michael J. Daniels and Robert E. Kass. Shrinkage estimators for covariance matrices. Biometrics, 57(4):1173–1184, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edwards:2001:LEC

[354] Lloyd J. Edwards, Paul W. Stewart, Keith E. Muller, and Ronald W. Helms. Linear equality constraints in the general linear mixed model. Biometrics, 57(4):1185–1190, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2001:MIM

[355] Kaifeng Lu and Anastasios A. Tsiatis. Multiple imputation methods for estimating regression coefficients in the competing risks model with missing cause of failure. Biometrics, 57(4):1191–1197, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fay:2001:SSA

[356] Michael P. Fay and Barry I. Graubard. Small-sample adjustments for Wald-type tests using sandwich estimators. Biometrics, 57(4):1198–1206, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anstrom:2001:UPS

[357] Kevin J. Anstrom and Anastasios A. Tsiatis. Utilizing propensity scores to estimate causal treatment effects with censored time-lagged data. Biometrics, 57(4):1207–1218, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peddada:2001:TOR

[358] Shyamal D. Peddada, Katherine E. Prescott, and Mark Conaway. Tests for order restrictions in binary data. Biometrics, 57(4):1219–1227, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Irizarry:2001:AHT

[359] Rafael A. Irizarry, Clarke Tankersley, Robert Frank, and Susan Flanders. Assessing homeostasis through circadian patterns. Biometrics, 57(4): 1228–1237, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lyles:2001:CTV

[360] Robert H. Lyles, Jovonne K. Williams, and Rutt Chuachoowong. Correlating two viral load assays with known detection limits. Biometrics, 57(4):1238–1244, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2001:MIA

[361] Wei Pan. A multiple imputation approach to regression analysis for doubly censored data with application to AIDS studies. Biometrics, 57 (4):1245–1250, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hutchinson:2001:PKA

[362] T. P. Hutchinson. Partial knowledge and answer-until-correct tasks in birds and humans. Biometrics, 57(4):1251–1252, December 2001. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bilder:2001:FOR

[363] Christopher R. Bilder and Thomas M. Loughin. On the first-order Rao– Scott correction of the Umesh–Loughin–Scherer statistic. Biometrics, 57(4):1253–1255, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rabe-Hesketh:2001:PMR

[364] S. Rabe-Hesketh and A. Skrondal. Parameterization of multivariate random effects models for categorical data. Biometrics, 57(4):1256–1263, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coull:2001:ARF

[365] Brent A. Coull and Alan Agresti. The authors replied as follows. Biometrics, 57(4):1263–1264, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2001:BRd

[366] M. S. Ridout. Book reviews. Biometrics, 57(4):1265–1278, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2001:A

[367] Anonymous. Acknowledgements. Biometrics, 57(4):1279–1280, December 2001. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2002:MCT

[368] D. Y. Lin, L. J. Wei, and Z. Ying. Model-checking techniques based on cumulative residuals. Biometrics, 58(1):1–12, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mallick:2002:SRM

[369] Bani Mallick, F. Owen Hoffman, and Raymond J. Carroll. Semiparametric regression modeling with mixtures of Berkson and classical error, with application to fallout from the Nevada Test Site. Biometrics, 58 (1):13–20, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frangakis:2002:PSC

[370] Constantine E. Frangakis and Donald B. Rubin. Principal stratification in causal inference. Biometrics, 58(1):21–29, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Parmigiani:2002:MOE

[371] Giovanni Parmigiani, Steven Skates, and Marvin Zelen. Modeling and optimization in early detection programs with a single exam. Biometrics, 58(1):30–36, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Faucett:2002:SAU

[372] Cheryl L. Faucett, Nathaniel Schenker, and Jeremy M. G. Taylor. Survival analysis using auxiliary variables via multiple imputation, with application to AIDS clinical trial data. Biometrics, 58(1):37–47, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lunceford:2002:ESD

[373] Jared K. Lunceford, Marie Davidian, and Anastasios A. Tsiatis. Estimation of survival distributions of treatment policies in two-stage randomization designs in clinical trials. Biometrics, 58(1):48–57, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Betensky:2002:TDB

[374] Rebecca A. Betensky and Dianne M. Finkelstein. Testing for dependence between failure time and visit compliance with interval-censored data. Biometrics, 58(1):58–63, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2002:ECP

[375] Wei Pan and Rick Chappell. Estimation in the Cox proportional hazards model with left-truncated and interval-censored data. Biometrics, 58(1): 64–70, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2002:PHM

[376] David B. Dunson and Donna D. Baird. A proportional hazards model for incidence and induced remission of disease. Biometrics, 58(1):71–78, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2002:BMM

[377] David B. Dunson and Gregg E. Dinse. Bayesian models for multivariate current status data with informative censoring. Biometrics, 58(1):79– 88, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2002:MRC

[378] Ying Kuen Cheung and Peter F. Thall. Monitoring the rates of composite events with censored data in Phase II clinical trials. Biometrics, 58(1): 89–97, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Herring:2002:FMM

[379] Amy H. Herring, Joseph G. Ibrahim, and Stuart R. Lipsitz. Frailty models with missing covariates. Biometrics, 58(1):98–109, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xue:2002:CCM

[380] Qian-Li Xue and Karen Bandeen-Roche. Combining complete multivariate outcomes with incomplete covariate information: a latent class approach. Biometrics, 58(1):110–120, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2002:FME

[381] Wensheng Guo. Functional mixed effects models. Biometrics, 58(1):121– 128, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2002:EPS

[382] Hao Zhang. On estimation and prediction for spatial generalized linear mixed models. Biometrics, 58(1):129–136, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

TenHave:2002:MEL

[383] Thomas R. Ten Have, Beth A. Reboussin, Michael E. Miller, and Allen Kunselman. Mixed effects logistic regression models for multiple longitudinal binary functional limitation responses with informative dropout and confounding by baseline outcomes. Biometrics, 58(1):137–144, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Foulkes:2002:CRB

[384] A. S. Foulkes and V. De Gruttola. Characterizing the relationship between HIV-1 genotype and phenotype: Prediction-based classification. Biometrics, 58(1):145–156, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fletcher:2002:GCP

[385] Scott D. Fletcher and Florence Jaffrezic. Generalized character process models: Estimating the genetic basis of traits that cannot be observed and that change with age or environmental conditions. Biometrics, 58 (1):157–162, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Satagopan:2002:TSD

[386] Jaya M. Satagopan, David A. Verbel, E. S. Venkatraman, Kenneth E. Offit, and Colin B. Begg. Two-stage designs for gene-disease association studies. Biometrics, 58(1):163–170, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stylianou:2002:DFU

[387] Mario Stylianou and Nancy Flournoy. Dose finding using the biased coin up-and-down design and isotonic regression. Biometrics, 58(1):171–177, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meester:2002:MRT

[388] Ronald Meester, Jan de Koning, Mart C. M. de Jong, and Odo Diekmann. Modeling and real-time prediction of classical swine fever epidemics. Biometrics, 58(1):178–184, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2002:ELE

[389] Alan Lee. Effect of list errors on the estimation of population size. Biometrics, 58(1):185–191, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yip:2002:UPR

[390] Paul S. F. Yip and Yan Wang. A unified parametric regression model for recapture studies with random removals in continuous time. Biometrics, 58(1):192–199, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bilder:2002:TCM

[391] Christopher R. Bilder and Thomas M. Loughin. Testing for conditional multiple marginal independence. Biometrics, 58(1):200–208, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dormer:2002:IED

[392] Allan Dormer and Guangyong Zou. Interval estimation for a difference between intraclass kappa statistics. Biometrics, 58(1):209–215, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

ONeill:2002:LTH

[393] Michael E. O’Neill and Ky L. Mathews. Levene tests of homogeneity of variance for general block and treatment designs. Biometrics, 58(1):216– 224, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pourahmadi:2002:DCL

[394] M. Pourahmadi and M. J. Daniels. Dynamic conditionally linear mixed models for longitudinal data. Biometrics, 58(1):225–231, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2002:UFH

[395] Yi Li, Rebecca A. Betensky, David N. Louis, and J. Gregory Cairncross. The use of frailty hazard models for unrecognized heterogeneity that interacts with treatment: Considerations of efficiency and power. Biometrics, 58(1):232–236, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2002:UNC

[396] Ying Kuen Cheung. On the use of nonparametric curves in Phase I trials with low toxicity tolerance. Biometrics, 58(1):237–240, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balding:2002:DDS

[397] David J. Balding. The DN a database search controversy. Biometrics, 58 (1):241–244, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2002:CFM

[398] John O’Quigley. Curve-free and model-based continual reassessment method designs. Biometrics, 58(1):245–249, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2002:EB

[399] Anonymous. To the Editor of Biometrics. Biometrics, 58(1):250, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2002:ARF

[400] Anonymous. The authors replied as follows. Biometrics, 58(1):250–251, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2002:BRa

[401] M. S. Ridout. Book reviews. Biometrics, 58(1):252–262, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2002:C

[402] Anonymous. Correction. Biometrics, 58(1):262, March 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2002:SEL

[403] Song Xi Chen, Paul S. F. Yip, and Yong Zhou. Sequential estimation in line transect surveys. Biometrics, 58(2):263–269, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Royle:2002:SMW

[404] J. Andrew Royle, Mark D. Koneff, and Ron E. Reynolds. Spatial modeling of wetland condition in the U.S. prairie pothole region. Biometrics, 58(2):270–279, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christensen:2002:BPS

[405] Ole F. Christensen and Rasmus Waagepetersen. Bayesian prediction of spatial count data using generalized linear mixed models. Biometrics, 58(2):280–286, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2002:MSS

[406] Yi Li and Louise Ryan. Modeling spatial survival data using semiparametric frailty models. Biometrics, 58(2):287–297, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finkelstein:2002:AFT

[407] Dianne M. Finkelstein, William B. Goggins, and David A. Schoenfeld. Analysis of failure time data with dependent interval censoring. Biometrics, 58(2):298–304, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2002:SAT

[408] Ronghui Xu and Sudeshna Adak. Survival analysis with time-varying regression effects using a tree-based approach. Biometrics, 58(2):305– 315, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lam:2002:MMS

[409] K. F. Lam, Y. W. Lee, and T. L. Leung. Modeling multivariate survival data by a semiparametric random effects proportional odds model. Biometrics, 58(2):316–323, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balshaw:2002:SMA

[410] Robert F. Balshaw and C. B. Dean. A semiparametric model for the analysis of recurrent-event panel data. Biometrics, 58(2):324–331, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rieger:2002:ACB

[411] Randall H. Rieger and Clarice R. Weinberg. Analysis of clustered binary outcomes using within-cluster paired resampling. Biometrics, 58(2):332– 341, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heagerty:2002:MTM

[412] Patrick J. Heagerty. Marginalized transition models and likelihood inference for longitudinal categorical data. Biometrics, 58(2):342–351, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2002:REL

[413] Haibo Zhou, Jianwei Chen, and Jianwen Cai. Random effects logistic regression analysis with auxiliary covariates. Biometrics, 58(2):352–360, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glidden:2002:RIE

[414] David V. Glidden. Robust inference for event probabilities with non-Markov event data. Biometrics, 58(2):361–368, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2002:MAS

[415] Kai Yu and Eleanor Feingold. Methods for analyzing the spatial distribution of chiasmata during meiosis based on recombination data. Biometrics, 58(2):369–377, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilbur:2002:VSH

[416] J. D. Wilbur, J. K. Ghosh, C. H. Nakatsu, S. M. Brouder, and R. W. Doerge. Variable selection in high-dimensional multivariate binary data with application to the analysis of microbial community DNA fingerprints. Biometrics, 58(2):378–386, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Biernacki:2002:GDR

[417] Christophe Biernacki, Farid Beninel, and Vincent Bretagnolle. A generalized discriminant rule when training population and test population differ on their descriptive parameters. Biometrics, 58(2):387–397, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Astuti:2002:TTC

[418] Erni Tri Astuti and Takashi Yanagawa. Testing trend for count data with extra-Poisson variability. Biometrics, 58(2):398–402, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mancuso:2002:AAE

[419] Jessica Y. Mancuso, Hongshik Ahn, James J. Chen, and James P. Mancuso. Age-adjusted exact trend tests in the event of rare occurrences. Biometrics, 58(2):403–412, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2002:SEL

[420] Haibo Zhou, M. A. Weaver, J. Qin, M. P. Longnecker, and M. C. Wang. A semiparametric empirical likelihood method for data from an outcomedependent sampling scheme with a continuous outcome. Biometrics, 58 (2):413–421, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peterson:2002:GAC

[421] John J. Peterson, Suntara Cahya, and Enrique del Castillo. A general approach to confidence regions for optimal factor levels of response surfaces. Biometrics, 58(2):422–431, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thach:2002:SDT

[422] Chau T. Thach and Lloyd D. Fisher. Self-designing two-stage trials to minimize expected costs. Biometrics, 58(2):432–438, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ng:2002:MPN

[423] Meei Pyng Ng. A modification of Peto’s nonparametric estimation of survival curves for interval-censored data. Biometrics, 58(2):439–442, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bacchetti:2002:GAM

[424] Peter Bacchetti and Christopher Quale. Generalized additive models with interval-censored data and time-varying covariates: Application to human immunodeficiency virus infection in hemophiliacs. Biometrics, 58(2):443–447, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Suh:2002:SML

[425] Eun-Young Suh and Daniel W. Schafer. Semiparametric maximum likelihood for nonlinear regression with measurement errors. Biometrics, 58 (2):448–453, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prescott:2002:SBA

[426] Gordon J. Prescott and Paul H. Garthwaite. A simple Bayesian analysis of misclassified binary data with a validation substudy. Biometrics, 58 (2):454–458, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2002:DUA

[427] Yuk W. Cheng and Anthony Y. C. Kuk. Determination of the unknown age at first capture of western rock lobsters (Panulims cygnus) by random effects model. Biometrics, 58(2):459–462, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2002:LAR

[428] Jianrong Wu, Guoyong Jiang, A. C. M. Wong, and Xiang Sun. Likelihood analysis for the ratio of means of two independent log–normal distributions. Biometrics, 58(2):463–469, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2002:BRb

[429] M. S. Ridout. Book reviews. Biometrics, 58(2):470–479, June 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2002:SIF

[430] Chang Yu and Daniel Zelterman. Statistical inference for familial disease clusters. Biometrics, 58(3):481–491, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Knorr-Held:2002:DMS

[431] Leonhard Knorr-Held, Günter Raßer, and Nikolaus Becker. Disease mapping of stage-specific cancer incidence data. Biometrics, 58(3):492–501, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:2002:ASD

[432] Joanna H. Shih and Nilanjan Chatterjee. Analysis of survival data from case-control family studies. Biometrics, 58(3):502–509, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2002:FMI

[433] Xuelin Huang and Robert A. Wolfe. A frailty model for informative censoring. Biometrics, 58(3):510–520, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shults:2002:UQL

[434] Justine Shults and Ardythe L. Morrow. Use of quasi-least squares to adjust for two levels of correlation. Biometrics, 58(3):521–530, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chao:2002:ENS

[435] Anne Chao and John Bunge. Estimating the number of species in a stochastic abundance model. Biometrics, 58(3):531–539, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Besbeas:2002:IMR

[436] P. Besbeas, S. N. Freeman, B. J. T. Morgan, and E. A. Catchpole. Integrating mark-recapture-recovery and census data to estimate animal abundance and demographic parameters. Biometrics, 58(3):540–547, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zucker:2002:SSR

[437] David M. Zucker and Jonathan Denne. Sample-size redetermination for repeated measures studies. Biometrics, 58(3):548–559, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2002:ADM

[438] Peter F. Thall, Lurdes Y. T. Inoue, and Thomas G. Martin. Adaptive decision making in a lymphocyte infusion trial. Biometrics, 58(3):560– 568, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Spiessens:2002:GSM

[439] Bart Spiessens, Emmanuel Lesaffre, Geert Verbeke, and KyungMann Kim. Group sequential methods for an ordinal logistic random-effects model under misspecification. Biometrics, 58(3):569–575, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2002:DBB

[440] Paul S. Albert, Lisa M. McShane, and Edward L. Korn. Design of a binary biomarker study from the results of a pilot study. Biometrics, 58(3):576–585, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barabesi:2002:RSS

[441] Lucio Barabesi and Caterina Pisani. Ranked set sampling for replicated sampling designs. Biometrics, 58(3):586–592, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kottas:2002:NBM

[442] Athanasios Kottas, Márcia D. Branco, and Alan E. Gelfand. A nonparametric Bayesian modeling approach for cytogenetic dosimetry. Biometrics, 58(3):593–600, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2002:EIS

[443] Lang Li, Morton B. Brown, Kyung-Hoon Lee, and Suneel Gupta. Estimation and inference for a spline-enhanced population pharmacokinetic model. Biometrics, 58(3):601–611, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tan:2002:SSI

[444] Ming Tan, Hong-Bin Fang, Guo-Liang Tian, and Peter J. Houghton. Small-sample inference for incomplete longitudinal data with truncation and censoring in tumor xenograft models. Biometrics, 58(3):612–620, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lipsitz:2002:PEL

[445] Stuart R. Lipsitz, Garrett M. Fitzmaurice, Joseph G. Ibrahim, Richard Gelber, and Steven Lipshultz. Parameter estimation in longitudinal studies with outcome-dependent follow-up. Biometrics, 58(3):621–630, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2002:LAM

[446] Paul S. Albert, Dean A. Follmann, Shaohua A. Wang, and Edward B. Suh. A latent autoregressive model for longitudinal binary data subject to informative missingness. Biometrics, 58(3):631–642, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bang:2002:MRC

[447] Heejung Bang and Anastasios A. Tsiatis. Median regression with censored cost data. Biometrics, 58(3):643–649, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tamhane:2002:ACC

[448] Ajit C. Tamhane and Brent R. Logan. Accurate critical constants for the one-sided approximate likelihood ratio test of a normal mean vector when the covariance matrix is estimated. Biometrics, 58(3):650–656, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McIntosh:2002:CSS

[449] Martin W. McIntosh and Margaret Sullivan Pepe. Combining several screening tests: Optimality of the risk score. Biometrics, 58(3):657–664, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2002:QCM

[450] Jing Qin, Marianne Berwick, Rosie Ashbolt, and Terry Dwyer. Quantifying the change of melanoma incidence by breslow thickness. Biometrics, 58(3):665–670, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2002:STE

[451] Ying Kuen Cheung and Rick Chappell. A simple technique to evaluate model sensitivity in the continual reassessment method. Biometrics, 58(3):671–674, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neuhaus:2002:ACL

[452] John M. Neuhaus. Analysis of clustered and longitudinal binary data subject to response misclassification. Biometrics, 58(3):675–683, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clegg:2002:EVD

[453] Limin X. Clegg, Mitchell H. Gail, and Eric J. Feuer. Estimating the variance of disease-prevalence estimates from population-based registries. Biometrics, 58(3):684–688, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kotz:2002:LEB

[454] Samuel Kotz, Saralees Nadarajah, and J. K. Lindsey. Letter to the Editor of Biometrics. Biometrics, 58(3):689–690, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2002:BRc

[455] M. S. Ridout. Book reviews. Biometrics, 58(3):691–698, September 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nguyen:2002:DME

[456] Danh V. Nguyen, A. Bulak Arpat, Naisyin Wang, and Raymond J. Carroll. DNA microarray experiments: Biological and technological aspects. Biometrics, 58(4):701–717, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

FitzGerald:2002:EGE

[457] Patrick E. B. FitzGerald. Extended generalized estimating equations for binary familial data with incomplete families. Biometrics, 58(4):718–726, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mills:2002:MSG

[458] J. E. Mills, C. A. Field, and D. J. Dupuis. Marginally specified generalized linear mixed models: a robust approach. Biometrics, 58(4):727–734, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ouwens:2002:MOD

[459] Mario J. N. M. Ouwens, Prans E. S. Tan, and Martijn P. F. Berger. Maximin D-optimal designs for longitudinal mixed effects models. Biometrics, 58(4):735–741, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2002:SLA

[460] Xiao Song, Marie Davidian, and Anastasios A. Tsiatis. A semiparametric likelihood approach to joint modeling of longitudinal and time-to-event data. Biometrics, 58(4):742–753, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maples:2002:TLP

[461] Jerry J. Maples, Susan A. Murphy, and William G. Axinn. Two-level proportional hazards models. Biometrics, 58(4):754–763, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2002:CCC

[462] Shou-En Lu and Mei-Cheng Wang. Cohort case-control design and analysis for clustered failure-time data. Biometrics, 58(4):764–772, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:2002:SIC

[463] Peter B. Gilbert, L. J. Wei, Michael R. Kosorok, and John D. Clemens. Simultaneous inferences on the contrast of two hazard functions with censored observations. Biometrics, 58(4):773–780, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2002:PPL

[464] Jie Huang and David Harrington. Penalized partial likelihood regression for right-censored data with bootstrap selection of the penalty parameter. Biometrics, 58(4):781–791, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Datta:2002:EIT

[465] Somnath Datta and Glen A. Satten. Estimation of integrated transition hazards and stage occupation probabilities for non-Markov systems under dependent censoring. Biometrics, 58(4):792–802, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2002:MPT

[466] Yue Wang and Jeremy M. G. Taylor. A measure of the proportion of treatment effect explained by a surrogate marker. Biometrics, 58(4): 803–812, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2002:BMI

[467] David B. Dunson and Donna D. Baird. Bayesian modeling of incidence and progression of disease from cross-sectional data. Biometrics, 58(4): 813–822, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Inoue:2002:SER

[468] Lurdes Y. T. Inoue, Peter F. Thall, and Donald A. Berry. Seamlessly expanding a randomized Phase II trial to Phase III. Biometrics, 58(4): 823–831, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeh:2002:SBW

[469] Judith Zeh, David Poole, Gary Miller, William Koski, Lisa Baraff, and David Hugh. Survival of bowhead whales, Balaena mysticetus, estimated from 1981–1998 photoidentification data. Biometrics, 58(4):832–840, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2002:MSI

[470] R. King and S. P. Brooks. Model selection for integrated recovery/ recapture data. Biometrics, 58(4):841–851, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yip:2002:RDU

[471] Paul S. F. Yip, K. S. Chan, and Emmy C. Y. Wan. Robust design and the use of auxiliary variables in estimating population size for an open population. Biometrics, 58(4):852–861, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pollard:2002:ALT

[472] J. H. Pollard, D. Palka, and S. T. Buckland. Adaptive line transect sampling. Biometrics, 58(4):862–870, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staudenmayer:2002:SRP

[473] John Staudenmayer and Donna Spiegelman. Segmented regression in the presence of covariate measurement error in main study/validation study designs. Biometrics, 58(4):871–877, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gustafson:2002:CEC

[474] Paul Gustafson and Nhu D. Le. Comparing the effects of continuous and discrete covariate mismeasurement, with emphasis on the dichotomization of mismeasured predictors. Biometrics, 58(4):878–887, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cerioli:2002:TMI

[475] Andrea Cerioli. Testing mutual independence between two discretevalued spatial processes: a correction to Pearson chi-squared. Biometrics, 58(4):888–897, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guthrie:2002:HAD

[476] Katherine A. Guthrie, Lianne Sheppard, and Jon Wakefield. A hierarchical aggregate data model with spatially correlated disease rates. Biometrics, 58(4):898–905, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

French:2002:BMT

[477] Jonathan L. French and Joseph G. Ibrahim. Bayesian methods for a three-state model for rodent carcinogenicity studies. Biometrics, 58(4): 906–916, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2002:UHC

[478] Yanping Ma, Jianhua Guo, Ning-Zhong Shi, and Man-Lai Tang. On the use of historical control information for trend test in carcinogenesis. Biometrics, 58(4):917–927, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glasbey:2002:ETP

[479] C. A. Glasbey and C. D. Robinson. Estimators of tissue proportions from X-ray CT images. Biometrics, 58(4):928–936, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:2002:SML

[480] Hans-Georg Müller, Chun-Lung Su, Stephen R. Dueker, Yumei Lin, Andrew Clifford, Bruce A. Buchholz, and John S. Vogel. Semiparametric modeling of labeled-cell kinetics, with application to isotope labeling of erythrocytes. Biometrics, 58(4):937–945, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsimikas:2002:PLI

[481] John V. Tsimikas, Ronald J. Bosch, Brent A. Coull, and Hammou El Barmi. Profile-likelihood inference for highly accurate diagnostic tests. Biometrics, 58(4):946–956, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2002:SSD

[482] Man-Lai Tang, Nian-Sheng Tang, Ivan Siu-Fung Chan, and Ben Ping-Shing Chan. Sample size determination for establishing equivalence/ noninferiority via ratio of two proportions in matched-pair design. Biometrics, 58(4):957–963, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nahhas:2002:RSS

[483] Ramzi W. Nahhas, Douglas A. Wolfe, and Haiying Chen. Ranked set sampling: Cost and optimal set size. Biometrics, 58(4):964–971, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2002:EUI

[484] Nian-Sheng Tang and Man-Lai Tang. Exact unconditional inference for risk ratio in a correlated 2 × 2 table with structural zero. Biometrics, 58(4):972–980, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2002:SMI

[485] Mei-Ling Ting Lee, Martha L. Bulyk, G. A. Whitmore, and George M. Church. A statistical model for investigating binding probabilities of DNA nucleotide sequences using microarrays. Biometrics, 58(4):981–988, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Birmingham:2002:PMM

[486] Jolene Birmingham and Garrett M. Fitzmaurice. A pattern-mixture model for longitudinal binary responses with nonignorable nonresponse. Biometrics, 58(4):989–996, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2002:FWL

[487] Lang Wu and Peter B. Gilbert. Flexible weighted log-rank tests optimal for detecting early and/or late survival differences. Biometrics, 58(4): 997–1004, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2002:VAS

[488] Thomas D. Cook. P -value adjustment in sequential clinical trials. Biometrics, 58(4):1005–1011, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barnhart:2002:WLS

[489] Huiman X. Barnhart and John M. Williamson. Weighted least-squares approach for comparing correlated kappa. Biometrics, 58(4):1012–1019, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barnhart:2002:OCC

[490] Huiman X. Barnhart, Michael Haber, and Jingli Song. Overall concordance correlation coefficient for evaluating agreement among multiple observers. Biometrics, 58(4):1020–1027, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dahm:2002:PMA

[491] P. Fred Dahm, Ann W. Olmsted, and Ira F. Greenbaum. Probability models and the applicability of statistical procedures in the identification of chromosomal fragile sites. Biometrics, 58(4):1028–1031, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See reply [492].

Hou:2002:RAA

[492] Chia-Ding Hou and John Jen Tai. Rationale and applicability of analysis methods for identification of chromosomal fragile sites — a reply to Drs Dahm et al. Biometrics, 58(4):1032–1033, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [491].

Lyles:2002:NEC

[493] Robert H. Lyles. A note on estimating crude odds ratios in case-control studies with differentially misclassified exposure. Biometrics, 58(4):1034– 1036, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morrissey:2002:E

[494] Mary Morrissey and Donna Spiegelman. To the Editor. Biometrics, 58(4):1036–1037, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2002:BR

[495] Anonymous. Book reviews. Biometrics, 58(4):1038–1048, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2002:A

[496] Anonymous. Acknowledgements. Biometrics, 58(4):1049–1050, December 2002. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Breslow:2003:SCM

[497] Norman E. Breslow. Are statistical contributions to medicine undervalued? Biometrics, 59(1):1–8, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wakefield:2003:SAE

[498] Jon Wakefield. Sensitivity analyses for ecological regression. Biometrics, 59(1):9–17, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stefanescu:2003:LIE

[499] Catalina Stefanescu and Bruce W. Turnbull. Likelihood inference for exchangeable binary data with varying cluster sizes. Biometrics, 59(1): 18–24, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2003:TCC

[500] Yi Li and Xihong Lin. Testing the correlation for clustered categorical and censored discrete time-to-event data when covariates are measured without/with errors. Biometrics, 59(1):25–35, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williamson:2003:MAC

[501] John M. Williamson, Somnath Datta, and Glen A. Satten. Marginal analyses of clustered data when cluster size is informative. Biometrics, 59(1):36–42, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pradel:2003:PGF

[502] Roger Pradel, Claire M. A. Wintrebert, and Olivier Gimenez. A proposal for a goodness-of-fit test to the arnason-Schwarz multisite capture– recapture model. Biometrics, 59(1):43–53, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barry:2003:ARR

[503] S. C. Barry, S. P. Brooks, E. A. Catchpole, and B. J. T. Morgan. The analysis of ring-recovery data using random effects. Biometrics, 59(1): 54–65, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lopes:2003:BMA

[504] Hedibert Freitas Lopes, Peter Müller, and Gary L. Rosner. Bayesian meta-analysis for longitudinal data models using multivariate mixture priors. Biometrics, 59(1):66–75, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gonen:2003:BMT

[505] Mithat Gönen, Peter H. Westfall, and Wesley O. Johnson. Bayesian multiple testing for two-sample multivariate endpoints. Biometrics, 59 (1):76–82, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2003:BMT

[506] David B. Dunson, Patricia Chulada, and Samuel J. Arbes, Jr. Bayesian modeling of time-varying and waning exposure effects. Biometrics, 59 (1):83–91, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenland:2003:GCP

[507] Sander Greenland. Generalized conjugate priors for Bayesian analysis of risk and survival regressions. Biometrics, 59(1):92–99, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loeys:2003:CPH

[508] T. Loeys and E. Goetghebeur. A causal proportional hazards estimator for the effect of treatment actually received in a randomized trial with allor-nothing compliance. Biometrics, 59(1):100–105, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brannath:2003:STN

[509] W. Brannath, P. Bauer, W. Maurer, and M. Posch. Sequential tests for noninferiority and superiority. Biometrics, 59(1):106–114, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joffe:2003:CCF

[510] Marshall M. Joffe. A case-control follow-up study for disease-specific mortality. Biometrics, 59(1):115–125, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [561].

Fu:2003:PEE

[511] Wenjiang J. Fu. Penalized estimating equations. Biometrics, 59(1):126– 132, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pepe:2003:SDE

[512] Margaret Sullivan Pepe, Gary Longton, Garnet L. Anderson, and Michel Schummer. Selecting differentially expressed genes from microarray experiments. Biometrics, 59(1):133–142, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2003:DRU

[513] Yinsheng Qu, Bao ling Adam, Mark Thornquist, John D. Potter, Mary Lou Thompson, Yutaka Yasui, John Davis, Paul F. Schellhammer, Lisa Cazares, MaryAnn Clements, George L. Wright Jr., and Ziding Feng. Data reduction using a discrete wavelet transform in discriminant analysis of very high dimensionality data. Biometrics, 59(1):143–151, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ko:2003:ECT

[514] Hyejin Ko, Joseph W. Hogan, and Kenneth H. Mayer. Estimating causal treatment effects from longitudinal HIV natural history studies using marginal structural models. Biometrics, 59(1):152–162, March 2003. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kosinski:2003:ANV

[515] Andrzej S. Kosinski and Huiman X. Barnhart. Accounting for nonignorable verification bias in assessment of diagnostic tests. Biometrics, 59 (1):163–171, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liquet:2003:BCE

[516] B. Liquet, C. Sakarovitch, and D. Commenges. Bootstrap choice of estimators in parametric and semiparametric families: an extension of EIC. Biometrics, 59(1):172–178, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Richardson:2003:MBI

[517] Barbra A. Richardson and James P. Hughes. Modeling breastmilk infectivity in HIV-1 infected mothers. Biometrics, 59(1):179–185, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aldrin:2003:CCS

[518] Magne Aldrin, Marit Holden, and Tore Schweder. Comment on Cowling’s “Spatial Methods for Line Transect Surveys”. Biometrics, 59(1):186–188, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Formann:2003:LCM

[519] Anton K. Formann. Latent class model diagnosis from a frequentist point of view. Biometrics, 59(1):189–196, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeger:2003:RLC

[520] Scott L. Zeger and Elizabeth S. Garrett. Rejoinder to “Latent Class Model Diagnosis from a Frequentist Point of View”. Biometrics, 59(1): 197–198, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:CEI

[521] Anonymous. Correction to “Estimating Intraclass Correlation for Binary Data” by M. S. Ridout, C. G. B. Demétrio, and D. Firth; 55, 137–148, March 1999. Biometrics, 59(1):199, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [3].

Nunez-Anton:2003:BRH

[522] V. Núñez-Antón. Book review: Huber-Carol, C., Balakrishnan, N., Nikulin, M. S. and Mesbah, M. (Editors). Goodness-of-Fit Tests and Model Validity. Birkhäuser, Boston, 2002. xxxiii + 507 pp. $89.95/Euro 136.00. ISBN 0-8176-4209-9. Biometrics, 59(1):200–201, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bowman:2003:BRG

[523] A. W. Bowman. Book review: Gu, C., Smoothing Spline ANOVA Models. Springer, New York, 2002. xiii + 289 pp. $79.95/£56.00. ISBN 0-387-95353-1. Biometrics, 59(1):201–202, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steel:2003:BRS

[524] M. F. J. Steel. Book review: Stulajter, F., Predictions in Time Series Using Regression Models. Springer, New York, 2002. xix + 231 pp. $69.95/£52.50. ISBN 0-387-95350-7. Biometrics, 59(1):202, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Medley:2003:BRD

[525] G. F. Medley. Book review: Dieckmann, U., Metz, J. A. J., Sabelis, M. W. And Sigmund, K. (Editors). Adaptive Dynamics of Infectious Diseases: In Pursuit of Virulence Management. Cambridge University Press, Cambridge, 2002. xvii + 532 pp. £55.00/$80.00. ISBN 0-521-78165-5. Biometrics, 59(1):202–203, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edwards:2003:BRI

[526] A. W. F. Edwards. Book review: Israel, G. and Gasca, A. M., The Biology of Numbers: The Correspondence of Vito Volterra on Mathematical Biology. Birkhäuser, Basel, 2002. ix + 405 pp. $89.95/Euro 74.00. ISBN 3-7643-6514-5. Biometrics, 59(1):203, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rudolfer:2003:BRZ

[527] S. M. Rudolfer. Book review: Zhou, X.-H., Obuchowski, N. A. and Mc-Clish, D. K., Statistical Methods in Diagnostic Medicine. Wiley, New York, 2002. xv + 437 pp. $94.95/£70.50. ISBN 0-471-34772-8. Biometrics, 59(1):203–204, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Woodward:2003:BRP

[528] P. Woodward. Book review: Press, S. J. and Tanur, J. M., The Subjectivity of Scientists and the Bayesian Approach. Wiley, New York, 2001. x + 274. $84.95/£59.50. ISBN 0-471-39685-0. Biometrics, 59(1):204–205, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Payne:2003:BRD

[529] R. W. Payne. Book review: Davies, C. S., Statistical Methods for the Analysis of Repeated Measurements. Springer, New York, 2002. xxiv + 415 pp. $79.95/£56.00. ISBN 0-387-95370-1. Biometrics, 59(1):205–206, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balding:2003:BRD

[530] D. J. Balding. Book review: Durrett, R., Probability Models for DNA Sequence Evolution. Springer, New York, 2002. viii + 240 pp. $69.95/ £52.50. ISBN 0-387-95435-X. Biometrics, 59(1):206, March 2003. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rothery:2003:BRG

[531] P. Rothery. Book review: Grafen, A. and Hails, R. Modern Statistics for the Life Sciences. Oxford University Press, Oxford, 2002. xv + 351 pp. £22.99/$34.95. ISBN 0-19-925231-9 (pbk). Biometrics, 59(1):206–207, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neuhauser:2003:BRH

[532] M. Neuhäuser. Book review: Hardy, I. C. W. (editor). Sex Ratios: Concepts and Research Methods. Cambridge University Press, Cambridge, 2002. xiii + 424 pp. £32.95/$48.00. ISBN 0-521-66578-7. Biometrics, 59 (1):207–208, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:BREa

[533] Anonymous. Brief reports by the Editor. Biometrics, 59(1):208–209, March 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carroll:2003:VAN

[534] Raymond J. Carroll. Variances are not always nuisance parameters. Biometrics, 59(2):211–220, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:2003:BSJ

[535] Elizabeth R. Brown and Joseph G. Ibrahim. A Bayesian semiparametric joint hierarchical model for longitudinal and survival data. Biometrics, 59(2):221–228, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hardwick:2003:DWD

[536] Janis Hardwick, Mary C. Meyer, and Quentin F. Stout. Directed walk designs for dose-response problems with competing failure modes. Biometrics, 59(2):229–236, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krishnamoorthy:2003:ICM

[537] K. Krishnamoorthy and Yong Lu. Inferences on the common mean of several normal populations based on the generalized variable method. Biometrics, 59(2):237–247, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nunez:2003:MDP

[538] Marian Núñez, Angel Villarroya, and José María Oller. Minimum distance probability discriminant analysis for mixed variables. Biometrics, 59(2):248–253, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Verbeke:2003:UST

[539] Geert Verbeke and Geert Molenberghs. The use of score tests for inference on variance components. Biometrics, 59(2):254–262, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tutz:2003:GSS

[540] Gerhard Tutz. Generalized semiparametrically structured ordinal models. Biometrics, 59(2):263–273, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brumback:2003:ISA

[541] Babette Brumback, Sander Greenland, Mary Redman, Nancy Kiviat, and Paula Diehr. The intensity-score approach to adjusting for confounding. Biometrics, 59(2):274–285, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2003:BIO

[542] David B. Dunson and Brian Neelon. Bayesian inference on orderconstrained parameters in generalized linear models. Biometrics, 59(2): 286–295, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2003:BLV

[543] David B. Dunson, M. Watson, and Jack A. Taylor. Bayesian latent variable models for median regression on multiple outcomes. Biometrics, 59(2):296–304, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

MacNab:2003:HBM

[544] Ying C. MacNab. Hierarchical Bayesian modeling of spatially correlated health service outcome and utilization rates. Biometrics, 59(2):305–315, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hodges:2003:PCA

[545] James S. Hodges, Bradley P. Carlin, and Qiao Fan. On the precision of the conditionally autoregressive prior in spatial models. Biometrics, 59(2):317–322, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kulldorff:2003:TBS

[546] Martin Kulldorff, Zixing Fang, and Stephen J. Walsh. A tree-based scan statistic for database disease surveillance. Biometrics, 59(2):323– 331, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ramsey:2003:PHH

[547] Fred L. Ramsey and Dale Usner. Persistence and heterogeneity in habitat selection studies using radio telemetry. Biometrics, 59(2):332–340, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2003:AMC

[548] Devin S. Johnson and Jennifer A. Hoeting. Autoregressive models for capture–recapture data: a Bayesian approach. Biometrics, 59(2):341– 350, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dorazio:2003:MME

[549] Robert M. Dorazio and J. Andrew Royle. Mixture models for estimating the size of a closed population when capture rates vary among individuals. Biometrics, 59(2):351–364, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2003:ESO

[550] Hsin-Chou Yang, Richard Huggins, and Austina S. S. Clark. Estimation of the size of an open population using local estimating equations II: a partially parametric approach. Biometrics, 59(2):365–374, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Filho:2003:PIB

[551] Júlio S. de S. Bueno Filho and Steven G. Gilmour. Planning incomplete block experiments when treatments are genetically related. Biometrics, 59(2):375–381, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tritchler:2003:TLC

[552] David Tritchler, Ying Liu, and Shafagh Fallah. A test of linkage for complex discrete and continuous traits in nuclear families. Biometrics, 59(2):382–392, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2003:IGG

[553] Yuqun Luo and Shili Lin. Information gain for genetic parameter estimation with incorporation of marker data. Biometrics, 59(2):393–401, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stallard:2003:DTD

[554] Nigel Stallard. Decision-theoretic designs for Phase II clinical trials allowing for competing studies. Biometrics, 59(2):402–409, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jansen:2003:LIA

[555] Ivy Jansen, Geert Molenberghs, Marc Aerts, Herbert Thijs, and Kristel Van Steen. A local influence approach applied to binary data from a psychiatric study. Biometrics, 59(2):410–419, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:2003:MOI

[556] Dean Follmann, Michael Proschan, and Eric Leifer. Multiple outputation: Inference for complex clustered data by averaging analyses from independent data. Biometrics, 59(2):420–429, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2003:CRM

[557] John O’Quigley and Xavier Paoletti. Continual reassessment method for ordered groups. Biometrics, 59(2):430–440, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mehrotra:2003:CNE

[558] Devan V. Mehrotra, Ivan S. F. Chan, and Roger L. Berger. A cautionary note on exact unconditional inference for a difference between two independent binomial proportions. Biometrics, 59(2):441–450, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:CEA

[559] Anonymous. Correction to “Estimation of absolute risk from nested casecontrol data” by B. Langholz and Ø. Borgan; 53, 767–774, June 1997. Biometrics, 59(2):451, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [2].

Anonymous:2003:CCA

[560] Anonymous. Correction to “Covariate adjustments of event histories estimated from Markov chains: The additive approach” by O. O. Aalen, Ø. Borgan, and H. Fekjær; 57, 993–1001, December 2001. Biometrics, 59(2):452–453, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [332].

Anonymous:2003:CCC

[561] Anonymous. Correction to “A Case-control Follow-up Study for Disease-Specific Mortality” by M. M. Joffe; 59, 115–125, March 2003. Biometrics, 59(2):454, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [510].

Ridout:2003:BRa

[562] M. S. Ridout. Book reviews. Biometrics, 59(2):455–465, June 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Becker:2003:EVE

[563] Niels G. Becker, Tom Britton, and Philip D. O’Neill. Estimating vaccine effects on transmission of infection from household outbreak data. Biometrics, 59(3):467–475, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morris:2003:BSE

[564] Jeffrey S. Morris, Keith A. Baggerly, and Kevin R. Coombes. Bayesian shrinkage estimation of the relative abundance of mRNA transcripts using SAGE. Biometrics, 59(3):476–486, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2003:DFT

[565] Peter F. Thall, Randall E. Millikan, Peter Mueller, and Sang-Joon Lee. Dose-finding with two agents in Phase I oncology trials. Biometrics, 59(3):487–496, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DiRienzo:2003:NCT

[566] A. G. DiRienzo. Nonparametric comparison of two survival-time distributions in the presence of dependent censoring. Biometrics, 59(3): 497–504, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hughes:2003:EAP

[567] James P. Hughes and Patricia Totten. Estimating the accuracy of polymerase chain reaction-based tests using endpoint dilution. Biometrics, 59(3):505–511, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Okamura:2003:AED

[568] Hiroshi Okamura, Toshihide Kitakado, Kazuhiko Hiramatsu, and Mitsuyo Mori. Abundance estimation of diving animals by the doubleplatform line transect method. Biometrics, 59(3):512–520, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2003:BAJ

[569] David B. Dunson, Zhen Chen, and Jean Harry. A Bayesian approach for joint modeling of cluster size and subunit-specific outcomes. Biometrics, 59(3):521–530, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comments [955] and rejoinder [956].

Gilbert:2003:SAA

[570] Peter B. Gilbert, Ronald J. Bosch, and Michael G. Hudgens. Sensitivity analysis for the assessment of causal vaccine effects on viral load in HIV vaccine trials. Biometrics, 59(3):531–541, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tadesse:2003:IDE

[571] Mahlet G. Tadesse, Joseph G. Ibrahim, and George L. Mutter. Identification of differentially expressed genes in high-density oligonucleotide arrays accounting for the quantification limits of the technology. Biometrics, 59(3):542–554, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Natarajan:2003:ESM

[572] Loki Natarajan, Charles C. Berry, and Christoph Gasche. Estimation of spontaneous mutation rates. Biometrics, 59(3):555–561, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morgan:2003:MYP

[573] B. J. T. Morgan, M. S. Ridout, and L. W. Ruddock. Models for yeast prions. Biometrics, 59(3):562–569, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2003:HRI

[574] Tianxi Cai and Rebecca A. Betensky. Hazard regression for intervalcensored data with penalized spline. Biometrics, 59(3):570–579, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comments [1917].

Jiroutek:2003:NMC

[575] Michael R. Jiroutek, Keith E. Muller, Lawrence L. Kupper, and Paul W. Stewart. A new method for choosing sample size for confidence intervalbased inferences. Biometrics, 59(3):580–590, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haines:2003:BOD

[576] Linda M. Haines, Inna Perevozskaya, and William F. Rosenberger. Bayesian optimal designs for Phase I clinical trials. Biometrics, 59(3): 591–600, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scharfstein:2003:GAS

[577] Daniel O. Scharfstein and Rafael A. Irizarry. Generalized additive selection models for the analysis of studies with potentially nonignorable missing outcome data. Biometrics, 59(3):601–613, September 2003. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dodd:2003:PAE

[578] Lori E. Dodd and Margaret S. Pepe. Partial AUC estimation and regression. Biometrics, 59(3):614–623, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xiong:2003:SCP

[579] Xiaoping Xiong, Ming Tan, and James Boyett. Sequential conditional probability ratio tests for normalized test statistic on information time. Biometrics, 59(3):624–631, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2003:AMM

[580] Yan Li, Daniel Zelterman, and Brian W. C. Forsyth. Analyzing multiply matched cohort studies with two different comparison groups: Application to pregnancy rates among HIV+ women. Biometrics, 59(3):632–639, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fewster:2003:SSP

[581] R. M. Fewster. A spatiotemporal stochastic process model for species spread. Biometrics, 59(3):640–649, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2003:BDA

[582] Timothy D. Johnson. Bayesian deconvolution analysis of pulsatile hormone concentration profiles. Biometrics, 59(3):650–660, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Braun:2003:PMR

[583] W. John Braun, Valentin Rousson, William A. Simpson, and Jennifer Prokop. Parametric modeling of reaction time experiment data. Biometrics, 59(3):661–669, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Paulino:2003:BRM

[584] Carlos Daniel Paulino, Paulo Soares, and John Neuhaus. Binomial regression with misclassification. Biometrics, 59(3):670–675, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yao:2003:SEF

[585] Fang Yao, Hans-Georg Müller, Andrew J. Clifford, Steven R. Dueker, Jennifer Follett, Yumei Lin, Bruce A. Buchholz, and John S. Vogel. Shrinkage estimation for functional principal component scores with application to the population kinetics of plasma folate. Biometrics, 59(3): 676–685, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:2003:BAJ

[586] Elizabeth R. Brown and Joseph G. Ibrahim. Bayesian approaches to joint cure-rate and longitudinal models with applications to cancer vaccine trials. Biometrics, 59(3):686–693, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jones:2003:BPS

[587] Beatrix Jones. Balancing population size and genetic information in parentage analysis studies. Biometrics, 59(3):694–700, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sooriyarachchi:2003:IDR

[588] Marina Roshini Sooriyarachchi, John Whitehead, Kim Bolland, and Anne Whitehead. Incorporating data received after a sequential trial has stopped into the final analysis: Implementation and comparison of methods. Biometrics, 59(3):701–709, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miglioretti:2003:LTR

[589] Diana L. Miglioretti. Latent transition regression for mixed outcomes. Biometrics, 59(3):710–720, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2003:GFM

[590] Debashis Ghosh. Goodness-of-fit methods for additive-risk models in tumorigenicity experiments. Biometrics, 59(3):721–726, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meester:2003:EVD

[591] Ronald Meester and Marjan Sjerps. The evidential value in the DNA database search controversy and the two-stain problem. Biometrics, 59(3):727–732, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wallenstein:2003:LEB

[592] Sylvan Wallenstein. Letter to the Editor of Biometrics. Biometrics, 59(3):733–734, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonaldi:2003:AR

[593] Christophe Bonaldi, Jean-Pierre Daurès, and Nicolas Molinari. Authors’ reply. Biometrics, 59(3):734, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Basford:2003:BR

[594] K. E. Basford. Book reviews: 1. Biometrics, 59(3):735–736, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heimann:2003:BR

[595] G. Heimann. Book reviews: 2. Biometrics, 59(3):736, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Billheimer:2003:BR

[596] D. Billheimer. Book reviews: 3. Biometrics, 59(3):736–737, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2003:BR

[597] J. C. Lee. Book reviews: 4. Biometrics, 59(3):737–738, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mead:2003:BR

[598] A. Mead. Book reviews: 5. Biometrics, 59(3):738–739, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2003:BRb

[599] M. S. Ridout. Book reviews: 6. Biometrics, 59(3):739, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:BREb

[600] Anonymous. Brief reports by the Editor. Biometrics, 59(3):740, September 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:EN

[601] Anonymous. Editor’s note. Biometrics, 59(3):740, September 2003. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dobson:2003:DJL

[602] Angela Dobson and Robin Henderson. Diagnostics for joint longitudinal and dropout time modeling. Biometrics, 59(4):741–751, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Apanasovich:2003:TSC

[603] Tatiyana V. Apanasovich, Simon Sheather, Joanne R. Lupton, Natasa Popovic, Nancy D. Turner, Robert S. Chapkin, Leslie A. Braby, and Raymond J. Carroll. Testing for spatial correlation in nonstationary binary data, with application to aberrant crypt foci in colon carcinogenesis. Biometrics, 59(4):752–761, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2003:RES

[604] Zhen Chen and David B. Dunson. Random effects selection in linear mixed models. Biometrics, 59(4):762–769, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burington:2003:FIG

[605] Bart E. Burington and Scott S. Emerson. Flexible implementations of group sequential stopping rules using constrained boundaries. Biometrics, 59(4):770–777, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Link:2003:DAS

[606] William A. Link, J. Andrew Royle, and Jeff S. Hatfield. Demographic analysis from summaries of an age-structured population. Biometrics, 59(4):778–785, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pledger:2003:OCR

[607] Shirley Pledger, Kenneth H. Pollock, and James L. Norris. Open capture– recapture models with heterogeneity: I. Cormack–Jolly–Seber model.

Biometrics, 59(4):786–794, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ekholm:2003:JRA

[608] Anders Ekholm, Jukka Jokinen, John W. McDonald, and Peter W. F. Smith. Joint regression and association modeling of longitudinal ordinal data. Biometrics, 59(4):795–803, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2003:SIM

[609] Yuedong Wang, Chunlei Ke, and Morton B. Brown. Shape-invariant modeling of circadian rhythms with random effects and smoothing spline ANOVA decompositions. Biometrics, 59(4):804–812, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2003:SSM

[610] Wenyang Zhang, Qiwei Yao, Howell Tong, and Nils Chr. Stenseth. Smoothing for spatiotemporal models and its application to modeling muskrat-mink interaction. Biometrics, 59(4):813–821, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kerr:2003:DCE

[611] M. Kathleen Kerr. Design considerations for efficient and effective microarray studies. Biometrics, 59(4):822–828, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:2003:MLD

[612] Jason Roy. Modeling longitudinal data with nonignorable dropouts using a latent dropout class model. Biometrics, 59(4):829–836, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2003:FML

[613] Wenqing He and Jerald F. Lawless. Flexible maximum likelihood methods for bivariate proportional hazards models. Biometrics, 59(4):837– 848, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carrasco:2003:EGC

[614] Josep L. Carrasco and Lluís Jover. Estimating the generalized concordance correlation coefficient through variance components. Biometrics, 59(4):849–858, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berhane:2003:ISB

[615] Kiros Berhane and Lisa A. Weissfeld. Inference in spline-based models for multiple time-to-event data, with applications to a breast cancer prevention trial. Biometrics, 59(4):859–868, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cadigan:2003:SET

[616] N. G. Cadigan and J. Brattey. Semiparametric estimation of tag loss and reporting rates for tag-recovery experiments using exact time-at-Liberty data. Biometrics, 59(4):869–876, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2003:SAR

[617] Debashis Ghosh and D. Y. Lin. Semiparametric analysis of recurrent events data in the presence of dependent censoring. Biometrics, 59(4): 877–885, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Green:2003:BHM

[618] Paul E. Green and Taesung Park. A Bayesian hierarchical model for categorical data with nonignorable nonresponse. Biometrics, 59(4):886– 896, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2003:MTG

[619] Paul S. Albert and Joanna H. Shih. Modeling tumor growth with random onset. Biometrics, 59(4):897–906, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hirotsu:2003:PAH

[620] Chihiro Hirotsu, Eri Ohta, Nobuyoshi Hirose, and Kenichiro Shimizu. Profile analysis of 24-hours measurements of blood pressure. Biometrics, 59(4):907–915, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2003:BIC

[621] David B. Dunson and Amy H. Herring. Bayesian inferences in the Cox model for order-restricted hypotheses. Biometrics, 59(4):916–923, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marques:2003:ICS

[622] Fernanda F. C. Marques and Stephen T. Buckland. Incorporating covariates into standard line transect analyses. Biometrics, 59(4):924–935, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dukic:2003:MAD

[623] V. Dukic and C. Gatsonis. Meta-analysis of diagnostic test accuracy assessment studies with varying number of thresholds. Biometrics, 59(4): 936–946, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meier:2003:DPH

[624] Amalia S. Meier, Barbra A. Richardson, and James P. Hughes. Discrete proportional hazards models for mismeasured outcomes. Biometrics, 59(4):947–954, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2003:MDB

[625] Jun Zhu, Jens C. Eickhoff, and Mark S. Kaiser. Modeling the dependence between number of trials and success probability in beta-binomial-Poisson mixture distributions. Biometrics, 59(4):955–961, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2003:BMA

[626] Chong Z. He. Bayesian modeling of age-specific survival in bird nesting studies under irregular visits. Biometrics, 59(4):962–973, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schweder:2003:AEM

[627] Tore Schweder. Abundance estimation from multiple photo surveys: Confidence distributions and reduced likelihoods for bowhead whales off Alaska. Biometrics, 59(4):974–983, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peterson:2003:ULC

[628] Derick R. Peterson, Hongwei Zhao, and Sara Eapen. Using local correlation in kernel-based smoothers for dependent data. Biometrics, 59(4): 984–991, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2003:PDM

[629] Debashis Ghosh. Penalized discriminant methods for the classification of tumors from gene expression data. Biometrics, 59(4):992–1000, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ivanova:2003:NDF

[630] Anastasia Ivanova. A new dose-finding design for bivariate outcomes. Biometrics, 59(4):1001–1007, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tian:2003:EPL

[631] L. Tian, W. Wang, and L. J. Wei. Estimating predictors for long- or short-term survivors. Biometrics, 59(4):1008–1015, December 2003. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2003:MHD

[632] Zudi Lu, Yer Van Hui, and Andy H. Lee. Minimum Hellinger distance estimation for finite mixtures of Poisson regression models and its applications. Biometrics, 59(4):1016–1026, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nam:2003:HST

[633] Jun mo Nam. Homogeneity score test for the intraclass version of the kappa statistics and sample-size determination in multiple or stratified studies. Biometrics, 59(4):1027–1035, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scheike:2003:EAC

[634] Thomas H. Scheike and Mei-Jie Zhang. Extensions and applications of the Cox–Aalen survival model. Biometrics, 59(4):1036–1045, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leon:2003:SET

[635] Selene Leon, Anastasios A. Tsiatis, and Marie Davidian. Semiparametric estimation of treatment effect in a pretest-posttest study. Biometrics, 59 (4):1046–1055, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Al-Saidy:2003:CBL

[636] Obaid M. Al-Saidy, Walter W. Piegorsch, R. Webster West, and Daniela K. Nitcheva. Confidence bands for low-dose risk estimation with quantal response data. Biometrics, 59(4):1056–1062, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dewanji:2003:ECR

[637] Anup Dewanji and Debasis Sengupta. Estimation of competing risks with general missing pattern in failure types. Biometrics, 59(4):1063–1070, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsai:2003:EFD

[638] Chen-An Tsai, Huey miin Hsueh, and James J. Chen. Estimation of false discovery rates in multiple testing: Application to gene microarray data. Biometrics, 59(4):1071–1081, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2003:BOM

[639] Xiaojie Zhou, Lawrence Joseph, David B. Wolfson, and Patrick Bélisle. A Bayesian A-optimal and model robust design criterion. Biometrics, 59(4):1082–1088, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosner:2003:ICE

[640] Bernard Rosner, Robert J. Glynn, and Mei-Ling Ting Lee. Incorporation of clustering effects for the Wilcoxon rank sum test: a largesample approach. Biometrics, 59(4):1089–1098, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2003:GFP

[641] Ying Zhang and Mortaza Jamshidian. The gamma-frailty Poisson model for the nonparametric estimation of panel count data. Biometrics, 59(4): 1099–1106, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shoukri:2003:ICE

[642] M. M. Shoukri, M. H. Asyali, and S. D. Walter. Issues of cost and efficiency in the design of reliability studies. Biometrics, 59(4):1107–1112, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hwang:2003:ECR

[643] Wen-Han Hwang and Steve Y. H. Huang. Estimation in capture– recapture models when covariates are subject to measurement errors.

Biometrics, 59(4):1113–1122, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Link:2003:NPS

[644] William A. Link. Nonidentifiability of population size from capture– recapture data with heterogeneous detection probabilities. Biometrics, 59(4):1123–1130, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wouters:2003:GEG

[645] Luc Wouters, Hinrich W. Göhlmann, Luc Bijnens, Stefan U. Kass, Geert Molenberghs, and Paul J. Lewi. Graphical exploration of gene expression data: a comparative study of three multivariate methods. Biometrics, 59(4):1131–1139, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stubbendick:2003:MLM

[646] Amy L. Stubbendick and Joseph G. Ibrahim. Maximum likelihood methods for nonignorable missing responses and covariates in random effects models. Biometrics, 59(4):1140–1150, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heinze:2003:ELR

[647] Georg Heinze, Michael Gnant, and Michael Schemper. Exact log-rank tests for unequal follow-up. Biometrics, 59(4):1151–1157, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2003:SRS

[648] Inyoung Kim, Noah D. Cohen, and Raymond J. Carroll. Semiparametric regression splines in matched case-control studies. Biometrics, 59(4): 1158–1169, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chan:2003:SAN

[649] Ivan S. F. Chan, Nian-Sheng Tang, Man-Lai Tang, and Ping-Shing Chan. Statistical analysis of noninferiority trials with a rate ratio in smallsample matched-pair designs. Biometrics, 59(4):1170–1177, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wienke:2003:BFM

[650] Andreas Wienke, Paul Lichtenstein, and Anatoli I. Yashin. A bivariate frailty model with a cure fraction for modeling familial correlations in diseases. Biometrics, 59(4):1178–1183, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chatterjee:2003:UBS

[651] Nilanjan Chatterjee and Joanna H. Shih. On use of bivariate survival models with cure fraction. Biometrics, 59(4):1184–1185, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laslett:2003:BSI

[652] Geoff M. Laslett, J. Paige Eveson, and Tom Polacheck. Biological and statistical issues in fitting growth curves to capture-recapture data. Biometrics, 59(4):1186–1187, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2003:RLE

[653] Yuk W. Cheng and Anthony Y. C. Kuk. Rejoinder to the letter to the Editors from G. M. Laslett, J. P. Eveson, and T. Polacheck. Biometrics, 59(4):1187–1189, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cottee:2003:BR

[654] Michaela J. Cottee. Book reviews: 1. Biometrics, 59(4):1190–1191, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walker:2003:BR

[655] S. G. Walker. Book reviews: 2. Biometrics, 59(4):1191, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Murray:2003:BR

[656] S. Murray. Book reviews: 3. Biometrics, 59(4):1192–1193, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Espejo:2003:BR

[657] M. Ruiz Espejo. Book reviews: 4. Biometrics, 59(4):1193, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Toutenburg:2003:BR

[658] H. Toutenburg. Book reviews: 5. Biometrics, 59(4):1193–1194, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kubler:2003:BR

[659] J. Kübler. Book reviews: 6. Biometrics, 59(4):1194–1195, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2003:BREc

[660] Anonymous. Brief reports by the Editor. Biometrics, 59(4):1195–1196, December 2003. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cullis:2004:RCE

[661] Brian Cullis, Daniel Commenges, Mike Kenward, and Xihong Lin. Report of the Co-Editors — 2003. Biometrics, 60(1):vii–ix, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2004:CEG

[662] Erning Li, Daowen Zhang, and Marie Davidian. Conditional estimation for generalized linear models when covariates are subject-specific parameters in a mixed model for longitudinal measurements. Biometrics, 60(1): 1–7, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2004:GLM

[663] Daowen Zhang. Generalized linear mixed models with varying coefficients for longitudinal data. Biometrics, 60(1):8–15, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hall:2004:MMN

[664] Daniel B. Hall and Michael Clutter. Multivariate multilevel nonlinear mixed effects models for timber yield predictions. Biometrics, 60(1):16– 24, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2004:BED

[665] Cong Han and Kathryn Chaloner. Bayesian experimental design for nonlinear mixed-effects models with application to HIV dynamics. Biometrics, 60(1):25–33, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:2004:BAA

[666] Debajyoti Sinha and Tapabrata Maiti. A Bayesian approach for the analysis of panel-count data with dependent termination. Biometrics, 60(1):34–40, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:2004:BSM

[667] Samiran Sinha, Bhramar Mukherjee, and Malay Ghosh. Bayesian semiparametric modeling for matched case-control studies with multiple disease states. Biometrics, 60(1):41–49, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oleson:2004:HBM

[668] Jacob J. Oleson and Chong Z. He. Hierarchical Bayesian modeling in dichotomous processes in the presence of nonresponse. Biometrics, 60 (1):50–59, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sisson:2004:BPE

[669] S. A. Sisson and M. A. Hurn. Bayesian point estimation of quantitative trait loci. Biometrics, 60(1):60–68, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Czika:2004:PMT

[670] Wendy Czika and B. S. Weir. Properties of the multiallelic trend test. Biometrics, 60(1):69–74, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leon:2004:FFD

[671] Larry F. León and Chih-Ling Tsai. Functional form diagnostics for Cox’s proportional hazards model. Biometrics, 60(1):75–84, March 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Larsen:2004:JAT

[672] Klaus Larsen. Joint analysis of time-to-event and multiple binary indicators of latent classes. Biometrics, 60(1):85–92, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Su:2004:MST

[673] Xiaogang Su and Juanjuan Fan. Multivariate survival trees: a maximum likelihood approach based on frailty models. Biometrics, 60(1):93–99, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rivest:2004:LMR

[674] Louis-Paul Rivest and Gaétan Daigle. Loglinear models for the robust design in mark-recapture experiments. Biometrics, 60(1):100–107, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Royle:2004:MME

[675] J. Andrew Royle. N-mixture models for estimating population size from spatially replicated counts. Biometrics, 60(1):108–115, March 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2004:EFA

[676] Jiahua Chen, Mary E. Thompson, and Changbao Wu. Estimation of fish abundance indices based on scientific research trawl surveys. Biometrics, 60(1):116–123, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wahed:2004:OES

[677] Abdus S. Wahed and Anastasios A. Tsiatis. Optimal estimator for the survival distribution and related quantities for treatment policies in twostage randomization designs in clinical trials. Biometrics, 60(1):124–133, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kosorok:2004:DAG

[678] Michael R. Kosorok, Shi Yuanjun, and David L. DeMets. Design and analysis of group sequential clinical trials with multiple primary endpoints. Biometrics, 60(1):134–145, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Copas:2004:BPB

[679] John Copas and Dan Jackson. A bound for publication bias based on the fraction of unpublished studies. Biometrics, 60(1):146–153, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scharfstein:2004:CBP

[680] Daniel O. Scharfstein, Charles F. Manski, and James C. Anthony. On the construction of bounds in prospective studies with missing ordinal outcomes: Application to the good behavior game trial. Biometrics, 60 (1):154–164, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2004:SSC

[681] Wesley O. Johnson, Chun-Lung Su, Ian A. Gardner, and Ronald Christensen. Sample size calculations for surveys to substantiate freedom of populations from infectious agents. Biometrics, 60(1):165–171, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Freedman:2004:NMD

[682] Laurence S. Freedman, Vitaly Fainberg, Victor Kipnis, Douglas Midthune, and Raymond J. Carroll. A new method for dealing with measurement error in explanatory variables of regression models. Biometrics, 60(1):172–181, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bryant:2004:SMC

[683] John Bryant and James J. Dignam. Semiparametric models for cumulative incidence functions. Biometrics, 60(1):182–190, March 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pinsky:2004:ELS

[684] Paul F. Pinsky. An early- and late-stage convolution model for disease natural history. Biometrics, 60(1):191–198, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yasui:2004:PSL

[685] Yutaka Yasui, Margaret Pepe, Li Hsu, Bao-Ling Adam, and Ziding Feng. Partially supervised learning using an EM-boosting algorithm. Biometrics, 60(1):199–206, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

MacEachern:2004:JPS

[686] Steven N. MacEachern, Elizabeth A. Stasny, and Douglas A. Wolfe. Judgement post-stratification with imprecise rankings. Biometrics, 60 (1):207–215, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Han:2004:EAD

[687] Karen E. Han, Paul J. Catalano, Pralay Senchaudhuri, and Cyrus Mehta. Exact analysis of dose response for multiple correlated binary outcomes. Biometrics, 60(1):216–224, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chan:2004:TCS

[688] K. S. Chan, H. Tong, and N. Chr. Stenseth. Testing for common structures in a panel of threshold models. Biometrics, 60(1):225–232, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2004:COO

[689] Paul R. Rosenbaum. The case-only odds ratio as a causal parameter. Biometrics, 60(1):233–240, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bilder:2004:TMI

[690] Christopher R. Bilder and Thomas M. Loughin. Testing for marginal independence between two categorical variables with multiple responses. Biometrics, 60(1):241–248, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pfeiffer:2004:MER

[691] Ruth M. Pfeiffer, Sam Mbulaiteye, and Eric Engels. A model to estimate risk of infection with human herpesvirus 8 associated with transfusion from cross-sectional data. Biometrics, 60(1):249–256, March 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2004:AEN

[692] Shou-En Lu, Yong Lin, and Wei-Chung Joe Shih. Analyzing excessive no changes in clinical trials with clustered data. Biometrics, 60(1):257– 267, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Banerjee:2004:PSC

[693] Sudipto Banerjee and Bradley P. Carlin. Parametric spatial cure rate models for interval-censored time-to-relapse data. Biometrics, 60(1):268– 275, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perlman:2004:NOS

[694] Michael D. Perlman and Lang Wu. A note on one-sided tests with multiple endpoints. Biometrics, 60(1):276–280, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1360].

Lin:2004:LPM

[695] Haiqun Lin, Charles E. McCulloch, and Robert A. Rosenheck. Latent pattern mixture models for informative intermittent missing data in longitudinal studies. Biometrics, 60(2):295–305, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Paik:2004:NMM

[696] Myunghee Cho Paik. Nonignorable missingness in matched case-control data analyses. Biometrics, 60(2):306–314, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2004:EMR

[697] Brent A. Johnson and Anastasios A. Tsiatis. Estimating mean response as a function of treatment duration in an observational study, where duration May be informatively censored. Biometrics, 60(2):315–323, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zucker:2004:IPH

[698] David M. Zucker and Donna Spiegelman. Inference for the proportional hazards model with misclassified discrete-valued covariates. Biometrics, 60(2):324–334, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Balasubramanian:2004:ATE

[699] Raji Balasubramanian and Stephen W. Lagakos. Analyzing time-to-event data in a clinical trial when an unknown proportion of subjects has experienced the event at entry. Biometrics, 60(2):335–343, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2004:ERE

[700] John O’Quigley and Loki Natarajan. Erosion of regression effect in a survival study. Biometrics, 60(2):344–351, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2004:SFP

[701] David B. Dunson and Zhen Chen. Selecting factors predictive of heterogeneity in multivariate event time data. Biometrics, 60(2):352–358, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McPeek:2004:BLU

[702] Mary Sara McPeek, Xiaodong Wu, and Carole Ober. Best linear unbiased allele-frequency estimation in complex pedigrees. Biometrics, 60(2):359– 367, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nicolae:2004:MRI

[703] Dan L. Nicolae and Augustine Kong. Measuring the relative information in allele-sharing linkage studies. Biometrics, 60(2):368–375, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kauermann:2004:MMD

[704] Göran Kauermann and Paul Eilers. Modeling microarray data using a threshold mixture model. Biometrics, 60(2):376–387, June 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dendukuri:2004:BSS

[705] Nandini Dendukuri, Elham Rahme, Patrick Bélisle, and Lawrence Joseph. Bayesian sample size determination for prevalence and diagnostic test studies in the absence of a gold standard test. Biometrics, 60(2):388–397, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neelon:2004:BIR

[706] Brian Neelon and David B. Dunson. Bayesian isotonic regression and trend analysis. Biometrics, 60(2):398–406, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gelman:2004:BAS

[707] Andrew Gelman, Ginger L. Chew, and Michael Shnaidman. Bayesian analysis of serial dilution assays. Biometrics, 60(2):407–417, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berry:2004:AMA

[708] Scott M. Berry and Donald A. Berry. Accounting for multiplicities in assessing drug safety: a three-level hierarchical mixture model. Biometrics, 60(2):418–426, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2004:CNR

[709] Paul S. Albert and Lori E. Dodd. A cautionary note on the robustness of latent class models for estimating diagnostic error without a gold standard. Biometrics, 60(2):427–435, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2004:CMM

[710] Richard J. Cook, Grace Y. Yi, Ker-Ai Lee, and Dafna D. Gladman. A conditional Markov model for clustered progressive multistate processes under incomplete observation. Biometrics, 60(2):436–443, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Altman:2004:AGF

[711] Rachel MacKay Altman. Assessing the goodness-of-fit of hidden Markov models. Biometrics, 60(2):444–450, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2004:PPM

[712] Lang Li, Xihong Lin, Morton B. Brown, Suneel Gupta, and Kyung-Hoon Lee. A population pharmacokinetic model with time-dependent covariates measured with errors. Biometrics, 60(2):451–460, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brumback:2004:SMF

[713] Lyndia C. Brumback and Mary J. Lindstrom. Self modeling with flexible, random time transformations. Biometrics, 60(2):461–470, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schoenberg:2004:TSS

[714] Frederic Paik Schoenberg. Testing separability in spatial-temporal marked point processes. Biometrics, 60(2):471–481, June 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2004:ATS

[715] Yong Lin and Weichung J. Shih. Adaptive two-stage designs for singlearm Phase IIA cancer clinical trials. Biometrics, 60(2):482–490, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kwong:2004:MTE

[716] Koon Shing Kwong, Siu Hung Cheung, and Wai Sum Chan. Multiple testing to establish superiority/equivalence of a new treatment compared with k standard treatments for unbalanced designs. Biometrics, 60(2): 491–498, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fernandez-Duran:2004:CDB

[717] J. J. Fernández-Durán. Circular distributions based on nonnegative trigonometric sums. Biometrics, 60(2):499–503, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OBrien:2004:CSC

[718] Sean M. O’Brien. Cutpoint selection for categorizing a continuous predictor. Biometrics, 60(2):504–509, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stanghellini:2004:MRS

[719] Elena Stanghellini and Peter G. M. van der Heijden. A multiple-record systems estimation method that takes observed and unobserved heterogeneity into account. Biometrics, 60(2):510–516, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krzanowski:2004:BMA

[720] W. J. Krzanowski. Biplots for multifactorial analysis of distance. Biometrics, 60(2):517–524, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gill:2004:SSI

[721] Paramjit S. Gill. Small-sample inference for the comparison of means of log–normal distributions. Biometrics, 60(2):525–527, June 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pepe:2004:APV

[722] Margaret Sullivan Pepe and Tianxi Cai. The analysis of placement values for evaluating discriminatory measures. Biometrics, 60(2):528–535, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Millar:2004:SBE

[723] Russell B. Millar. Sensitivity of Bayes estimators to hyper-parameters with an application to maximum yield from fisheries. Biometrics, 60 (2):536–542, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Larsen:2004:MRT

[724] David R. Larsen and Paul L. Speckman. Multivariate regression trees for analysis of abundance data. Biometrics, 60(2):543–549, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2004:CIR

[725] Man-Lai Tang, Nian-Sheng Tang, and Vincent J. Carey. Confidence interval for rate ratio in a 2 × 2 table with structural zero: an application in assessing false-negative rate ratio when combining two diagnostic tests. Biometrics, 60(2):550–555, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lui:2004:ARF

[726] Kung-Jong Lui. The author replied as follows. Biometrics, 60(2):555, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2004:GRS

[727] You-Gan Wang, Zehua Chen, and Jianbin Liu. General ranked set sampling with cost considerations. Biometrics, 60(2):556–561, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bhattacharya:2004:BR

[728] R. N. Bhattacharya. Book reviews: 1. Biometrics, 60(2):562–563, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Didelez:2004:BRa

[729] V. Didelez. Book reviews: 2. Biometrics, 60(2):563, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donev:2004:BR

[730] A. Donev. Book reviews: 3. Biometrics, 60(2):564, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Homering:2004:BR

[731] M. Homering. Book reviews: 4. Biometrics, 60(2):564–565, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoyle:2004:BR

[732] R. H. Hoyle. Book reviews: 5. Biometrics, 60(2):565–566, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Katz:2004:BR

[733] R. W. Katz. Book reviews: 6. Biometrics, 60(2):566–567, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pohlabeln:2004:BR

[734] H. Pohlabeln. Book reviews: 8. Biometrics, 60(2):567–568, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Osius:2004:BRa

[735] G. Osius. Book reviews: 7. Biometrics, 60(2):567–577, June 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Verbeke:2004:BR

[736] G. Verbeke and K. U. Leuven. Book reviews: 9. Biometrics, 60(2):568– 569, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wild:2004:BR

[737] P. Wild. Book reviews: 10. Biometrics, 60(2):569–570, June 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2004:BREa

[738] Anonymous. Brief report by the Editor. Biometrics, 60(2):570–572, June 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boys:2004:BAD

[739] Richard J. Boys and Daniel A. Henderson. A Bayesian approach to DNA sequence segmentation. Biometrics, 60(3):573–581, September 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [740, 741, 742, 743, 918, 919] and rejoinder [744].

Lawrence:2004:DBA

[740] Charles Lawrence and Ivan Auger. Discussions on “A Bayesian Approach to DNA Sequence Segmentation”. Biometrics, 60(3):581–582, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [739, 744].

Gupta:2004:DBA

[741] Mayetri Gupta and Jun S. Liu. Discussions on “A Bayesian Approach to DNA Sequence Segmentation”. Biometrics, 60(3):582–583, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [739, 744].

Speed:2004:DBA

[742] Terry Speed. Discussions on “A Bayesian Approach to DNA Sequence Segmentation”. Biometrics, 60(3):583–584, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [739, 744].

Thorne:2004:DBA

[743] Jeffrey L. Thorne. Discussions on “A Bayesian Approach to DNA Sequence Segmentation”. Biometrics, 60(3):584–585, September 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [739, 744].

Anonymous:2004:RBA

[744] Anonymous. Rejoinder to “A Bayesian Approach to DNA Sequence Segmentation”. Biometrics, 60(3):585–588, September 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [739, 740, 741, 742, 743].

Satagopan:2004:TSD

[745] Jaya M. Satagopan, E. S. Venkatraman, and Colin B. Begg. Two-stage designs for gene-disease association studies with sample size constraints. Biometrics, 60(3):589–597, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peng:2004:EGL

[746] Yahong Peng, Roderick J. A. Little, and Trivellore E. Raghunathan. An extended general location model for causal inferences from data subject to noncompliance and missing values. Biometrics, 60(3):598–607, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duchateau:2004:PPL

[747] Luc Duchateau and Paul Janssen. Penalized partial likelihood for frailties and smoothing splines in time to first insemination models for dairy cows. Biometrics, 60(3):608–614, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Whitaker:2004:IVC

[748] H. J. Whitaker and C. P. Farrington. Infections with varying contact rates: Application to varicella. Biometrics, 60(3):615–623, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2004:MLA

[749] Sik-Yum Lee and Xin-Yuan Song. Maximum likelihood analysis of a general latent variable model with hierarchically mixed data. Biometrics, 60(3):624–636, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2004:RAD

[750] Liuquan Sun, Yang jin Kim, and Jianguo Sun. Regression analysis of doubly censored failure time data using the additive hazards model. Biometrics, 60(3):637–643, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2004:SMA

[751] Molin Wang, John M. Williamson, and Susan Redline. A semiparametric method for analyzing matched case-control family studies with a continuous outcome and proband sampling. Biometrics, 60(3):644–650, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2004:ECR

[752] Jian-Lun Xu, Richard M. Fagerstrom, Philip C. Prorok, and Barnett S. Kramer. Estimating the cumulative risk of a false-positive test in a repeated screening program. Biometrics, 60(3):651–660, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conaway:2004:DSM

[753] Mark R. Conaway, Stephanie Dunbar, and Shyamal D. Peddada. Designs for single- or multiple-agent Phase I trials. Biometrics, 60(3):661–669, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2004:EGP

[754] You-Gan Wang. Estimation of growth parameters from multiplerecapture data. Biometrics, 60(3):670–675, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2004:BMM

[755] D. B. Dunson, C. Holloman, C. Calder, and L. H. Gunn. Bayesian modeling of multiple lesion onset and growth from interval-censored data. Biometrics, 60(3):676–683, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2004:DFB

[756] Peter F. Thall and John D. Cook. Dose-finding based on efficacytoxicity trade-offs. Biometrics, 60(3):684–693, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wellek:2004:TEC

[757] Stefan Wellek. Tests for establishing compatibility of an observed genotype distribution with Hardy-Weinberg equilibrium in the case of a biallelic locus. Biometrics, 60(3):694–703, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Siannis:2004:APM

[758] Fotios Siannis. Applications of a parametric model for informative censoring. Biometrics, 60(3):704–714, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brannath:2004:OCE

[759] Werner Brannath and Peter Bauer. Optimal conditional error functions for the control of conditional power. Biometrics, 60(3):715–723, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2004:PAM

[760] Ariel Alonso, Geert Molenberghs, Tomasz Burzykowski, Didier Renard, Helena Geys, Ziv Shkedy, Fabián Tibaldi, José Cortiñas Abrahantes, and Marc Buyse. Prentice’s approach and the meta-analytic paradigm: a reflection on the role of statistics in the evaluation of surrogate endpoints. Biometrics, 60(3):724–728, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2004:FMQ

[761] Rongling Wu, Chang-Xing Ma, Min Lin, Zuoheng Wang, and George Casella. Functional mapping of quantitative trait loci underlying growth trajectories using a transform-both-sides logistic model. Biometrics, 60(3):729–738, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OBrien:2004:BML

[762] Sean M. O’Brien and David B. Dunson. Bayesian multivariate logistic regression. Biometrics, 60(3):739–746, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2004:SFM

[763] Lei Liu, Robert A. Wolfe, and Xuelin Huang. Shared frailty models for recurrent events and a terminal event. Biometrics, 60(3):747–756, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marques:2004:PCB

[764] Tiago A. Marques. Predicting and correcting bias caused by measurement error in line transect sampling using multiplicative error models. Biometrics, 60(3):757–763, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aalen:2004:DAM

[765] Odd O. Aalen, Johan Fosen, Harald Weedon-Fekjær, Ørnulf Borgan, and Einar Husebye. Dynamic analysis of multivariate failure time data. Biometrics, 60(3):764–773, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Delongchamp:2004:MTS

[766] Robert R. Delongchamp, John F. Bowyer, James J. Chen, and Ralph L. Kodell. Multiple-testing strategy for analyzing cDNA array data on gene expression. Biometrics, 60(3):774–782, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McNamee:2004:TPS

[767] R. McNamee. Two-phase sampling for simultaneous prevalence estimation and case detection. Biometrics, 60(3):783–792, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greven:2004:PMS

[768] Sonja Greven, A. John Bailer, Lawrence L. Kupper, Keith E. Muller, and Jeremy L. Craft. A parametric model for studying organism fitness using step-stress experiments. Biometrics, 60(3):793–799, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2004:CSS

[769] Qi Jiang, Steven Snapinn, and Boris Iglewicz. Calculation of sample size in survival trials: The impact of informative noncompliance. Biometrics, 60(3):800–806, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zou:2004:CIE

[770] Guangyong Zou and Allan Donner. Confidence interval estimation of the intraclass correlation coefficient for binary outcome data. Biometrics, 60(3):807–811, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sha:2004:BVS

[771] Naijun Sha, Marina Vannucci, Mahlet G. Tadesse, Philip J. Brown, Ilaria Dragoni, Nick Davies, Tracy C. Roberts, Andrea Contestabile, Mike Salmon, Chris Buckley, and Francesco Falciani. Bayesian variable selection in multinomial probit models to identify molecular signatures of disease stage. Biometrics, 60(3):812–819, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2004:MAI

[772] Richard J. Cook, Leilei Zeng, and Grace Y. Yi. Marginal analysis of incomplete longitudinal binary data: a cautionary note on LOCF imputation. Biometrics, 60(3):820–828, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ganju:2004:SUA

[773] Jitendra Ganju. Some unexamined aspects of analysis of covariance in pretest-posttest studies. Biometrics, 60(3):829–833, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holt:2004:LEB

[774] John Holt and Gerarda Darlington. Letter to the Editor of Biometrics. Biometrics, 60(3):834–835, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Indrayan:2004:LEB

[775] Abhaya Indrayan. Letter to the Editor of Biometrics. Biometrics, 60(3): 836–837, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunstan:2004:ARF

[776] Frank Dunstan. The author replied as follows. Biometrics, 60(3):837, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bammann:2004:BR

[777] K. Bammann. Book reviews: 1. Biometrics, 60(3):838–839, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berke:2004:BR

[778] O. Berke. Book reviews: 2. Biometrics, 60(3):839–840, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fahrmeir:2004:BR

[779] L. Fahrmeir. Book reviews: 3. Biometrics, 60(3):840, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fraga:2004:BR

[780] S. Fraga. Book reviews: 4. Biometrics, 60(3):840–841, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hand:2004:BR

[781] D. J. Hand. Book reviews: 5. Biometrics, 60(3):841, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Held:2004:BR

[782] L. Held. Book reviews: 6. Biometrics, 60(3):841–842, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lebreton:2004:BR

[783] J.-D. Lebreton. Book reviews: 7. Biometrics, 60(3):842–843, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2004:BREb

[784] Anonymous. Brief reports by the Editor. Biometrics, 60(3):843–844, September 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2004:VSM

[785] Ariel Alonso, Helena Geys, Geert Molenberghs, Michael G. Kenward, and Tony Vangeneugden. Validation of surrogate markers in multiple randomized clinical trials with repeated measurements: Canonical correlation approach. Biometrics, 60(4):845–853, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hogan:2004:MVC

[786] Joseph W. Hogan, Xihong Lin, and Benjamin Herman. Mixtures of varying coefficient models for longitudinal data with discrete or continuous nonignorable dropout. Biometrics, 60(4):854–864, December 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

MacNab:2004:EBD

[787] Ying C. MacNab, Patrick J. Farrell, Paul Gustafson, and Sijin Wen. Estimation in Bayesian disease mapping. Biometrics, 60(4):865–873, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2004:EMS

[788] Xiao Song and Margaret Sullivan Pepe. Evaluating markers for selecting a Patient’s treatment. Biometrics, 60(4):874–883, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2004:MBM

[789] Zengri Wang and Thomas A. Louis. Marginalized binary mixed-effects models with covariate-dependent random effects and likelihood inference. Biometrics, 60(4):884–891, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ratcliffe:2004:JML

[790] Sarah J. Ratcliffe, Wensheng Guo, and Thomas R. Ten Have. Joint modeling of longitudinal and survival data via a common frailty. Biometrics, 60(4):892–899, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conn:2004:GMA

[791] Paul B. Conn, William L. Kendall, and Michael D. Samuel. A general model for the analysis of mark-resight, mark-recapture, and bandrecovery data under tag loss. Biometrics, 60(4):900–909, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2004:EPE

[792] Yi Cheng and Yu Shen. Estimation of a parameter and its exact confidence interval following sequential sample size reestimation trials. Biometrics, 60(4):910–918, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donner:2004:MSA

[793] Allan Donner, Neil Klar, and Guangyong Zou. Methods for the statistical analysis of binary data in split-cluster designs. Biometrics, 60(4):919– 925, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Herring:2004:MEB

[794] Amy H. Herring, David B. Dunson, and Nancy Dole. Modeling the effects of a bidirectional latent predictor from multivariate questionnaire data. Biometrics, 60(4):926–935, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsu:2004:SEM

[795] Li Hsu, Lu Chen, Malka Gorfine, and Kathleen Malone. Semiparametric estimation of marginal hazard function from case-control family studies. Biometrics, 60(4):936–944, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghidey:2004:SRE

[796] Wendimagegn Ghidey, Emmanuel Lesaffre, and Paul Eilers. Smooth random effects distribution in a linear mixed model. Biometrics, 60(4): 945–953, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reilly:2004:UIC

[797] Cavan Reilly, Phillip Price, Andrew Gelman, and Scott A. Sandgathe. Using image and curve registration for measuring the goodness of fit of spatial and temporal predictions. Biometrics, 60(4):954–964, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsui:2004:ATR

[798] Shigeyuki Matsui. Analysis of times to repeated events in two-arm randomized trials with noncompliance and dependent censoring. Biometrics, 60(4):965–976, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ducharme:2004:STG

[799] Gilles R. Ducharme and Bénédicte Fontez. A smooth test of goodness-of-fit for growth curves and monotonic nonlinear regression models. Biometrics, 60(4):977–986, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greene:2004:MEC

[800] Wendy F. Greene and Jianwen Cai. Measurement error in covariates in the marginal hazards model for multivariate failure time data. Biometrics, 60(4):987–996, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2004:ERA

[801] Zehua Chen and You-Gan Wang. Efficient regression analysis with ranked-set sampling. Biometrics, 60(4):997–1004, December 2004. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2004:CSL

[802] Hyune-Ju Kim, Michael P. Fay, Binbing Yu, Michael J. Barrett, and Eric J. Feuer. Comparability of segmented line regression models. Biometrics, 60(4):1005–1014, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2004:SSP

[803] Jianwen Cai and Donglin Zeng. Sample size/power calculation for casecohort studies. Biometrics, 60(4):1015–1024, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Petzold:2004:SLM

[804] Max Petzold, Christian Sonesson, Eva Bergman, and Helle Kieler. Surveillance in longitudinal models: Detection of intrauterine growth restriction. Biometrics, 60(4):1025–1033, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schmid:2004:BES

[805] Volker Schmid and Leonhard Held. Bayesian extrapolation of space-time trends in cancer registry data. Biometrics, 60(4):1034–1042, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yasui:2004:ECI

[806] Yutaka Yasui, Ziding Feng, Paula Diehr, Dale McLerran, Shirley A. A. Beresford, and Charles E. McCulloch. Evaluation of communityintervention trials via generalized linear mixed models. Biometrics, 60 (4):1043–1052, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2004:EB

[807] Anonymous. To the Editors of Biometrics. Biometrics, 60(4):1053–1054, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Natarajan:2004:ARF

[808] Loki Natarajan and Charles C. Berry. The authors replied as follows. Biometrics, 60(4):1054–1055, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Casella:2004:BR

[809] George Casella. Book reviews: 1. Biometrics, 60(4):1056–1057, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Osius:2004:BRb

[810] Gerhard Osius. Book reviews: 4. Biometrics, 60(4):1056–1057, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Didelez:2004:BRb

[811] Vanessa Didelez. Book reviews: 2. Biometrics, 60(4):1057, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lausen:2004:BR

[812] Berthold Lausen. Book reviews: 3. Biometrics, 60(4):1058, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2004:BREc

[813] Anonymous. Brief reports by the Editor. Biometrics, 60(4):1059–1060, December 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wagenmakers:2004:NNB

[814] Eric-Jan Wagenmakers, Simon Farrell, and Roger Ratcliff. Naïve nonparametric bootstrap model weights are biased. Biometrics, 63(2):281– 283, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [815].

Buckland:2004:REJ

[815] S. T. Buckland, K. P. Burnham, and N. H. Augustin. Rejoinder to the Editors from E.-J. Wagenmakers, S. Farrell, and R. Ratcliff. Biometrics, 63(2):283, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [814].

Julious:2004:LE

[816] Steven Julious. Letter to the Editors. Biometrics, 63(2):284, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zucker:2004:RSJ

[817] David M. Zucker and Jonathan Denne. Rejoinder to steven Julious’s letter to the Editors. Biometrics, 63(2):284–285, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Argac:2004:CV

[818] D. Argac. Components of variance. Biometrics, 63(2):286–287, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edwards:2004:SAG

[819] D. Edwards. Statistical analysis of gene expression microarray data. Biometrics, 63(2):287–289, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huesler:2004:SMP

[820] J. Huesler. Stochastic musings. perspectives from the pioneers of the late 20th century. Biometrics, 63(2):289, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mansmann:2004:BRC

[821] U. Mansmann. Book review: Case Studies in Bayesian Statistics (Volume IV). Biometrics, 63(2):289–290, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mueller:2004:SSM

[822] U. U. Mueller. The stress-strength model and its generalizations. theory and applications. Biometrics, 63(2):290–291, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Newman:2004:MMS

[823] S. C. Newman. Modern medical statistics. a practical guide. Biometrics, 63(2):291, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schill:2004:ALD

[824] W. Schill. Applied longitudinal data analysis for epidemiology: a practical guide. Biometrics, 63(2):291–292, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2004:BREd

[825] Anonymous. Brief report by the Editor. Biometrics, 63(2):292–293, March 2004. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cullis:2005:RCE

[826] Brian Cullis, Laurence Freedman, Mike Kenward, and Xihong Lin. Report of the Co-Editors — 2004. Biometrics, 61(1):vii–ix, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Molenberghs:2005:BBB

[827] Geert Molenberghs. Biometry, biometrics, biostatistics, bioinformatics, …, bio-X. Biometrics, 61(1):1–9, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tseng:2005:TCR

[828] George C. Tseng and Wing H. Wong. Tight clustering: a resamplingbased approach for identifying stable and tight patterns in data. Biometrics, 61(1):10–16, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2005:IPL

[829] Jie Huang and David Harrington. Iterative partial least squares with right-censored data analysis: a comparison to other dimension reduction techniques. Biometrics, 61(1):17–24, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tadesse:2005:WTB

[830] Mahlet G. Tadesse, Joseph G. Ibrahim, Marina Vannucci, and Robert Gentleman. Wavelet thresholding with Bayesian false discovery rate control. Biometrics, 61(1):25–35, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fuentes:2005:MES

[831] Montserrat Fuentes and Adrian E. Raftery. Model evaluation and spatial interpolation by Bayesian combination of observations with outputs from numerical models. Biometrics, 61(1):36–45, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Link:2005:MAA

[832] William A. Link and Richard J. Barker. Modeling association among demographic parameters in analysis of open population capture–recapture data. Biometrics, 61(1):46–54, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2005:GSP

[833] Stuart G. Baker, Paul Lichtenstein, Jaakko Kaprio, and Niels Holm. Genetic susceptibility to prostate, breast, and colorectal cancer among Nordic twins. Biometrics, 61(1):55–63, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1916].

Brown:2005:FBS

[834] Elizabeth R. Brown, Joseph G. Ibrahim, and Victor DeGruttola. A flexible B-spline model for multiple longitudinal biomarkers and survival. Biometrics, 61(1):64–73, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gelman:2005:MIM

[835] Andrew Gelman, Iven Van Mechelen, Geert Verbeke, Daniel F. Heitjan, and Michel Meulders. Multiple imputation for model checking: Completed-data plots with missing and latent data. Biometrics, 61(1): 74–85, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eng:2005:SSF

[836] Kevin Hasegawa Eng and Michael R. Kosorok. A sample size formula for the supremum log-rank statistic. Biometrics, 61(1):86–91, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heagerty:2005:SMP

[837] Patrick J. Heagerty and Yingye Zheng. Survival model predictive accuracy and ROC curves. Biometrics, 61(1):92–105, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:2005:TST

[838] Peter B. Gilbert, A. J. Rossini, and Raj Shankarappa. Two-sample tests for comparing intra-individual genetic sequence diversity between populations. Biometrics, 61(1):106–117, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Atkinson:2005:BAB

[839] Anthony C. Atkinson and Atanu Biswas. Bayesian adaptive biased-coin designs for clinical trials with normal responses. Biometrics, 61(1):118– 125, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunson:2005:BIP

[840] David B. Dunson and Joseph B. Stanford. Bayesian inferences on predictors of conception probabilities. Biometrics, 61(1):126–133, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sutherland:2005:MLM

[841] Jason Sutherland and Carl James Schwarz. Multi-list methods using incomplete lists in closed populations. Biometrics, 61(1):134–140, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhai:2005:MCM

[842] J. Zhai and R. W. Morris. A Markov chain model for animal estrous cycling data. Biometrics, 61(1):141–150, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yin:2005:QRM

[843] Guosheng Yin and Jianwen Cai. Quantile regression models with multivariate failure time data. Biometrics, 61(1):151–161, March 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dueck:2005:REM

[844] Amylou Dueck and Sharon Lohr. Robust estimation of multivariate covariance components. Biometrics, 61(1):162–169, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2005:SEP

[845] Y. Q. Chen, N. P. Jewell, X. Lei, and S. C. Cheng. Semiparametric estimation of proportional mean residual life model in presence of censoring. Biometrics, 61(1):170–178, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Saha:2005:BCM

[846] Krishna Saha and Sudhir Paul. Bias-corrected maximum likelihood estimator of the negative binomial dispersion parameter. Biometrics, 61 (1):179–185, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2005:GCA

[847] Gang Zheng, Boris Freidlin, Zhaohai Li, and Joseph L. Gastwirth. Genomic control for association studies under various genetic models. Biometrics, 61(1):186–192, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Crespi:2005:QMC

[848] Catherine M. Crespi, William G. Cumberland, and Sally Blower. A queueing model for chronic recurrent conditions under panel observation. Biometrics, 61(1):193–198, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hyrien:2005:SMA

[849] Ollivier Hyrien, Margot Mayer-Pröschel, Mark Noble, and Andrei Yakovlev. A stochastic model to analyze clonal data on multi-type cell populations. Biometrics, 61(1):199–207, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yin:2005:CBS

[850] Guosheng Yin and Joseph G. Ibrahim. A class of Bayesian shared gamma frailty models with multivariate failure time data. Biometrics, 61(1):208– 216, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2005:TDD

[851] Kai Wang and Anastasia Ivanova. Two-dimensional dose finding in discrete dose space. Biometrics, 61(1):217–222, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klein:2005:RMC

[852] John P. Klein and Per Kragh Andersen. Regression modeling of competing risks data based on pseudovalues of the cumulative incidence function. Biometrics, 61(1):223–229, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frigessi:2005:BPD

[853] Arnoldo Frigessi, Marit Holden, Clare Marshall, Hildegunn Viljugrein, Nils Chr. Stenseth, Lars Holden, Vladimir Ageyev, and Nikolay L. Klassovskiy. Bayesian population dynamics of interacting species: Great gerbils and fleas in Kazakhstan. Biometrics, 61(1):230–238, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosner:2005:BMC

[854] Gary L. Rosner. Bayesian monitoring of clinical trials with failure-time endpoints. Biometrics, 61(1):239–245, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2005:AEC

[855] Paul R. Rosenbaum. Attributable effects in case^{2}-studies. Biometrics, 61(1):246–253, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2005:CMS

[856] Gang Zheng and Zehua Chen. Comparison of maximum statistics for hypothesis testing when a nuisance parameter is present only under the alternative. Biometrics, 61(1):254–258, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2005:AAM

[857] Yingye Zheng, William E. Barlow, and Gary Cutter. Assessing accuracy of mammography in the presence of verification bias and intrareader correlation. Biometrics, 61(1):259–268, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hellmich:2005:CPM

[858] M. Hellmich and W. Lehmacher. Closure procedures for monotone bifactorial dose–response designs. Biometrics, 61(1):269–276, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nitcheva:2005:MAI

[859] Daniela K. Nitcheva, Walter W. Piegorsch, R. Webster West, and Ralph L. Kodell. Multiplicity-adjusted inferences in risk assessment: Benchmark analysis with quantal response data. Biometrics, 61(1):277– 286, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lyles:2005:EMT

[860] Robert H. Lyles, John M. Williamson, Hung-Mo Lin, and Charles M. Heilig. Extending McNemar’s test: Estimation and inference when paired binary outcome data are misclassified. Biometrics, 61(1):287–294, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vangeneugden:2005:ACG

[861] Tony Vangeneugden, Annouschka Laenen, Helena Geys, Didier Renard, and Geert Molenberghs. Applying concepts of generalizability theory on clinical trial data to investigate sources of variation and their impact on reliability. Biometrics, 61(1):295–304, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hayden:2005:ETC

[862] Douglas Hayden, Donna K. Pauler, and David Schoenfeld. An estimator for treatment comparisons among survivors in randomized trials. Biometrics, 61(1):305–310, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:CK

[863] Anonymous. Correction of Kodlin (1967). Biometrics, 61(1):311, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1].

Arminger:2005:BR

[864] Gerhard Arminger. Book reviews: 1. Biometrics, 61(1):312–313, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bammann:2005:BR

[865] Karin Bammann. Book reviews: 2. Biometrics, 61(1):313–314, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bernardo:2005:BR

[866] José M. Bernardo. Book reviews: 3. Biometrics, 61(1):314, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boehning:2005:BR

[867] Dankmar Boehning. Book reviews: 4. Biometrics, 61(1):314–315, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brueckner:2005:BR

[868] Andreas Brueckner. Book reviews: 5. Biometrics, 61(1):315–316, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kohl:2005:BR

[869] Jochen Kohl and Arndt Von Haeseler. Book reviews: 6. Biometrics, 61(1):316, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Palta:2005:BR

[870] Mari Palta. Book reviews: 7. Biometrics, 61(1):316–317, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peplies:2005:BR

[871] Jenny Peplies. Book reviews: 8. Biometrics, 61(1):317–318, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pohlabeln:2005:BR

[872] Hermann Pohlabeln. Book reviews: 9. Biometrics, 61(1):318–319, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steinberg:2005:BR

[873] David M. Steinberg. Book reviews: 10. Biometrics, 61(1):319–320, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wackernagel:2005:BR

[874] Hans Wackernagel. Book reviews: 11. Biometrics, 61(1):320, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yanagawa:2005:BR

[875] Takashi Yanagawa. Book reviews: 12. Biometrics, 61(1):320–321, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ziegler:2005:BR

[876] Andreas Ziegler. Book reviews: 13. Biometrics, 61(1):321–322, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zimmermann:2005:BR

[877] Karl-Heinz Zimmermann. Book reviews: 14. Biometrics, 61(1):322–323, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREa

[878] Anonymous. Brief reports by the Editor. Biometrics, 61(1):323–324, March 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OMalley:2005:LMT

[879] A. James O’Malley and Sharon-Lise T. Normand. Likelihood methods for treatment noncompliance and subsequent nonresponse in randomized trials. Biometrics, 61(2):325–334, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Braun:2005:DMT

[880] Thomas M. Braun, Zheng Yuan, and Peter F. Thall. Determining a maximum-tolerated schedule of a cytotoxic agent. Biometrics, 61(2): 335–343, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bekele:2005:BAJ

[881] B. Nebiyou Bekele and Yu Shen. A Bayesian approach to jointly modeling toxicity and biomarker expression in a Phase I/II dose-finding trial. Biometrics, 61(2):343–354, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miller:2005:VEC

[882] Frank Miller. Variance estimation in clinical studies with interim sample size reestimation. Biometrics, 61(2):355–361, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yin:2005:ADE

[883] Guosheng Yin and Yu Shen. Adaptive design and estimation in randomized clinical trials with correlated observations. Biometrics, 61(2): 362–369, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

LeBlanc:2005:ARG

[884] Michael LeBlanc, James Moon, and John Crowley. Adaptive risk group refinement. Biometrics, 61(2):370–378, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2005:PCS

[885] Yingye Zheng and Patrick J. Heagerty. Partly conditional survival models for longitudinal data. Biometrics, 61(2):379–391, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2005:NEB

[886] Chiung-Yu Huang and Mei-Cheng Wang. Nonparametric estimation of the bivariate recurrence time distribution. Biometrics, 61(2):392–402, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yin:2005:GCB

[887] Guosheng Yin and Joseph G. Ibrahim. A general class of Bayesian survival models with zero and nonzero cure fractions. Biometrics, 61(2): 403–412, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2005:EVF

[888] You-Gan Wang and Xu Lin. Effects of variance-function misspecification in analysis of longitudinal data. Biometrics, 61(2):413–421, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morton:2005:EBV

[889] Richard Morton and Jordan M. Howarth. Estimation of breeding values using selected pedigree records. Biometrics, 61(2):422–432, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2005:ESA

[890] Chang Xuan Mao, Robert K. Colwell, and Jing Chang. Estimating the species accumulation curve using mixtures. Biometrics, 61(2):433–441, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pradel:2005:MEM

[891] Roger Pradel. Multievent: an extension of multistate capture–recapture models to uncertain states. Biometrics, 61(2):442–447, June 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Calinski:2005:AME

[892] T. Caliński, S. Czajka, Z. Kaczmarek, P. Krajewski, and W. Pilarczyk. Analyzing multi-environment variety trials using randomization-derived mixed models. Biometrics, 61(2):448–455, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2005:STC

[893] Jing Qin and Denis H. Y. Leung. A semiparametric two-component “compound” mixture model and its application to estimating malaria attributable fractions. Biometrics, 61(2):456–464, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2005:LAQ

[894] K. Wang. A likelihood approach for quantitative-trait-locus mapping with selected pedigrees. Biometrics, 61(2):465–473, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2005:FEA

[895] Zehua Chen. The full EM algorithm for the MLEs of QTL effects and positions and their estimated variances in multiple-interval mapping. Biometrics, 61(2):474–480, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goncalves:2005:SAP

[896] Luzia Gonçalves, Líbia Zé-Zé, Hildete Prisco Pinheiro, and Maria Antónia Amaral-Turkman. Statistical aspects in physical mapping application to the genome of o. oeni strain GM. Biometrics, 61(2):481–487, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tadesse:2005:BEV

[897] Mahlet G. Tadesse, Joseph G. Ibrahim, Robert Gentleman, Sabina Chiaretti, Jerome Ritz, and Robin Foa. Bayesian error-in-variable survival model for the analysis of GeneChip arrays. Biometrics, 61(2):488– 497, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2005:IVS

[898] Xiaowei Yang, Thomas R. Belin, and W. John Boscardin. Imputation and variable selection in linear regression models with missing covariates. Biometrics, 61(2):498–506, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cantoni:2005:VSM

[899] Eva Cantoni, Joanna Mills Flemming, and Elvezio Ronchetti. Variable selection for marginal longitudinal generalized linear models. Biometrics, 61(2):507–514, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:2005:FPB

[900] Alan Agresti and Yongyi Min. Frequentist performance of Bayesian confidence intervals for comparing proportions in 2 × 2 contingency tables. Biometrics, 61(2):515–523, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2005:ETS

[901] Ying Kuen Cheung. Exact two-sample inference with missing data. Biometrics, 61(2):524–531, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2005:AOT

[902] Peng Huang, Barbara C. Tilley, Robert F. Woolson, and Stuart Lipsitz. Adjusting O’Brien’s test to control Type I error for the generalized nonparametric Behrens–Fisher problem. Biometrics, 61(2):532–539, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ivanova:2005:CTM

[903] Anastasia Ivanova, Bahjat F. Qaqish, and Michael J. Schell. Continuous toxicity monitoring in Phase II trials in oncology. Biometrics, 61(2):540– 545, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gillen:2005:NVU

[904] Daniel L. Gillen and Scott S. Emerson. A note on P -values under group sequential testing and nonproportional hazards. Biometrics, 61(2):546– 551, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yin:2005:BCR

[905] Guosheng Yin. Bayesian cure rate frailty models with application to a root canal therapy study. Biometrics, 61(2):552–558, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsuura:2005:MLE

[906] Masaaki Matsuura and Shinto Eguchi. Modeling late entry bias in survival analysis. Biometrics, 61(2):559–566, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jiang:2005:SAS

[907] Hongyu Jiang, Jason P. Fine, and Rick Chappell. Semiparametric analysis of survival data with left truncation and dependent right censoring. Biometrics, 61(2):567–575, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nan:2005:VCC

[908] Bin Nan, Xihong Lin, Lynda D. Lisabeth, and Siobán D. Harlow. A varying-coefficient Cox model for the effect of age at a marker event on age at menopause. Biometrics, 61(2):576–583, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pfeiffer:2005:SCC

[909] Ruth M. Pfeiffer and Nilanjan Chatterjee. On a supplemented casecontrol design. Biometrics, 61(2):584–590, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2005:EBM

[910] Jing X. Zhao, Andrea S. Foulkes, and Edward I. George. Exploratory Bayesian model selection for serial genetics data. Biometrics, 61(2):591– 599, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2005:NER

[911] Xiao-Hua Zhou, Pete Castelluccio, and Chuan Zhou. Nonparametric estimation of ROC curves in the absence of a gold standard. Biometrics, 61(2):600–609, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clough:2005:QUA

[912] H. E. Clough, D. Clancy, P. D. O’Neill, S. E. Robinson, and N. P. French. Quantifying uncertainty associated with microbial count data: a Bayesian approach. Biometrics, 61(2):610–616, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Banerjee:2005:GDC

[913] Sudipto Banerjee. On geodetic distance computations in spatial modeling. Biometrics, 61(2):617–625, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jung:2005:BTO

[914] Byoung Cheol Jung, Myoungshic Jhun, and Jae Won Lee. Bootstrap tests for overdispersion in a zero-inflated Poisson regression model. Biometrics, 61(2):626–628, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2005:RBT

[915] Martin Ridout, John P. Hinde, and Clarice G. B. Demétrio. Rejoinder to “Bootstrap Tests for Overdispersion in a Zero-Inflated Poisson Regression Model”. Biometrics, 61(2):628–629, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pittelkow:2005:UPC

[916] Yvonne Pittelkow and Susan R. Wilson. Use of principal component analysis and the GE-biplot for the graphical exploration of gene expression data. Biometrics, 61(2):630–632, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See rejoinder [917].

Wouters:2005:RUP

[917] Luc Wouters, Hinrich W. Göhlmann, Luc Bijnens, Stefan U. Kass, Geert Molenberghs, and Paul J. Lewi. Rejoinder to use of principal component analysis and the GE-biplot for the graphical exploration of gene expression data. Biometrics, 61(2):632–634, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [916].

Booth:2005:DBA

[918] Hilary S. Booth, Conrad J. Burden, John H. Maindonald, Lucia Santoso, Matthew J. Wakefield, and Susan R. Wilson. Discussion of “A Bayesian Approach to DNA Sequence Segmentation”. Biometrics, 61(2):635–637, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [739, 744].

Boys:2005:DBA

[919] R. J. Boys and D. A. Henderson. Discussion of “A Bayesian Approach to DNA Sequence Segmentation”. Biometrics, 61(2):637–639, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [739, 744].

Vines:2005:BR

[920] S. Karen Vines. Book reviews: 1. Biometrics, 61(2):640, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Timmer:2005:BR

[921] Antje Timmer. Book reviews: 2. Biometrics, 61(2):640–641, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edwards:2005:BR

[922] David George Edwards. Book reviews: 3. Biometrics, 61(2):641–643, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Singer:2005:BR

[923] Judith D. Singer. Book reviews: 4. Biometrics, 61(2):643, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Didelez:2005:BR

[924] Vanessa Didelez. Book reviews: 5. Biometrics, 61(2):643–644, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baltas:2005:BR

[925] George Baltas. Book reviews: 6. Biometrics, 61(2):644–645, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREb

[926] Anonymous. Brief reports by the Editor. Biometrics, 61(2):645–646, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:RAa

[927] Anonymous. Regular articles. Biometrics, 61(2):647–652, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:AT

[928] Anonymous. Abstract translations. Biometrics, 61(2):647–655, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:CFa

[929] Anonymous. Consultant’s forum. Biometrics, 61(2):652–655, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:RRa

[930] Anonymous. Reader reaction. Biometrics, 61(2):655, June 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cowen:2005:CRS

[931] Laura Cowen and Carl J. Schwarz. Capture–recapture studies using radio telemetry with premature radio-tag failure. Biometrics, 61(3):657–664, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Suchard:2005:MEB

[932] Marc A. Suchard, Robert E. Weiss, and Janet S. Sinsheimer. Models for estimating Bayes factors with applications to phylogeny and tests of monophyly. Biometrics, 61(3):665–673, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2005:GLL

[933] J. Zhu, J. C. Eickhoff, and P. Yan. Generalized linear latent variable models for repeated measures of spatially correlated multivariate data. Biometrics, 61(3):674–683, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2005:REF

[934] You-Gan Wang, Xu Lin, and Min Zhu. Robust estimating functions and bias correction for longitudinal data analysis. Biometrics, 61(3):684–691, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comments [1919] and rejoinder [1920].

Cook:2005:RTT

[935] Richard J. Cook, Wei Wei, and Grace Y. Yi. Robust tests for treatment effects based on censored recurrent event data observed over multiple periods. Biometrics, 61(3):692–701, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2005:CSA

[936] Xiao Song and Yijian Huang. On corrected score approach for proportional hazards model with covariate measurement error. Biometrics, 61(3):702–714, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andrei:2005:SGS

[937] Adin-Cristian Andrei and Susan Murray. Simultaneous group sequential analysis of rank-based and weighted Kaplan–Meier tests for paired censored survival data. Biometrics, 61(3):715–720, September 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2005:PSM

[938] Bo Lu. Propensity score matching with time-dependent covariates. Biometrics, 61(3):721–728, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Naskar:2005:SMM

[939] Malay Naskar, Kalyan Das, and Joseph G. Ibrahim. A semiparametric mixture model for analyzing clustered competing risks data. Biometrics, 61(3):729–737, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bretz:2005:CMC

[940] F. Bretz, J. C. Pinheiro, and M. Branson. Combining multiple comparisons and modeling techniques in dose-response studies. Biometrics, 61(3):738–748, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2005:RAS

[941] John O’Quigley. Retrospective analysis of sequential dose-finding designs. Biometrics, 61(3):749–756, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yu:2005:QRA

[942] Zi-Fan Yu and Paul J. Catalano. Quantitative risk assessment for multivariate continuous outcomes with application to neurotoxicology: The bivariate case. Biometrics, 61(3):757–766, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2005:BAG

[943] Lan Huang, Ming-Hui Chen, and Joseph G. Ibrahim. Bayesian analysis for generalized linear models with nonignorably missing covariates. Biometrics, 61(3):767–780, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Golightly:2005:BIS

[944] A. Golightly and D. J. Wilkinson. Bayesian inference for stochastic kinetic models using a diffusion approximation. Biometrics, 61(3):781–788, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diao:2005:SMM

[945] Guoqing Diao and D. Y. Lin. Semiparametric methods for mapping quantitative trait loci with censored data. Biometrics, 61(3):789–798, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Briggs:2005:ASY

[946] William Briggs and David Ruppert. Assessing the skill of yes/no predictions. Biometrics, 61(3):799–807, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thas:2005:STZ

[947] Olivier Thas and J. C. W. Rayner. Smooth tests for the zero-inflated Poisson distribution. Biometrics, 61(3):808–815, September 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsui:2005:SAR

[948] Shigeyuki Matsui. Stratified analysis in randomized trials with noncompliance. Biometrics, 61(3):816–823, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2005:LME

[949] Liang Li, Jun Shao, and Mari Palta. A longitudinal measurement error model with a semicontinuous covariate. Biometrics, 61(3):824–830, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buonaccorsi:2005:EMB

[950] John P. Buonaccorsi, Petter Laake, and Marit B. Veierød. On the effect of misclassification on bias of perfectly measured covariates in regression. Biometrics, 61(3):831–836, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:2005:MCL

[951] Jason Roy and Xihong Lin. Missing covariates in longitudinal data with informative dropouts: Bias analysis and inference. Biometrics, 61(3): 837–846, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graubard:2005:SEA

[952] Barry I. Graubard and Thomas R. Fears. Standard errors for attributable risk for simple and complex sample designs. Biometrics, 61(3):847–855, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fewster:2005:LTS

[953] Rachel M. Fewster, Jeffrey L. Laake, and Stephen T. Buckland. Line transect sampling in small and large regions. Biometrics, 61(3):856–859, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Melville:2005:RLT

[954] G. J. Melville and A. H. Welsh. Rejoinder to “Line Transect Sampling in Small and Large Regions”. Biometrics, 61(3):859–861, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gueorguieva:2005:CAJ

[955] Ralitza V. Gueorguieva. Comments about joint modeling of cluster size and binary and continuous subunit-specific outcomes. Biometrics, 61(3): 862–866, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [569] and rejoinder [956].

Chen:2005:RCA

[956] Zhen Chen and David B. Dunson. Rejoinder to “Comments about Joint Modeling of Cluster Size and Binary and Continuous Subunit-Specific Outcomes”. Biometrics, 61(3):866–867, September 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [569, 955].

Pledger:2005:PMM

[957] Shirley Pledger. The performance of mixture models in heterogeneous closed population capture–recapture. Biometrics, 61(3):868–873, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dorazio:2005:RPM

[958] Robert M. Dorazio and J. Andrew Royle. Rejoinder to “The Performance of Mixture Models in Heterogeneous Closed Population Capture– Recapture”. Biometrics, 61(3):874–876, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2005:BAB

[959] Jing Cao and Chong Z. He. Bias adjustment in Bayesian estimation of bird nest age-specific survival rates. Biometrics, 61(3):877–878, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2005:LE

[960] Paul S. Albert. Letter to the Editor. Biometrics, 61(3):879–880, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:2005:RLE

[961] Weichung Joe Shih, Shou-En Lu, and Yong Lin. Rejoinder to the letter to the Editor from P. S. Albert. Biometrics, 61(3):880–881, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maller:2005:BR

[962] Ross A. Maller. Book reviews: 1. Biometrics, 61(3):882–883, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gentle:2005:BR

[963] James E. Gentle. Book reviews: 2. Biometrics, 61(3):883–884, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hand:2005:BR

[964] David J. Hand. Book reviews: 3. Biometrics, 61(3):884, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Abrams:2005:BR

[965] Peter A. Abrams. Book reviews: 4. Biometrics, 61(3):884–885, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:2005:BR

[966] Richard O. Gilbert. Book reviews: 5. Biometrics, 61(3):885–886, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gurrin:2005:BR

[967] Lyle Gurrin. Book reviews: 6. Biometrics, 61(3):886–887, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shoukri:2005:BR

[968] Mohamed M. Shoukri. Book reviews: 7. Biometrics, 61(3):887–888, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Claustriaux:2005:BR

[969] Jean-Jacques Claustriaux. Book reviews: 8. Biometrics, 61(3):888, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schill:2005:BR

[970] Walter Schill. Book reviews: 9. Biometrics, 61(3):888, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREc

[971] Anonymous. Brief reports by the Editor. Biometrics, 61(3):888–889, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BRa

[972] Anonymous. Book reviews: 11. Biometrics, 61(3):889, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BRb

[973] Anonymous. Book reviews: 12. Biometrics, 61(3):889, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BRc

[974] Anonymous. Book reviews: 13. Biometrics, 61(3):890, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:RAb

[975] Anonymous. Regular articles. Biometrics, 61(3):891–895, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:CFb

[976] Anonymous. Consultant’s forum. Biometrics, 61(3):895–896, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:RRb

[977] Anonymous. Reader reaction. Biometrics, 61(3):896–897, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:C

[978] Anonymous. Correspondence. Biometrics, 61(3):897, September 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prentice:2005:SIA

[979] Ross L. Prentice, Mary Pettinger, and Garnet L. Anderson. Statistical issues arising in the Women’s Health Initiative. Biometrics, 61(4):899– 911, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [980, 981, 982, 983, 984, 985, 986, 987].

Carroll:2005:DSI

[980] Raymond J. Carroll. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4):911–912, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [979].

Day:2005:DSI

[981] N. E. Day. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4):912–914, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [979].

DeMets:2005:DSI

[982] David L. DeMets. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4):914–918, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [979].

Freedman:2005:DSI

[983] David A. Freedman and Diana B. Petitti. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4): 918–920, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [979].

Greenland:2005:DSI

[984] Sander Greenland. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4):920–921, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [979].

Hernan:2005:DSI

[985] Miguel A. Hernán, James M. Robins, and Luis A. García Rodríguez. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4):922–930, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [979].

Thomas:2005:DSI

[986] Duncan C. Thomas. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4):930–933, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See

Tsiatis:2005:DSI

[987] Anastasios A. Tsiatis and Marie Davidian. Discussion on “Statistical Issues Arising in the Women’s Health Initiative”. Biometrics, 61(4): 933–935, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [979].

Prentice:2005:R

[988] Ross L. Prentice, Mary Pettinger, and Garnet L. Anderson. Rejoinder. Biometrics, 61(4):935–941, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeLong:2005:ETW

[989] Elizabeth R. DeLong, Laura P. Coombs, T. Bruce Ferguson Jr., and Eric D. Peterson. The evaluation of treatment when center-specific selection criteria vary with respect to patient risk. Biometrics, 61(4):942–949, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jin:2005:GHM

[990] Xiaoping Jin, Bradley P. Carlin, and Sudipto Banerjee. Generalized hierarchical multivariate CAR models for areal data. Biometrics, 61(4): 950–961, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bang:2005:DRE

[991] Heejung Bang and James M. Robins. Doubly robust estimation in missing data and causal inference models. Biometrics, 61(4):962–973, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1571].

Wang:2005:GMH

[992] Molin Wang and John M. Williamson. Generalization of the Mantel– Haenszel estimating function for sparse clustered binary data. Biometrics, 61(4):973–981, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pfeiffer:2005:CCD

[993] Ruth M. Pfeiffer, Louise Ryan, Augusto Litonjua, and David Pee. A case-cohort design for assessing covariate effects in longitudinal studies. Biometrics, 61(4):982–991, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shen:2005:NEA

[994] Yu Shen and Xuelin Huang. Nonparametric estimation of asymptomatic duration from a randomized prospective cancer screening trial. Biometrics, 61(4):992–999, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2005:GFM

[995] Zhiying Pan and D. Y. Lin. Goodness-of-fit methods for generalized linear mixed models. Biometrics, 61(4):1000–1009, December 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2005:MAB

[996] Hsin-Chou Yang and Anne Chao. Modeling animals’ behavioral response by Markov chain models for capture–recapture experiments. Biometrics, 61(4):1010–1017, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hans:2005:BIU

[997] Chris Hans and David B. Dunson. Bayesian inferences on umbrella orderings. Biometrics, 61(4):1018–1026, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoff:2005:SCB

[998] Peter D. Hoff. Subset clustering of binary sequences, with an application to genomic abnormality data. Biometrics, 61(4):1027–1036, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harezlak:2005:IPP

[999] Jaroslaw Harezlak, Louise M. Ryan, Jay N. Giedd, and Nicholas Lange. Individual and population penalized regression splines for accelerated longitudinal designs. Biometrics, 61(4):1037–1048, December 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Larsen:2005:CPH

[1000] Klaus Larsen. The Cox proportional hazards model with a continuous latent variable measured by multiple binary indicators. Biometrics, 61 (4):1049–1055, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2005:MBI

[1001] Dongfeng Wu, Gary L. Rosner, and Lyle Broemeling. MLE and Bayesian inference of age-dependent sensitivity and transition probability in periodic screening. Biometrics, 61(4):1056–1063, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:2005:TVF

[1002] Hans-Georg Müller and Ying Zhang. Time-varying functional regression for predicting remaining lifetime distributions from longitudinal trajectories. Biometrics, 61(4):1064–1075, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pang:2005:SRM

[1003] Zhen Pang and Anthony Y. C. Kuk. A shared response model for clustered binary data in developmental toxicity studies. Biometrics, 61(4): 1076–1084, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yip:2005:SME

[1004] Paul S. F. Yip, Hua-Zhen Lin, and Liqun Xi. A semiparametric method for estimating population size for capture–recapture experiments with random covariates in continuous time. Biometrics, 61(4):1085–1092, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dorazio:2005:IRB

[1005] Robert M. Dorazio, Howard L. Jelks, and Frank Jordan. Improving removal-based estimates of abundance by sampling a population of spatially distinct subpopulations. Biometrics, 61(4):1093–1101, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:2005:CLP

[1006] Jeremy M. G. Taylor, Yue Wang, and Rodolphe Thiébaut. Counterfactual links to the proportion of treatment effect explained by a surrogate marker. Biometrics, 61(4):1102–1111, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Faddy:2005:MDB

[1007] M. J. Faddy and D. M. Smith. Modeling the dependence between the number of trials and the success probability in binary trials. Biometrics, 61(4):1112–1114, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2005:ACR

[1008] Paul S. Albert and Sally Hunsberger. On analyzing circadian rhythms data using nonlinear mixed models with harmonic terms. Biometrics, 61(4):1115–1120, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2005:RAC

[1009] Yuedong Wang, Chunlei Ke, and Morton B. Brown. Rejoinder to “On Analyzing Circadian Rhythms Data Using Nonlinear Mixed Models with Harmonic Terms”. Biometrics, 61(4):1120–1122, December 2005. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Novikov:2005:MCP

[1010] Ilya Novikov, Bernice Oberman, and Laurence Freedman. Modification of the computational procedure in Parker and Bregman’s method of calculating sample size from matched case-control studies with a dichotomous exposure. Biometrics, 61(4):1123–1126, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1012].

Parker:2005:RMC

[1011] Robert A. Parker. Rejoinder to “Modification of the Computational Procedure in Parker and Bregman’s Method of Calculating Sample Size from Matched Case-Control Studies with a Dichotomous Exposure”. Biometrics, 61(4):1126–1127, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:CMC

[1012] Anonymous. Correction to “Modification of the computational procedure in Parker and Bregman’s method of calculating sample size from matched case-control studies with a dichotomous exposure” by I. Novikov, B. Oberman, and L. Freedman; 61, 1123–1126, December 2005. Biometrics, 61(4):1128, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1010].

Vanclay:2005:BR

[1013] Jerry K. Vanclay. Book reviews: 1. Biometrics, 61(4):1129–1130, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Al-Saleh:2005:BR

[1014] Mohammad Fraiwan Al-Saleh. Book reviews: 2. Biometrics, 61(4):1130, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roehmel:2005:BR

[1015] Joachim Roehmel. Book reviews: 3. Biometrics, 61(4):1130–1131, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dey:2005:BR

[1016] Aloke Dey. Book reviews: 4. Biometrics, 61(4):1131–1132, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Woolson:2005:BR

[1017] Robert F. Woolson. Book reviews: 5. Biometrics, 61(4):1132, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cody:2005:BR

[1018] Ronald Cody. Book reviews: 6. Biometrics, 61(4):1132–1133, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berke:2005:BR

[1019] Olaf Berke. Book reviews: 7. Biometrics, 61(4):1133–1134, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hayter:2005:BR

[1020] Anthony Hayter. Book reviews: 8. Biometrics, 61(4):1134–1135, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beirlant:2005:BR

[1021] Jan Beirlant. Book reviews: 9. Biometrics, 61(4):1135, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morton:2005:BR

[1022] Richard Morton. Book reviews: 10. Biometrics, 61(4):1135–1136, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hanley:2005:BR

[1023] James A. Hanley. Book reviews: 11. Biometrics, 61(4):1136–1137, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Foraita:2005:BR

[1024] Ronja Foraita. Book reviews: 12. Biometrics, 61(4):1137, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsu:2005:BR

[1025] John S. J. Hsu. Book reviews: 13. Biometrics, 61(4):1137–1138, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREd

[1026] Anonymous. Brief reports by the Editor: 1. Biometrics, 61(4):1138–1139, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREe

[1027] Anonymous. Brief reports by the Editor: 2. Biometrics, 61(4):1139, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREf

[1028] Anonymous. Brief reports by the Editor: 3. Biometrics, 61(4):1139, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREg

[1029] Anonymous. Brief reports by the Editor: 4. Biometrics, 61(4):1139–1140, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREh

[1030] Anonymous. Brief reports by the Editor: 5. Biometrics, 61(4):1140, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREi

[1031] Anonymous. Brief reports by the Editor: 6. Biometrics, 61(4):1140, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:BREj

[1032] Anonymous. Brief reports by the Editor: 7. Biometrics, 61(4):1140, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:DP

[1033] Anonymous. Discussion paper. Biometrics, 61(4):1142, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:RAc

[1034] Anonymous. Regular articles. Biometrics, 61(4):1142–1146, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2005:RRc

[1035] Anonymous. Reader reaction. Biometrics, 61(4):1146–1147, December 2005. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gottardo:2006:BRI

[1036] Raphael Gottardo, Adrian E. Raftery, Ka Yee Yeung, and Roger E. Bumgarner. Bayesian robust inference for differential gene expression in microarrays with multiple samples. Biometrics, 62(1):10–18, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lewin:2006:BMD

[1037] Alex Lewin, Sylvia Richardson, Clare Marshall, Anne Glazier, and Tim Aitman. Bayesian modeling of differential gene expression. Biometrics, 62(1):10–18, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kendziorski:2006:SME

[1038] C. M. Kendziorski, M. Chen, M. Yuan, H. Lan, and A. D. Attie. Statistical methods for expression quantitative trait loci (eQTL) mapping. Biometrics, 62(1):19–27, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2006:HBA

[1039] Jinbo Chen and Nilanjan Chatterjee. Haplotype-based association analysis in cohort and nested case-control studies. Biometrics, 62(1):28–35, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chatterjee:2006:CCC

[1040] Nilanjan Chatterjee, Zeynep Kalaylioglu, Joanna H. Shih, and Mitchell H. Gail. Case-control and case-only designs with genotype and family history data: Estimating relative risk, residual familial aggregation, and cumulative risk. Biometrics, 62(1):36–48, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huber:2006:MCA

[1041] Mark Huber, Yuguo Chen, Ian Dinwoodie, Adrian Dobra, and Mike Nicholas. Monte Carlo algorithms for Hardy-Weinberg proportions. Biometrics, 62(1):49–53, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bagiella:2006:ECD

[1042] Emilia Bagiella. Estimating and comparing the distribution of onset of disease from prevalence case-control data. Biometrics, 62(1):54–60, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brookmeyer:2006:HTE

[1043] Ron Brookmeyer and Xiaojun You. A hypothesis test for the end of a common source outbreak. Biometrics, 62(1):61–65, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bertail:2006:IGS

[1044] Patrice Bertail and Jessica Tressou. Incomplete generalized U-statistics for food risk assessment. Biometrics, 62(1):66–74, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carroll:2006:SUM

[1045] Raymond J. Carroll, Douglas Midthune, Laurence S. Freedman, and Victor Kipnis. Seemingly unrelated measurement error models, with application to nutritional epidemiology. Biometrics, 62(1):75–84, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kuchenhoff:2006:GMD

[1046] Helmut Küchenhoff, Samuel M. Mwalili, and Emmanuel Lesaffre. A general method for dealing with misclassification in regression: The misclassification SIMEX. Biometrics, 62(1):85–96, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Royle:2006:SOM

[1047] J. Andrew Royle. Site occupancy models with heterogeneous detection probabilities. Biometrics, 62(1):97–102, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:2006:CRS

[1048] E. R. Williams, J. A. John, and D. Whitaker. Construction of resolvable spatial row-column designs. Biometrics, 62(1):103–108, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kneib:2006:SAR

[1049] Thomas Kneib and Ludwig Fahrmeir. Structured additive regression for categorical space-time data: a mixed model approach. Biometrics, 62 (1):109–118, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guan:2006:AIS

[1050] Yongtao Guan, Michael Sherman, and James A. Calvin. Assessing isotropy for spatial point processes. Biometrics, 62(1):119–125, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guan:2006:TIB

[1051] Yongtao Guan. Tests for independence between marks and points of a marked point process. Biometrics, 62(1):126–134, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haas:2006:ENS

[1052] Peter J. Haas, Yushan Liu, and Lynne Stokes. An estimator of number of species from quadrat sampling. Biometrics, 62(1):135–141, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonner:2006:ECJ

[1053] S. J. Bonner and C. J. Schwarz. An extension of the Cormack–Jolly– Seber model for continuous covariates with application to Microtus pennsylvanicus. Biometrics, 62(1):142–149, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2006:URS

[1054] Haiying Chen, Elizabeth A. Stasny, and Douglas A. Wolfe. Unbalanced ranked set sampling for estimating a population proportion. Biometrics, 62(1):150–158, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lai:2006:FMH

[1055] Tze Leung Lai, Mei-Chiung Shih, and Samuel Po-Shing Wong. Flexible modeling via a hybrid estimation scheme in generalized mixed models for longitudinal data. Biometrics, 62(1):159–167, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilkins:2006:HMN

[1056] Kenneth J. Wilkins and Garrett M. Fitzmaurice. A hybrid model for nonignorable dropout in longitudinal binary responses. Biometrics, 62 (1):168–176, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2006:SMM

[1057] Qingxia Chen and Joseph G. Ibrahim. Semiparametric models for missing covariate and response data in regression models. Biometrics, 62(1):177– 184, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosner:2006:WSR

[1058] Bernard Rosner, Robert J. Glynn, and Mei-Ling T. Lee. The Wilcoxon signed rank test for paired comparisons of clustered data. Biometrics, 62 (1):185–192, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2006:GAC

[1059] Peter F. Thall, Leiko H. Wooten, and Elizabeth J. Shpall. A geometric approach to comparing treatments for rapidly fatal diseases. Biometrics, 62(1):193–201, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2006:ARM

[1060] Shuangge Ma, Michael R. Kosorok, and Jason P. Fine. Additive risk models for survival data with high-dimensional covariates. Biometrics, 62(1):202–210, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2006:FIS

[1061] R. King, S. P. Brooks, B. J. T. Morgan, and T. Coulson. Factors influencing Soay sheep survival: a Bayesian analysis. Biometrics, 62(1):211–220, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pepe:2006:CPC

[1062] Margaret Sullivan Pepe, Tianxi Cai, and Gary Longton. Combining predictors for classification using the area under the receiver operating characteristic curve. Biometrics, 62(1):221–229, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2006:DPH

[1063] Yu-Chieh Yang, Anna Liu, and Yuedong Wang. Detecting pulsatile hormone secretions using nonlinear mixed effects partial spline models. Biometrics, 62(1):230–238, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krzanowski:2006:SMS

[1064] W. J. Krzanowski. Sensitivity in metric scaling and analysis of distance. Biometrics, 62(1):239–244, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anderson:2006:DBT

[1065] Marti J. Anderson. Distance-based tests for homogeneity of multivariate dispersions. Biometrics, 62(1):245–253, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jacqmin-Gadda:2006:RCM

[1066] Hélène Jacqmin-Gadda, Daniel Commenges, and Jean-François Dartigues. Random changepoint model for joint modeling of cognitive decline and dementia. Biometrics, 62(1):254–260, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2006:MER

[1067] Li C. Liu and Donald Hedeker. A mixed-effects regression model for longitudinal multivariate ordinal data. Biometrics, 62(1):261–268, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Congdon:2006:MFM

[1068] Peter Congdon. A model framework for mortality and health data classified by age, area, and time. Biometrics, 62(1):269–278, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2006:ATD

[1069] Yingye Zheng, Tianxi Cai, and Ziding Feng. Application of the timedependent ROC curves for prognostic accuracy with multiple biomarkers.

Biometrics, 62(1):279–287, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choudhary:2006:AAU

[1070] Pankaj K. Choudhary and Hon Keung Tony Ng. Assessment of agreement under nonstandard conditions using regression models for mean and variance. Biometrics, 62(1):288–296, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2006:QPC

[1071] John O’Quigley and Philippe Flandre. Quantification of the Prentice criteria for surrogate endpoints. Biometrics, 62(1):297–300, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Metcalfe:2006:LEB

[1072] Chris Metcalfe. Letter to the Editor of Biometrics. Biometrics, 62(1): 301, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schill:2006:BRC

[1073] Walter Schill. Book review: Categorical Data Analysis Using the SAS System, by M. E. Stokes, C. S. Davis and G. G. Koch. Biometrics, 62 (1):302–303, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Behrens:2006:BRH

[1074] Thomas Behrens. Book review: A History of Epidemiologic Methods and Concepts, edited by A. Morabia. Biometrics, 62(1):303, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

May:2006:BRA

[1075] Susanne May. Book review: Advanced Excel(R) for Scientific Data Analysis, by R. De Levie. Biometrics, 62(1):303–304, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2006:BRM

[1076] Annie Qu. Book review: Mixed Models. Theory and Applications, by E. Demidenko. Biometrics, 62(1):304–305, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burger:2006:BRG

[1077] Reinhard Bürger. Book review: Genetic Structure and Selection in Subdivided Populations, by F. Rousset. Biometrics, 62(1):305–306, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ades:2006:BRB

[1078] Tony Ades. Book review: Bayesian Approaches to Clinical Trials and Health-Care Evaluation, by D. J. Spiegelhalter, K. R. Abrams and J. P. Miles. Biometrics, 62(1):306–307, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buzas:2006:BRM

[1079] Jeffrey S. Buzas. Book review: Measurement Error and Misclassification in Statistics and Epidemiology: Impacts and Bayesian Adjustments, by P. Gustafson. Biometrics, 62(1):307–308, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berke:2006:BRP

[1080] Olaf Berke. Book review: A Primer of Ecological Statistics, by H. J. Gotelli and A. M. Ellison. Biometrics, 62(1):308, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dahlhaus:2006:BRD

[1081] Rainer Dahlhaus. Book review: Diagnostic Checks in Time Series, by W. K. Li. Biometrics, 62(1):308–309, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Duembgen:2006:BRE

[1082] Lutz Duembgen. Book review: Exploring Multivariate Data with the Forward Search, by A. Atkinson, M. Riani and A. Cerioli. Biometrics, 62(1):309, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kotz:2006:BRH

[1083] Samuel Kotz. Book review: Handbook of Beta Distribution and Its Applications, edited by A. K. Gupta and S. Nadarajah. Biometrics, 62 (1):309–310, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BREa

[1084] Anonymous. Brief reports by the Editor. Biometrics, 62(1):310, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRU

[1085] Anonymous. Book reviews: Using R for Introductory Statistics, by J. Verzani and An R and S-Plus Companion to Multivariate Analysis by B. S. Everitt. Biometrics, 62(1):311, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRF

[1086] Anonymous. Book review: Francis Galton. Pioneer of Heredity and Biometry, by M. Bulmer. Biometrics, 62(1):311, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRSa

[1087] Anonymous. Book review: Statistics without Math, by W. E. Magnusson and G. Mourao. Biometrics, 62(1):311–312, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRD

[1088] Anonymous. Book review: Data Mining-Next Generation Challenges and Future Directions, edited by H. Kargupta, A. Joshi, K. Sivakumar and Y. Yesha. Biometrics, 62(1):312, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BREb

[1089] Anonymous. Book review: Essentials of Statistical Inference, by G. A. Young and R. L. Smith. Biometrics, 62(1):312, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRSb

[1090] Anonymous. Book review: Survival Analysis: a Self-Learning Text, by D. G. Kleinbaum and M. Klein. Biometrics, 62(1):312, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:RAa

[1091] Anonymous. Regular articles. Biometrics, 62(1):313–320, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:CFa

[1092] Anonymous. Consultant’s forum. Biometrics, 62(1):320–321, March 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:RRa

[1093] Anonymous. Reader reaction. Biometrics, 62(1):321, March 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilmour:2006:RSD

[1094] Steven G. Gilmour. Response surface designs for experiments in bioprocessing. Biometrics, 62(2):323–331, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shepherd:2006:SAC

[1095] Bryan E. Shepherd, Peter B. Gilbert, Yannis Jemiai, and Andrea Rotnitzky. Sensitivity analyses comparing outcomes only existing in a subset selected post-randomization, conditional on covariates, with application to HIV vaccine trials. Biometrics, 62(2):332–342, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2006:PCI

[1096] Fan Li and Constantine E. Frangakis. Polydesigns and causal inference. Biometrics, 62(2):343–351, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Elliott:2006:POA

[1097] Michael R. Elliott, Marshall M. Joffe, and Zhen Chen. A potential outcomes approach to developmental toxicity analyses. Biometrics, 62(2): 352–360, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chao:2006:ABS

[1098] Anne Chao, Robin L. Chazdon, Robert K. Colwell, and Tsung-Jen Shen. Abundance-based similarity indices and their estimation when there are unseen species in samples. Biometrics, 62(2):361–371, June 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borchers:2006:AUH

[1099] D. L. Borchers, J. L. Laake, C. Southwell, and C. G. M. Paxton. Accommodating unmodeled heterogeneity in double-observer distance sampling surveys. Biometrics, 62(2):372–378, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2006:QIF

[1100] Annie Qu and Runze Li. Quadratic inference functions for varyingcoefficient models with longitudinal data. Biometrics, 62(2):379–391, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Welham:2006:ALD

[1101] Sue J. Welham, Brian R. Cullis, Michael G. Kenward, and Robin Thompson. The analysis of longitudinal data using mixed model L-splines. Biometrics, 62(2):392–401, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2006:SIT

[1102] Wenqin Pan, Xihong Lin, and Donglin Zeng. Structural inference in transition measurement error models for longitudinal data. Biometrics, 62(2):402–412, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2006:HBM

[1103] Yangxin Huang, Dacheng Liu, and Hulin Wu. Hierarchical Bayesian methods for estimation of parameters in a longitudinal HIV dynamic system. Biometrics, 62(2):413–423, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fieuws:2006:PFM

[1104] Steffen Fieuws and Geert Verbeke. Pairwise fitting of mixed models for the joint modeling of multivariate longitudinal profiles. Biometrics, 62 (2):424–431, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chi:2006:JMM

[1105] Yueh-Yun Chi and Joseph G. Ibrahim. Joint models for multivariate longitudinal and multivariate survival data. Biometrics, 62(2):432–445, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2006:BCS

[1106] Bo Cai and David B. Dunson. Bayesian covariance selection in generalized linear mixed models. Biometrics, 62(2):446–457, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2006:RAD

[1107] Yang-Jin Kim. Regression analysis of doubly censored failure time data with frailty. Biometrics, 62(2):458–464, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosychuk:2006:SEC

[1108] Rhonda J. Rosychuk, Carolyn Huston, and Narasimha G. N. Prasad. Spatial event cluster detection using a compound Poisson distribution. Biometrics, 62(2):465–470, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ollier:2006:OTD

[1109] S. Ollier, P. Couteron, and D. Chessel. Orthonormal transform to decompose the variance of a life-history trait across a phylogenetic tree. Biometrics, 62(2):471–477, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Katki:2006:EMF

[1110] Hormuzd A. Katki. Effect of misreported family history on Mendelian mutation prediction models. Biometrics, 62(2):478–487, June 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Neuhaus:2006:FSA

[1111] J. M. Neuhaus, A. J. Scott, and C. J. Wild. Family-specific approaches to the analysis of case-control family data. Biometrics, 62(2):488–494, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Swartz:2006:SSG

[1112] Michael D. Swartz, Marek Kimmel, Peter Mueller, and Christopher I. Amos. Stochastic search gene suggestion: a Bayesian hierarchical model for gene mapping. Biometrics, 62(2):495–503, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2006:IGA

[1113] Dong Wang and Dan Nettleton. Identifying genes associated with a quantitative trait or quantitative trait locus via selective transcriptional profiling. Biometrics, 62(2):504–514, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2006:BMM

[1114] Chuan Zhou and Jon Wakefield. A Bayesian mixture model for partitioning gene expression data. Biometrics, 62(2):515–525, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2006:CRM

[1115] Li-Xuan Qin and Steven G. Self. The clustering of regression models method with applications in gene expression data. Biometrics, 62(2):

526–533, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hong:2006:FHM

[1116] F. Hong and H. Li. Functional hierarchical models for identifying genes with different time-course expression profiles. Biometrics, 62(2):534– 544, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2006:NIL

[1117] Peter X.-K. Song, Xin Gao, Rui Liu, and Wen Le. Nonparametric inference for local extrema with application to oligonucleotide microarray data in yeast genome. Biometrics, 62(2):545–554, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2006:RSM

[1118] Shuangge Ma, Michael R. Kosorok, Jian Huang, Hehuang Xie, Liliana Manzella, and Marcelo Bento Soares. Robust semiparametric microarray normalization and significance analysis. Biometrics, 62(2):555–561, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2006:RAR

[1119] Lanju Zhang and William F. Rosenberger. Response-adaptive randomization for clinical trials with continuous outcomes. Biometrics, 62(2): 562–569, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2006:EIC

[1120] Hongkun Wang and Hongwei Zhao. Estimating incremental costeffectiveness ratios and their confidence intervals with differentially censored data. Biometrics, 62(2):570–575, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dinh:2006:NSM

[1121] Phillip Dinh and Xiao-Hua Zhou. Nonparametric statistical methods for cost-effectiveness analyses. Biometrics, 62(2):576–588, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Randolph:2006:MPM

[1122] T. W. Randolph and Y. Yasui. Multiscale processing of mass spectrometry data. Biometrics, 62(2):589–597, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Betensky:2006:HVV

[1123] Rebecca A. Betensky, Caprice K. Christian, Michael L. Gustafson, Jennifer Daley, and Michael J. Zinner. Hospital volume versus outcome: an unusual example of bivariate association. Biometrics, 62(2):598–604, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonzo:2006:NDE

[1124] Todd A. Alonzo and John M. Kittelson. A novel design for estimating relative accuracy of screening tests when complete disease verification is not feasible. Biometrics, 62(2):605–612, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2006:ELI

[1125] Gengsheng Qin and Xiao-Hua Zhou. Empirical likelihood inference for the area under the ROC curve. Biometrics, 62(2):613–622, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gray:2006:LEB

[1126] Robert Gray. Letter to the Editor of Biometrics. Biometrics, 62(2): 623, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2006:ARF

[1127] Zhen Chen and David B. Dunson. The authors replied as follows. Biometrics, 62(2):623–624, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Augustin:2006:BRI

[1128] Thomas Augustin and Carolin Strobl. Book review: Interactive Statistics for the Behavioral Sciences, Edited by P. Williams. Biometrics, 62(2): 625–626, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bammann:2006:IBA

[1129] Karin Bammann. An introduction to bioinformatics algorithms. Biometrics, 62(2):626, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Behrens:2006:PBM

[1130] Thomas Behrens. Publication bias in meta-analyses: Prevention, assessment and adjustments. Biometrics, 62(2):626–627, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Blair:2006:HE

[1131] Aaron Blair. Handbook of epidemiology. Biometrics, 62(2):627–628, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Karlsson:2006:MDL

[1132] Andreas Karlsson. Models for discrete longitudinal data. Biometrics, 62 (2):628, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klaassen:2006:NSM

[1133] Chris A. J. Klaassen. Nonparametric and semiparametric models. Biometrics, 62(2):628–629, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kreiss:2006:NTS

[1134] Jens-Peter Kreiss. Nonlinear time series-nonparametric and parametric methods. Biometrics, 62(2):629–630, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kuchenhoff:2006:MTP

[1135] Helmut Küchenhoff. Measurement theory and practice: The world through qualification. Biometrics, 62(2):630–631, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Langner:2006:SAT

[1136] Ingo Langner. Survival analysis: Techniques for censored and truncated data. Biometrics, 62(2):631, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reineke:2006:DAT

[1137] Achim Reineke. Data analysis tools for DNA microarrays. Biometrics, 62(2):632, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:2006:IRE

[1138] Peter J. Smith. Introduction to rare event simulation. Biometrics, 62 (2):632–633, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Trommer:2006:RMD

[1139] Joerg Trommer. Resampling methods for dependent data. Biometrics, 62(2):633–634, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loughin:2006:BRE

[1140] T. M. Loughin. Brief reports by the Editor. Biometrics, 62(2):634, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRSc

[1141] Anonymous. Book review: Statistical Analysis of Medical Data Using SAS, by Der, G. and Everitt, B. S., Taylor & Francis/CRC Press, Boca Raton, Florida, 2006. vi + 428 pp. US$79.95/£39.99, ISBN 1-58488-469-X. Biometrics, 62(2):634–635, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:RAb

[1142] Anonymous. Regular articles. Biometrics, 62(2):636–643, June 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:CFb

[1143] Anonymous. Consultant’s forum. Biometrics, 62(2):643, June 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ji:2006:CBT

[1144] Hongkai Ji and Wing Hung Wong. Computational biology: Toward deciphering gene regulatory information in mammalian genomes. Biometrics, 62(3):645–663, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burman:2006:FDS

[1145] Carl-Fredrik Burman and Christian Sonesson. Are flexible designs sound? Biometrics, 62(3):664–669, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jennison:2006:D

[1146] Christopher Jennison and Bruce W. Turnbull. Discussions. Biometrics, 62(3):670–673, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proschan:2006:D

[1147] Michael A. Proschan. Discussions. Biometrics, 62(3):674–676, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bauer:2006:D

[1148] P. Bauer. Discussions. Biometrics, 62(3):676–678, September 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frisen:2006:D

[1149] Marianne Frisén. Discussions. Biometrics, 62(3):678–680, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burman:2006:R

[1150] Carl-Fredrik Burman and Christian Sonesson. Rejoinder. Biometrics, 62(3):680–683, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huggins:2006:SEA

[1151] Richard Huggins. Semiparametric estimation of animal abundance using capture–recapture data from open populations. Biometrics, 62(3):684– 690, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gimenez:2006:SRC

[1152] O. Gimenez, C. Crainiceanu, C. Barbraud, S. Jenouvrier, and B. J. T. Morgan. Semiparametric regression in capture–recapture modeling. Biometrics, 62(3):691–698, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cowen:2006:JSM

[1153] Laura Cowen and Carl J. Schwarz. The Jolly–Seber model with tag loss. Biometrics, 62(3):699–705, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dupuis:2006:BES

[1154] Jérôme A. Dupuis and Jean Joachim. Bayesian estimation of species richness from quadrat sampling data in the presence of prior information.

Biometrics, 62(3):706–712, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cadigan:2006:LID

[1155] N. G. Cadigan. Local influence diagnostics for quasi-likelihood and lognormal estimates of a biological reference point from some fish stock and recruitment models. Biometrics, 62(3):713–720, September 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Newton:2006:STH

[1156] Michael A. Newton, Linda Clipson, Andrew T. Thliveris, and Richard B. Halberg. A statistical test of the hypothesis that polyclonal intestinal tumors arise by random collision of initiated clones. Biometrics, 62(3): 721–727, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stewart:2006:IEG

[1157] William C. L. Stewart and Elizabeth A. Thompson. Improving estimates of genetic maps: a maximum likelihood approach. Biometrics, 62(3): 728–734, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2006:PNF

[1158] Baolin Wu, Zhong Guan, and Hongyu Zhao. Parametric and nonparametric FDR estimation revisited. Biometrics, 62(3):735–744, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gurrin:2006:UBM

[1159] Lyle C. Gurrin, John B. Carlin, Jonathan A. C. Sterne, Gillian S. Dite, and John L. Hopper. Using bivariate models to understand betweenand within-cluster regression coefficients, with application to twin data. Biometrics, 62(3):745–751, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chervoneva:2006:GAT

[1160] Inna Chervoneva, Boris Iglewicz, and Terry Hyslop. A general approach for two-stage analysis of multilevel clustered non-Gaussian data. Biometrics, 62(3):752–759, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gustafson:2006:PRC

[1161] Paul Gustafson and Sander Greenland. The performance of random coefficient regression in accounting for residual confounding. Biometrics, 62(3):760–768, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vittinghoff:2006:COA

[1162] E. Vittinghoff and D. C. Bauer. Case-only analysis of treatment-covariate interactions in clinical trials. Biometrics, 62(3):769–776, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yin:2006:BDF

[1163] Guosheng Yin, Yisheng Li, and Yuan Ji. Bayesian dose-finding in Phase I/II clinical trials using toxicity and efficacy odds ratios. Biometrics, 62(3):777–787, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baker:2006:UJI

[1164] Rose Baker and Dan Jackson. Using journal impact factors to correct for the publication bias of medical studies. Biometrics, 62(3):785–792, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2006:PSE

[1165] Jianhua Z. Huang and Linxu Liu. Polynomial spline estimation and inference of proportional hazards regression models with flexible relative risk form. Biometrics, 62(3):793–802, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2006:SSB

[1166] Jiang Lin, Daowen Zhang, and Marie Davidian. Smoothing spline-based score tests for proportional hazards models. Biometrics, 62(3):803–812, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2006:REA

[1167] Jian Huang, Shuangge Ma, and Huiliang Xie. Regularized estimation in the accelerated failure time model with high-dimensional covariates. Biometrics, 62(3):813–820, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tian:2006:APO

[1168] Lu Tian and Stephen Lagakos. Analysis of a partially observed binary covariate process and a censored failure time in the presence of truncation and competing risks. Biometrics, 62(3):821–828, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Datta:2006:NES

[1169] Somnath Datta and Rajeshwari Sundaram. Nonparametric estimation of stage occupation probabilities in a multistage model with current status data. Biometrics, 62(3):829–837, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2006:EDR

[1170] Ruiguang Song, John M. Karon, Edward White, and Gary Goldbaum. Estimating the distribution of a renewal process from times at which events from an independent process are detected. Biometrics, 62(3): 838–846, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fay:2006:EAA

[1171] Michael P. Fay, Ram C. Tiwari, Eric J. Feuer, and Zhaohui Zou. Estimating average annual percent change for disease rates without assuming constant change. Biometrics, 62(3):847–854, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fuentes:2006:SAB

[1172] Montserrat Fuentes, Hae-Ryoung Song, Sujit K. Ghosh, David M. Holland, and Jerry M. Davis. Spatial association between speciated fine particles and mortality. Biometrics, 62(3):855–863, September 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Banerjee:2006:CSM

[1173] Sudipto Banerjee and Gregg A. Johnson. Coregionalized single- and multiresolution spatially varying growth curve modeling with application to weed growth. Biometrics, 62(3):864–876, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2006:MTS

[1174] Yujun Wu, Marc G. Genton, and Leonard A. Stefanski. A multivariate two-sample mean test for small sample size and missing data. Biometrics, 62(3):877–885, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wright:2006:OED

[1175] Stephen E. Wright and A. John Bailer. Optimal experimental design for a nonlinear response in environmental toxicology. Biometrics, 62(3): 886–892, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mehrotra:2006:CEM

[1176] Devan V. Mehrotra, Xiaoming Li, and Peter B. Gilbert. A comparison of eight methods for the dual-endpoint evaluation of efficacy in a proof-of-concept HIV vaccine trial. Biometrics, 62(3):893–900, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2006:ECE

[1177] Xuelin Huang, Janice N. Cormier, and Peter W. T. Pisters. Estimation of the causal effects on survival of two-stage nonrandomized treatment sequences for recurrent diseases. Biometrics, 62(3):901–909, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schaubel:2006:SSM

[1178] Douglas E. Schaubel, Robert A. Wolfe, and Friedrich K. Port. A sequential stratification method for estimating the effect of a time-dependent experimental treatment in observational studies. Biometrics, 62(3):910– 917, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ninan:2006:EMS

[1179] Ipe Ninan, Ottavio Arancio, and Daniel Rabinowitz. Estimation of the mean from sums with unknown numbers of summands. Biometrics, 62(3):918–920, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [1258].

Tutz:2006:GAM

[1180] Gerhard Tutz and Harald Binder. Generalized additive modeling with implicit variable selection by likelihood-based boosting. Biometrics, 62(4):961–971, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Claeskens:2006:VSL

[1181] Gerda Claeskens, Christophe Croux, and Johan Van Kerckhoven. Variable selection for logistic regression using a prediction-focused information criterion. Biometrics, 62(4):972–979, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goeman:2006:GFT

[1182] Jelle J. Goeman and Saskia le Cessie. A goodness-of-fit test for multinomial logistic regression. Biometrics, 62(4):980–985, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kong:2006:GRS

[1183] Maiying Kong and J. Jack Lee. A generalized response surface model with varying relative potency for assessing drug interaction. Biometrics, 62(4):986–995, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lam:2006:SAZ

[1184] K. F. Lam, Hongqi Xue, and Yin Bun Cheung. Semiparametric analysis of zero-inflated count data. Biometrics, 62(4):996–1003, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Naskar:2006:SAT

[1185] Malay Naskar and Kalyan Das. Semiparametric analysis of two-level bivariate binary data. Biometrics, 62(4):1004–1013, December 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proust:2006:NML

[1186] Cécile Proust, Hélène Jacqmin-Gadda, Jeremy M. G. Taylor, Julien Ganiayre, and Daniel Commenges. A nonlinear model with latent process for cognitive evolution using multivariate longitudinal data. Biometrics, 62(4):1014–1024, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wood:2006:LRS

[1187] Simon N. Wood. Low-rank scale-invariant tensor product smooths for generalized additive mixed models. Biometrics, 62(4):1025–1036, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hsieh:2006:JMS

[1188] Fushing Hsieh, Yi-Kuan Tseng, and Jane-Ling Wang. Joint modeling of survival and longitudinal data: Likelihood approach revisited. Biometrics, 62(4):1037–1043, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pennell:2006:BSD

[1189] Michael L. Pennell and David B. Dunson. Bayesian semiparametric dynamic frailty models for multiple event time data. Biometrics, 62(4): 1044–1052, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2006:IMQ

[1190] Mengling Liu, Wenbin Lu, and Yongzhao Shao. Interval mapping of quantitative trait loci for time-to-event data with the proportional hazards mixture cure model. Biometrics, 62(4):1053–1061, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Houseman:2006:FSP

[1191] E. Andrés Houseman, Brent A. Coull, and Rebecca A. Betensky. Featurespecific penalized latent class analysis for genomic data. Biometrics, 62 (4):1062–1070, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vincent:2006:QGI

[1192] Quentin Vincent, Alexandre Alcaïs, Andrea Alter, Erwin Schurr, and Laurent Abel. Quantifying genomic imprinting in the presence of linkage. Biometrics, 62(4):1071–1080, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2006:MEP

[1193] Yuanjia Wang, Ruth Ottman, and Daniel Rabinowitz. A method for estimating penetrance from families sampled for linkage analysis. Biometrics, 62(4):1081–1088, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2006:UAS

[1194] Ming Yuan and Christina Kendziorski. A unified approach for simultaneous gene clustering and differential expression identification. Biometrics, 62(4):1089–1098, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2006:SVA

[1195] Debashis Ghosh. Shrunken p-values for assessing differential expression with applications to genomic data analysis. Biometrics, 62(4):1099–1106, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gort:2006:FLD

[1196] Gerrit Gort, Wim J. M. Koopman, and Alfred Stein. Fragment length distributions and collision probabilities for AFLP markers. Biometrics, 62(4):1107–1115, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moskvina:2006:SBH

[1197] Valentina Moskvina and Karl Michael Schmidt. Susceptibility of biallelic haplotype and genotype frequencies to genotyping error. Biometrics, 62 (4):1116–1123, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2006:SAH

[1198] H. Zhang, G. Zheng, and Z. Li. Statistical analysis for haplotype-based matched case-control studies. Biometrics, 62(4):1124–1131, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bukszar:2006:OTS

[1199] József Bukszár and Edwin J. C. G. van den Oord. Optimization of two-stage genetic designs where data are combined using an accurate and efficient approximation for Pearson’s statistic. Biometrics, 62(4): 1132–1137, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2006:CCD

[1200] Shou-En Lu and Joanna H. Shih. Case-cohort designs and analysis for clustered failure time data. Biometrics, 62(4):1138–1148, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2006:SEL

[1201] Xiaofei Wang and Haibo Zhou. A semiparametric empirical likelihood method for biased sampling schemes with auxiliary covariates. Biometrics, 62(4):1149–1160, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:2006:ADA

[1202] Dean Follmann. Augmented designs to assess immune response in vaccine trials. Biometrics, 62(4):1161–1169, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lekone:2006:SIS

[1203] Phenyo E. Lekone and Bärbel F. Finkenstädt. Statistical inference in a stochastic epidemic SEIR model with control intervention: Ebola as a case study. Biometrics, 62(4):1170–1177, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staudenmayer:2006:MER

[1204] John Staudenmayer and John P. Buonaccorsi. Measurement error in a random walk model with applications to population dynamics. Biometrics, 62(4):1178–1189, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2006:MEM

[1205] Aiyi Liu, Enrique F. Schisterman, and Chengqing Wu. Multistage evaluation of measurement error in a reliability study. Biometrics, 62(4): 1190–1196, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2006:ERS

[1206] Brian J. Reich, James S. Hodges, and Vesna Zadnik. Effects of residual smoothing on the posterior of the fixed effects in disease-mapping models. Biometrics, 62(4):1197–1206, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohning:2006:ETC

[1207] Dankmar Böhning and Ronny Kuhnert. Equivalence of truncated count mixture distributions and mixtures of truncated count distributions. Biometrics, 62(4):1207–1215, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridout:2006:NAM

[1208] M. S. Ridout, D. J. Cole, B. J. T. Morgan, L. J. Byrne, and M. F. Tuite. New approximations to the Malthusian parameter. Biometrics, 62(4): 1216–1223, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:2006:AWS

[1209] Steven K. Thompson. Adaptive Web sampling. Biometrics, 62(4):1224– 1234, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ridall:2006:MUN

[1210] P. Gareth Ridall, Anthony N. Pettitt, Robert D. Henderson, and Pamela A. McCombe. Motor unit number estimation — a Bayesian approach. Biometrics, 62(4):1235–1250, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosner:2006:ERS

[1211] Bernard Rosner, Robert J. Glynn, and Mei-Ling T. Lee. Extension of the rank sum test for clustered data: Two-group comparisons with group membership defined at the subunit level. Biometrics, 62(4):1251–1259, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2006:SLS

[1212] Jiameng Zhang and Daniel F. Heitjan. A simple local sensitivity analysis tool for nonignorable coarsening: Application to dependent censoring. Biometrics, 62(4):1260–1268, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Asmussen:2006:BRB

[1213] Søren Asmussen. Book review: Branching Processes: Variation, Growth and Extinction of Populations, edited by P. Haccou, P. Jagers, and V. A. Vatutin. Biometrics, 62(4):1269–1270, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Doerge:2006:BRB

[1214] R. W. Doerge. Book review: Bioinformatics and Computational Biology Solutions Using R and Bioconductor, edited by Gentleman, R., Carey, V., Huber, W., Irizarry, R., and Dudoit, S. Biometrics, 62(4):1270–1271, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Elliott:2006:BRR

[1215] Michael Elliott. Book review: Regression Methods in Biostatistics: Linear, Logistic, Survival, and Repeated Measures Models, Edited by Vittinghoff, E., Glidden, D. V., Shiboski, S. C., and McCulloch, C. E. Biometrics, 62(4):1271–1272, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerow:2006:BRS

[1216] Ken Gerow. Book review: Statistics Explained: an Introductory Guide for Life Scientists, Edited by Killup, S. Biometrics, 62(4):1272, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Horgan:2006:BRS

[1217] Graham Horgan. Book review: Stereology for Statisticians, Edited by Baddeley, A. and Vedel Jensen, E. B. Biometrics, 62(4):1272–1273, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loughin:2006:BRS

[1218] Thomas M. Loughin. Book review: SAS(R) for Mixed Models, 2nd edition, edited by Littell, R. C., Milliken, G. A., Stroup, W. W., Wolfinger, R. D., and Schabenberger, O. Biometrics, 62(4):1273–1274, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pigeot:2006:BRD

[1219] Iris Pigeot. Book review: Data Monitoring in Clinical Trials. A Case Studies Approach, edited by Demets, D. L., Furberg, C. D., and Friedman, L. M. Biometrics, 62(4):1274–1275, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pontius:2006:BRD

[1220] Jeffrey S. Pontius. Book review: Design of Experiments for Agriculture and the Natural Sciences, 2nd edition, Edited by Hoshmand A. R. Biometrics, 62(4):1275, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Espejo:2006:BRS

[1221] M. Ruiz Espejo and F. J. Montero. Book review: Statistical Demography and Forecasting, edited by Alho, J. M. and Spencer, B. D. Biometrics, 62(4):1275–1276, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sandercock:2006:BRH

[1222] Brett K. Sandercock. Book review: Handbook of Capture–Recapture Analysis, Edited by Amstrup, S. C., Mcdonald, T. L., and Manly, B. F. J. Biometrics, 62(4):1276–1277, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sargent:2006:BRH

[1223] Daniel Sargent, Sumithra Mandrekar, and Ann Oberg. Book review: Handbook of Statistics in Clinical Oncology, 2nd edition, Edited by Crowley, J. and Ankerst, D. P. Biometrics, 62(4):1277–1278, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2006:BRE

[1224] Haiyan Wang. Book review: Extending the Linear Model with R: Generalized Linear, Mixed Effects and Nonparametric Regression Models, Edited by Faraway, J. J. Biometrics, 62(4):1278, December 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRHa

[1225] Anonymous. Book review: A Handbook of Statistical Analyses Using R, Edited by Everitt, B. S. and Hothorn, T. Biometrics, 62(4):1279, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRSd

[1226] Anonymous. Book review: Screening: Methods for Experimentation in Industry, Drug Discovery, and Genetics, Edited by Dean, A. and Lewis, S. Biometrics, 62(4):1279, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRHb

[1227] Anonymous. Book review: Handbook of Univariate and Multivariate Data Analyses and Interpretation with SPSS, edited by Ho, R. Biometrics, 62(4):1279–1280, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRC

[1228] Anonymous. Book review: Cancer Bioinformatics: From Therapy Design to Treatment, Edited by Nagl S. Biometrics, 62(4):1280, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:RAc

[1229] Anonymous. Regular articles. Biometrics, 62(4):1282–1290, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:CFc

[1230] Anonymous. Consultant’s forum. Biometrics, 62(4):1290, December 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klingenberg:2006:MEM

[1231] Bernhard Klingenberg and Alan Agresti. Multivariate extensions of Mc-Nemar’s Test. Biometrics, 63(2):921–928, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2006:ACP

[1232] Mimi Y. Kim, Xiaonan Xue, and Yunling Du. Approaches for calculating power for case-cohort studies. Biometrics, 63(2):929–933, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2006:ARF

[1233] Jianwen Cai and Donglin Zeng. The authors replied as follows. Biometrics, 63(2):933, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holzmann:2006:ICR

[1234] Hajo Holzmann, Axel Munk, and Walter Zucchini. On identifiability in capture–recapture models. Biometrics, 63(2):934–936, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Link:2006:ARF

[1235] William A. Link. The author replied as follows. Biometrics, 63(2):936– 939, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alpizar-Jara:2006:BRA

[1236] Russell Alpizar-Jara. Book review: Advanced Distance Sampling: Estimating Abundance of Biological Populations, Edited by S. T. Buckland, D. R. Anderson, K. P. Burnham, J. L. Laake, D. L. Borchers, and L. Thomas. Biometrics, 63(2):940–941, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andersen:2006:FMM

[1237] Per Kragh Andersen. Flowgraph models for multistate time-to-event data. Biometrics, 63(2):941–942, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bailey:2006:SAG

[1238] Trevor Bailey. Spatial analysis: a guide for ecologists. Biometrics, 63(2): 942–943, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bammann:2006:SMT

[1239] Karin Bammann. Statistical models: Theory and practice. Biometrics, 63(2):943, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fahrmeir:2006:HSS

[1240] Ludwig Fahrmeir. Handbook of statistics 21: Stochastic processes: Modelling and simulation. Biometrics, 63(2):943–944, September 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graur:2006:SMM

[1241] Dan Graur and Dror Berel. Statistical methods in molecular evolution. Biometrics, 63(2):944, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gueorguieva:2006:MLD

[1242] Ralitza Gueorguieva. Modeling longitudinal data. Biometrics, 63(2): 944–945, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hauschke:2006:BSC

[1243] Dieter Hauschke. Bioequivalence and statistics in clinical pharmacology. Biometrics, 63(2):945–946, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Karlsson:2006:NSE

[1244] Andreas Karlsson. The nature of scientific evidence: Statistical, philosophical and empirical considerations. Biometrics, 63(2):946–947, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mortensen:2006:CCP

[1245] Uwe Mortensen. Cognition and chance — the psychology of probabilistic reasoning. Biometrics, 63(2):947–948, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rassler:2006:ABM

[1246] Susanne Rässler. Applied Bayesian modeling and causal inference from incomplete-data perspectives. Biometrics, 63(2):948, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanHouwelingen:2006:ESE

[1247] Hans C. Van Houwelingen. The evaluation of surrogate endpoints. Biometrics, 63(2):948–949, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schill:2006:BRA

[1248] Walter Schill. Book review: Advanced Log–Linear Models Using SAS, by D. Zelterman, SAS Institute, Cary, North Carolina, 2002 . vii + 189 pp . US$34.95, ISBN 1-59047-080-X. Biometrics, 63(2):949–950, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tu:2006:NMC

[1249] Dongsheng Tu. Nonparametric Monte-Carlo tests and their applications. Biometrics, 63(2):950–951, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BREc

[1250] Anonymous. Brief reports by the Editor. Biometrics, 63(2):951, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:MDD

[1251] Anonymous. Models for discrete data, revised edition. Biometrics, 63 (2):951, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:MLM

[1252] Anonymous. Multilevel and longitudinal modeling using Stata. Biometrics, 63(2):951, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:DP

[1253] Anonymous. Discussion paper. Biometrics, 63(2):952, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:EIP

[1254] Anonymous. Editors’ invited paper. Biometrics, 63(2):952, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:RAd

[1255] Anonymous. Regular articles. Biometrics, 63(2):952–958, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:CFd

[1256] Anonymous. Consultant’s forum. Biometrics, 63(2):958–959, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:RRb

[1257] Anonymous. Reader reaction. Biometrics, 63(2):959, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rabinowitz:2007:CEM

[1258] Daniel Rabinowitz. Correction to “Estimation of the Mean from Sums with Unknown Numbers of Summands,” by I. Ninan, O. Arancio, and D. Rabinowitz; 62, 918–920, September 2006. Biometrics, 60(1):623, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1179].

Wakefield:2007:BRS

[1259] Jon Wakefield. Book review: Statistical Analysis of Environmental Space-Time Processes, edited by N. D. Le and J. V. Zidek. Biometrics, 62(3): 624–625, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:2007:BRC

[1260] A. W. Kemp. Book reviews: Computer-Intensive Methods of Data Analysis in Biology, edited by D. A. Roff. Randomization, Bootstrap, and Monte Carlo Methods in Biology (3rd edition) edited by B. F. J. Manly. Biometrics, 62(3):625–626, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gurka:2007:BRA

[1261] Matthew J. Gurka. Book review: Applied Mixed Models in Medicine, 2nd edition edited by H. Brown and R. Prescott. Biometrics, 62(3):626– 627, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wolkenhauer:2007:BRS

[1262] Olaf Wolkenhauer. Book review: Stochastic Modeling for Systems Biology, edited by D. J. Wilkinson. Biometrics, 62(3):627–628, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2007:BRK

[1263] Zhen Zhang. Book review: Knowledge Discovery in Proteomics, edited by I. Jurisica and D. Wigle. Biometrics, 62(3):628, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Suling:2007:BRC

[1264] Marc Suling. Book review: Computational Genome Analysis: an Introduction, by R. C. Deonier, S. Tavaré, and M. S. Waterman. Biometrics, 62(3):628–629, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kochar:2007:BRS

[1265] Subhash Kochar. Book review: Stochastic Orders, by M. Shaked and J. G. Shantikumar. Biometrics, 62(3):629, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRMa

[1266] Anonymous. Book review: Measurement Error in Nonlinear Models: a Modern Perspective, 2nd edition by R. J. Carroll, D. Ruppert, L. A. Stefansky, and C. M. Crainiceanu. Biometrics, 62(3):629–630, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRBa

[1267] Anonymous. Book review: Basic Statistics and Pharmaceutical Statistical Applications, 2nd edition by J. E. De Muth. Biometrics, 62(3): 630, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRHa

[1268] Anonymous. Book review: A Handbook of Statistical Analyses Using Stata, 4th edition by S. Rabe-Hesketh and B. S. Everitt. Biometrics, 62 (3):630, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRIa

[1269] Anonymous. Book review: Introduction to Randomized Controlled Clinical Trials, 2nd edition by J. N. S. Matthews. Biometrics, 62(3):630, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRPa

[1270] Anonymous. Book review: A Pocket Guide to Epidemiology, by D. G. Kleinbaum, K. M. Sullivan, and N. D. Barker. Biometrics, 62(3):631, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRSa

[1271] Anonymous. Book review: Survival and Event History Analysis, edited by P. K. Andersen and N. Keiding. Biometrics, 62(3):631, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:IAL

[1272] Anonymous. IBS 60th anniversary lecture. Biometrics, 62(3):632, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:RAa

[1273] Anonymous. Regular articles. Biometrics, 62(3):632–639, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:CFa

[1274] Anonymous. Consultant’s forum. Biometrics, 62(3):639–640, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Louis:2007:OFH

[1275] Thomas A. Louis. Our future as history. Biometrics, 63(1):1–9, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Keles:2007:MMG

[1276] Sündüz Keles. Mixture modeling for genome-wide localization of transcription factors. Biometrics, 63(1):10–21, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2007:MBI

[1277] Nancy R. Zhang and David O. Siegmund. A modified Bayes information criterion with applications to the analysis of comparative genomic hybridization data. Biometrics, 63(1):22–32, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sterrett:2007:ILT

[1278] Andrew Sterrett and Fred A. Wright. Inferring the location of tumor suppressor genes by modeling frequency of allelic loss. Biometrics, 63 (1):33–40, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2007:ADG

[1279] Jianhua Hu and Fred A. Wright. Assessing differential gene expression with small sample sizes in oligonucleotide arrays using a mean-variance model. Biometrics, 63(1):41–49, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2007:EQN

[1280] Jianhua Hu and Xuming He. Enhanced quantile normalization of microarray data to reduce loss of information in gene expression profiles.

Biometrics, 63(1):50–59, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2007:JMP

[1281] Chengcheng Hu and Victor De Gruttola. Joint modeling of progression of HIV resistance mutations measured with uncertainty and failure time data. Biometrics, 63(1):60–68, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gustafson:2007:RMF

[1282] Paul Gustafson. On robustness and model flexibility in survival analysis: Transformed hazard models and average effects. Biometrics, 63(1):69– 77, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ye:2007:SAC

[1283] Yining Ye, John D. Kalbfleisch, and Douglas E. Schaubel. Semiparametric analysis of correlated recurrent and terminal events. Biometrics, 63(1):78–87, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hanson:2007:BSP

[1284] Timothy Hanson and Mingan Yang. Bayesian semiparametric proportional odds models. Biometrics, 63(1):88–95, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peng:2007:RMS

[1285] Limin Peng and Jason P. Fine. Regression modeling of semicompeting risks data. Biometrics, 63(1):96–108, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2007:SSS

[1286] Lan Huang, Martin Kulldorff, and David Gregorio. A spatial scan statistic for survival data. Biometrics, 63(1):109–118, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tango:2007:CMA

[1287] Toshiro Tango. A class of multiplicity adjusted tests for spatial clustering based on case-control point data. Biometrics, 63(1):119–127, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haneuse:2007:HMC

[1288] Sebastien J-P. A. Haneuse and Jonathan C. Wakefield. Hierarchical models for combining ecological and case-control data. Biometrics, 63(1):128– 136, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gorfine:2007:LME

[1289] Malka Gorfine, Nurit Lipshtat, Laurence S. Freedman, and Ross L. Prentice. Linear measurement error models with restricted sampling. Biometrics, 63(1):137–142, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sugar:2007:LRE

[1290] Elizabeth A. Sugar, Ching-Yun Wang, and Ross L. Prentice. Logistic regression with exposure biomarkers and flexible measurement error. Biometrics, 63(1):143–151, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2007:MCR

[1291] Tianxi Cai and Yingye Zheng. Model checking for ROC regression analysis. Biometrics, 63(1):152–163, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2007:NEC

[1292] Ying Guo and Amita K. Manatunga. Nonparametric estimation of the concordance correlation coefficient under univariate censoring. Biometrics, 63(1):164–172, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2007:CRM

[1293] Z. Yuan, R. Chappell, and H. Bailey. The continual reassessment method for multiple toxicity grades: a Bayesian quasi-likelihood approach. Biometrics, 63(1):173–179, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2007:SME

[1294] Ariel Alonso and Geert Molenberghs. Surrogate marker evaluation from an information theory perspective. Biometrics, 63(1):180–186, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jackson:2007:AIP

[1295] Dan Jackson. Assessing the implications of publication bias for two popular estimates of between-study variance in meta-analysis. Biometrics, 63(1):187–193, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

May:2007:NTT

[1296] Susanne May and Victor DeGruttola. Nonparametric tests for two-group comparisons of dependent observations obtained at varying time points. Biometrics, 63(1):194–200, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2007:MHS

[1297] Anna Liu and Yuedong Wang. Modeling of hormone secretion-generating mechanisms with splines: a pseudo-likelihood approach. Biometrics, 63 (1):201–208, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2007:LRS

[1298] Haonan Wang and M. Giovanna Ranalli. Low-rank smoothing splines on complicated domains. Biometrics, 63(1):209–217, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pang:2007:TMC

[1299] Zhen Pang and Anthony Y. C. Kuk. Test of marginal compatibility and smoothing methods for exchangeable binary data with unequal cluster sizes. Biometrics, 63(1):218–227, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xi:2007:ULB

[1300] Liqun Xi, Paul S. F. Yip, and Ray Watson. A unified likelihood-based approach for estimating population size in continuous-time capture– recapture experiments with frailty. Biometrics, 63(1):228–236, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dryden:2007:SAU

[1301] Ian L. Dryden, Jonathan D. Hirst, and James L. Melville. Statistical analysis of unlabeled point sets: Comparing molecules in chemoinformatics. Biometrics, 63(1):237–251, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Waagepetersen:2007:EFA

[1302] Rasmus Plenge Waagepetersen. An estimating function approach to inference for inhomogeneous Neyman-Scott processes. Biometrics, 63(1): 252–258, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Datta:2007:PPS

[1303] Susmita Datta, Jennifer Le-Rademacher, and Somnath Datta. Predicting patient survival from microarray data by accelerated failure time modeling using partial least squares and LASSO. Biometrics, 63(1):259– 271, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ladouceur:2007:RPE

[1304] Martin Ladouceur, Elham Rahme, Christian A. Pineau, and Lawrence Joseph. Robustness of prevalence estimates derived from misclassified data from administrative databases. Biometrics, 63(1):272–279, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:2007:SBI

[1305] Peter Müller, Fernando A. Quintana, and Gary L. Rosner. Semiparametric Bayesian inference for multilevel repeated measurement data. Biometrics, 63(1):280–289, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Assuncao:2007:NTS

[1306] Renato Assunção and Alexandra Maia. A note on testing separability in spatial-temporal marked point processes. Biometrics, 63(1):290–294, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schoenberg:2007:ARF

[1307] Frederic Paik Schoenberg. The authors replied as follows. Biometrics, 63 (1):294–295, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Senn:2007:DNS

[1308] Stephen Senn. Drawbacks to noninteger scoring for ordered categorical data. Biometrics, 63(1):296–298, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berger:2007:ARF

[1309] Vance W. Berger. The authors replied as follows. Biometrics, 63(1):298– 299, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Billheimer:2007:BRF

[1310] D. Billheimer. Book review: Functional Data Analysis, 2nd edition edited by J. O. Ramsay and B. W. Silverman. Biometrics, 63(1):300–301, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Park:2007:BRB

[1311] Taesung Park. Book review: Bayesian Models for Categorical Data, edited by Congdon, P. Biometrics, 63(1):301–302, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mehrotra:2007:BRE

[1312] Devan V. Mehrotra. Book review: Exact Analysis of Discrete Data, edited by Hirji, K. F. Biometrics, 63(1):302–303, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Naranjo:2007:BRR

[1313] J. Naranjo. Book review: Robust Statistical Methods with R, edited by Jureckova, A. and Picek, J. Biometrics, 63(1):303, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Njuho:2007:BRD

[1314] Peter M. Njuho. Book review: Dose Finding in Drug Development, edited by Naitee T. Biometrics, 63(1):303–304, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Therneau:2007:BRD

[1315] Terry Therneau. Book review: Dynamic Regression Models for Survival Data, edited by Martinussen, T. and Scheike, T. H. Biometrics, 63(1): 304, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berman:2007:BRS

[1316] Nancy G. Berman. Book reviews: Statistical Evidence in Medical Trials: What Do the Data Really Tell Us?, edited by Simon, S. D.; Statistical Reasoning in Medicine: The Intuitive P -value Primer edited by Moye, L. A. Biometrics, 63(1):305–306, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRMb

[1317] Anonymous. Book review: Multivariable Analysis: a Practical Guide for Clinicians, 2nd edition, edited by Katz, M. H. Study Design and Statistical Analysis: a Practical Guide for Clinicians edited by Katz, M. H. Biometrics, 63(1):306, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRSb

[1318] Anonymous. Book review: Survival Analysis: a Practical Approach, 2nd Edition edited by Machin, D., Cheung, Y. B., and Parmar, M. K. Biometrics, 63(1):306–307, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRL

[1319] Anonymous. Book review: Linear Model Theory: Univariate, Multivariate, and Mixed Models, edited by Muller, K. E. and Stewart, P. W. Biometrics, 63(1):307, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRIb

[1320] Anonymous. Book review: Introduction to Randomized Controlled Clinical Trials, 2nd Edition edited by Matthews, J. N. S. Biometrics, 63 (1):307–308, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:PA

[1321] Anonymous. Presidential address. Biometrics, 63(1):309, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:RAb

[1322] Anonymous. Regular articles. Biometrics, 63(1):309–315, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:CFb

[1323] Anonymous. Consultant’s forum. Biometrics, 63(1):315–316, March 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:RRa

[1324] Anonymous. Reader reaction. Biometrics, 63(1):316, March 2007. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stigler:2007:PIB

[1325] Stephen M. Stigler. The pedigree of the International Biometric Society. Biometrics, 63(2):317–321, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schildcrout:2007:MMM

[1326] Jonathan S. Schildcrout and Patrick J. Heagerty. Marginalized models for moderate to long series of longitudinal binary response data. Biometrics, 63(2):322–331, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2007:PAL

[1327] Yingye Zheng and Patrick J. Heagerty. Prospective accuracy for longitudinal markers. Biometrics, 63(2):332–341, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2007:SIS

[1328] Wei Liu and Lang Wu. Simultaneous inference for semiparametric nonlinear mixed-effects models with covariate measurement errors and missing responses. Biometrics, 63(2):342–350, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2007:TSF

[1329] Daowen Zhang, Xihong Lin, and MaryFran Sowers. Two-stage functional mixed models for evaluating the effect of longitudinal covariate profiles on a scalar outcome. Biometrics, 63(2):351–362, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2007:JAL

[1330] Mengling Liu and Zhiliang Ying. Joint analysis of longitudinal data with informative right censoring. Biometrics, 63(2):363–371, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hudgens:2007:NEJ

[1331] Michael G. Hudgens, Marloes H. Maathuis, and Peter B. Gilbert. Nonparametric estimation of the joint distribution of a survival time subject to interval censoring and a continuous mark variable. Biometrics, 63 (2):372–380, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Herring:2007:BMM

[1332] Amy H. Herring and Juan Yang. Bayesian modeling of multiple episode occurrence and severity with a terminating event. Biometrics, 63(2):381– 388, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2007:JFM

[1333] Xuelin Huang and Lei Liu. A joint frailty model for survival and gap times between recurrent events. Biometrics, 63(2):389–397, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andrei:2007:RMM

[1334] Adin-Cristian Andrei and Susan Murray. Regression models for the mean of the quality-of-life-adjusted restricted survival time using pseudoobservations. Biometrics, 63(2):398–404, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandel:2007:TGF

[1335] Micha Mandel and Rebecca A. Betensky. Testing goodness of fit of a uniform truncation model. Biometrics, 63(2):405–412, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2007:SOR

[1336] Hua Yun Chen. A semiparametric odds ratio model for measuring association. Biometrics, 63(2):413–421, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lokhnygina:2007:CRM

[1337] Yuliya Lokhnygina and Jeffrey D. Helterbrand. Cox regression methods for two-stage randomization designs. Biometrics, 63(2):422–428, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2007:PPI

[1338] Xuelin Huang, Swati Biswas, Yasuhiro Oki, Jean-Pierre Issa, and Donald A. Berry. A parallel Phase I/II clinical trial design for combination therapies. Biometrics, 63(2):429–436, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mattei:2007:APS

[1339] A. Mattei and F. Mealli. Application of the principal stratification approach to the faenza randomized experiment on breast self-examination.

Biometrics, 63(2):437–446, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moodie:2007:DOD

[1340] Erica E. M. Moodie, Thomas S. Richardson, and David A. Stephens. Demystifying optimal dynamic treatment regimes. Biometrics, 63(2): 447–455, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2007:SAE

[1341] Paul R. Rosenbaum. Sensitivity analysis for m-estimates, tests, and confidence intervals in matched observational studies. Biometrics, 63 (2):456–464, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fay:2007:AVS

[1342] Michael P. Fay, M. Elizabeth Halloran, and Dean A. Follmann. Accounting for variability in sample size estimation with applications to nonadherence and estimation of variance and effect size. Biometrics, 63 (2):465–474, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Henmi:2007:CIV

[1343] Masayuki Henmi, John B. Copas, and Shinto Eguchi. Confidence intervals and P -values for meta-analysis with publication bias. Biometrics, 63(2):475–482, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ruppert:2007:EIV

[1344] David Ruppert, Dan Nettleton, and J. T. Gene Hwang. Exploring the information in p-values for the analysis and planning of multiple-test experiments. Biometrics, 63(2):483–495, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chi:2007:BHM

[1345] Yueh-Yun Chi, Joseph G. Ibrahim, Anika Bissahoyo, and David W. Threadgill. Bayesian hierarchical modeling for time course microarray experiments. Biometrics, 63(2):496–504, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mardia:2007:PBM

[1346] Kanti V. Mardia, Charles C. Taylor, and Ganesh K. Subramaniam. Protein bioinformatics and mixtures of bivariate von Mises distributions for angular data. Biometrics, 63(2):505–512, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2007:EBM

[1347] Shizhong Xu. An empirical Bayes method for estimating epistatic effects of quantitative trait loci. Biometrics, 63(2):513–521, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Begg:2007:STC

[1348] Colin B. Begg, Kevin H. Eng, and Amanda J. Hummer. Statistical tests for clonality. Biometrics, 63(2):522–530, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Serban:2007:MMP

[1349] Nicoleta Serban. MICE: Multiple-peak identification, characterization, and estimation. Biometrics, 63(2):531–539, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2007:SCD

[1350] Andrea J. Cook, Diane R. Gold, and Yi Li. Spatial cluster detection for censored outcome data. Biometrics, 63(2):540–549, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Diggle:2007:SOA

[1351] P. J. Diggle, V. Gómez-Rubio, P. E. Brown, A. G. Chetwynd, and S. Gooding. Second-order analysis of inhomogeneous spatial point processes using case-control data. Biometrics, 63(2):550–557, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hooten:2007:HSM

[1352] Mevin B. Hooten, Christopher K. Wikle, Robert M. Dorazio, and J. Andrew Royle. Hierarchical spatiotemporal matrix models for characterizing invasions. Biometrics, 63(2):558–567, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartolucci:2007:CLM

[1353] Francesco Bartolucci and Fulvia Pennoni. A class of latent Markov models for capture–recapture data allowing for time, heterogeneity, and behavior effects. Biometrics, 63(2):568–578, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fernandez-Duran:2007:MCL

[1354] J. J. Fernández-Durán. Models for circular-linear and circular-circular data constructed from circular distributions based on nonnegative trigonometric sums. Biometrics, 63(2):579–585, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sexton:2007:BRT

[1355] Joseph Sexton and Petter Laake. Boosted regression trees with errors in variables. Biometrics, 63(2):586–592, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2007:REM

[1356] Paul S. Albert. Random effects modeling approaches for estimating ROC curves from repeated ordinal tests without a gold standard. Biometrics, 63(2):593–602, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nakas:2007:RGA

[1357] Christos T. Nakas and Todd A. Alonzo. ROC graphs for assessing the ability of a diagnostic marker to detect three disease classes with an umbrella ordering. Biometrics, 63(2):603–609, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Comulada:2007:MBD

[1358] W. Scott Comulada and Robert E. Weiss. On models for binomial data with random numbers of trials. Biometrics, 63(2):610–617, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morgan:2007:NSL

[1359] Byron J. T. Morgan, David J. Revell, and Stephen N. Freeman. A note on simplifying likelihoods for site occupancy models. Biometrics, 63(2): 618–621, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Perlman:2007:CNO

[1360] Michael D. Perlman and Lang Wu. Correction to “A Note on One-Sided Tests with Multiple Endpoints,” by M. D. Perlman and L. Wu; 60, 276– 280, March 2004. Biometrics, 63(2):622, June 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [694].

Frangakis:2007:PSD

[1361] Constantine E. Frangakis, Donald B. Rubin, Ming-Wen An, and Ellen MacKenzie. Principal stratification designs to estimate input data missing due to death. Biometrics, 63(3):641–649, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robins:2007:D

[1362] James Robins, Andrea Rotnitzky, and Stijn Vansteelandt. Discussions. Biometrics, 63(3):650–653, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

TenHave:2007:D

[1363] Tom Ten Have. Discussions. Biometrics, 63(3):654–655, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xie:2007:D

[1364] Yu Xie and Susan Murphy. Discussions. Biometrics, 63(3):655–658, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frangakis:2007:R

[1365] Constantine E. Frangakis, Donald B. Rubin, Ming-Wen An, and Ellen MacKenzie. Rejoinder. Biometrics, 63(3):658–662, September 2007. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cong:2007:MAC

[1366] Xiuyu J. Cong, Guosheng Yin, and Yu Shen. Marginal analysis of correlated failure time data with informative cluster sizes. Biometrics, 63(3): 663–672, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shih:2007:AFT

[1367] Joanna H. Shih and Shou-En Lu. Analysis of failure time data with multilevel clustering, with application to the child vitamin A intervention trial in Nepal. Biometrics, 63(3):673–680, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2007:MPA

[1368] You-Gan Wang and Yuning Zhao. A modified pseudolikelihood approach for analysis of longitudinal data. Biometrics, 63(3):681–689, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kinney:2007:FRE

[1369] Satkartar K. Kinney and David B. Dunson. Fixed and random effects selection in linear and logistic models. Biometrics, 63(3):690–698, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:2007:SSD

[1370] Anindya Roy, Dulal K. Bhaumik, Subhash Aryal, and Robert D. Gibbons. Sample size determination for hierarchical longitudinal designs with differential attrition rates. Biometrics, 63(3):699–707, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huggins:2007:AAS

[1371] Richard M. Huggins, Peter Hall, Paul S. F. Yip, and Quang M. Bui. Applications of additive semivarying coefficient models: Monthly suicide data from Hong Kong. Biometrics, 63(3):708–713, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shi:2007:GPF

[1372] J. Q. Shi, B. Wang, R. Murray-Smith, and D. M. Titterington. Gaussian process functional regression modeling for batch data. Biometrics, 63(3): 714–723, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bigelow:2007:BAR

[1373] Jamie L. Bigelow and David B. Dunson. Bayesian adaptive regression splines for hierarchical data. Biometrics, 63(3):724–732, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:2007:HSR

[1374] Li-Jung Liang and Robert E. Weiss. A hierarchical semiparametric regression model for combining HIV-1 phylogenetic analyses using iterative reweighting algorithms. Biometrics, 63(3):733–741, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Healy:2007:CCS

[1375] Brian Healy, Victor De Gruttola, and Marcello Pagano. Combining crosssectional and prospective data methods to improve transition parameter estimation for characterizing the accumulation of HIV-1 drug resistance mutations. Biometrics, 63(3):742–750, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2007:CMM

[1376] Shuangge Ma and Jian Huang. Combining multiple markers for classification using ROC. Biometrics, 63(3):751–757, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Picard:2007:SCM

[1377] F. Picard, S. Robin, E. Lebarbier, and J.-J. Daudin. A segmentation/ clustering model for the analysis of array CGH data. Biometrics, 63(3): 758–766, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qiu:2007:SSC

[1378] Jing Qiu and J. T. Gene Hwang. Sharp simultaneous confidence intervals for the means of selected populations with application to microarray data analysis. Biometrics, 63(3):767–776, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See comment [1744].

Wang:2007:MAM

[1379] Zailong Wang, Shili Lin, Magdalena Popesco, and Andrej Rotter. Modeling and analysis of multi-library, multi-group SAGE data with application to a study of mouse cerebellum. Biometrics, 63(3):777–786, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2007:CCD

[1380] Ming Zheng, Leah O. Barrera, Bing Ren, and Ying Nian Wu. ChIP-chip: Data, model, and analysis. Biometrics, 63(3):787–796, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gupta:2007:GHM

[1381] Mayetri Gupta. Generalized hierarchical Markov models for the discovery of length-constrained sequence features from genome tiling arrays. Biometrics, 63(3):797–805, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

vandeWiel:2007:EFD

[1382] Mark A. van de Wiel and Kyung In Kim. Estimating the false discovery rate using nonparametric deconvolution. Biometrics, 63(3):806–815, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Khanin:2007:SRT

[1383] R. Khanin, V. Vinciotti, V. Mersinias, C. P. Smith, and E. Wit. Statistical reconstruction of transcription factor activity using Michaelis–Menten kinetics. Biometrics, 63(3):816–823, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2007:BMP

[1384] Inyoung Kim, Yin Liu, and Hongyu Zhao. Bayesian methods for predicting interacting protein pairs using domain information. Biometrics, 63(3):824–833, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mukherjee:2007:SBA

[1385] Bhramar Mukherjee, Li Zhang, Malay Ghosh, and Samiran Sinha. Semiparametric Bayesian analysis of case-control data under conditional geneenvironment independence. Biometrics, 63(3):834–844, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2007:SCI

[1386] Yuguo Chen, Ian H. Dinwoodie, and Brenda MacGibbon. Sampling for conditional inference on case-control data. Biometrics, 63(3):845–855, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fan:2007:DPC

[1387] Shenghua K. Fan and You-Gan Wang. Designs for Phase I clinical trials with multiple courses of subjects at different doses. Biometrics, 63(3): 856–864, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2007:SMR

[1388] Yanqiong Zhang, William F. Rosenberger, and Robert T. Smythe. Sequential monitoring of randomization tests: Stratified randomization. Biometrics, 63(3):865–872, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2007:BIL

[1389] Dongfeng Wu, Gary L. Rosner, and Lyle D. Broemeling. Bayesian inference for the lead time in periodic cancer screening. Biometrics, 63(3): 873–880, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nathoo:2007:MMS

[1390] F. Nathoo and C. B. Dean. A mixed mover-stayer model for spatiotemporal two-state processes. Biometrics, 63(3):881–891, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Oman:2007:ASD

[1391] Samuel D. Oman, Victoria Landsman, Yohay Carmel, and Ronen Kadmon. Analyzing spatially distributed binary data using independentblock estimating equations. Biometrics, 63(3):892–900, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2007:MBM

[1392] Zhuoxin Sun, Ori Rosen, and Allan R. Sampson. Multivariate Bernoulli mixture models with application to postmortem tissue studies in schizophrenia. Biometrics, 63(3):901–909, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sutherland:2007:MPE

[1393] Jason M. Sutherland, Carl James Schwarz, and Louis-Paul Rivest. Multilist population estimation with incomplete and partial stratification. Biometrics, 63(3):910–916, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2007:EPS

[1394] Ying Xu, Liping Liu, Na You, Hungyu Pan, and Paul Yip. Estimating population size for a continuous time frailty model with covariates in a capture–recapture study. Biometrics, 63(3):917–921, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Storvik:2007:DDS

[1395] Geir Storvik and Thore Egeland. The DNA database search controversy revisited: Bridging the Bayesian-frequentist gap. Biometrics, 63(3):922– 925, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

TenHave:2007:CMA

[1396] Thomas R. Ten Have, Marshall M. Joffe, Kevin G. Lynch, Gregory K. Brown, Stephen A. Maisto, and Aaron T. Beck. Causal mediation analyses with rank preserving models. Biometrics, 63(3):926–934, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:2007:CTB

[1397] Bing Lu, John S. Preisser, Bahjat F. Qaqish, Chirayath Suchindran, Shrikant I. Bangdiwala, and Mark Wolfson. A comparison of two biascorrected covariance estimators for generalized estimating equations. Biometrics, 63(3):935–941, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fitzmaurice:2007:NPT

[1398] Garrett M. Fitzmaurice, Stuart R. Lipsitz, and Joseph G. Ibrahim. A note on permutation tests for variance components in multilevel generalized linear mixed models. Biometrics, 63(3):942–946, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2007:IAE

[1399] Paul S. Albert. Imputation approaches for estimating diagnostic accuracy for multiple tests from partially verified designs. Biometrics, 63(3):947– 957, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qu:2007:QES

[1400] Yongming Qu and Michael Case. Quantifying the effect of the surrogate marker by information gain. Biometrics, 63(3):958–960, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Alonso:2007:ARF

[1401] A. Alonso, G. Molenberghs, T. Burzykowski, D. Renard, H. Geys, Z. Shkedy, F. Tibaldi, J. Cortinas Abrahantes, and M. Buyse. The authors replied as follows. Biometrics, 63(3):960–962, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Begg:2007:LEB

[1402] Colin B. Begg. Letter to the Editor of Biometrics. Biometrics, 63(3): 964–965, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chatterjee:2007:ARF

[1403] Nilanjan Chatterjee, Zeynep Kalayliouglu, Joanna H. Shih, and Mitchell Gail. The authors replied as follows. Biometrics, 63(3):965–966, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Janssen:2007:BRL

[1404] Arnold Janssen. Book review: Laws of Small Numbers: Extremes and Rare Events, 2nd revised and extended edition edited by M. Falk, J. Huesler, and R.-D. Reiss. Biometrics, 63(3):967–968, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bailey:2007:BRS

[1405] T. C. Bailey. Book review: Statistical Methods in Spatial Epidemiology, 2nd edition by A. B. Lawson. Biometrics, 63(3):968–969, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2007:BRS

[1406] X. Joan Hu. Book review: Statistical Monitoring of Clinical Trials: a Unified Approach, by M. A. Proschan, K. K. G. Lan, and J. T. Wittes. Biometrics, 63(3):969, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chow:2007:BRB

[1407] Shein-Chung Chow. Book review: Bioequivalence Studies in Drug Development, Methods and Applications, by D. Hauschke, V. Steinijans, and I. Pigeot. Biometrics, 63(3):969–970, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Biswas:2007:BRT

[1408] Atanu Biswas. Book review: The Theory of Response-Adaptive Randomization in Clinical Trials, by F. Hu and W. F. Rosenberger. Biometrics, 63(3):970–971, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bohning:2007:BRF

[1409] Dankmar Böhning. Book review: Finite Mixture and Markov Switching Models, by S. Frühwirth-Schnatter. Biometrics, 63(3):971–972, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Flora:2007:BRI

[1410] David B. Flora. Book review: An Introduction to Latent Variable Growth Curve Modeling: Concepts, Issues, and Applications, 2nd edition by T. E. Duncan, S. C. Duncan, and L. A. Strycker. Biometrics, 63(3):972–973, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McLeod:2007:BRM

[1411] Robert G. McLeod. Book review: Modern Experimental Design, by T. P. Ryan. Biometrics, 63(3):973–974, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gomez:2007:BRS

[1412] Guadalupe Gómez. Book review: The Statistical Analysis of Interval-Censored Failure Time Data, by J. Sun. Biometrics, 63(3):974, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwarz:2007:BRD

[1413] Carl James Schwarz. Book review: Data Analysis and Graphics using R, 2nd edition, by J. Maindonald and J. Braun. Biometrics, 63(3):974–975, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joseph:2007:BRB

[1414] Lawrence Joseph. Book review: Bayesian Core: a Practical Approach to Computational Bayesian Statistics, by J. M. Marin and C. P. Robert. Biometrics, 63(3):975–976, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRBb

[1415] Anonymous. Book review: Bayesian Statistical Modelling 2nd edition by P. Congdon. Biometrics, 63(3):976–977, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRSc

[1416] Anonymous. Book review: Structural Equation Modeling and Natural Systems, by J. B. Grace. Biometrics, 63(3):977, September 2007. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRPb

[1417] Anonymous. Book reviews: Principles of Statistical Inference, by D. R. Cox. Modes of Parametric Statistical Inference by S. Geisser. Biometrics, 63(3):977–978, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:DP

[1418] Anonymous. Discussion paper. Biometrics, 63(3):979, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:RAc

[1419] Anonymous. Regular articles. Biometrics, 63(3):979–986, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:CFc

[1420] Anonymous. Consultant’s forum. Biometrics, 63(3):986–987, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:RRb

[1421] Anonymous. Reader reaction. Biometrics, 63(3):987, September 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buckland:2007:LTM

[1422] S. T. Buckland, D. L. Borchers, A. Johnston, P. A. Henrys, and T. A. Marques. Line transect methods for plant surveys. Biometrics, 63(4): 989–998, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rivest:2007:AEC

[1423] Louis-Paul Rivest and Sophie Baillargeon. Applications and extensions of Chao’s moment estimator for the size of a closed population. Biometrics, 63(4):999–1006, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mazzetta:2007:STP

[1424] Chiara Mazzetta, Steve Brooks, and Stephen N. Freeman. On smoothing trends in population index modeling. Biometrics, 63(4):1007–1014, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dupuis:2007:BAM

[1425] Jerome A. Dupuis and Carl James Schwarz. A Bayesian approach to the multistate Jolly–Seber capture–recapture model. Biometrics, 63(4): 1015–1022, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cole:2007:AEG

[1426] D. J. Cole, M. S. Ridout, B. J. T. Morgan, L. J. Byrne, and M. F. Tuite. Approximations for expected generation number. Biometrics, 63(4):1023–1030, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yan:2007:DNC

[1427] Mingjin Yan and Keying Ye. Determining the number of clusters using the weighted gap statistic. Biometrics, 63(4):1031–1037, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Litiere:2007:TTI

[1428] Saskia Litière, Ariel Alonso, and Geert Molenberghs. Type I and Type II error under random-effects misspecification in generalized linear mixed models. Biometrics, 63(4):1038–1044, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Copt:2007:RAT

[1429] Samuel Copt and Stephane Heritier. Robust alternatives to the F -test in mixed linear models based on MM-estimates. Biometrics, 63(4):1045– 1052, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2007:RAP

[1430] Jianguo Sun, Xingwei Tong, and Xin He. Regression analysis of panel count data with dependent observation times. Biometrics, 63(4):1053– 1059, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2007:CMM

[1431] Keunbaik Lee and Michael J. Daniels. A class of Markov models for longitudinal ordinal data. Biometrics, 63(4):1060–1067, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2007:JMP

[1432] Erning Li, Naisyin Wang, and Nae-Yuh Wang. Joint models for a primary endpoint and multiple longitudinal covariate processes. Biometrics, 63 (4):1068–1078, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2007:SRM

[1433] Dawei Liu, Xihong Lin, and Debashis Ghosh. Semiparametric regression of multidimensional genetic pathway data: Least-squares kernel machines and linear mixed models. Biometrics, 63(4):1079–1088, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Owzar:2007:CAD

[1434] Kouros Owzar, Sin-Ho Jung, and Pranab Kumar Sen. A copula approach for detecting prognostic genes associated with survival outcome in microarray studies. Biometrics, 63(4):1089–1098, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2007:CLM

[1435] Jinbo Chen and Carmen Rodriguez. Conditional likelihood methods for haplotype-based association analysis using matched case-control data. Biometrics, 63(4):1099–1107, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gelfond:2007:PME

[1436] Jonathan A. L. Gelfond, Joseph G. Ibrahim, and Fei Zou. Proximity model for expression quantitative trait loci (eQTL) detection. Biometrics, 63(4):1108–1116, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bochkina:2007:TPP

[1437] N. Bochkina and S. Richardson. Tail posterior probability for inference in pairwise and multiclass gene expression data. Biometrics, 63(4):1117– 1125, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2007:NBE

[1438] Yongqiang Tang, Subhashis Ghosal, and Anindya Roy. Nonparametric Bayesian estimation of positive false discovery rates. Biometrics, 63(4): 1126–1134, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ninomiya:2007:CTM

[1439] Yoshiyuki Ninomiya and Hironori Fujisawa. A conservative test for multiple comparison based on highly correlated test statistics. Biometrics, 63(4):1135–1142, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Strassburger:2007:OMC

[1440] K. Strassburger, F. Bretz, and H. Finner. Ordered multiple comparisons with the best and their applications to dose-response studies. Biometrics, 63(4):1143–1151, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2007:CGT

[1441] Hae-Young Kim, Michael G. Hudgens, Jonathan M. Dreyfuss, Daniel J. Westreich, and Christopher D. Pilcher. Comparison of group testing algorithms for case identification in the presence of test error. Biometrics, 63(4):1152–1163, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:2007:CIU

[1442] Paul R. Rosenbaum. Confidence intervals for uncommon but dramatic responses to treatment. Biometrics, 63(4):1164–1171, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2007:PSM

[1443] Ying Yuan and Roderick J. A. Little. Parametric and semiparametric model-based estimates of the finite population mean for two-stage cluster samples with item nonresponse. Biometrics, 63(4):1172–1180, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2007:EPC

[1444] Ying Huang, Margaret Sullivan Pepe, and Ziding Feng. Evaluating the predictiveness of a continuous marker. Biometrics, 63(4):1181–1188, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gupte:2007:MMI

[1445] N. Gupte, R. Brookmeyer, R. Bollinger, and G. Gray. Modeling maternal-infant HIV transmission in the presence of breastfeeding with an imperfect test. Biometrics, 63(4):1189–1197, December 2007. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guedj:2007:MLE

[1446] J. Guedj, R. Thiébaut, and D. Commenges. Maximum likelihood estimation in dynamical models of HIV. Biometrics, 63(4):1198–1206, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2007:APH

[1447] Timothy D. Johnson. Analysis of pulsatile hormone concentration profiles with nonconstant basal concentration: a Bayesian approach. Biometrics, 63(4):1207–1217, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2007:AAV

[1448] Lan Wang and Xiao-Hua Zhou. Assessing the adequacy of variance function in heteroscedastic regression models. Biometrics, 63(4):1218–1225, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2007:SUM

[1449] Yehua Li, Annamaria Guolo, F. Owen Hoffman, and Raymond J. Carroll. Shared uncertainty in measurement error problems, with application to Nevada test site fallout data. Biometrics, 63(4):1226–1236, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dahlberg:2007:PHC

[1450] Suzanne E. Dahlberg and Molin Wang. A proportional hazards cure model for the analysis of time to event with frequently unidentifiable causes. Biometrics, 63(4):1237–1244, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2007:ESS

[1451] Jianliang Dai, Li Li, Sangkyu Kim, Beth Kimball, S. Michal Jazwinski, Jonathan Arnold, and for the Georgia Centenarian Study. Exact sample size needed to detect dependence in 2 × 2 × 2 tables. Biometrics, 63(4): 1245–1252, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2007:TVC

[1452] Duncan Lee and Gavin Shaddick. Time-varying coefficient models for the analysis of air pollution and health outcome data. Biometrics, 63(4): 1253–1261, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shin:2007:JIV

[1453] Yongyun Shin and Stephen W. Raudenbush. Just-identified versus overidentified two-level hierarchical linear models with missing data. Biometrics, 63(4):1262–1268, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Houseman:2007:PIR

[1454] E. Andrés Houseman, Carmen Marsit, Margaret Karagas, and Louise M. Ryan. Penalized item response theory models: Application to epigenetic alterations in bladder cancer. Biometrics, 63(4):1269–1277, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jeske:2007:TPO

[1455] Daniel R. Jeske, Guangyao Wang, and Milton E. McGiffen, Jr. Testing a partial ordering of population means with application to inference about growth habits of cowpea genotypes. Biometrics, 63(4):1278–1282, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gerds:2007:ETM

[1456] Thomas A. Gerds and Martin Schumacher. Efron-type measures of prediction error for survival analysis. Biometrics, 63(4):1283–1287, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2007:PCC

[1457] Jianwen Cai and Donglin Zeng. Power calculation for case-cohort studies with nonrare events. Biometrics, 63(4):1288–1295, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Booth:2007:BRG

[1458] James G. Booth. Book review: Generalized Linear Models with Random Effects: Unified Analysis via H-likelihood, by Y. Lee, J. A. Nelder, and Y. Pawitan. Biometrics, 63(4):1296–1297, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2007:BRL

[1459] Youngjo Lee. Book review: Linear and Generalized Linear Models and their Applications, by J. Jiang. Biometrics, 63(4):1297–1298, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stasinopoulos:2007:BRG

[1460] Mikis Stasinopoulos. Book review: Generalized Additive Models: an Introduction with R, by S. N. Wood. Biometrics, 63(4):1298–1299, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Walls:2007:BRL

[1461] Theodore A. Walls and Moon-Ho Ringo Ho. Book review: Longitudinal Models in the Behavioral and Related Sciences, edited by K. Van Montfort, J. Oud, and A. Satorra. Biometrics, 63(4):1299–1300, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pigeot:2007:BRS

[1462] Iris Pigeot. Book review: Statistical Monitoring of Clinical Trials. Fundamentals for Investigators, by L. A. Moyé. Biometrics, 63(4):1300–1301, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buncher:2007:BRP

[1463] C. Ralph Buncher. Book review: Pharmaceutical Statistics Using SAS — A Practical Guide, edited by A. Dmitrienko, C. Chuang-Stein and R. D’Agostino. Biometrics, 63(4):1301, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:2007:BRS

[1464] John O’Quigley. Book review: Statistical Methods for Dose-Finding Experiments, edited by S. Chevret. Biometrics, 63(4):1301–1302, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hammitt:2007:BRS

[1465] James K. Hammitt. Book review: Statistical Analysis of Cost-Effectiveness Data, by A. R. Willan, and A. H. Briggs. Biometrics, 63(4):1302–1303, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mengersen:2007:BRS

[1466] Kerrie Mengersen. Book review: Structural Equation Modeling: a Bayesian Approach, by S.-Y. Lee. Biometrics, 63(4):1303, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mielke:2007:BRR

[1467] Paul W. Mielke, Jr. Book review: Randomization Tests, by E. S. Edgington and P. Onghena. Biometrics, 63(4):1303–1304, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Suchard:2007:BRA

[1468] Marc A. Suchard. Book review: Analysis of Phylogenetics and Evolution with R, by E. Paradis. Biometrics, 63(4):1304–1305, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meyer:2007:BRR

[1469] Mary Meyer. Book review: Restricted Parameter Space Estimation Problems, by C. Van Eeden. Biometrics, 63(4):1305–1306, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRHb

[1470] Anonymous. Book review: Handbook of Regression and Modeling: Applications for the Clinical and Pharmaceutical Sciences, by D. S. Paulson. Biometrics, 63(4):1306, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRHc

[1471] Anonymous. Book review: Handbook of Parametric and Nonparametric Statistical Procedures, 4th edition by D. J. Sheskin. Biometrics, 63(4): 1306, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRMc

[1472] Anonymous. Book review: Matched Sampling for Causal Effects, by D. R. Rubin. Biometrics, 63(4):1306–1307, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:BRIc

[1473] Anonymous. Book review: An Introduction to Categorical Data Analysis, 2nd edition by A. Agresti. Biometrics, 63(4):1307, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:AT

[1474] Anonymous. Acknowledgement of translators. Biometrics, 63(4):1308, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:RAd

[1475] Anonymous. Regular articles. Biometrics, 63(4):1309–1315, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:CFd

[1476] Anonymous. Consultant’s forum. Biometrics, 63(4):1315–1316, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2007:AR

[1477] Anonymous. Acknowledgement of referees. Biometrics, 63(4):1317–1318, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Freedman:2008:RE

[1478] Laurence Freedman, Geert Molenberghs, Naisyin Wang, David Zucker, and Marie Davidian. Report of the Editors — 2007. Biometrics, 64(1): vii–x, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Janes:2008:MSC

[1479] Holly Janes and Margaret S. Pepe. Matching in studies of classification accuracy: Implications for analysis, efficiency, and assessment of incremental value. Biometrics, 64(1):1–9, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Telesca:2008:ELT

[1480] Donatello Telesca, Ruth Etzioni, and Roman Gulati. Estimating lead time and overdiagnosis associated with PSA screening from prostate cancer incidence trends. Biometrics, 64(1):10–19, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Young:2008:RAD

[1481] Jessica G. Young, Nicholas P. Jewell, and Steven J. Samuels. Regression analysis of a disease onset distribution using diagnosis data. Biometrics, 64(1):20–28, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2008:EEP

[1482] Peng Zhang, Peter X.-K. Song, Annie Qu, and Tom Greene. Efficient estimation for patient-specific rates of disease progression using nonnormal linear mixed models. Biometrics, 64(1):29–38, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2008:WRR

[1483] You-Gan Wang and Yudong Zhao. Weighted rank regression for clustered data analysis. Biometrics, 64(1):39–45, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ninomiya:2008:SMD

[1484] Yoshiyuki Ninomiya and Atsushi Yoshimoto. Statistical method for detecting structural change in the growth process. Biometrics, 64(1):46– 53, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:2008:BMS

[1485] Wesley K. Thompson and Ori Rosen. A Bayesian model for sparse functional data. Biometrics, 64(1):54–63, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baladandayuthapani:2008:BHS

[1486] Veerabhadran Baladandayuthapani, Bani K. Mallick, Mee Young Hong, Joanne R. Lupton, Nancy D. Turner, and Raymond J. Carroll. Bayesian hierarchical spatially correlated functional data analysis with application to colon carcinogenesis. Biometrics, 64(1):64–73, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2008:BAC

[1487] L. Sun and M. K. Clayton. Bayesian analysis of crossclassified spatial data with autocorrelation. Biometrics, 64(1):74–84, March 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2008:EEE

[1488] C. Y. Wang, Yijian Huang, Edward C. Chao, and Marjorie K. Jeffcoat. Expected estimating equations for missing data, measurement error, and misclassification, with application to longitudinal nonignorable missing data. Biometrics, 64(1):85–95, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beunckens:2008:LCM

[1489] Caroline Beunckens, Geert Molenberghs, Geert Verbeke, and Craig Mallinckrodt. A latent-class mixture model for incomplete longitudinal Gaussian data. Biometrics, 64(1):96–105, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinha:2008:AMC

[1490] Samiran Sinha and Tapabrata Maiti. Analysis of matched case-control data in presence of nonignorable missing exposure. Biometrics, 64(1): 106–114, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bondell:2008:SRS

[1491] Howard D. Bondell and Brian J. Reich. Simultaneous regression shrinkage, variable selection, and supervised clustering of predictors with OS-CAR. Biometrics, 64(1):115–123, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2008:SIR

[1492] Lexin Li and Xiangrong Yin. Sliced inverse regression with regularizations. Biometrics, 64(1):124–131, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2008:DPB

[1493] Sijian Wang, Bin Nan, Ji Zhu, and David G. Beer. Doubly penalized Buckley-James method for survival data with high-dimensional covariates. Biometrics, 64(1):132–140, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2008:PHR

[1494] Debashis Ghosh. Proportional hazards regression for cancer studies. Biometrics, 64(1):141–148, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2008:SIS

[1495] Debashis Ghosh. Semiparametric inference for surrogate endpoints with bivariate censored data. Biometrics, 64(1):149–156, March 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jeong:2008:NIM

[1496] Jong-Hyeon Jeong, Sin-Ho Jung, and Joseph P. Costantino. Nonparametric inference on median residual life function. Biometrics, 64(1):157– 163, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Martinussen:2008:MRE

[1497] Torben Martinussen, Odd O. Aalen, and Thomas H. Scheike. The Mizon– Richard encompassing test for the Cox and Aalen additive hazards models. Biometrics, 64(1):164–171, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2008:CRA

[1498] Bingshu E. Chen, Joan L. Kramer, Mark H. Greene, and Philip S. Rosenberg. Competing risks analysis of correlated failure time data. Biometrics, 64(1):172–179, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lakhal:2008:ESA

[1499] Lajmi Lakhal, Louis-Paul Rivest, and Belkacem Abdous. Estimating survival and association in a semicompeting risks model. Biometrics, 64 (1):180–188, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brittain:2008:DCK

[1500] Erica Brittain, Dean Follmann, and Song Yang. Dynamic comparison of Kaplan–Meier proportions: Monitoring a randomized clinical trial with a long-term binary endpoint. Biometrics, 64(1):189–197, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:2008:GST

[1501] Peter B. Gilbert, Chunyuan Wu, and David V. Jobes. Genome scanning tests for comparing amino acid sequences between groups. Biometrics, 64(1):198–207, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jarrett:2008:DAT

[1502] Richard G. Jarrett and Katya Ruggiero. Design and analysis of two-phase experiments for gene expression microarrays — Part I. Biometrics, 64 (1):208–216, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matsui:2008:SSC

[1503] Shigeyuki Matsui, Shu Zeng, Takeharu Yamanaka, and John Shaughnessy. Sample size calculations based on ranking and selection in microarray experiments. Biometrics, 64(1):217–226, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Korn:2008:NCN

[1504] Edward L. Korn and Boris Freidlin. A note on controlling the number of false positives. Biometrics, 64(1):227–231, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Polley:2008:TSD

[1505] Mei-Yin Polley and Ying Kuen Cheung. Two-stage designs for dosefinding trials with a biologic endpoint using stepwise tests. Biometrics, 64(1):232–241, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xi:2008:MCP

[1506] Liqun Xi, Ray Watson, and Paul S. F. Yip. The minimum capture proportion for reliable estimation in capture–recapture models. Biometrics, 64(1):242–249, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Briggs:2008:SPG

[1507] William M. Briggs and Russell Zaretzki. The skill plot: a graphical technique for evaluating continuous diagnostic tests. Biometrics, 64(1): 250–256, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pepe:2008:D

[1508] Margaret Sullivan Pepe. Discussions. Biometrics, 64(1):256–258, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hand:2008:C

[1509] David J. Hand. Comment. Biometrics, 64(1):259, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Briggs:2008:R

[1510] William M. Briggs and Russell Zaretzki. Rejoinder. Biometrics, 64(1): 259–261, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zimmerman:2008:EIS

[1511] Dale L. Zimmerman. Estimating the intensity of a spatial point process from locations coarsened by incomplete geocoding. Biometrics, 64(1): 262–270, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nathoo:2008:SMT

[1512] F. S. Nathoo and C. B. Dean. Spatial multistate transitional models for longitudinal event data. Biometrics, 64(1):271–279, March 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rabe-Hesketh:2008:BMT

[1513] S. Rabe-Hesketh, A. Skrondal, and H. K. Gjessing. Biometrical modeling of twin and family data using standard mixed model software. Biometrics, 64(1):280–288, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Midthune:2008:BRT

[1514] Douglas Midthune, Victor Kipnis, Laurence S. Freedman, and Raymond J. Carroll. Binary regression in truncated samples, with application to comparing dietary instruments in a large prospective study. Biometrics, 64(1):289–298, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhou:2008:IBD

[1515] Yinghui Zhou, John Whitehead, Pasi Korhonen, and Mika Mustonen. Implementation of a Bayesian design in a dose-escalation study of an experimental agent in healthy volunteers. Biometrics, 64(1):299–308, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gumpertz:2008:BRV

[1516] Marcia Gumpertz. Book review: Variations on Split Plot and Split Block Experiment Designs, by W. T. Federer and F. King. Biometrics, 64(1): 309–310, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwarz:2008:BRAa

[1517] Carl James Schwarz. Book review: Analysing Ecological Data, by A. F. Zuur, E. N. Ieno, and G. M. Smith. Biometrics, 64(1):310–311, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drake:2008:BRN

[1518] John M. Drake. Book review: Niche Modeling: Predictions from Statistical Distributions, by D. Stockwell. Biometrics, 64(1):311–312, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brendel:2008:BRD

[1519] Volker Brendel. Book review: DNA, Words and Models-Statistics of Exceptional Words, by S. Robin, F. Rodolphe, and S. Schbath. Biometrics, 64(1):312–313, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holford:2008:BRP

[1520] Nick Holford. Book review: Pharmacometrics: The Science of Quantitative Pharmacology, by E. I. Ette and P. J. Williams. Biometrics, 64(1): 313, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Behr:2008:BRN

[1521] Sigrid Behr. Book review: New Drug Development: Design, Methodology, and Analysis, by J. R. Turner. Biometrics, 64(1):313–314, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Friede:2008:BRA

[1522] Tim Friede. Book review: Adaptive Design Methods in Clinical Trials, by S.-C. Chow and M. Chang. Biometrics, 64(1):314–315, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Austin:2008:BRG

[1523] Matt Austin. Book review: Graphics of Large Datasets: Visualizing a Million, by A. Unwin, M. Theus, and H. Hofmann. Biometrics, 64 (1):315–316, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hand:2008:BRD

[1524] David J. Hand. Book review: Data Mining: Methods and Models, by D. T. Larose. Biometrics, 64(1):316, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gupta:2008:BRS

[1525] Ramesh C. Gupta. Book review: Stochastic Ageing and Dependence for Reliability, by C.-D. Lai and M. Xie. Biometrics, 64(1):317, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loughin:2008:BRN

[1526] Thomas M. Loughin. Book review: The Nature of Statistical Evidence, by B. Thompson. Biometrics, 64(1):317–318, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRIa

[1527] Anonymous. Book review: Introduction to Bioinformatics, by A. Tramontano. Biometrics, 64(1):318, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRAa

[1528] Anonymous. Book review: Applied Nonparametric Statistical Methods, by P. Sprent and N. C. Smeeton. Biometrics, 64(1):318–319, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRR

[1529] Anonymous. Book review: Reliability, Life Testing, and Prediction of Service Lives for Engineers and Scientists, by S. C. Saunders. Biometrics, 64(1):319, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BMa

[1530] Anonymous. Biometric methodology. Biometrics, 64(1):320–326, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BPa

[1531] Anonymous. Biometric practice. Biometrics, 64(1):326–327, March 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2008:RBM

[1532] Yang Yang and Victor DeGruttola. Resampling-based multiple testing methods with covariate adjustment: Application to investigation of antiretroviral drug susceptibility. Biometrics, 64(2):329–336, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Griffin:2008:DAA

[1533] B. A. Griffin and S. W. Lagakos. Design and analysis of arm-in-cage experiments: Inference for three-state progressive disease models with common periodic observation times. Biometrics, 64(2):337–344, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Du:2008:JPS

[1534] Juan Du and Steven N. MacEachern. Judgement post-stratification for designed experiments. Biometrics, 64(2):345–354, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2008:NME

[1535] Xinlei Wang, Johan Lim, and Lynne Stokes. A nonparametric mean estimator for judgment poststratified data. Biometrics, 64(2):355–363, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Royle:2008:MIE

[1536] J. Andrew Royle. Modeling individual effects in the Cormack–Jolly– Seber model: a state-space formulation. Biometrics, 64(2):364–370, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

You:2008:PSE

[1537] Na You and Chang Xuan Mao. Population size estimation in a two-list surveillance system with a discrete covariate. Biometrics, 64(2):371– 376, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borchers:2008:SEM

[1538] D. L. Borchers and M. G. Efford. Spatially explicit maximum likelihood methods for capture–recapture studies. Biometrics, 64(2):377–385, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2008:MME

[1539] Sandra J. Lee and Marvin Zelen. Mortality modeling of early detection programs. Biometrics, 64(2):386–395, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kong:2008:SRS

[1540] Maiying Kong and J. Jack Lee. A semiparametric response surface model for assessing drug interaction. Biometrics, 64(2):396–405, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2008:TSS

[1541] John T. Chen. A two-stage stepwise estimation procedure. Biometrics, 64(2):406–412, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pennell:2008:NBT

[1542] Michael L. Pennell and David B. Dunson. Nonparametric Bayes testing of changes in a response distribution with an ordinal predictor. Biometrics, 64(2):413–423, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choi:2008:BAL

[1543] Taeryon Choi, Mark J. Schervish, Ketra A. Schmitt, and Mitchell J. Small. A Bayesian approach to a logistic regression model with incomplete information. Biometrics, 64(2):424–430, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2008:CMB

[1544] Zheng Yuan and Debashis Ghosh. Combining multiple biomarker models in logistic regression. Biometrics, 64(2):431–439, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2008:VSM

[1545] Sijian Wang and Ji Zhu. Variable selection for model-based highdimensional clustering and its application to microarray data. Biometrics, 64(2):440–448, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2008:EQA

[1546] Huixia Wang and Xuming He. An enhanced quantile approach for assessing differential gene expressions. Biometrics, 64(2):449–457, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vansteelandt:2008:TEG

[1547] Stijn Vansteelandt, Dawn L. DeMeo, Jessica Lasky-Su, Jordan W. Smoller, Amy J. Murphy, Matt McQueen, Kady Schneiter, Juan C. Celedon, Scott T. Weiss, Edwin K. Silverman, and Christoph Lange. Testing and estimating gene-environment interactions in family-based association studies. Biometrics, 64(2):458–467, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gottardo:2008:FPB

[1548] Raphael Gottardo, Wei Li, W. Evan Johnson, and X. Shirley Liu. A flexible and powerful Bayesian hierarchical model for ChIP-Chip experiments. Biometrics, 64(2):468–478, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morris:2008:BAM

[1549] Jeffrey S. Morris, Philip J. Brown, Richard C. Herrick, Keith A. Baggerly, and Kevin R. Coombes. Bayesian analysis of mass spectrometry proteomic data using wavelet-based functional mixed models. Biometrics, 64(2):479–489, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Apanasovich:2008:ACF

[1550] Tatiyana V. Apanasovich, David Ruppert, Joanne R. Lupton, Natasa Popovic, Nancy D. Turner, Robert S. Chapkin, and Raymond J. Carroll. Aberrant crypt foci and semiparametric modeling of correlated binary data. Biometrics, 64(2):490–500, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Datta:2008:SRT

[1551] Somnath Datta and Glen A. Satten. A signed-rank test for clustered data. Biometrics, 64(2):501–507, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xiang:2008:MHD

[1552] Liming Xiang, Kelvin K. W. Yau, Yer Van Hui, and Andy H. Lee. Minimum Hellinger distance estimation for k-component Poisson mixture with random effects. Biometrics, 64(2):508–518, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2008:VCM

[1553] Liang Li and Tom Greene. Varying coefficients model with measurement error. Biometrics, 64(2):519–526, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Albert:2008:MLB

[1554] Paul S. Albert. Modeling longitudinal biomarker data from multiple assays that have different known detection limits. Biometrics, 64(2): 527–537, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roy:2008:GCP

[1555] Jason Roy and Michael J. Daniels. A general class of pattern mixture models for nonignorable dropout with many possible dropout times. Biometrics, 64(2):538–545, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ding:2008:MLD

[1556] Jimin Ding and Jane-Ling Wang. Modeling longitudinal data with nonparametric multiplicative random effects jointly with survival data. Biometrics, 64(2):546–556, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2008:SAJ

[1557] Xiao Song and C. Y. Wang. Semiparametric approaches for joint modeling of longitudinal and survival data with time-varying coefficients. Biometrics, 64(2):557–566, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2008:SSR

[1558] Min Zhang and Marie Davidian. “Smooth” semiparametric regression analysis for arbitrarily censored time-to-event data. Biometrics, 64(2): 567–576, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bacchetti:2008:SDS

[1559] Peter Bacchetti, Charles E. McCulloch, and Mark R. Segal. Simple, defensible sample sizes based on cost efficiency. Biometrics, 64(2):577– 585, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hanley:2008:D

[1560] James A. Hanley and Stanley H. Shapiro. Discussions. Biometrics, 64 (2):586–587, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muller:2008:D

[1561] Peter Müller and Gary L. Rosner. Discussions. Biometrics, 64(2):587– 589, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Simon:2008:D

[1562] Richard Simon. Discussions. Biometrics, 64(2):589–591, June 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bacchetti:2008:R

[1563] Peter Bacchetti, Charles E. McCulloch, and Mark R. Segal. Rejoinder. Biometrics, 64(2):592–594, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morita:2008:DES

[1564] Satoshi Morita, Peter F. Thall, and Peter Müller. Determining the effective sample size of a parametric prior. Biometrics, 64(2):595–602, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schoop:2008:QPP

[1565] R. Schoop, E. Graf, and M. Schumacher. Quantifying the predictive performance of prognostic models for censored survival data with timedependent covariates. Biometrics, 64(2):603–610, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rizopoulos:2008:TPJ

[1566] Dimitris Rizopoulos, Geert Verbeke, Emmanuel Lesaffre, and Yves Vanrenterghem. A two-part joint model for the analysis of survival and longitudinal binary data with excess zeros. Biometrics, 64(2):611–619, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Salway:2008:GGL

[1567] Ruth Salway and Jon Wakefield. Gamma generalized linear models for pharmacokinetic data. Biometrics, 64(2):620–626, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hedeker:2008:AME

[1568] Donald Hedeker, Robin J. Mermelstein, and Hakan Demirtas. An application of a mixed-effects location scale model for analysis of ecological momentary assessment (EMA) data. Biometrics, 64(2):627–634, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dorazio:2008:MUS

[1569] Robert M. Dorazio, Bhramar Mukherjee, Li Zhang, Malay Ghosh, Howard L. Jelks, and Frank Jordan. Modeling unobserved sources of heterogeneity in animal abundance using a Dirichlet process prior. Biometrics, 64(2):635–644, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanderWeele:2008:SAD

[1570] Tyler J. VanderWeele. Sensitivity analysis: Distributional assumptions and confounding assumptions. Biometrics, 64(2):645–649, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bang:2008:CDR

[1571] Heejung Bang and James M. Robins. Correction to “Doubly Robust Estimation in Missing Data and Causal Inference Models,” by H. Bang and J. M. Robins; 61, 962–972, December 2005. Biometrics, 64(2):650, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [991].

Bailer:2008:BRU

[1572] A. John Bailer. Book review: Useless Arithmetic: Why Environmental Scientists Can’t Predict the Future, by Pilkey, O. H. and Pilkey-Jarvis, L. Biometrics, 64(2):651–652, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Barabesi:2008:BRS

[1573] Lucio Barabesi. Book review: Sampling Strategies for Natural Resources and the Environment, by Gregoire, T. G. and Valentine, H. T. Biometrics, 64(2):652–653, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Banerjee:2008:BRM

[1574] Sudipto Banerjee. Book review: Model-Based Geostatistics, by Diggle, P. J. and Ribeiro, P. J. Biometrics, 64(2):653–654, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Paradis:2008:BRR

[1575] Emmanuel Paradis. Book review: Reconstructing Evolution: New Mathematical and Computational Advances, edited by Gascuel, O. and Steel, M. Biometrics, 64(2):654, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rao:2008:BRB

[1576] J. Sunil Rao. Book review: Bayesian Inference for Gene Expression and Proteomics, by Do, K.-A., Müller, P., and Vannucci, M. Biometrics, 64 (2):654–655, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hubert:2008:BRD

[1577] Mia Hubert. Book review: Data Depth: Robust Multivariate Analysis, Computational Geometry and Applications, by Liu, R., Serfling, R., and Souvaine, D. L. Biometrics, 64(2):655–656, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Celeux:2008:BRD

[1578] Gilles Celeux. Book review: Data Clustering: Theory Algorithms and Applications, by Gan, G., Chaoqun, M. A., and Wu, J. Biometrics, 64 (2):656–657, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ryan:2008:BRR

[1579] Louise M. Ryan. Book review: Reliability and Risk: a Bayesian Perspective, by Singpurwalla, N. D. Biometrics, 64(2):657–658, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chappell:2008:BRT

[1580] Rick Chappell. Book review: Textbook of Clinical Trials, by Machin, D., Day, S., and Green, S. Biometrics, 64(2):658, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steiner:2008:BRS

[1581] Stefan Steiner. Book review: Statistical Development of Quality in Medicine, by Winkel, P. and Zhang, N. F. Biometrics, 64(2):658–659, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2008:BRM

[1582] Liansheng Tang. Book review: Model-Based Geostatistics, by Gönen, M. Biometrics, 64(2):659, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Waldorp:2008:BRA

[1583] Lourens Waldorp. Book review: Applied Asymptotics: Case Studies in Small Sample Statistics, by Brazzale, A. R., Davison, A. C., and Reid, N. Biometrics, 64(2):660, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRAb

[1584] Anonymous. Book review: Analysis of Correlated Data with SAS and R, 3rd edition. by Shoukri, M. M. and Chaudhary, M. A. Biometrics, 64(2):660–661, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRE

[1585] Anonymous. Book review: Elementary Bayesian Statistics, by Moyé, L. A. Biometrics, 64(2):661, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRIb

[1586] Anonymous. Book review: Information Criteria and Statistical Modeling, by Konishi, S. and Kitagawa, G. Biometrics, 64(2):661, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRMa

[1587] Anonymous. Book review: Modelling for Field Biologists and Other Interesting People, by Kokko, H. Biometrics, 64(2):662, June 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BMb

[1588] Anonymous. Biometric methodology. Biometrics, 64(2):663–669, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BPb

[1589] Anonymous. Biometric practice. Biometrics, 64(2):669, June 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:CF

[1590] Anonymous. Consultant’s forum. Biometrics, 64(2):669–671, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:RRa

[1591] Anonymous. Reader reaction. Biometrics, 64(2):671, June 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lobach:2008:HBR

[1592] Iryna Lobach, Raymond J. Carroll, Christine Spinka, Mitchell H. Gail, and Nilanjan Chatterjee. Haplotype-based regression analysis and inference of case-control studies with unphased genotypes and measurement errors in environmental exposures. Biometrics, 64(3):673–684, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mukherjee:2008:EGE

[1593] Bhramar Mukherjee and Nilanjan Chatterjee. Exploiting geneenvironment independence for analysis of case-control studies: an empirical Bayes-type shrinkage estimator to trade-off between bias and efficiency. Biometrics, 64(3):685–694, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2008:BDE

[1594] Zhihong Cai, Manabu Kuroki, Judea Pearl, and Jin Tian. Bounds on direct effects in the presence of confounded intermediate variables. Biometrics, 64(3):695–701, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanderWeele:2008:SBU

[1595] Tyler J. VanderWeele. The sign of the bias of unmeasured confounding. Biometrics, 64(3):702–706, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2008:IEI

[1596] Min Zhang, Anastasios A. Tsiatis, and Marie Davidian. Improving efficiency of inferences in randomized clinical trials using auxiliary covariates. Biometrics, 64(3):707–715, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lloyd:2008:NEM

[1597] Chris J. Lloyd. A new exact and more powerful unconditional test of no treatment effect from binary matched pairs. Biometrics, 64(3):716–723, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wei:2008:ECT

[1598] Guanghui Wei and Douglas E. Schaubel. Estimating cumulative treatment effects in the presence of nonproportional hazards. Biometrics, 64(3):724–732, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Logan:2008:CTP

[1599] Brent R. Logan, John P. Klein, and Mei-Jie Zhang. Comparing treatments in the presence of crossing survival curves: an application to bone marrow transplantation. Biometrics, 64(3):733–740, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2008:RCA

[1600] Rui Song, Michael R. Kosorok, and Jianwen Cai. Robust covariateadjusted log-rank statistics and corresponding sample size formula for recurrent events data. Biometrics, 64(3):741–750, September 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nielsen:2008:CMN

[1601] J. D. Nielsen and C. B. Dean. Clustered mixed nonhomogeneous Poisson process spline models for the analysis of recurrent event panel data. Biometrics, 64(3):751–761, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Elashoff:2008:JML

[1602] Robert M. Elashoff, Gang Li, and Ning Li. A joint model for longitudinal measurements and survival data in the presence of multiple failure types. Biometrics, 64(3):762–771, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goetgeluk:2008:CGE

[1603] Sylvie Goetgeluk and Stijn Vansteelandt. Conditional generalized estimating equations for the analysis of clustered and longitudinal data. Biometrics, 64(3):772–780, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ma:2008:IKL

[1604] Yan Ma, Wan Tang, Changyong Feng, and Xin M. Tu. Inference for kappas for longitudinal study data: Applications to sexual health research. Biometrics, 64(3):781–789, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reich:2008:MLS

[1605] Brian J. Reich and James S. Hodges. Modeling longitudinal spatial periodontal data: a spatially adaptive model with tools for specifying priors and checking fit. Biometrics, 64(3):790–799, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guan:2008:KTS

[1606] Yongtao Guan. A KPSS test for stationarity for spatial point processes. Biometrics, 64(3):800–806, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zucchini:2008:MTS

[1607] Walter Zucchini, David Raubenheimer, and Iain L. MacDonald. Modeling time series of animal behavior by means of a latent-state model with feedback. Biometrics, 64(3):807–815, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2008:BEC

[1608] R. King and S. P. Brooks. On the Bayesian estimation of a closed population size in the presence of heterogeneity and model uncertainty. Biometrics, 64(3):816–824, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Branscum:2008:BNM

[1609] Adam J. Branscum and Timothy E. Hanson. Bayesian nonparametric meta-analysis using Polya tree mixture models. Biometrics, 64(3):825– 833, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Massonnet:2008:FCS

[1610] Goele Massonnet, Paul Janssen, and Tomasz Burzykowski. Fitting conditional survival models to meta-analytic data by using a transformation toward mixed-effects models. Biometrics, 64(3):834–842, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hubbard:2008:MNM

[1611] R. A. Hubbard, L. Y. T. Inoue, and J. R. Fann. Modeling nonhomogeneous Markov processes via time transformation. Biometrics, 64(3): 843–850, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Drovandi:2008:MMP

[1612] C. C. Drovandi and A. N. Pettitt. Multivariate Markov process models for the transmission of methicillin-resistant staphylococcus aureus in a hospital Ward. Biometrics, 64(3):851–859, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2008:OOT

[1613] Alex R. Cook, Gavin J. Gibson, and Christopher A. Gilligan. Optimal observation times in experimental epidemic processes. Biometrics, 64(3): 860–868, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lam:2008:TCF

[1614] K. F. Lam, J. V. Deshpande, E. H. Y. Lau, U. V. Naik-Nimbalkar, P. S. F. Yip, and Ying Xu. A test for constant fatality rate of an emerging epidemic: With applications to severe acute respiratory syndrome in Hong Kong and Beijing. Biometrics, 64(3):869–876, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peng:2008:SCI

[1615] Jianan Peng, Chu-In Charles Lee, Karelyn A. Davis, and Weizhen Wang. Stepwise confidence intervals for monotone dose-response studies. Biometrics, 64(3):877–885, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ding:2008:BOD

[1616] Meichun Ding, Gary L. Rosner, and Peter Müller. Bayesian optimal design for Phase II screening trials. Biometrics, 64(3):886–894, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vexler:2008:MLR

[1617] Albert Vexler, Aiyi Liu, Ekaterina Eliseeva, and Enrique F. Schisterman. Maximum likelihood ratio tests for comparing the discriminatory ability of biomarkers subject to limit of detection. Biometrics, 64(3):895–903, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2008:ESI

[1618] Runze Li and Lei Nie. Efficient statistical inference procedures for partially nonlinear models and their applications. Biometrics, 64(3):904– 911, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Deschepper:2008:TDP

[1619] E. Deschepper, O. Thas, and J. P. Ottoy. Tests and diagnostic plots for detecting lack-of-fit for circular-linear regression models. Biometrics, 64(3):912–920, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xie:2008:VSP

[1620] Benhuai Xie, Wei Pan, and Xiaotong Shen. Variable selection in penalized model-based clustering via regularization on grouped parameters. Biometrics, 64(3):921–930, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2008:OPD

[1621] Sang Han Lee, Johan Lim, Marina Vannucci, Eva Petkova, Maurice Preter, and Donald F. Klein. Order-preserving dimension reduction procedure for the dominance of two mean curves with application to tidal volume curves. Biometrics, 64(3):931–939, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheung:2008:SSB

[1622] Ying Kuen Cheung. Simple sequential boundaries for treatment selection in multi-armed randomized clinical trials with a control. Biometrics, 64(3):940–949, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2008:ALD

[1623] Lei Liu, Xuelin Huang, and John O’Quigley. Analysis of longitudinal data in the presence of informative observational times and a dependent terminal event, with application to medical cost data. Biometrics, 64(3): 950–958, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2008:EPP

[1624] Jiguo Cao, Gregor F. Fussmann, and James O. Ramsay. Estimating a predator-prey dynamical model with the parameter cascades method.

Biometrics, 64(3):959–967, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2008:GFA

[1625] Devin S. Johnson, Dana L. Thomas, Jay M. Ver Hoef, and Aaron Christ. A general framework for the analysis of animal resource selection from telemetry data. Biometrics, 64(3):968–976, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2008:NPS

[1626] Chang Xuan Mao. On the nonidentifiability of population sizes. Biometrics, 64(3):977–979, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holzmann:2008:ARF

[1627] Hajo Holzmann and Axel Munk. The authors replied as follows. Biometrics, 64(3):979–981, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bernard-Michel:2008:NSI

[1628] C. Bernard-Michel, L. Gardes, and S. Girard. A note on sliced inverse regression with regularizations. Biometrics, 64(3):982–984, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2008:ARF

[1629] Lexin Li and Xiangrong Yin. The authors replied as follows. Biometrics, 64(3):984–986, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2008:BRS

[1630] Chiung-Yu Huang and Xianghua Luo. Book review: The Statistical Analysis of Recurrent Events, by Cook, R. J. and Lawless, J. F. Biometrics, 64(3):987–988, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2008:BRI

[1631] Changyong Feng. Book review: Introduction to Empirical Processes and Semiparametric Inference, by Kosorok, M. R. Biometrics, 64(3):988–989, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dickey:2008:BRU

[1632] James M. Dickey. Book review: Uncertain Judgements: Eliciting Experts’ Probabilities, by O’Hagan, A, Buck, C. E., Daneshkhah, A., Eiser, J. R., Garthwaite, P. H., Jenkinson, D. J., Oakley, J. E., and Rakow, T. Biometrics, 64(3):989–990, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Farewell:2008:BRM

[1633] Daniel Farewell. Book review: Models for Intensive Longitudinal Data, edited by Walls, T. A. and Schafer, J. L. Biometrics, 64(3):990–991, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartell:2008:BRQ

[1634] Scott M. Bartell. Book review: Quantitative Environmental Risk Analysis for Human Health, by Fjeld, R. A., Eisenberg, N. A., and Compton, K. L. Biometrics, 64(3):991–992, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schulte:2008:BRR

[1635] Oliver Schulte. Book review: Reliable Reasoning: Induction and Statistical Learning Theory, by Harman, G. and Kulkarni, S. Biometrics, 64(3): 992–993, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2008:BRM

[1636] Carol Y. Lin. Book review: Modeling Infectious Diseases in Humans and Animals, by Keeling, M. J. and Rohani, P. Biometrics, 64(3):993, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Leese:2008:BRC

[1637] Morven Leese. Book review: Cluster and Classification Techniques for the Biosciences, by Fielding, A. H. Biometrics, 64(3):993–994, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frey:2008:BRM

[1638] Jesse Frey. Book review: Modern Applied U-Statistics, by Kowalski, J. and Tu, X. M. Biometrics, 64(3):994–995, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bammann:2008:BRH

[1639] Karin Bammann, Ronja Foraita, and Frauke Günther. Book review: Handbook of Statistical Genetics, edited by Balding, D. J., Bishop, M., and Cannings, C. Biometrics, 64(3):995, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sloan:2008:BRQ

[1640] Jeff A. Sloan, Amylou Dueck, Rui Qin, Wenting Wu, Pamela J. Atherton, Paul Novotny, Heshan Liu, Kelli N. Burger, Angelina D. Tan, Daniel Szydlo, Victor M. Johnson, Sara J. Felten, Xinghua Zhao, and Brent Diekmann. Book review: Quality of Life: The Assessment, Analysis, and Interpretation of Patient-Reported Outcomes, by Fayers, P. M. and Machin, D. Biometrics, 64(3):996, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwarz:2008:BRAb

[1641] Carl James Schwarz. Book review: Advanced Distance Sampling: Estimating Abundance of Biological Populations, by Buckland, S. T., Anderson, D. R., Burnham, K. P., Laake, J. L., Borchers, C. L., and Thomas, L. Biometrics, 64(3):997, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRC

[1642] Anonymous. Book review: The Construction of Optimal Stated Choice Experiments: Theory and Methods, by Street, D. J. and Burgess, L. Biometrics, 64(3):997–998, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRAc

[1643] Anonymous. Book review: Asymptotic Theory of Statistics and Probability, by Dasgupta, A. Biometrics, 64(3):998, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRD

[1644] Anonymous. Book review: Design and Analysis of Experiments, Volume 1: Introduction to Experimental Design, by Hinkelmann, K. and Kempthorne, O. Biometrics, 64(3):998, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRIc

[1645] Anonymous. Book review: Introduction to Bayesian Statistics, by Bolstad, W. M. Biometrics, 64(3):998, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRB

[1646] Anonymous. Book review: Bootstrap Methods: a Guide for Practitioners and Researchers, by Chernick, M. R. Biometrics, 64(3):998–999, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BMc

[1647] Anonymous. Biometric methodology. Biometrics, 64(3):1000–1007, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BPc

[1648] Anonymous. Biometric practice. Biometrics, 64(3):1007–1008, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:RRb

[1649] Anonymous. Reader reaction. Biometrics, 64(3):1008, September 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2008:ICE

[1650] Debashis Ghosh, Moulinath Banerjee, and Pinaki Biswas. Inference for constrained estimation of tumor size distributions. Biometrics, 64(4): 1009–1017, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ostrovnaya:2008:CPT

[1651] Irina Ostrovnaya, Venkatraman E. Seshan, and Colin B. Begg. Comparison of properties of tests for assessing tumor clonality. Biometrics, 64(4):1018–1022, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krafty:2008:VCM

[1652] Robert T. Krafty, Phyllis A. Gimotty, David Holtz, George Coukos, and Wensheng Guo. Varying coefficient model with unknown within-subject covariance for analysis of tumor growth curves. Biometrics, 64(4):1023– 1031, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2008:STM

[1653] Haiqun Lin, Zhenchao Guo, Peter N. Peduzzi, Thomas M. Gill, and Heather G. Allore. A semiparametric transition model with latent traits for longitudinal multistate data. Biometrics, 64(4):1032–1042, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Escaramis:2008:DSD

[1654] Geòrgia Escaramís, Josep L. Carrasco, and Carlos Ascaso. Detection of significant disease risks using a spatial conditional autoregressive model. Biometrics, 64(4):1043–1053, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staniswalis:2008:IMC

[1655] Joan G. Staniswalis. Incorporating marginal covariate information in a nonparametric regression model for a sample of R×C tables. Biometrics, 64(4):1054–1061, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Claeskens:2008:VSI

[1656] Gerda Claeskens and Fabrizio Consentino. Variable selection with incomplete covariate data. Biometrics, 64(4):1062–1069, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sun:2008:TET

[1657] Yanqing Sun, Seunggeun Hyun, and Peter Gilbert. Testing and estimation of time-varying cause-specific hazard ratios with covariate adjustment. Biometrics, 64(4):1070–1079, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Peng:2008:NTC

[1658] L. Peng and J. P. Fine. Nonparametric tests for continuous covariate effects with multistate survival data. Biometrics, 64(4):1080–1089, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2008:RSA

[1659] Xuelin Huang and Nan Zhang. Regression survival analysis with an assumed copula for dependent censoring: a sensitivity analysis approach.

Biometrics, 64(4):1090–1099, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2008:MMB

[1660] Richard J. Cook, Leilei Zeng, and Ker-Ai Lee. A multistate model for bivariate interval-censored failure time data. Biometrics, 64(4):1100– 1109, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schacht:2008:NNA

[1661] Alexander Schacht, Kris Bogaerts, Erich Bluhmki, and Emmanuel Lesaffre. A new nonparametric approach for baseline covariate adjustment for two-group comparative studies. Biometrics, 64(4):1110–1116, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pei:2008:SFQ

[1662] Lixia Pei and Michael D. Hughes. A statistical framework for quantile equivalence clinical trials with application to pharmacokinetic studies that bridge from HIV-infected adults to children. Biometrics, 64(4): 1117–1125, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:2008:PSD

[1663] Peter F. Thall, Hoang Q. Nguyen, and Elihu H. Estey. Patient-specific dose finding based on bivariate outcomes and covariates. Biometrics, 64(4):1126–1136, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tang:2008:NSG

[1664] Liansheng Tang, Scott S. Emerson, and Xiao-Hua Zhou. Nonparametric and semiparametric group sequential methods for comparing accuracy of diagnostic tests. Biometrics, 64(4):1137–1145, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilbert:2008:ECP

[1665] Peter B. Gilbert and Michael G. Hudgens. Evaluating candidate principal surrogate endpoints. Biometrics, 64(4):1146–1154, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2008:RBU

[1666] Zhenqiu Liu and Ming Tan. ROC-based utility function maximization for feature selection and classification with applications to high-dimensional protease data. Biometrics, 64(4):1155–1161, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bogdan:2008:EMB

[1667] Małgorzata Bogdan, Florian Frommlet, Przemysław Biecek, Riyan Cheng, Jayanta K. Ghosh, and R. W. Doerge. Extending the modified Bayesian information criterion (mBIC) to dense markers and multiple interval mapping. Biometrics, 64(4):1162–1169, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Conn:2008:BAW

[1668] Paul B. Conn, Duane R. Diefenbach, Jeffrey L. Laake, Mark A. Ternent, and Gary C. White. Bayesian analysis of wildlife age-at-harvest data. Biometrics, 64(4):1170–1177, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chao:2008:TSP

[1669] Anne Chao, Lou Jost, S. C. Chiang, Y.-H. Jiang, and Robin L. Chazdon. A two-stage probabilistic approach to multiple-community similarity indices. Biometrics, 64(4):1178–1186, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

King:2008:ACC

[1670] R. King, S. P. Brooks, and T. Coulson. Analyzing complex capture– recapture data in the presence of individual and temporal covariates and model uncertainty. Biometrics, 64(4):1187–1195, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miller:2008:FSC

[1671] Timothy J. Miller and Per K. Andersen. A finite-state continuoustime approach for inferring regional migration and mortality rates from archival tagging and conventional tag-recovery experiments. Biometrics, 64(4):1196–1206, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guolo:2008:FAM

[1672] A. Guolo. A flexible approach to measurement error correction in casecontrol studies. Biometrics, 64(4):1207–1214, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Benjamini:2008:SPC

[1673] Yoav Benjamini and Ruth Heller. Screening for partial conjunction hypotheses. Biometrics, 64(4):1215–1222, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2008:BEI

[1674] Bo Hu, Yuan Ji, and Kam-Wah Tsui. Bayesian estimation of inverse dose response. Biometrics, 64(4):1223–1230, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:2008:BRE

[1675] Grace Hui-Min Wu, Shu-Hui Chang, and Tony Hsiu-Hsi Chen. A Bayesian random-effects Markov model for tumor progression in women with a family history of breast cancer. Biometrics, 64(4):1231–1237, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ye:2008:SML

[1676] Wen Ye, Xihong Lin, and Jeremy M. G. Taylor. Semiparametric modeling of longitudinal measurements and time-to-event data — a two-stage regression calibration approach. Biometrics, 64(4):1238–1246, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2008:WLM

[1677] Zhiguo Li, Peter Gilbert, and Bin Nan. Weighted likelihood method for grouped survival data in case-cohort studies with application to HIV vaccine trials. Biometrics, 64(4):1247–1255, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heo:2008:SPS

[1678] Moonseong Heo and Andrew C. Leon. Statistical power and sample size requirements for three level hierarchical cluster randomized trials. Biometrics, 64(4):1256–1262, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hothorn:2008:GMS

[1679] Torsten Hothorn and Achim Zeileis. Generalized maximally selected statistics. Biometrics, 64(4):1263–1269, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:2008:SCI

[1680] Alan Agresti, Matilde Bini, Bruno Bertaccini, and Euijung Ryu. Simultaneous confidence intervals for comparing binomial parameters. Biometrics, 64(4):1270–1275, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zheng:2008:AOC

[1681] Gang Zheng. Analysis of ordered categorical data: Two scoreindependent approaches. Biometrics, 64(4):1276–1279, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2008:CTC

[1682] Yi Li and Ram C. Tiwari. Comparing trends in cancer rates across overlapping regions. Biometrics, 64(4):1280–1286, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Commenges:2008:ADS

[1683] Daniel Commenges and Benoit Liquet. Asymptotic distribution of score statistics for spatial cluster detection with censored data. Biometrics, 64(4):1287–1289, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cook:2008:ARF

[1684] Andrea J. Cook and Yi Li. The authors replied as follow. Biometrics, 64(4):1289–1292, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tuyl:2008:IPC

[1685] Frank Tuyl, Richard Gerlach, and Kerrie Mengersen. Inference for proportions in a 2 × 2 contingency table: HPD or not HPD? Biometrics, 64(4):1293–1295, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:2008:ARF

[1686] Alan Agresti and Yongyi Min. The authors replied as follow. Biometrics, 64(4):1295–1296, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2008:LEB

[1687] Changyong Feng. Letter to the Editor of Biometrics. Biometrics, 64 (4):1297, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2008:ARF

[1688] Bingshu E. Chen, Joan L. Kramer, Mark H. Greene, and Philip S. Rosenberg. The authors replied as follows. Biometrics, 64(4):1297–1298, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Topaz:2008:BRC

[1689] Chad Topaz. Book review: Continuum Modeling in the Physical Sciences, by Van Groesen, E. and Molenaar, J. Biometrics, 64(4):1299–1300, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gueron:2008:BRR

[1690] Shay Gueron. Book review: Random Fragmentation and Coagulation Processes, by Bertoin, J. Biometrics, 64(4):1300–1301, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2008:BRC

[1691] Mei-Jie Zhang. Book review: Continuous Semi-Markov Process, by Harlamov, B. Biometrics, 64(4):1301, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Emerson:2008:BRI

[1692] John W. Emerson. Book review: Interactive and Dynamic Graphics for Data Analysis: With R and GGobi, by Cook, D. and Swayne, D. Biometrics, 64(4):1301–1303, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bai:2008:BRI

[1693] Z. D. Bai. Book review: Inference and Prediction in Large Dimensions, by Bosc, D. and Blanke, D. Biometrics, 64(4):1303–1304, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vahl:2008:BRS

[1694] Christopher I. Vahl. Book review: Statistical Design, by Casella, G. Biometrics, 64(4):1304–1305, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kachman:2008:BRI

[1695] Stephen D. Kachman. Book reviews: Introduction to Mixed Modelling: Beyond Regression and Analysis of Variance, by Galwey, N. W. Linear Mixed Models: a Practical Guide Using Statistical Software by West, B. T., Welch, K. B., and Galecki, A. T. Biometrics, 64(4):1305, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Woolson:2008:BRS

[1696] Robert F. Woolson. Book review: Statistical Advances in the Biomedical Sciences: Clinical Trials, Epidemiology, Survival Analysis, and Bioinformatics, edited by Biswas, A., Datta, S., Fine, J. P., and Segal, M. R. Biometrics, 64(4):1305–1306, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Day:2008:BRI

[1697] Simon Day. Book review: Introduction to Statistical Methods for Clinical Trials, by Cook, T. D. and DeMets, D. L. Biometrics, 64(4):1306–1307, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glueck:2008:BRS

[1698] D. H. Glueck. Book review: Sample Size Calculations in Clinical Research, 2nd edition by Chow, S.-C., Shao, J., and Wang, H. Biometrics, 64(4):1307–1308, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gentle:2008:BRS

[1699] James E. Gentle. Book review: Simulation and the Monte Carlo Method, 2nd edition by Rubinsteina, R. Y. and Kroese, D. P. Biometrics, 64(4): 1308, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tseng:2008:BRG

[1700] George C. Tseng. Book review: Genomic Signal Processing, by Shmulevich, I. and Dougherty, E. R. Biometrics, 64(4):1308–1309, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRMb

[1701] Anonymous. Book review: Model Based Inference in the Life Sciences: a Primer on Evidence, by Anderson, D. A. Biometrics, 64(4):1309, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRMc

[1702] Anonymous. Book review: Medical Statistics: a Textbook for the Health Sciences, 4th edition by Campbell, M. J., Machin, D., and Walters, S. J.

Biometrics, 64(4):1309–1310, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRH

[1703] Anonymous. Book review: Handbook of Hidden Markov Models in Bioinformatics, by Gollery, M. Biometrics, 64(4):1310, December 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRMd

[1704] Anonymous. Book review: Multilevel and Longitudinal Modeling Using Stata, 2nd edition by Rabe-Hesketh, S. and Skrondal, A. Biometrics, 64 (4):1310, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BMd

[1705] Anonymous. Biometric methodology. Biometrics, 64(4):1312–1318, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BPd

[1706] Anonymous. Biometric practice. Biometrics, 64(4):1318–1320, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:RRc

[1707] Anonymous. Reader reaction. Biometrics, 64(4):1320, December 2008. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Molenberghs:2009:RE

[1708] Geert Molenberghs, Naisyin Wang, David Zucker, and Marie Davidian. Report of the Editors — 2008. Biometrics, 65(1):vii–x, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mead:2009:BIS

[1709] Andrew Mead. Biometry and the IBS — strength through diversity. Biometrics, 65(1):1–8, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2009:IRL

[1710] Shili Lin and Jie Ding. Integration of ranked lists via cross entropy Monte Carlo with applications to mRNA and microRNA studies. Biometrics, 65(1):9–18, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

vanWieringen:2009:NTD

[1711] Wessel N. van Wieringen and Mark A. van de Wiel. Nonparametric testing for DNA copy number induced differential mRNA gene expression. Biometrics, 65(1):19–29, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yang:2009:NFM

[1712] Jie Yang, Rongling Wu, and George Casella. Nonparametric functional mapping of quantitative trait loci. Biometrics, 65(1):30–39, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tai:2009:GRU

[1713] Yu Chuan Tai and Terence P. Speed. On gene ranking using replicated microarray time course data. Biometrics, 65(1):40–51, March 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qin:2009:RML

[1714] Guo You Qin and Zhong Yi Zhu. Robustified maximum likelihood estimation in generalized partial linear mixed model for longitudinal data. Biometrics, 65(1):52–59, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Song:2009:JRA

[1715] Peter X.-K. Song, Mingyao Li, and Ying Yuan. Joint regression analysis of correlated data using Gaussian copulas. Biometrics, 65(1):60–68, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marshall:2009:DAL

[1716] Guillermo Marshall, Rolando De la Cruz-Mesía, Fernando A. Quintana, and Anna E. Barón. Discriminant analysis for longitudinal data with multiple continuous responses and possibly missing data. Biometrics, 65(1):69–80, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsonaka:2009:SPS

[1717] Roula Tsonaka, Geert Verbeke, and Emmanuel Lesaffre. A semiparametric shared parameter model to handle nonmonotone nonignorable missingness. Biometrics, 65(1):81–87, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:2009:MIM

[1718] L. Taylor and X. H. Zhou. Multiple imputation methods for treatment noncompliance and nonresponse in randomized clinical trials. Biometrics, 65(1):88–95, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2009:EIC

[1719] Jing Cheng. Estimation and inference for the causal effect of receiving treatment on a multinomial outcome. Biometrics, 65(1):96–103, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sanchez:2009:RBD

[1720] B. N. Sánchez, E. A. Houseman, and L. M. Ryan. Residual-based diagnostics for structural equation models. Biometrics, 65(1):104–115, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cho:2009:BCI

[1721] Hyunsoon Cho, Joseph G. Ibrahim, Debajyoti Sinha, and Hongtu Zhu. Bayesian case influence diagnostics for survival models. Biometrics, 65 (1):116–124, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guo:2009:MAM

[1722] Ying Guo and Amita K. Manatunga. Measuring agreement of multivariate discrete survival times using a modified weighted kappa coefficient. Biometrics, 65(1):125–134, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kong:2009:CCA

[1723] Lan Kong and Jianwen Cai. Case-cohort analysis with accelerated failure time model. Biometrics, 65(1):135–142, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Frydman:2009:NEM

[1724] Halina Frydman and Michael Szarek. Nonparametric estimation in a Markov “illness–death” process from interval censored observations with missing intermediate transition status. Biometrics, 65(1):143–151, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chiang:2009:EOC

[1725] Chin-Tsang Chiang and Shr-Yan Huang. Estimation for the optimal combination of markers without modeling the censoring distribution. Biometrics, 65(1):152–158, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Qiao:2009:AWL

[1726] Xingye Qiao and Yufeng Liu. Adaptive weighted learning for unbalanced multicategory classification. Biometrics, 65(1):159–168, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bondell:2009:SFS

[1727] Howard D. Bondell and Brian J. Reich. Simultaneous factor selection and collapsing levels in ANOVA. Biometrics, 65(1):169–177, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dai:2009:SEE

[1728] James Y. Dai, Michael LeBlanc, and Charles Kooperberg. Semiparametric estimation exploiting covariate independence in two-phase randomized trials. Biometrics, 65(1):178–187, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosner:2009:PSS

[1729] B. Rosner and R. J. Glynn. Power and sample size estimation for the Wilcoxon rank sum test with application to comparisons of c statistics from alternative prediction models. Biometrics, 65(1):188–197, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bornkamp:2009:BNE

[1730] Björn Bornkamp and Katja Ickstadt. Bayesian nonparametric estimation of continuous monotone functions with applications to dose-response analysis. Biometrics, 65(1):198–205, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Calder:2009:RSM

[1731] C. A. Calder, P. F. Craigmile, and J. Zhang. Regional spatial modeling of topsoil geochemistry. Biometrics, 65(1):206–215, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pawlas:2009:PSS

[1732] Zbynek Pawlas, Jens R. Nyengaard, and Eva B. Vedel Jensen. Particle sizes from sectional data. Biometrics, 65(1):216–224, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fewster:2009:EER

[1733] Rachel M. Fewster, Stephen T. Buckland, Kenneth P. Burnham, David L. Borchers, Peter E. Jupp, Jeffrey L. Laake, and Len Thomas. Estimating the encounter rate variance in distance sampling. Biometrics, 65(1):225– 236, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McClintock:2009:EAU

[1734] Brett T. McClintock, Gary C. White, Michael F. Antolin, and Daniel W. Tripp. Estimating abundance using mark-resight when sampling is with replacement or the number of marked individuals is unknown. Biometrics, 65(1):237–246, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bandos:2009:AUF

[1735] Andriy I. Bandos, Howard E. Rockette, Tao Song, and David Gur. Area under the free-response ROC curve (FROC) and a related summary index. Biometrics, 65(1):247–256, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Crepet:2009:UEL

[1736] Amélie Crépet, Hugo Harari-Kermadec, and Jessica Tressou. Using empirical likelihood to combine data: Application to food risk assessment. Biometrics, 65(1):257–266, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Royle:2009:ACR

[1737] J. Andrew Royle. Analysis of capture–recapture models with individual covariates using data augmentation. Biometrics, 65(1):267–274, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OHara:2009:ERB

[1738] R. B. O’Hara, S. Lampila, and M. Orell. Estimation of rates of births, deaths, and immigration from mark-recapture data. Biometrics, 65(1): 275–281, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Welty:2009:BDL

[1739] L. J. Welty, R. D. Peng, S. L. Zeger, and F. Dominici. Bayesian distributed lag models: Estimating effects of particulate matter air pollution on daily mortality. Biometrics, 65(1):282–291, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miao:2009:DEM

[1740] Hongyu Miao, Carrie Dykes, Lisa M. Demeter, and Hulin Wu. Differential equation modeling of HIV viral fitness experiments: Model identification, model selection, and multimodel inference. Biometrics, 65 (1):292–300, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

An:2009:NDS

[1741] Ming-Wen An, Constantine E. Frangakis, Beverly S. Musick, and Constantin T. Yiannoutsos. The need for double-sampling designs in survival studies: an application to monitor PEPFAR. Biometrics, 65(1):301–306, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ivanova:2009:DFC

[1742] Anastasia Ivanova and Se Hee Kim. Dose finding for continuous and ordinal outcomes with a monotone objective function: a unified approach. Biometrics, 65(1):307–315, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Proschan:2009:CT

[1743] Michael A. Proschan and Martha Nason. Conditioning in 2 × 2 tables. Biometrics, 65(1):316–322, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Glimm:2009:LEC

[1744] Ekkehard Glimm. Letter to the Editor “Comment on Qiu and Hwang (2007), Biometrics 63, 767–776”. Biometrics, 65(1):323–324, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1378].

Qiu:2009:ARF

[1745] Jing Qiu and J. T. Gene Hwang. The authors replied as follows. Biometrics, 65(1):324–325, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rotnitzky:2009:BRS

[1746] Andrea Rotnitzky. Book review: Semiparametric Theory and Missing Data, by Tsiatis, A. A. Biometrics, 65(1):326–328, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heitjan:2009:BRM

[1747] Daniel F. Heitjan. Book review: Missing Data in Longitudinal Studies: Strategies for Bayesian Modeling and Sensitivity Analysis, by Daniels, M. J. and Hogan, J. W. Biometrics, 65(1):328, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gustafson:2009:BRB

[1748] Paul Gustafson. Book review: Bayesian Biostatistics and Diagnostic Medicine, by Broemeling, L. D. Biometrics, 65(1):328–329, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DAgostino:2009:BRS

[1749] Ralph B. D’Agostino, Jr. Book review: Statistics in the Pharmaceutical Industry, 3rd edition edited by Buncher, C. R. and Tsay, J.-Y. Biometrics, 65(1):329–331, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:2009:BRI

[1750] Yulan Liang. Book review: Introduction to Machine Learning and Bioinformatics, by Mitra, S., Datta, S., Perkins, T., and Michailidis, G. Biometrics, 65(1):331, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhao:2009:BRS

[1751] Hongyu Zhao. Book review: The Statistics of Gene Mapping, by Siegmund, D. and Yakir, B. Biometrics, 65(1):331–332, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Siegmund:2009:BRD

[1752] Kimberly D. Siegmund. Book review: DNA Methylation Microarrays: Experimental Design and Statistical Analysis, by Wang, S.-C. and Petronis, A. Biometrics, 65(1):332–333, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heller:2009:BRM

[1753] Ruth Heller. Book review: Multiple Testing Procedures with Applications to Genomics, by Dudoit, S. and van der Laan, M. J. Biometrics, 65(1): 333–334, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:2009:BRS

[1754] Wesley K. Thompson. Book review: The Statistical Analysis of Functional MRI Data, by Lazar, N. A. Biometrics, 65(1):334–335, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Campbell:2009:BRS

[1755] Dave Campbell. Book review: Simulation and Inference for Stochastic Differential Equations with R Examples, by Iacus, S. M. Biometrics, 65 (1):335–336, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gomes:2009:BRN

[1756] M. Ivette Gomes. Book review: Nonparametric Analysis of Univariate Heavy-Tailed Data: Research and Practice, by Markovich, N. Biometrics, 65(1):336, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRAa

[1757] Anonymous. Book review: Analysis of Variance and Covariance: How to Choose and Construct Models for the Life Sciences, by Doncaster, C. P. and Davey, A. J. H. Biometrics, 65(1):337, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:2009:BRT

[1758] Timothy D. Johnson. Book review: Time Series Analysis with Applications in R, 2nd edition by Cryer, J. D. and Chan, K.-S. Biometrics, 65(1):337, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRCa

[1759] Anonymous. Book review: Computational Statistics Handbook with MATLAB(R), 2nd edition by Martinez, W. L. and Martinez, A. R. Biometrics, 65(1):338, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRMa

[1760] Anonymous. Book review: Medical Biostatistics, 2nd edition by Indrayan, A. Biometrics, 65(1):338, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRMb

[1761] Anonymous. Book review: Models for Probability and Statistical Inference: Theory and Applications, by Stapleton, J. H. Biometrics, 65(1): 338, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRCb

[1762] Anonymous. Book review: Computational Methods in Biomedical Research, edited by Khattree, R. and Naik, D. N. Biometrics, 65(1):338– 339, March 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maitra:2009:CPS

[1763] Ranjan Maitra and Ivan P. Ramler. Clustering in the presence of scatter. Biometrics, 65(2):341–352, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2009:CDB

[1764] Tonglin Zhang and Ge Lin. Cluster detection based on spatial associations and iterated residuals in generalized linear mixed models. Biometrics, 65(2):353–360, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2009:DRE

[1765] Xianzheng Huang. Diagnosis of random-effect model misspecification in generalized linear mixed models for binary response. Biometrics, 65 (2):361–368, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Saville:2009:TRE

[1766] Benjamin R. Saville and Amy H. Herring. Testing random effects in the linear mixed model using approximate Bayes factors. Biometrics, 65(2):369–376, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:2009:JMA

[1767] Yu Liang, Wenbin Lu, and Zhiliang Ying. Joint modeling and analysis of longitudinal data with informative observation times. Biometrics, 65 (2):377–384, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cheng:2009:NAA

[1768] Yu Cheng, Jason P. Fine, and Michael R. Kosorok. Nonparametric association analysis of exchangeable clustered competing risks data. Biometrics, 65(2):385–393, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2009:REA

[1769] T. Cai, J. Huang, and L. Tian. Regularized estimation for the accelerated failure time model. Biometrics, 65(2):394–404, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kang:2009:MHR

[1770] Sangwook Kang and Jianwen Cai. Marginal hazards regression for retrospective studies within cohort with possibly correlated failure time data. Biometrics, 65(2):405–414, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

French:2009:MMR

[1771] Benjamin French and Patrick J. Heagerty. Marginal mark regression analysis of recurrent marked point process data. Biometrics, 65(2):415– 422, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Henrys:2009:ICI

[1772] P. A. Henrys and P. E. Brown. Inference for clustered inhomogeneous spatial point processes. Biometrics, 65(2):423–430, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yan:2009:PFT

[1773] Jun Yan and Jian Huang. Partly functional temporal process regression with semiparametric profile estimating functions. Biometrics, 65(2):431– 440, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finley:2009:HSM

[1774] Andrew O. Finley, Sudipto Banerjee, Patrik Waldmann, and Tore Ericsson. Hierarchical spatial modeling of additive and dominance genetic variance for large spatial trial datasets. Biometrics, 65(2):441–451, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klingenberg:2009:TMH

[1775] B. Klingenberg, A. Solari, L. Salmaso, and F. Pesarin. Testing marginal homogeneity against stochastic order in multivariate ordinal data. Biometrics, 65(2):452–462, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:2009:EMF

[1776] Hye-Seung Lee, Myunghee Cho Paik, and Joseph H. Lee. Estimating a multivariate familial correlation using joint models for canonical correlations: Application to memory score analysis from familial Hispanic Alzheimer’s disease study. Biometrics, 65(2):463–469, June 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2009:MGL

[1777] Zehua Chen and Jianbin Liu. Mixture generalized linear models for multiple interval mapping of quantitative trait loci in experimental crosses. Biometrics, 65(2):470–477, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2009:MEH

[1778] Ying Yuan and Roderick J. A. Little. Mixed-effect hybrid models for longitudinal data with nonignorable dropout. Biometrics, 65(2):478– 486, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yuan:2009:MAS

[1779] Ying Yuan and Roderick J. A. Little. Meta-analysis of studies with missing data. Biometrics, 65(2):487–496, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Egleston:2009:ESA

[1780] Brian L. Egleston, Daniel O. Scharfstein, and Ellen MacKenzie. On estimation of the survivor average causal effect in observational studies when important confounders are missing due to death. Biometrics, 65 (2):497–504, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lin:2009:NMC

[1781] Julia Y. Lin, Thomas R. Ten Have, and Michael R. Elliott. Nested Markov compliance class model in the presence of time-varying noncompliance. Biometrics, 65(2):505–513, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sjolander:2009:SAP

[1782] Arvid Sjölander, Keith Humphreys, Stijn Vansteelandt, Rino Bellocco, and Juni Palmgren. Sensitivity analysis for principal stratum direct effects, with an application to a study of physical activity and coronary heart disease. Biometrics, 65(2):514–520, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2009:ASS

[1783] Debashis Ghosh. On assessing surrogacy in a single trial setting using a semicompeting risks paradigm. Biometrics, 65(2):521–529, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joffe:2009:RCF

[1784] Marshall M. Joffe and Tom Greene. Related causal frameworks for surrogate outcomes. Biometrics, 65(2):530–538, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brannath:2009:ECB

[1785] Werner Brannath, Cyrus R. Mehta, and Martin Posch. Exact confidence bounds following adaptive group sequential tests. Biometrics, 65(2):539– 546, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2009:CMM

[1786] Chang Xuan Mao and Na You. On comparison of mixture models for closed population capture–recapture studies. Biometrics, 65(2):547–553, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ward:2009:POD

[1787] Gill Ward, Trevor Hastie, Simon Barry, Jane Elith, and John R. Leathwick. Presence-only data and the EM algorithm. Biometrics, 65(2):554– 563, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2009:WWT

[1788] Lan Wang and Runze Li. Weighted Wilcoxon-type smoothly clipped absolute deviation method. Biometrics, 65(2):564–571, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Newman:2009:MCI

[1789] Ken B. Newman, Carmen Fernández, Len Thomas, and Stephen T. Buckland. Monte Carlo inference for state-space models of wild animal populations. Biometrics, 65(2):572–583, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Feng:2009:ATD

[1790] Rui Feng, Gongfu Zhou, Meizhuo Zhang, and Heping Zhang. Analysis of twin data using SAS. Biometrics, 65(2):584–589, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gregoire:2009:REM

[1791] Timothy G. Gregoire and Christian Salas. Ratio estimation with measurement error in the auxiliary variate. Biometrics, 65(2):590–598, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Langholz:2009:SMA

[1792] Bryan Langholz, Duncan C. Thomas, Marilyn Stovall, Susan A. Smith, John D. Boice Jr., Roy E. Shore, Leslie Bernstein, Charles F. Lynch, Xinbo Zhang, and Jonine L. Bernstein. Statistical methods for analysis of radiation effects with tumor and dose location-specific information with application to the WECARE study of asynchronous contralateral breast cancer. Biometrics, 65(2):599–608, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ritz:2009:FRA

[1793] Christian Ritz and Jens C. Streibig. Functional regression analysis of fluorescence curves. Biometrics, 65(2):609–617, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yi:2009:MRM

[1794] Grace Y. Yi and Wenqing He. Median regression models for longitudinal data with dropouts. Biometrics, 65(2):618–625, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kneib:2009:VSM

[1795] Thomas Kneib, Torsten Hothorn, and Gerhard Tutz. Variable selection and model choice in geoadditive regression models. Biometrics, 65(2): 626–634, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shardell:2009:CSS

[1796] Michelle D. Shardell and Samer S. El-Kamary. Calculating sample size for studies with expected all-or-none nonadherence and selection bias. Biometrics, 65(2):635–639, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Little:2009:CME

[1797] Roderick J. Little, Qi Long, and Xihong Lin. A comparison of methods for estimating the causal effect of a treatment in randomized clinical trials subject to noncompliance. Biometrics, 65(2):640–649, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carlson:2009:BAM

[1798] Nichole E. Carlson, Timothy D. Johnson, and Morton B. Brown. A Bayesian approach to modeling associations between pulsatile hormones. Biometrics, 65(2):650–659, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hoef:2009:BRA

[1799] Jay M. Ver Hoef. Book review: Applied Spatial Data Analysis with R. by Bivand, R. S., Pebesma, E. J., and Gomez-Rubio, V. Biometrics, 65(2):660–661, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Law:2009:BRB

[1800] J. Law. Book review: Bayesian Disease Mapping: Hierarchical Modeling in Spatial Epidemiology, by Lawson, A. B. Biometrics, 65(2):661–662, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pigeot:2009:BRD

[1801] Iris Pigeot. Book review: Disease Surveillance: a Public Health Informatics Approach, edited by Lombardo, J. S. and Buckeridge, D. L. Biometrics, 65(2):662, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Valsecchi:2009:BRS

[1802] M. G. Valsecchi. Book review: Survival Analysis for Epidemiologic and Medical Research, by Selvin, S. Biometrics, 65(2):663, June 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grambsch:2009:BRS

[1803] Patricia Grambsch. Book review: Survival and Event History Analysis: a Process Point of View, by Aalen, O. O., Borgan, O., and Gjessing, H. K. Biometrics, 65(2):663–665, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:2009:BRN

[1804] Haonan Wang. Book review: Nonlinear Dimensionality Reduction, by Lee, J. A. and Verleysen, M. Biometrics, 65(2):665, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Loughin:2009:BRM

[1805] Thomas M. Loughin. Book review: Model Selection and Model Averaging, by Claeskens, G. and Hjort, N. L. Biometrics, 65(2):666, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kaizar:2009:BRM

[1806] Eloise Kaizar. Book review: Meta-Analysis of Binary Data Using Profile Likelihood, by Böhning, D., Kuhnert, R., and Rattanasiri, S. Biometrics, 65(2):666–667, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morris:2009:BRW

[1807] Jeffrey S. Morris. Book review: Wavelet Methods in Statistics with R, by Nason, G. P. Biometrics, 65(2):667–668, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hu:2009:BRA

[1808] Feifang Hu. Book review: Adaptive Design Theory and Implementation Using SAS and R, by Chang, M. Biometrics, 65(2):668–669, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fewster:2009:BRE

[1809] Rachel M. Fewster. Book review: Ecological Models and Data in R, by Bolker, B. M. Biometrics, 65(2):669–670, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coppi:2009:BRA

[1810] Renato Coppi. Book review: Applied Multiway Data Analysis, by Kroonenberg, P. M. Biometrics, 65(2):670–671, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRSa

[1811] Anonymous. Book review: Sampling of Populations: Methods and Applications, 4th edition by Levy, P. S. and Lemeshow, S. Biometrics, 65(2): 671, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRAb

[1812] Anonymous. Book review: Applied Survival Analysis: Regression Modeling of Time-to-Event Data, 2nd edition by Hosmer, D. W., Lemeshow, S., and May, S. Biometrics, 65(2):671–672, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRAc

[1813] Anonymous. Book review: Advances in Mathematical and Statistical Modeling, edited by Arnold, B. C., Balakrishnan, N., Sarabia, J. M. and Mínguez, R. Biometrics, 65(2):672, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRSb

[1814] Anonymous. Book review: SAS for Data Analysis: Intermediate Statistical Methods, by Marasinghe, M. G. and Kennedy, W. J. Biometrics, 65 (2):672, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRI

[1815] Anonymous. Book review: An Introduction to Generalized Linear Models, 3rd edition by Dobson, A. J. and Barnett, A. G. Biometrics, 65 (2):672–673, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRD

[1816] Anonymous. Book review: Design and Analysis of Bioavailability and Bioequivalence Studies, 3rd edition, by Chow, S.-C. and Liu, J.-P. Biometrics, 65(2):673, June 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2009:IEC

[1817] Hua Chen, Zhi Geng, and Xiao-Hua Zhou. Identifiability and estimation of causal effects in randomized trials with noncompliance and completely nonignorable missing data. Biometrics, 65(3):675–682, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See discussion [1818].

Small:2009:DIE

[1818] Dylan S. Small and Jing Cheng. Discussion of “Identifiability and Estimation of Causal Effects in Randomized Trials with Noncompliance and Completely Nonignorable Missing Data”. Biometrics, 65(3):682–686, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1817].

Vansteelandt:2009:D

[1819] Stijn Vansteelandt. Discussions. Biometrics, 65(3):686–689, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2009:R

[1820] Hua Chen, Zhi Geng, and Xiao-Hua Zhou. Rejoinder. Biometrics, 65(3): 689–691, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Boos:2009:FFV

[1821] Dennis D. Boos, Leonard A. Stefanski, and Yujun Wu. Fast FSR variable selection with applications to clinical trials. Biometrics, 65(3):692–700, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maugis:2009:VSC

[1822] Cathy Maugis, Gilles Celeux, and Marie-Laure Martin-Magniette. Variable selection for clustering with Gaussian mixture models. Biometrics, 65(3):701–709, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2009:LVM

[1823] Xianzheng Huang and Joshua M. Tebbs. On latent-variable model misspecification in structural measurement error models for binary response. Biometrics, 65(3):710–718, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2009:LMR

[1824] Xianzheng Huang, Leonard A. Stefanski, and Marie Davidian. Latentmodel robustness in joint models for a primary endpoint and a longitudinal process. Biometrics, 65(3):719–727, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2009:EST

[1825] Wenqin Pan, Donglin Zeng, and Xihong Lin. Estimation in semiparametric transition measurement error models for longitudinal data. Biometrics, 65(3):728–736, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2009:SJM

[1826] Liang Li, Bo Hu, and Tom Greene. A semiparametric joint model for longitudinal and survival data with application to hemodialysis study. Biometrics, 65(3):737–745, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeng:2009:STM

[1827] Donglin Zeng and D. Y. Lin. Semiparametric transformation models with random effects for joint analysis of recurrent and terminal events. Biometrics, 65(3):746–752, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pan:2009:FED

[1828] Qing Pan and Douglas E. Schaubel. Flexible estimation of differences in treatment-specific recurrent event means in the presence of a terminating event. Biometrics, 65(3):753–761, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeIorio:2009:BNN

[1829] Maria De Iorio, Wesley O. Johnson, Peter Müller, and Gary L. Rosner. Bayesian nonparametric nonproportional hazards survival modeling. Biometrics, 65(3):762–771, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scarpa:2009:BHF

[1830] Bruno Scarpa and David B. Dunson. Bayesian hierarchical functional data analysis via contaminated informative priors. Biometrics, 65(3): 772–780, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Senturk:2009:CAC

[1831] Damla Sentürk, Danh V. Nguyen, Flora Tassone, Randi J. Hagerman, Raymond J. Carroll, and Paul J. Hagerman. Covariate adjusted correlation analysis with application to FMR1 premutation female carrier data. Biometrics, 65(3):781–792, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Telesca:2009:DEN

[1832] Donatello Telesca, Lurdes Y. T. Inoue, Mauricio Neira, Ruth Etzioni, Martin Gleave, and Colleen Nelson. Differential expression and network inferences through functional data modeling. Biometrics, 65(3):793–804, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xiao:2009:IDD

[1833] Guanghua Xiao, Cavan Reilly, and Arkady B. Khodursky. Improved detection of differentially expressed genes through incorporation of gene locations. Biometrics, 65(3):805–814, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Scherag:2009:FDG

[1834] André Scherag, Johannes Hebebrand, Helmut Schäfer, and Hans-Helge Müller. Flexible designs for genomewide association studies. Biometrics, 65(3):815–821, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tzeng:2009:GTS

[1835] Jung-Ying Tzeng, Daowen Zhang, Sheng-Mao Chang, Duncan C. Thomas, and Marie Davidian. Gene-trait similarity regression for multimarker-based association analysis. Biometrics, 65(3):822–832, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wright:2009:IGU

[1836] Janine A. Wright, Richard J. Barker, Matthew R. Schofield, Alain C. Frantz, Andrea E. Byrom, and Dianne M. Gleeson. Incorporating genotype uncertainty into mark-recapture-type models for estimating abundance using DNA samples. Biometrics, 65(3):833–840, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sutherland:2009:MMC

[1837] Jason M. Sutherland, Pete Castelluccio, and Carl James Schwarz. A multilevel model for continuous time population estimation. Biometrics, 65(3):841–849, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cangelosi:2009:MBS

[1838] Amanda R. Cangelosi and Mevin B. Hooten. Models for bounded systems with continuous dynamics. Biometrics, 65(3):850–856, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kaiser:2009:EDD

[1839] Mark S. Kaiser and Petruta C. Caragea. Exploring dependence with data on spatial lattices. Biometrics, 65(3):857–865, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yin:2009:LCT

[1840] Guosheng Yin and Ying Yuan. A latent contingency table approach to dose finding for combinations of two agents. Biometrics, 65(3):866–875, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ji:2009:ARM

[1841] Yuan Ji and B. Nebiyou Bekele. Adaptive randomization for multiarm comparative clinical trials based on joint efficacy/toxicity outcomes. Biometrics, 65(3):876–884, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Follmann:2009:CLT

[1842] Dean Follmann, Michael P. Fay, and Michael Proschan. Chop-lump tests for vaccine trials. Biometrics, 65(3):885–893, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tian:2009:ISW

[1843] L. Tian, T. Cai, and L. J. Wei. Identifying subjects who benefit from additional information for better prediction of the outcome variables. Biometrics, 65(3):894–902, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:2009:TDA

[1844] Hae-Young Kim and Michael G. Hudgens. Three-dimensional arraybased group testing algorithms. Biometrics, 65(3):903–910, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhang:2009:EPS

[1845] Guangyu Zhang and Roderick Little. Extensions of the penalized spline of propensity prediction method of imputation. Biometrics, 65(3):911–918, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zhu:2009:AET

[1846] Hongtu Zhu, Haibo Zhou, Jiahua Chen, Yimei Li, Jeffrey Lieberman, and Martin Styner. Adjusted exponentially tilted likelihood with applications to brain morphology. Biometrics, 65(3):919–927, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hooker:2009:FFD

[1847] Giles Hooker. Forcing function diagnostics for nonlinear dynamics. Biometrics, 65(3):928–936, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenblum:2009:URM

[1848] Michael Rosenblum and Mark J. van der Laan. Using regression models to analyze randomized trials: Asymptotically valid hypothesis tests despite incorrectly specified models. Biometrics, 65(3):937–945, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

He:2009:CIC

[1849] Yulei He and Alan M. Zaslavsky. Combining information from cancer registry and medical records data to improve analyses of adjuvant cancer therapies. Biometrics, 65(3):946–952, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buzoianu:2009:OBD

[1850] Manuela Buzoianu and Joseph B. Kadane. Optimal Bayesian design for patient selection in a clinical study. Biometrics, 65(3):953–961, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Millar:2009:CHB

[1851] Russell B. Millar. Comparison of hierarchical Bayesian models for overdispersed count data using DIC and Bayes’ factors. Biometrics, 65(3):962–969, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luo:2009:BIS

[1852] Sheng Luo, Ciprian M. Crainiceanu, Thomas A. Louis, and Nilanjan Chatterjee. Bayesian inference for smoking cessation with a latent cure state. Biometrics, 65(3):970–978, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kawaguchi:2009:AEB

[1853] Atsushi Kawaguchi, Gary G. Koch, and Ratna Ramaswamy. Applications of extensions of bivariate rank sum statistics to the crossover design to compare two treatments through four sequence groups. Biometrics, 65(3):979–988, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hosmer:2009:BRM

[1854] David W. Hosmer. Book review: Multivariable Model-building: a Pragmatic Approach to Regression Analysis Based on Fractional Polynomials for Modelling Continuous Variables, by Royston, P. and Sauerbrei, W. Biometrics, 65(3):989–990, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghosh:2009:BRM

[1855] Debashis Ghosh. Book review: Modern Multivariate Statistical Techniques: Regression, Classification, and Manifold Learning, by Izenman, A. J. Biometrics, 65(3):990–991, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Srivastava:2009:BRP

[1856] Muni S. Srivastava. Book review: A Primer on Linear Models, by Monahan, J. F. Biometrics, 65(3):991–992, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schwarz:2009:BRM

[1857] Carl James Schwarz. Book review: Mixed Effects Models and Extensions in Ecology with R, by Zuur, A. F., Ieno, E. N., Walker, N. J., Saveliev, A. A., and Smith, G. M. Biometrics, 65(3):992–993, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rice:2009:BRR

[1858] Kenneth Rice. Book review: Random Effect and Latent Variable Model Selection, edited by Dunson, D. B. Biometrics, 65(3):993, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dubnicka:2009:BRM

[1859] Suzanne R. Dubnicka. Book review: Missing Data in Clinical Studies, by Molenberghs, G. and Kenward, M. G. Biometrics, 65(3):993–994, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nielsen:2009:BRF

[1860] Jason D. Nielsen. Book review: Functional and Operatorial Statistics, edited by Dabo-Niang, S. and Ferraty, F. Biometrics, 65(3):994–995, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moller:2009:BRS

[1861] Jesper Møller. Book review: Statistical Analysis and Modelling of Spatial Point Patterns, by Illian, J., Penttinen, A., Stoyan, H., and Stoyan, D. Biometrics, 65(3):995, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christensen:2009:BRS

[1862] Ole F. Christensen. Book review: Solved Problem in Geostatistics, by Leuangthong, O., Khan, K. D., and Deutsch, C. V. Biometrics, 65(3): 995–996, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buckeridge:2009:BRS

[1863] David Buckeridge. Book review: Statistical Methods for Environmental Epidemiology with R: a Case Study in Air Pollution and Health, by Peng, R. D. and Dominici, F. Biometrics, 65(3):996–997, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heiser:2009:BRC

[1864] Willem Heiser. Book review: Correspondence Analysis in Practice, by Greenacre, M. Biometrics, 65(3):997–998, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanderWeele:2009:BRI

[1865] Tyler J. VanderWeele. Book review: Introduction to Statistical Mediation Analysis, by MacKinnon, D. P. Biometrics, 65(3):998–1000, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRE

[1866] Anonymous. Book review: The EM Algorithm and Extensions, by McLachlan, G. J. and Krishnan, T. Biometrics, 65(3):1000, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRP

[1867] Anonymous. Book review: Probability Models for DNA Sequence Evolution, by Durrett, R. Biometrics, 65(3):1000, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRB

[1868] Anonymous. Book review: Bayesian Methods for Data Analysis, by Carlin, B. P. and Louis, T. A. Biometrics, 65(3):1000–1001, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRCc

[1869] Anonymous. Book review: Cancer Mortality and Morbidity Patterns in the U.S. Population: an Interdisciplinary Approach, by Manton, K. G., Akushevich, I., and Kravchenko, J. Biometrics, 65(3):1001, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRN

[1870] Anonymous. Book review: Nonlinear Regression with R, by Ritz, C. and Streibig, J. C. Biometrics, 65(3):1001, September 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kipnis:2009:MDE

[1871] Victor Kipnis, Douglas Midthune, Dennis W. Buckman, Kevin W. Dodd, Patricia M. Guenther, Susan M. Krebs-Smith, Amy F. Subar, Janet A. Tooze, Raymond J. Carroll, and Laurence S. Freedman. Modeling data with excess zeros and measurement error: Application to evaluating relationships between episodically consumed foods and health outcomes. Biometrics, 65(4):1003–1010, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xie:2009:LMD

[1872] Minge Xie, Qiankun Sun, and Joseph Naus. A latent model to detect multiple clusters of varying sizes. Biometrics, 65(4):1011–1020, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pang:2009:SBD

[1873] Herbert Pang, Tiejun Tong, and Hongyu Zhao. Shrinkage-based diagonal discriminant analysis and its applications in high-dimensional data. Biometrics, 65(4):1021–1029, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2009:BVS

[1874] Wei Chen, Debashis Ghosh, Trivellore E. Raghunathan, and Daniel J. Sargent. Bayesian variable selection with joint modeling of categorical and survival outcomes: an application to individualizing chemotherapy treatment in advanced colorectal cancer. Biometrics, 65(4):1030–1040, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2009:MIS

[1875] Lei Xu, Timothy D. Johnson, Thomas E. Nichols, and Derek E. Nee. Modeling inter-subject variability in fMRI activation location: a Bayesian hierarchical spatial model. Biometrics, 65(4):1041–1051, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cao:2009:MAN

[1876] Jing Cao, Chong Z. He, Kimberly M. Suedkamp Wells, Joshua J. Millspaugh, and Mark R. Ryan. Modeling age and nest-specific survival using a hierarchical Bayesian approach. Biometrics, 65(4):1052–1062, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mao:2009:CSA

[1877] Chang Xuan Mao and Jun Li. Comparing species assemblages via species accumulation curves. Biometrics, 65(4):1063–1067, December 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yap:2009:NML

[1878] John Stephen Yap, Jianqing Fan, and Rongling Wu. Nonparametric modeling of longitudinal covariance structure in functional mapping of quantitative trait loci. Biometrics, 65(4):1068–1077, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thurston:2009:BMM

[1879] Sally W. Thurston, David Ruppert, and Philip W. Davidson. Bayesian models for multiple outcomes nested in domains. Biometrics, 65(4):1078– 1086, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gelfond:2009:BHM

[1880] Jonathan A. L. Gelfond, Mayetri Gupta, and Joseph G. Ibrahim. A Bayesian hidden Markov model for Motif discovery through joint modeling of genomic sequence and ChIP-Chip data. Biometrics, 65(4):1087– 1095, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Li:2009:THW

[1881] Yan Li and Barry I. Graubard. Testing Hardy-Weinberg equilibrium and homogeneity of Hardy-Weinberg disequilibrium using complex survey data. Biometrics, 65(4):1096–1104, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2009:FMB

[1882] Lu Chen, Li Hsu, and Kathleen Malone. A frailty-model-based approach to estimating the age-dependent penetrance function of candidate genes using population-based case-control study designs: an application to data on the BRCA1 gene. Biometrics, 65(4):1105–1114, December 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Joo:2009:RGW

[1883] Jungnam Joo, Minjung Kwak, Kwangmi Ahn, and Gang Zheng. A robust genome-wide scan statistic of the Wellcome Trust Case-Control Consortium. Biometrics, 65(4):1115–1122, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Costain:2009:BPM

[1884] Deborah A. Costain. Bayesian partitioning for modeling and mapping spatial case-control data. Biometrics, 65(4):1123–1132, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huang:2009:PRM

[1885] Y. Huang and M. S. Pepe. A parametric ROC model-based approach for evaluating the predictiveness of continuous markers in case-control studies. Biometrics, 65(4):1133–1144, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2009:PLC

[1886] Huiping Xu and Bruce A. Craig. A probit latent class model with general correlation structures for evaluating accuracy of diagnostic tests. Biometrics, 65(4):1145–1155, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sexton:2009:SAB

[1887] Joseph Sexton and Petter Laake. Stochastic approximation boosting for incomplete data problems. Biometrics, 65(4):1156–1163, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shi:2009:LIG

[1888] Xiaoyan Shi, Hongtu Zhu, and Joseph G. Ibrahim. Local influence for generalized linear models with missing covariates. Biometrics, 65(4): 1164–1174, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Xu:2009:ELP

[1889] Lei Xu and Jun Shao. Estimation in longitudinal or panel data models with random-effect-based missing responses. Biometrics, 65(4):1175– 1183, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2009:EPL

[1890] Yanyan Liu, Haibo Zhou, and Jianwen Cai. Estimated pseudopartiallikelihood method for correlated failure time data with auxiliary covariates. Biometrics, 65(4):1184–1193, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DiRienzo:2009:FRM

[1891] A. Gregory DiRienzo. Flexible regression model selection for survival probabilities: With application to AIDS. Biometrics, 65(4):1194–1202, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jung:2009:RQR

[1892] Sin-Ho Jung, Jong-Hyeon Jeong, and Hanna Bandos. Regression on quantile residual life. Biometrics, 65(4):1203–1212, December 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ozturk:2009:ECV

[1893] Omer Ozturk and N. Balakrishnan. An exact control-versus-treatment comparison test based on ranked set samples. Biometrics, 65(4):1213– 1222, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hudgens:2009:AVE

[1894] Michael G. Hudgens and Peter B. Gilbert. Assessing vaccine effects in repeated low-dose challenge experiments. Biometrics, 65(4):1223–1232, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Slaughter:2009:BLV

[1895] James C. Slaughter, Amy H. Herring, and John M. Thorp. A Bayesian latent variable mixture model for longitudinal fetal growth. Biometrics, 65(4):1233–1242, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:2009:BWS

[1896] Shengde Liang, Sudipto Banerjee, and Bradley P. Carlin. Bayesian wombling for spatial point processes. Biometrics, 65(4):1243–1253, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Czado:2009:PMA

[1897] Claudia Czado, Tilmann Gneiting, and Leonhard Held. Predictive model assessment for count data. Biometrics, 65(4):1254–1261, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Maiti:2009:BRS

[1898] Tapabrata Maiti and Vivek Pradhan. Bias reduction and a solution for separation of logistic regression with missing covariates. Biometrics, 65 (4):1262–1269, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:2009:GTR

[1899] Peng Chen, Joshua M. Tebbs, and Christopher R. Bilder. Group testing regression models with fixed and random effects. Biometrics, 65(4):1270– 1278, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liu:2009:ANN

[1900] W. Liu, F. Bretz, A. J. Hayter, and H. P. Wynn. Assessing nonsuperiority, noninferiority, or equivalence when comparing two regression models over a restricted covariate region. Biometrics, 65(4):1279–1287, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kwak:2009:RTT

[1901] Minjung Kwak, Jungnam Joo, and Gang Zheng. A robust test for twostage design in genome-wide association studies. Biometrics, 65(4):1288– 1295, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeSantis:2009:LCM

[1902] Stacia M. DeSantis, E. Andrés Houseman, Brent A. Coull, David N. Louis, Gayatry Mohapatra, and Rebecca A. Betensky. A latent class model with hidden Markov dependence for array CGH data. Biometrics, 65(4):1296–1305, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goad:2009:BRD

[1903] Carla L. Goad. Book review: Design of Comparative Experiments (Cambridge Series in Statistical and Probabilistic Mathematics) by Bailey, R. A. Biometrics, 65(4):1306–1307, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Campbell:2009:BRC

[1904] Michael J. Campbell. Book review: Cluster Randomised Trials, by Hayes, R. J. and Moulton, L. H. Biometrics, 65(4):1307–1308, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cai:2009:BRB

[1905] Bo Cai. Book review: Bayesian Evaluation of Informative Hypotheses, edited by Hoijtink, H., Klugkist, I., and Boelen, P. A. Biometrics, 65(4): 1308–1309, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hamilton:2009:BRB

[1906] Cody Hamilton. Book review: Bayesian Methods for Measures of Agreement, by Broemeling, L. Biometrics, 65(4):1309, December 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yasui:2009:BRS

[1907] Yutaka Yasui and Xiaoming Wang. Book review: Statistical Learning from a Regression Perspective, by Berk, R. A. Biometrics, 65(4):1309– 1310, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ugarte:2009:BRS

[1908] Maria Dolores Ugarte. Book review: Statistical Detection and Surveillance of Geographic Clusters, by Rogerson, P. and Yamada, I. Biometrics, 65(4):1310–1311, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shepherd:2009:BRG

[1909] Bryan E. Shepherd. Book review: Global Sensitivity Analysis. The Primer, by Saltelli, A., Ratto, M., Andres, T., Campolongo, F., Cariboni, J., Gatelli, D., Saisana, M., And Tarantola, S. Biometrics, 65(4): 1311–1312, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Guedon:2009:BRS

[1910] Yann Guédon. Book review: Semi-Markov Chains and Hidden Semi-Markov Models toward Applications: Their Use in Reliability and DNA Analysis, by Barbu, V. S. and Limnios, N. Biometrics, 65(4):1312, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRH

[1911] Anonymous. Book review: Handbook of Statistical Analyses Using SAS, 3rd edition by Der, G. and Everitt, B. S. Biometrics, 65(4):1312–1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRRa

[1912] Anonymous. Book review: Recent Advances in Linear Models and Related Areas: Essays in Honour of Helge Toutenburg, edited by Shalabh and Heumann, C. Biometrics, 65(4):1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRRb

[1913] Anonymous. Book review: R for SAS and SPSS Users, by Muenchen, R. A. Biometrics, 65(4):1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRSc

[1914] Anonymous. Book review: Software for Data Analysis: Programming with R, by Chambers, J. M. Biometrics, 65(4):1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRMc

[1915] Anonymous. Book review: Medical Statistics from A to Z, 2nd edition by Everitt, B. Biometrics, 65(4):1313–1314, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:CGS

[1916] Anonymous. Correction to “Genetic Susceptibility to Prostate, Breast and Colorectal Cancer among Nordic Twins,” by Stuart G. Baker, Paul Lichtenstein, Jaakko Kaprio and Niels Holm; 61, 55–63, December 2005. Biometrics, 68(3):993, September 2012. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [833].

Bove:2013:CCB

[1917] Daniel Sabanés Bové and Leonhard Held. Comment on Cai and Betensky (2003), On the Poisson Approximation for Hazard Regression. Biometrics, 69(3):795, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [574, 1918].

Cai:2013:RSB

[1918] Tianxi Cai and Rebecca Betensky. Reply to Sabanés Bové and Held’s “Comment on Cai and Betensky (2003), On the Poisson Approximation for Hazard Regression”. Biometrics, 69(3):796, September 2013. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [1917].

Lunardon:2020:CWA

[1919] Nicola Lunardon and Giovanna Menardi. Comment on “Wang et al. (2005), Robust estimating functions and bias correction for longitudinal data analysis”. Biometrics, 76(3):1040–1042, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [934].

Wang:2020:RCW

[1920] You-Gan Wang, Xu Lin, and Min Zhu. Rejoinder to “Comment on ‘Wang et al. (2005), Robust estimating functions and bias correction for longitudinal data analysis’ by Nicola Lunardon and Giovanna Menardi”. Biometrics, 76(3):1043–1044, September 2020. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [934].